### 11 May 2026 v1

# Well-posedness of a generalized Stokes operator on domains with cylindrical ends via layer-potentials

Abstract.
We study the generalized Stokes operator
$\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}\mathrel{\,:=\,}\left(\begin{array}[]{ccc}\boldsymbol{L}+V&\nabla\\ \nabla^{*}&-V_{0}\end{array}\right)$
on a domain with straight cylindrical ends $\Omega$ using the method of layer potentials on a larger manifold with straight cylindrical ends $M$, $\Omega\subset M$. The operator $\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}$ recovers the classical Stokes operator $\boldsymbol{\Xi}_{0,0}$ when the potentials $V$ and $V_{0}$ vanish. Under suitable positivity assumptions at infinity on $V$ and $V_{0}$, we prove that $\boldsymbol{\Xi}$ is Fredholm on $L^{2}(M;TM\oplus\mathbb{C})$. This allows us then to define the single- and double-layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$. Under further positivity assumptions at infinity, we prove that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are also Fredholm. Under slightly stronger assumptions on $V$ and $V_{0}$ (including positivity everywhere), we prove the invertibility of the operators $\boldsymbol{\Xi}$, $\boldsymbol{S}$, and $\frac{1}{2}+\boldsymbol{K}$. As it is well-known, the invertibility of these operators leads to well-posedness results for the associated (linear) Stokes boundary value problem with Dirichlet boundary conditions on domains with straight cylindrical ends. The proofs of these results required us to develop many other related tools. In particular, we develop an “algebra tool kit” to deal with limit and jump relations of layer potential operators, in general, and on manifolds with straight cylindrical ends, in particular. We do that first on $\mathbb{R}^{n}$, then on (possibly) non-compact manifolds, and then we deal with the specific case of manifolds with straight cylindrical ends. We use these results to study the limit and jump relations of the layer potential operators associated to our generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ on manifolds with cylindrical ends. We also develop Green formulas and energy estimates for our generalized Stokes operator $\boldsymbol{\Xi}$ on manifolds with straight cylindrical ends, which requires, in particular, a careful geometric study of the related differential operators, such as the deformation operator $\operatorname{Def}\,$. We also use suitable classes of pseudodifferential operators on manifolds with straight cylindrical ends that were studied in some previous papers of ours (including “The Stokes operator on manifolds with cylindrical ends,” J. Diff. Equations, 2024). For completeness, we include a review of the needed results on these pseudodifferential operators. As an application of all these results, we prove the well-posedness result for the Dirichlet problem for the generalized Navier-Stokes system with small data on a domain with cylindrical ends. We expect our results to have applications to the study of Navier-Stokes equations on domains with conical points.
Key words and phrases:
The Stokes operator; manifolds with straight cylindrical ends; Sobolev spaces; Stokes layer potentials; Deformation operator; Dirichlet problem.
2000 Mathematics Subject Classification:
Primary 35R01; Secondary 76M
M.K. has been partially supported by AGC35124/31.10.2018. V.N. has been partially supported by ANR-14-CE25-0012-01.
1. Introduction
We prove the well-posedness of the Dirichlet boundary value problem for a generalized Stokes system (1.4) on a domain with straight cylindrical ends $\Omega$ using the method of layer potentials. In the process, we obtain many other results of independent interest on the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (1.3), on related differential operators, and on their associated layer potentials. We carefully develop the needed concepts and methods and prove in detail our results. Before stating in more detail our main results, let us introduce the setting of our paper.
1.1. The setting of our well-posedness results
We will work on manifolds and domains with straight cylindrical ends (see the next figure, Figure 1, as well as Figure 2 at the beginning of Section 6).
Figure 1. A boundaryless manifold with straight cylindrical ends.
More precisely, by a manifold with straight cylindrical ends $M$ we will mean a non-compact Riemannian manifold without boundary (i.e., boundaryless) isometric to one of the form
(1.1) $M\,=\,M_{0}\cup\big[M^{\prime}\times(-\infty,R_{M}]\big]\,$,
where
  • $(M_{0},g_{0})$ is a compact Riemannian manifold,
  • $M^{\prime}:=\partial M_{0}\neq\emptyset$ has induced metric $g_{\partial}$ and is identified with $M^{\prime}\times\{0\}$, and
  • the cylindrical end $M^{\prime}\times(-\infty,R_{M})$ is endowed with the product metric $g=g_{\partial}+(dt)^{2}$, $t\in(-\infty,R_{M})$,
(see, for instance, [63, 67, 68]). There is no loss of generality to assume that $M$ is connected, so we will do that throughout the paper, in order to simplify our statements. (Also, we may assume $R_{M}=0$, which we will often do.)
Second, $\Omega$ will be a smooth domain with straight cylindrical ends in $M$, in the sense that
(1.2) $\Omega\cap\big[M^{\prime}\times(-\infty,R_{\Omega})\big]\,=\,\Omega^{\prime}\times(-\infty,R_{\Omega})$
(up to an isometry), for some $R_{\Omega}\leq R_{M}$, and its boundary $\Gamma:=\partial\Omega$ is a smooth manifold. (So $\Gamma$ will also be a manifold with straight cylindrical ends, see Figure 2 at the beginning of Section 6.) Again, there is no loss of generality to assume that $\Omega$ is connected, so we will do that throughout the paper.
We will also assume throughout this paper that $\Omega$ is on one side of its boundary, meaning that $\Gamma$ is the boundary of $\Omega_{-}:=M\smallsetminus\overline{\Omega}$ as well (see Assumption 2.7). This assumption is useful when studying boundary value problems on $\Omega$ using layer potentials. This implies that $\Omega^{\prime}$ is also on one side of its boundary $\Gamma^{\prime}:=\partial\Omega^{\prime}\subset M^{\prime}$.
The generalized Stokes operators $\boldsymbol{\Xi}_{V,V_{0}}$ are defined as follows (see Section 2 for more details, as well as for the concepts not recalled here). Let
$\operatorname{Def}\,:\mathcal{C}^{\infty}(M;TM)\to\mathcal{C}^{\infty}(M;T^{*}M\otimes T^{*}M)$
be the deformation operator, (see formula (2.1) and the discussion around it for more details), and $\boldsymbol{L}:=2\operatorname{Def}^{*}\,\operatorname{Def}\,$ be the “Deformation Laplacian.” Let $\operatorname{End}(E)$ be the set of linear maps $E\to E$ (the endomorphisms of $E$). For suitable “potentials” $V\in\mathcal{C}^{\infty}(M;\operatorname{End}(TM))$ and $V_{0}\in\mathcal{C}^{\infty}(M)$, we consider the operator
(1.3) $\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}\mathrel{\,:=\,}\left(\begin{array}[]{ccc}\boldsymbol{L}+V&\nabla\\ \nabla^{*}&-V_{0}\end{array}\right):\mathcal{C}^{\infty}(M;TM\oplus\mathbb{C})\to\mathcal{C}^{\infty}(M;TM\oplus\mathbb{C})\,.$
Let
$U\mathrel{\,:=\,}\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)\in L^{2}(\Omega;TM\oplus\mathbb{C})\,$,
so that $\boldsymbol{u}$ is the “vector part” of $U$. We then consider the Dirichlet boundary value problem
(1.4) $\begin{cases}\boldsymbol{\Xi}U\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}U\,=\,0&\mbox{ in }\Omega\\ \boldsymbol{u}\,=\,\boldsymbol{h}&\mbox{ on }\partial\Omega\,.\end{cases}$
When $V$ and $V_{0}$ vanish identically on $M$, this is nothing but the classical Dirichlet problem for the Stokes operator.
1.2. Main results
Our main results are for manifolds with straight cylindrical ends $M$ and for compatible open subsets $\Omega\subset M$ (with boundary $\Gamma$, as above) or for closed manifolds (i.e., compact, smooth, boundaryless manifolds). The case of closed manifolds was treated in great detail in [69], so we will largely ignore this case in this paper (but we will recall the results for the closed case in Subsection 12.2 following [69]). Let us thus assume in this introduction that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a manifold with straight cylindrical ends (see 1.1 or Definition 3.11); this will be the case for the most part of our paper. Our first main results are to prove the invertibility on $\Gamma$ of the single layer and double layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$, associated to suitable generalizations $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ of the Stokes operator $\boldsymbol{\Xi}_{0,0}$, which will be defined shortly, where $V$ and $V_{0}$ are suitable non-negative functions, see Theorems 10.4 and 10.8. (We usually denote the identity operator $\mathbb{I}$ with 1, we thus write $\frac{1}{2}+\boldsymbol{K}$ instead of $\frac{1}{2}\mathbb{I}+\boldsymbol{K}$.) We also obtain, for suitable potentials $V$ and $V_{0}$ satisfying slightly weaker assumptions, that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are Fredholm.
In order to define our layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$, we first prove that $\boldsymbol{\Xi}$ is Fredholm under suitable positivity assumptions at infinity for $V$ and $V_{0}$. We prove that it ($\boldsymbol{\Xi}$) is even invertible if the positivity conditions hold everywhere (Theorem 8.4). We use these results to define the single and double layer potential operators $\mathcal{S}_{\rm{ST}}$, $\boldsymbol{S}$, $\mathcal{D}_{\rm{ST}}$, and $\boldsymbol{K}$ associated to the Dirichlet boundary value problem for $\boldsymbol{\Xi}$. We prove that they satisfy the usual mapping and “jump” properties and that they have the usual principal symbols (Theorem 8.16; but in order to obtain this result, we develop an extensive theory: energy and representation formulas, geometric representations and formulas for our operators, a “took-kit” for proving jump relations, and many other results of independent interest). Then, using the jump relations and principal symbol properties of the layer potential operators, we prove that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are invertible. As is well known, this result implies the solvability of the Dirichlet problem (1.4), Theorem 1.1 next, which is the main result of this paper. We write $v\succ 0$ for a non-negative function that does not vanish identically on any connected component of its domain (see Definition 2.13). The space $\mathcal{C}^{\infty}_{\operatorname{inv}}(M)$ consists of functions (or sections of a vector bundle in case of $\mathcal{C}^{\infty}_{\operatorname{inv}}(M;E)$) that are translation invariant in a neighborhood of invinity (Definition 3.14)
Theorem 1.1.
Let $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ be a manifold with straight cylindrical ends, let $\Omega\subset M$ be a compatible smooth domain (i.e., with straight cylindrical ends, Equation (1.2)) that is on one side of its boundary $\Gamma:=\partial\Omega\neq\emptyset$ (Assumption 2.7), and let $m\in\mathbb{Z}_{+}$. Let $V\in\mathcal{C}^{\infty}_{\operatorname{inv}}(M;TM)$ and $V_{0}\in\mathcal{C}^{\infty}_{\operatorname{inv}}(M)$ be such that:
  1. (i)
    $V\geq 0$ and $V_{0}\geq 0$ on $M$,
  2. (ii)
    $V_{0}\succ 0$ on $\Omega$,
  3. (iii)
    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, and
  4. (iv)
    if $\Omega_{0}$ is a connected component of $M^{\prime}$ contained in $\Omega^{\prime}$, then $\color[rgb]{1,0,1}\widehat{V}\not\equiv 0$ on $\Omega_{0}$.
Then, there exists $C_{m}\geq 0$ such that, for any $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$, the homogeneous Dirichlet problem (11.1) has a unique solution $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ and this solution satisfies
$\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}+\|p\|_{H^{m}(\Omega)}\leq C_{m}\|\boldsymbol{f}\|_{H^{m+1/2}(\Gamma;TM)}\,.$
The proof of this theorem is provided in Subsection 11.1. At this time, a proof in general of this theorem using energy methods (i.e., the Riesz representation formula) seems out of reach. We also obtain applications of this well-posedness results to the Dirichlet-to-Neumann operator and to the solvability of the stationary Navier-Stokes system on domains with straight cylindrical ends of small dimension.
We have outlined here the most important results. However, in this paper, we obtain many other results of independent interest that lead to the main results. These other results are discussed in much more detail in the part on the contents of the paper, which is next.
1.3. Contents of the paper
Our paper consists of 12.1 sections (including an introduction), each divided into subsections. The assumptions of this paper (especially on the manifold $M$ on which we work) change from section to section. These assumptions are always stated at the beginning of that section and are reminded regularly, especially at the beginning of the subsections and in the statement of the most important results. (Note, however, that, beginning with Section 5, we work almost exclusively on manifolds with straight cylindrical ends.)
The first section is this introduction. Section 2 is organized as follows. Its first subsection, Subsection 2.1, is devoted to the review of some classical differential operators that appear in the definition and in the study of the Stokes operator $\boldsymbol{\Xi}=\boldsymbol{\Xi}_{V,V_{0}}$ on a Riemannian manifold (recall that $\boldsymbol{\Xi}_{V,V_{0}}$ was introduced in Equation (1.3)), such as the deformation operator $\operatorname{Def}\,$ and the “traction” $\boldsymbol{T}_{\boldsymbol{\nu}}$. In Subsection 2.2, we recall an abstract integration by parts formula, Proposition 2.9. This general formula is used in Subsection 2.3, where we obtain Green-type formulas for the Stokes operator $\boldsymbol{\Xi}$, as well as some representation formulas and energy estimates. In this section, we work on a general Riemannian manifold $M$.
The third section is devoted to some basic preliminary results such as the concepts of traces and normal lateral limits, some useful Fourier transforms, the basic definitions of symbols and pseudodifferential operators. In particular, they explain in detail the main “Fourier transform calculation” that is the prototypical “jump/limit result”, see Lemma 3.3 and Corollary 3.4. These two results ultimately are responsible and explain all the other jump/limit results. These results are formulated on $\mathbb{R}^{n}$. However, in the last subsection of this section, we work on manifolds with straight cylindrical ends. In this subsection, we recall the basic definitions related to manifolds with cylindrical ends and their compatible vector bundles, differential operators, and Sobolev spaces. Although the results of this section are classical, our presentation is new. (More references and more related background material can be found in our book [63].)
In the forth section, we discuss the jump relations for pseudodifferential operators $a(x,D)$ at the hyperplane $\{x_{n}=0\}$ in $\mathbb{R}^{n}$. This is a long and technical section in which we perform in great detail all the necessary local calculations needed to obtain the jump and limit relations that are so important for the study of layer potentials. We first establish the jump relations for operators of order m<1m<-1 (the “jump” in this case is zero, so these are rather “limit” results). We prove our results for operators with Kohn-Nirenberg type symbols. We then establish the jump relations and the other usual results for operators of order 1-1 with homogeneous, odd principal symbol. The presentation expands and details the one in our book [63]. For instance, in this paper, we do not restrict ourselves to classical pseudodifferential operators (however, for operators of order 1-1, we give our results directly in the case of an odd symbol, the only one needed in our paper). Our presentation in this section provides a complete, yet concise, new introduction to the subject of limit and jump relations for potential operators on a half-space. They are the first step in the development of an “algebra tool kit” to study the limit/jump relations. These results, together with those in Section 6 provide a short–but complete–introduction to jump relations on manifolds with cylindrical ends. (Results for manifolds with bounded geometry can be found in [69].)
The setting of our paper, up to this point, was on $\mathbb{R}^{n}$ or on a general Riemannian manifold (except Subsection 3.4, where we defined manifolds with straight cylindrical ends). Beginning with the next section, Section 5, however, we will work almost exclusively on manifolds with straight cylindrical ends. (Sometimes, we need to consider the particular case of straight cylinders.) In particular, Section 5 is devoted to pseudodifferential operators on manifolds with straight cylindrical ends. Here we introduce the main two classes of pseudodifferential operators used in this paper. The differential operators considered in this paper will belong to the first class, while their parametrices and layer potential operators will belong to the second of these classes. More precisely, in the first half of this section, we recall the definition and main properties of the $\operatorname{inv}$-calculus of translation invariant operators in a neighborhood of infinity, and, in the second half we recall the definition and the main properties of the $\operatorname{ess}$-calculus introduced in [67] (see also [63, 68]). This calculus contains the $\operatorname{inv}$-calculus, but it has further important properties due to the fact that it contains the parametrices of its operators (i.e., it is spectrally invariant). Since this material is well-known, the proofs are short. However, all the necessary definitions and properties are discussed in great detail. In particular, we introduce the “limit operator” ${\mathcal{R}_{\infty}}(P)$ associated to an operator $P$ in any of our two calculi. The limit operator ${\mathcal{R}_{\infty}}(P)$ is a translation invariant operator on the infinite cylinder $M^{\prime}\times\mathbb{R}$ associated to our manifold with straight cylindrical ends $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$. The concept of limit operator is crucial for the study of operators on manifolds with cylindrical ends. We also introduce the indicial family $\widehat{P}(\tau)$, which is the Fourier transform of ${\mathcal{R}_{\infty}}(P)$. See [47, 84, 94, 124] for the closely related $b$-calculus.
In the sixth section, we extend the “jump/limit’ results of the forth section to the case of Riemannian manifolds with straight cylindrical ends. We thus study here the normal lateral limits on manifolds with straight cylindrical ends. To this end, we develop an “algebra tool kit” to study the limit/jump relations. We prove, for example, that normal lateral limits are compatible with limits at infinity (i.e., with the limit operator map $\mathcal{R}_{\infty}$}), a result that was not published before in a refereed journal (to our knowledge), see Theorem 6.10. This result is key in the study of the layer potentials in the remaining sections. In order to define the normal lateral limits on manifolds, we need to develop some geometric tools, including the normal geodesics, and to make some corresponding geometric assumptions, which we prove are satisfied for our manifolds with cylindrical ends. Our generalized Stokes operator $\boldsymbol{\Xi}$ is a differential operator of order two, but is not elliptic in the usual sense. In order to use the theory of elliptic operators, we recall the definition of Douglis-Nirenberg-operators and ellipticity in this setting in the last subsection of section. The normal lateral limit results extend right away to Douglis-Nirenberg-operators. The content of this section is motivated by the need to establish the Douglis-Nirenberg-ellipticity of the Stokes operator $\boldsymbol{\Xi}$. Our results extend those in [63, 69].
The seventh section is devoted to the study of some of the main technical results on a cylinder. We begin by studying the form of our differential operators ($\boldsymbol{\Xi}$, $\operatorname{Def}\,$, and $\nabla$) on a cylinder $\mathfrak{S}\times\mathbb{R}$ (which is a particular case of a manifold with cylindrical ends), where $\mathfrak{S}$ is a boundaryless, smooth, compact manifold (i.e., a closed manifold). Then we extend the results of Section 2 on Green formulas and energy estimates to the cylinder $\mathfrak{S}\times\mathbb{R}$. We also extend the Green formulas and the energy estimates of Section 2 to the indicial family $\widehat{\boldsymbol{\Xi}}(\tau)$ of $\boldsymbol{\Xi}$. As an application, we obtain criteria for the invertibility of the translation invariant, generalized Stokes operator $\boldsymbol{\Xi}$ on the cylinder $\mathfrak{S}\times\mathbb{R}$, under suitable positivity assumptions on $V$ and $V_{0}$ (Assumption 8.1).
In Section 8, we begin the study of the main operator of this paper, namely, the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (see Equation (1.3)) acting on a manifold with straight cylindrical ends. (Recall that, if $V$ and $V_{0}$ vanish, then $\boldsymbol{\Xi}$ becomes the usual Stokes operator.) We first obtain Fredholm and invertibility results for the Stokes operator $\boldsymbol{\Xi}$ under suitable positivity conditions on the potentials $V$ and $V_{0}$ (Assumption 8.1 and, respectively, Assumption 8.3). These results are interesting in themselves. Moreover, they play a crucial role in our paper, as these Fredholm and invertibility results are used then (in this section) to define the various layer potentials for the Stokes operator. In the last subsection of this section, we prove the jump and limit properties of these potentials. (These results were proved for the usual Stokes operator on Lipschitz domains in [100], but our method is different.) The proof of the jumpt/limit relations is based on the “algebra tool kit” developed for this purpose in the fourth and sixth sections of this paper.
In the ninth section, we specialize again to the case of a cylinder $M=\mathfrak{S}\times\mathbb{R}$ with an invariant metric and study the Fourier transforms (or indicial operators) of the corresponding layer potential operators. In combination with the Green formulas on cylinders, this allows us to establish the invertibility of the main layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ on a cylinder. This result is interesting in itself and is then used to study these operators on a general manifold with straight cylindrical ends.
The tenth section is devoted to Fredholmness and invertibility of the Stokes layer potential operators when $M$ is a general manifold with straight cylindrical ends, depending on the properties of our potentials $V$ and $V_{0}$. Thus, under suitable positivity assumptions on $V$ and $V_{0}$, we prove the invertibility of the single layer potential operator $\boldsymbol{S}$ and of the operator $\frac{1}{2}+\boldsymbol{K}$, where $\boldsymbol{K}$ is the corresponding double layer potential operator, on Sobolev spaces on the boundary of a domain with straight cylindrical ends. These are some of the main results of this paper and rely on the results developed in the previous sections. (The presentation in this paper is such as to lead as quickly as possible to this invertibility result, while preserving clarity and completeness, but sacrificing occasionally the generality.)
In the eleventh section, we prove the main well-posedness result of this paper, that is, the well-posedness of the Dirichlet problem for the generalized Stokes system, Theorem 1.1, assuming that $\Omega$ is a smooth domain with straight cylindrical ends in a connected manifold with straight cylindrical ends $M$ and suitable positivity assumptions on $V$ and $V_{0}$. Our proof uses layer potentials. This well-posedness result is useful to define the Dirichlet-to-Neumann operator for the generalized Stokes system and to prove that the conormal derivative of the Stokes double layer potential has no jump across the boundary of the domain $\Omega$, Theorem 11.6.
The twelfth (and last) section is devoted to applications and extensions. As an application, we obtain an existence and uniqueness result for the Dirichlet problem for the generalized Navier-Stokes system (12.4) in a smooth domain $\Omega$ with straight cylindrical ends assuming $k\in\mathbb{Z}_{+}$, $\operatorname{dim}\,M=n<2(k+3)$, that the given data is in $H^{k}(\Omega;TM)\oplus H^{k+3/2}(\partial\Omega;TM)$ and is sufficiently small in a sense that is described in detail in Theorem 12.3. The proof is based on the well-posedness of the Dirichlet problem for the non-homogeneous generalized Stokes system, Theorem 11.4, and a fixed point theorem. In the last subsection of this section (and of this paper), we state the well-posedness of the Dirichlet problem for the generalized Stokes system on a closed manifold [69]. We also explain what modifications need to be brought to the statements of our results for manifolds with straight cylindrical ends so that they are valid also in the case of closed manifolds. These resuls were proved in [69] (for closed manifolds), so in this paper we content ourselves with reminding a few of these results.
1.4. Overview of the existing literature
The layer potentials method is a classical method for solving elliptic boundary value problems [32, 35, 37, 39, 45, 54, 55, 59, 63, 76, 87, 96, 117, 134]. It plays a crucial role in the analysis of elliptic boundary value problems in various function spaces, their main advantage consisting in the fact that they provide explicit formulas for the the solutions and that they provide a convenient way of extending the results to $L^{p}$-type Sobolev spaces. There are also results that can be proved at this time only using layer potential methods. An example is provided by our well-posedness theorem, Theorem 1.1. More on the history of the method of layer potentials can be found, for example, in [2, 43, 45, 54, 59, 76, 87, 115, 135].
Let us now review some of the more specialized results in the subject, most of them from research papers devoted to the analysis of the Stokes and Navier-Stokes equations on domains, in Euclidean spaces, or on Riemannian manifolds. Fabes, Kenig and Verchota [43] initiated the study of boundary value problems for the Stokes system in Lipschitz domains of $\mathbb{R}^{n}$ with data in $L^{2}$-based Sobolev spaces and obtained various mapping properties for the corresponding layer potentials. They also obtained well-posedness results for the Dirichlet and Neumann problems in $L^{p}$-based spaces, with $p$ in a neighborhood of $2$, by combining Rellich type formulas with layer potential approaches. Mitrea and Wright [100] used layer potential methods and extended the results in [43] to a variety of $L^{p}$, Sobolev, Bessel potential, and Besov spaces. They also obtained well-posedness results for the main boundary value problems for the constant coefficient Stokes system in arbitrary Lipschitz domains in Euclidean spaces, together with the determination of the optimal range for $p$. The authors of [61] obtained mapping properties for the constant-coefficient Stokes and Brinkman layer potential operators in standard and weighted Sobolev spaces on $\mathbb{R}^{3}$. Shen [127] used a layer potential approach and obtained resolvent estimates in $L^{p}$ spaces for the Stokes operator in Lipschitz domains in the Euclidean spaces. Layer potential theoretical methods have been also combined with fixed point theorems in the analysis of boundary value problems for (linear and nonlinear) elliptic systems with nonlinear boundary conditions. For instance, the authors in [62] used a layer potential approach and the Leray-Schauder fixed point theorem to show the existence result in $L^{p}$, Sobolev, and Besov spaces for the constant-coefficient Stokes and Brinkman systems with nonlinear Neumann-transmission conditions. Girault and Sequeira [46] used a variational approach to prove the well-posedness of the exterior Dirichlet problem for the Stokes system on exterior Lipschitz domains in $\mathbb{R}^{n}$, $n=2,3$. Galdi [45, Chapters 5, 11, 12] used variational approaches in the analysis of steady Stokes and Navier-Stokes flows in exterior domains. Medková [90] used an integral equation method to analyze the Dirichlet problem for the Brinkman and Darcy-Forchheimer-Brinkman systems in Sobolev spaces on unbounded domains with compact Lipschitz boundary in $\mathbb{R}^{n}$. Exterior Dirichlet and transmission problems for anisotropic Stokes and Navier-Stokes systems with $L_{\infty}$ variable coefficients have been studied in [65] and [63, Chapters 5-7] using a variational approach and in [92]. Other boundary problems for the Stokes system in Sobolev spaces on Lipschitz domains via layer potential theoretical methods have also been studied in [88, 89, 91, 115, 118, 136].
An important paper is Costabel’s contribution [35] that establishes many very useful properties of layer potentials using elementary methods. Chandler-Wilde et al. [31, 32] studied boundary integral equations on locally-dilation-invariant Lipschitz domains and fractal screens. Buffa, Hiptmair, Petersdorff, and Schwab [28] studied Maxwell transmission problems in Lipschitz domains by using boundary element methods. Labarca-Figueroa and Hiptmair [75] used coupled boundary and volume integral equations in the study of electromagnetic scattering. Singularly perturbed boundary value problems based on a functional analytic approach have been studied by M. Dalla Riva, M. Lanza de Cristoforis, and P. Musolino [37]. This approach, which also uses a layer potential analysis, is very useful in the study of various linear nonlinear elliptic problems. A nonvariational form of the acoustic single layer potential has been obtained by Lanza de Cristoforis [78]. A nonvariational form of the Neumann problem for Hölder continuous harmonic functions, based also on a layer potential analysis, has been obtained in [77]. Ando, Kang, Miyanishi, and Putinar have obtained Carleman factorizations of layer potentials [15]. The recent book [96] of D. Mitrea, I. Mitrea and M. Mitrea is a rigorous interplay between Harmonic Analysis, Geometric Measure Theory, Function Space Theory, and Partial Differential Equations, and provide a powerful tool in the analysis of boundary problems for complex coefficient elliptic systems in various geometric settings, including the class of Lipschitz domains. The theory of Fredholm and layer potential operators in Euclidean spaces and in Riemannian manifolds plays an important role in their book [96] (and also in our book [63]).
The layer potential methods are also very useful in the analysis of elliptic boundary value problems on manifolds. Next we provide a brief overview of some of the important contributions related especially to the Stokes and Navier-Stokes equations in the setting of compact or non-compact Riemannian manifolds. Mitrea and Taylor [99] and Dindos̆ and Mitrea [39] obtained the well-posedness of the Dirichlet problem for the Stokes and Navier-Stokes systems with smooth coefficients on Lipschitz domains in closed Riemannian manifolds and in $L^{p}$, Sobolev and Besov spaces, by using the mapping properties of Stokes layer potentials, (see also [69] for similar results for a modified Stokes operator). In [64], the authors studied transmission problems for the smooth coefficients Navier-Stokes type equations in Lipschitz domains on closed Riemannian manifolds by using a layer potential approach combined with a fixed point theorem. Mitrea and Monniaux [97] studied the analyticity of the semigroup generated by the Stokes operator equipped with Neumann-type boundary conditions on $L^{p}$-spaces on Lipschitz domains of closed Riemannian manifolds. In [70], the authors proved the well-posedness of Dirichlet problems for the Stokes and Navier-Stokes systems with $L^{\infty}$-variable coefficients in $L^{2}$-based Sobolev spaces in Lipschitz domains on closed Riemannian manifolds. They used a variational approach to define the Newtonian and layer potentials for the non-smooth coefficients Stokes system on Lipschitz surfaces in closed Riemannian manifolds. Benavides, Nochetto, and Shakipov [18] used a variational approach and studied the well-posedness and regularity in $L^{p}$-based Sobolev spaces, $p\in(1,\infty)$, for the weak formulations of the stationary tangent Stokes and tangent Navier-Stokes systems on a compact and connected $d$-dimensional manifold without boundary of class $C^{m}$, $m\geq 2$, embedded in $\mathbb{R}^{d+1}$, in terms of the regularity of the source terms and the manifold.
The analysis on non-compact manifolds poses more challenges than that the one on compact manifolds. A significant issue in this sense is the construction of the classes of pseudodifferential operators. In this regard, we refer to the contributions of Ammann, Lauter and Nistor [11, 12], Mazzeo and Melrose [84], Melrose [93], Melrose and Mendoza [95], Nistor [103], Nistor, Weinstein, and Xu [104] Schrohe [120, 121], Schulze [122, 123], Schulze and Wong [124], Shubin [128], to mention just a few among very many others.
Another important issue in the analysis on non-compact manifolds is finding the “right definition” of the relevant function spaces. In this paper, as in [66], we have defined our function spaces by using connections. See also [3, 4, 119]. Amann [4] and Ammann, Große, Nistor [6, 8] studied function spaces on manifolds with bounded geometry and obtained well-posedness results for the Laplace operator. Using the uniform Shapiro-Lopatinski conditions, Große and Nistor [49] obtained regularity results for elliptic boundary value problems on manifolds with bounded geometry. Lie manifolds are special manifolds with bounded geometry [10, 11], which allow the implementation of the layer potential methods. See also [101, 102].
The layer potential theory for the Laplace operator on manifolds with cylindrical ends was initiated by Lewis and Parenti [81] and further extended by Mitrea and Nistor [98]. Manifolds with cylindrical ends are very useful in the study of boundary value problems on manifolds with conical points. See, for example, the papers [10, 27, 33, 38, 47, 71, 72, 73, 79, 82, 85, 94, 124], among many others devoted to the subject. The reason is provided by the fact that a domain with a conical point is mapped via the Kondratiev transformation [71] to a manifold with cylindrical ends. This fact was exploited by Qiao and Nistor [111] to obtain a solvability result for the Dirichlet problem for the Laplace operator in an open straight cone in $\mathbb{R}^{n}$, $n\geq 3$, in the setting of Kondratiev-Babuška weighted (Sobolev) spaces. Roidos, Schrohe and Seiler [113] studied the $H_{\infty}$-calculus in $L^{p}$-based Sobolev spaces for parameter-elliptic boundary value problems on manifolds with conical singularities at the boundary. As an application they treated the porous medium equation on manifolds with conical singularities (see also [112] and the references therein). Mitrea and Nistor [98] and Kohr, Nistor and Wendland [63, 67, 68] developed an essentially translation invariant pseudodifferential calculus on manifolds with cylindrical ends, which is very useful to provide the invertibility and structure of the Stokes operator and, as a consequence, the construction and the invertibility the Stokes layer potential operators on manifolds with cylindrical ends. Examples of domains with conical points are the polygonal domains and some polyhedral domains. Manifolds with cylindrical ends are particular cases of Lie manifolds [11, 29] and of manifolds with bounded geometry [6, 7, 8] (see also the references therein). The papers by Carvalho and Qiao [30], Qiao and Li [110], Qiao [109], Putinar and Perfect [105, 106] contain results on layer potential on polyhedral domains. See also [13, 14].
In [48], Große, Kohr, and Nistor have obtained the $L^{2}$-unique continuation property for the deformation operator $\operatorname{Def}\,$ on manifolds with bounded geometry, a result that was used for the analysis of the Stokes various classes of manifolds, including manifolds with cylindrical ends, in [63, 67, 68]. Manifolds with cylindrical ends appear in may applications that involve “wave-guides” or “layers” (see e.g. [19, 20, 21, 22, 23, 24, 25, 42, 63, 68, 74, 137]). Let us finally add that the main operator in the structure of the Stokes and Navier-Stokes equations on Riemannian manifolds is the deformation Laplacian $\boldsymbol{L}:=2\operatorname{Def}\,^{*}\operatorname{Def}\,$. In their groundbreaking paper [40], Ebin and Marsden described the deformation Laplacian as the right Laplace type operator to describe the Navier-Stokes equation on compact Riemannian manifolds. Pierfelice [108] also used the deformation-Laplacian for the definition and the analysis of the Navier-Stokes equation on a non-compact Riemannian manifold with negative Ricci curvature. This is the choice that we also adopt in the description of our Stokes operator $\boldsymbol{\Xi}$ of Equation (1.3). Our generalized Stokes operator is closely related to the Brinkman operator. The Brinkman flow was studied by Karageorghis, Lesnic, and Marin in [57, 58], and Bacuta, Hassell, Hsiao, and Sayas in [26]. A class of solutions to the Navier-Stokes equations for flows in plane-parallel channels and related singular perturbation problems were studied by Mazzucato and Taylor [86].
1.5. Acknowledgements
We thank Massimo Lanza de Cristoforis, Dorina, Irina, and Marius Mitrea, Sergey Mikhailov, and Mihai Putinar for useful discussions.
2. Green formulas for the Stokes operator
In this section we introduce the basic differential operators used for the study of our (generalized) Stokes operators and establish Green formulas and energy estimates for the Stokes operator. More precisely, in the first part of this section, Subsection 2.1, we introduce and study the deformation operator $\operatorname{Def}\,$, Equation (2.1). The deformation operator appears in the definition of the the Stokes operator $\boldsymbol{\Xi}=\boldsymbol{\Xi}_{V,V_{0}}$ introduced in Equation (1.3). We also introduce a few other basic differential operators needed for the study of the Stokes operator. In the second subsection, we recall an abstract integration by parts formulas, Proposition 2.9. The last subsection, Subsection 2.3 establishes the needed Green function formulas for the Stokes operator, as well as some related energy estimates and representation formulas.
The material contained in this section is thus mostly background material, thus we also introduce here other needed concepts or notation, mostly related to the differential operators appearing in the study of the Stokes operator. The results of this section, while mostly known, are sometimes hard to find in the literature. We thus provide here a complete and largely self-contained presentation of these results (a different approach is contained in Section A.3.3 of our book [63]).
In this section, we assume that $(M,g)$ is a Riemannian manifold, occasionally with some additional properties. More precisely, in the third subsection of this section, we assume that $M$ has bounded geometry. Recall that a manifold with bounded geometry is one that has positive injectivity radius and all the covariant derivatives of its curvature are bounded. (See [48] for references and for more on manifolds with bounded geometry. We note that the results of this section will be used only when $M$ has straight cylindrical ends, and that the manifolds with straight cylindrical ends have bounded geometry.)
2.1. Differential operators
We now introduce the basic differential operators used for the study of our (generalized) Stokes operators. If $E\to M$ is a (smooth) vector bundle, then $\mathcal{C}^{\infty}(M;E)$ denotes the set of smooth sections of $E$, whereas $\mathcal{C}^{\infty}_{c}(M;E)\subset\mathcal{C}^{\infty}(M;E)$ denotes the subset of compactly supported (smooth) sections of $E$. When necessary, the real vector bundles and their sections will be complexified. All our vector bundles will be smooth. If $E=\mathbb{C}$, it will be omitted from the notation, thus $\mathcal{C}^{\infty}(M)$ denotes the set of smooth functions on $M$.
One of the most basic differential operators is the derivation
$X:\mathcal{C}^{\infty}(M)\to\mathcal{C}^{\infty}(M)$
defined by a smooth vector field $X$. (Recall that a smooth vector field is a section of the tangent bundle $TM\to M$, whose dual is the cotangent bundle to $M$ and is denoted $T^{*}M$.) Another very basic differential operators is then the Levi-Civita connection
$\nabla^{LC}:\mathcal{C}^{\infty}(M;TM)\to\mathcal{C}^{\infty}(M;T^{*}M\otimes TM)\,.$
(Recall that the Levi-Civita connection $\nabla^{LC}$ is the unique torsion-free, metric preserving connection on $TM$.) The Levi-Civita connection on $TM$ then induces connections on the tensor bundles $T^{\otimes k}M\otimes T^{*\otimes l}M$, also be denoted $\nabla^{LC}$. If $X$ is a smooth vector field on $M$, then contraction with $X$ (on the first component tensor) defines a map $\iota_{X}:\mathcal{C}^{\infty}(M;T^{*}M\otimes TM)\to\mathcal{C}^{\infty}(M;TM)$ by $\iota(X)(\omega\otimes Y):=\omega(X)Y$, where $\omega\in\mathcal{C}^{\infty}(M;T^{*}M)$ (i.e., it is a 1-form) and $Y\in\mathcal{C}^{\infty}(M;TM)$ (i.e., $Y$ is a smooth vector field). This yields the differential operator
$\nabla_{X}^{LC}\mathrel{\,:=\,}\iota_{X}\circ\nabla^{LC}:\mathcal{C}^{\infty}(M;TM)\to\mathcal{C}^{\infty}(M;TM)$
that will often be used below. We obtain similar maps $\nabla_{X}^{LC}\mathrel{\,:=\,}\iota_{X}\circ\nabla^{LC}:\mathcal{C}^{\infty}(M;T^{*\otimes j}M\otimes T^{\otimes k}M)\to\mathcal{C}^{\infty}(M;T^{*\otimes j}M\otimes T^{\otimes k}M)$ between sections of tensor bundles.
We shall often use the musical isomorphism $\sharp:T^{*}M\to TM$ defined by the metric $g$, explicitly, if $X$ and $Y$ are vector fields on $M$, then
$\sharp:T^{*}M\to TM\ \mbox{ satisfies }\ \langle X^{\sharp},Y\rangle\mathrel{\,:=\,}X^{\sharp}(Y)\mathrel{\,:=\,}g(X,Y)\,.$
The inverse of $\sharp$ is denoted by the same symbol: $\sharp:TM\to T^{*}M$. One should be careful not to confuse $\nabla^{LC}X\in\mathcal{C}^{\infty}(M;T^{*}M\otimes TM)$, $X\in\mathcal{C}^{\infty}(M;TM)$ (a smooth vector field), with the gradient $\nabla f:=(df)^{\sharp}\in\mathcal{C}^{\infty}(M;TM)$, $f\in\mathcal{C}^{\infty}(M)$, a smooth function.
A crucial role in this paper will be played by the deformation operator
$\operatorname{Def}\,:\mathcal{C}^{\infty}(M;TM)\to\mathcal{C}^{\infty}(M;T^{*}M\otimes T^{*}M)\,,\quad\operatorname{Def}\,(X)\mathrel{\,:=\,}\frac{1}{2}\mathcal{L}_{X}g\,.$
Let $X,Y$, and $Z$ be smooth vector fields on $M$ and let $X\cdot Y=g(X,Y)$ denote the scalar product induced by the metric $g$ on $M$. The following alternative defining formula of $\operatorname{Def}\,$ will be more useful:
(2.1) $\operatorname{Def}\,(X)(Y,Z)\,=\,\langle\operatorname{Def}\,(X),Y\otimes Z\rangle\,=\,\frac{1}{2}\big[(\nabla_{Y}^{LC}X)\cdot Z+(\nabla_{Z}^{LC}X)\cdot Y\big]\,.$
To introduce other needed operators, let $\boldsymbol{\nu}$ be a fixed smooth vector field on $M$ (in our applications, $\boldsymbol{\nu}$ will extend the outer unit normal vector field to a domain $\Omega$ on which our boundary value problems are formulated; this explains why, in this paper, we require $\Omega$ to be on one side of its boundary, that is, to satisfy Assumption 2.7). The chosen vector field $\boldsymbol{\nu}$ induces a map
$\boldsymbol{\nu}\otimes\sharp:T^{*}M\otimes T^{*}M\to TM\,,\quad(\boldsymbol{\nu}\otimes\sharp)(\xi\otimes\eta)=\xi(\boldsymbol{\nu})\eta^{\sharp}\,.$
The map $\sharp\otimes\boldsymbol{\nu}:T^{*}M\otimes T^{*}M\to TM$ is defined analogously. We then let
(2.2) $\begin{gathered}\boldsymbol{D}_{\boldsymbol{\nu}}:\mathcal{C}^{\infty}(M;TM)\to\mathcal{C}^{\infty}(M;TM)\\ \boldsymbol{D}_{\boldsymbol{\nu}}X\mathrel{\,:=\,}\frac{1}{2}(\boldsymbol{\nu}\otimes\sharp+\sharp\otimes\boldsymbol{\nu})\,\operatorname{Def}\,(X)\,=\,\big\langle\,\operatorname{Def}\,(X)\,,\,\boldsymbol{\nu}\otimes 1\,\big\rangle^{\sharp}\,.\end{gathered}$
The operator $\boldsymbol{D}_{\boldsymbol{\nu}}$ and its (formal) adjoint $\boldsymbol{D}_{\boldsymbol{\nu}}^{*}$ will play an important role in what follows. (In the last formula, $\langle\ \cdot\ ,\,\boldsymbol{\nu}\otimes 1\,\rangle:T^{*}M\otimes T^{*}M\to\mathbb{C}\otimes T^{*}M=T^{*}M$ is the contraction with $\boldsymbol{\nu}$ on the first variable.) This allows us to introduce the operator
[…]
where $U:=(\boldsymbol{u}\ \ p)^{\top}$ is a smooth, compactly supported section of $TM\oplus\mathbb{C}$ (see also formula (2.2)). The formula $\boldsymbol{D}_{\boldsymbol{\nu}}X:=(\operatorname{Def}\,(X)\boldsymbol{\nu}\otimes 1)^{\sharp}$ is also used [39, 99, 134]. We shall also write
(2.4) $\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U\mathrel{\,:=\,}\left(\begin{array}[]{c}{-2\boldsymbol{D}_{\boldsymbol{\nu}}(\boldsymbol{u})+p\boldsymbol{\nu}}\\ {0}\end{array}\right)\,=\,\left(\begin{array}[]{cc}-2\boldsymbol{D}_{\boldsymbol{\nu}}&\boldsymbol{\nu}\\ 0&0\end{array}\right)U\,.$
Let $V,V_{0}:M\to[0,\infty)$ be measurable functions (usually smooth). Recall from the introduction that the deformation Laplacian is the second order differential operator $\boldsymbol{L}\mathrel{\,:=\,}2\operatorname{Def}^{*}\,\operatorname{Def}\,$. The operator
$\boldsymbol{L}_{V}\mathrel{\,:=\,}2\operatorname{Def}^{*}\,\operatorname{Def}\,+V$
will be called the perturbed deformation Laplacian. Also, recall that the generalized Stokes operator (Equation (1.3)) is the operator
$\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}\mathrel{\,:=\,}\left(\begin{array}[]{ccc}\boldsymbol{L}_{V}&\nabla\\ \nabla^{*}&-V_{0}\end{array}\right)\in\operatorname{End}(\mathcal{C}^{\infty}_{c}(M;TM\oplus\mathbb{C}))\,.$
(Recall that for a module $E$ over some ring $R$, $\operatorname{End}(E)$ denotes the set of $R$-linear maps $E\to E$, that is, the set of endomorphisms of $E$.)
We now study some of the properties of these operators. Then a direct calculation gives right away the following result.
Lemma 2.1.
We have
$\boldsymbol{T}_{\boldsymbol{\nu}}^{*}(\boldsymbol{u})\,=\,\left(\begin{array}[]{c}-2\boldsymbol{D}_{\boldsymbol{\nu}}^{*}\boldsymbol{u}\\ \boldsymbol{\nu}\cdot\boldsymbol{u}\end{array}\right)\,=\,\left(\begin{array}[]{c}-2\boldsymbol{D}_{\boldsymbol{\nu}}^{*}\\ \boldsymbol{\nu}^{\sharp}\end{array}\right)\,\boldsymbol{u}\,=\,\left(\begin{array}[]{cc}-2\boldsymbol{D}_{\boldsymbol{\nu}}^{*}&0\\ \boldsymbol{\nu}^{\sharp}&0\end{array}\right)U\,.$
In this paper, $\imath$ will denote the imaginary unit: $\imath^{2}=-1$. The principal symbols of the operators of interest can be determined from the functorial properties of the principal symbol using also the following formula (stationary phase approximation [52, 54, 130]).
Lemma 2.2.
Let $\phi:M\to\mathbb{R}$ be a smooth function and $P$ be a first order differential operator $\mathcal{C}^{\infty}_{c}(M;E)\to\mathcal{C}^{\infty}_{c}(M;F)$, where $E,F\to M$ are smooth vector bundles. Then
$\sigma_{1}(P;d\phi(x))u(x)\,=\,\lim_{t\to\infty}t^{-1}e^{-\imath t\phi(x)}\big[P(e^{\imath t\phi}u)\big](x)\ \mbox{ if }\ d\phi(x)\neq 0\,.$
The principal symbol will be discussed in general in Section 4, see Definition 3.7. For more details on the principal symbol, see the above mentioned textbooks [52, 54, 130] or [63] (or any basic monograph on pseudodifferential operators).
We shall use repeatedly in what follows the following notation.
Notation 2.3.
Let $\mathfrak{W}$ be a vector space and let $v\in\mathfrak{W}$ and $\xi\in\mathfrak{W}^{*}$. We shall then regard $v\otimes\xi\in\mathfrak{W}\otimes\mathfrak{W}^{*}$ as an endomorphism $v\otimes\xi\in\operatorname{End}(\mathfrak{W})\simeq\mathfrak{W}\otimes\mathfrak{W}^{*}$ via the formula $(v\otimes\xi)x:=\xi(x)v$, for any $x\in\mathfrak{W}$. In particular, if $\mathfrak{W}$ is a real vector space with an inner product denoted $\cdot$ and $\sharp:\mathfrak{W}\to\mathfrak{W}^{*}$ is the musical isomorphism (i.e., $v^{\sharp}(w):=v\cdot w$), then, for all $v,w,x\in\mathfrak{W}$, we have
[…]
where, for the last formula, we have used the fact that $\mathfrak{W}$ is a real vector space.
For any vector space or vector bundle $E$, we let $S^{2}E\subset E\otimes E$ be the symmetric part of the tensor product $E\otimes E$. The following formulas are well known, see [39, 99, 130]. We state them here for further use and we prove them for completeness.
Proposition 2.4.
Let $X$, $Y$, and $Z$ be three smooth vector fields on $M=M^{\prime}\times\mathbb{R}$. Then
  1. (i)
    $\sigma_{1}(\operatorname{Def}\,;\xi)X\,=\,\frac{\imath}{2}\big[\xi\otimes X^{\sharp}+X^{\sharp}\otimes\xi]\in S^{2}T^{*}M\subset T^{*\otimes 2}M$ and
  2. (ii)
    $\sigma_{1}(\operatorname{Def}^{*}\,;\xi)\big(Y^{\sharp}\otimes Z^{\sharp}\big)\,=\,-\frac{\imath}{2}\big[\xi(Y)Z+\xi(Z)Y\big]\,.$
Note that, in the last equation, $\sigma_{1}(\operatorname{Def}^{*}\,;\xi)$ was tacitly extended to act on all tensors $T^{*}M\otimes T^{*}M$ (not just on $S^{2}T^{*}M\subset T^{*\otimes 2}M$). Nevertheless, this extension is canonical, in the sense that it vanishes on anti-symmetric tensors:
$\sigma_{1}(\operatorname{Def}^{*}\,;\xi)\big(Y^{\sharp}\otimes Z^{\sharp}-Z^{\sharp}\otimes Y^{\sharp}\big)\,=\,-\frac{\imath}{2}\big[\xi(Y)Z+\xi(Z)Y-\xi(Z)Y-\xi(Y)Z\big]\,=\,0\,.$
Proof.
Let $x\in M$ and $\phi:M\to\mathbb{R}$ be a smooth function such that $d\phi(x)\neq 0$. Let us replace $X$ with $e^{\imath t\phi}X$ in Equation (2.1). Lemma 2.2 then gives:
$\langle\sigma_{1}(\operatorname{Def}\,;d\phi)X,Y\otimes Z\rangle$ $\,=\,\lim_{t\to\infty}\frac{e^{-\imath t\phi}}{2t}\big[Z\cdot\nabla_{Y}^{LC}(e^{\imath t\phi}X)+Y\cdot\nabla_{Z}^{LC}(e^{\imath t\phi}X)\big]$
$\,=\,\frac{\imath}{2}\big[d\phi(Y)X\cdot Z+d\phi(Z)X\cdot Y\big]$
$\,=\,\frac{\imath}{2}\langle d\phi\otimes X^{\sharp}+X^{\sharp}\otimes d\phi,Y\otimes Z\rangle\,.$
This proves (i). From this, using that the principal symbol is stable for adjoints (i.e., $\sigma_{m}(P^{*};\xi)=\sigma_{m}(P;\xi)^{*}$), we obtain
$(\sigma_{1}(\operatorname{Def}^{*};\xi)\big(Y^{\sharp}\otimes Z^{\sharp}\big),X)$ $\,=\,(Y^{\sharp}\otimes Z^{\sharp},\sigma_{1}(\operatorname{Def}\,;\xi)X)$
$\,=\,\big(Y^{\sharp}\otimes Z^{\sharp},\frac{\imath}{2}\big[\xi\otimes X^{\sharp}+X^{\sharp}\otimes\xi]\big)$
$\,=\,-\frac{\imath}{2}\big[\xi(Y)X\cdot Z+\xi(Z)X\cdot Y\big]$
$\,=\,-\frac{\imath}{2}\big(\xi(Y)Z+\xi(Z)Y,X\big)\,.$
This proves also (ii) and completes the proof. ∎
For any first order differential operator $P:\mathcal{C}^{\infty}(M;E)\to\mathcal{C}^{\infty}(M;F)$, we let
(2.5) $\partial_{\boldsymbol{\nu}}^{P}\mathrel{\,:=\,}-\imath\sigma_{1}(P;\boldsymbol{\nu}^{\sharp})\,$,
where $\boldsymbol{\nu}$ is the chosen vector field (in this section, this vector field is only required to be smooth, but, beginning with the next subsection, we will require $\boldsymbol{\nu}$ to be outer unit normal at the boundary of our given domain $\Omega$). This definition is motivated by the abstract integration formula of the following subsection.
The next formulas are an immediate consequence of the formulas of Propositions 2.9 and 2.4. (See Section A.3.4 from [63] for more details. Also, recall the formula $v\otimes w^{\sharp}x\mathrel{\,:=\,}(w\cdot x)v$ from Notation (2.3).)
Corollary 2.5.
The deformation operator $\operatorname{Def}\,$ of Equation (2.1) satisfies the following formulas:
  1. (i)
    $\sigma_{1}(\operatorname{Def}\,;\xi)=\frac{\imath}{2}\big[\xi\otimes\sharp+\sharp\otimes\xi\big]\in\operatorname{Hom}(TM;T^{*\otimes 2}M)$;
  2. (ii)
    $\partial_{\boldsymbol{\nu}}^{\operatorname{Def}\,}:=-\imath\sigma_{1}(\operatorname{Def}\,;\boldsymbol{\nu}^{\sharp})=\frac{1}{2}\big[\boldsymbol{\nu}^{\sharp}\otimes\sharp+\sharp\otimes\boldsymbol{\nu}^{\sharp}\big]\in\operatorname{Hom}(TM;T^{*\otimes 2}M)$;
  3. (iii)
    $\sigma_{1}(\operatorname{Def}^{*}\,;\xi)=-\frac{\imath}{2}\big[\xi^{\sharp}\otimes\sharp+\sharp\otimes\xi^{\sharp}\big]\in\operatorname{Hom}(T^{*\otimes 2}M;TM)$;
  4. (iv)
    $\partial_{\boldsymbol{\nu}}^{\operatorname{Def}^{*}\,}:=-\imath\sigma_{1}(\operatorname{Def}\,^{*};\boldsymbol{\nu}^{\sharp})=-\frac{1}{2}\big[\boldsymbol{\nu}\otimes\sharp+\sharp\otimes\boldsymbol{\nu}\big]\in\operatorname{Hom}(T^{*\otimes 2}M;TM)$;
Proof.
The first point is nothing but a reformulation of the first point of the last proposition. The second point follows right away from the first point. The third point point is obtained by applying again Proposition 2.9. Indeed, the point (ii) of that proposition, namely, $\sigma_{1}(\operatorname{Def}^{*}\,;\xi)\big(Y^{\sharp}\otimes Z^{\sharp}\big)\,=\,-\frac{\imath}{2}\big[\xi(Y)Z+\xi(Z)Y\big]$ gives right away
$\frac{\imath}{2}\big[\xi^{\sharp}\otimes\sharp+\sharp\otimes\xi^{\sharp}\big]\big(Y^{\sharp}\otimes Z^{\sharp}\big)\mathrel{\,:=\,}-\frac{\imath}{2}\xi(Y)Z+\xi(Z)Y\\ \,=\,\sigma_{1}(\operatorname{Def}^{*};\xi)\big(Y^{\sharp}\otimes Z^{\sharp}\big)\,.$
The last point follows from (iii) just proved using also that $\sharp$ is involutive, that is, $\sharp^{2}=id$. ∎
We also obtain the following slightly more difficult formulas.
Corollary 2.6.
The deformation operator $\operatorname{Def}\,$ and the differential operator $\boldsymbol{D}_{\boldsymbol{\nu}}$ of Equation (2.2) satisfy the following formulas:
  1. (i)
    $\partial_{\boldsymbol{\nu}}^{\operatorname{Def}^{*}\,}\operatorname{Def}\,=-\boldsymbol{D}_{\boldsymbol{\nu}}$;
  2. (ii)
    $\sigma_{1}(\boldsymbol{D}_{\boldsymbol{\nu}};\xi)=\frac{\imath}{2}\big[\xi(\boldsymbol{\nu})+\xi^{\sharp}\otimes\boldsymbol{\nu}^{\sharp}\big]$;
  3. (iii)
    $\sigma_{1}(\boldsymbol{D}_{\boldsymbol{\nu}}^{*};\xi)=-\frac{\imath}{2}\big[\xi(\boldsymbol{\nu})+\boldsymbol{\nu}\otimes\xi\big]$; and
  4. (iv)
    $\sigma_{2}(\operatorname{Def}^{*}\,\operatorname{Def}\,;\xi)=\frac{1}{2}(|\xi|^{2}+\xi^{\sharp}\otimes\xi)$.
Proof.
The point (i) follows directly from $\boldsymbol{D}_{\boldsymbol{\nu}}\mathrel{\,:=\,}\frac{1}{2}(\boldsymbol{\nu}\otimes\sharp+\sharp\otimes\boldsymbol{\nu})\,\operatorname{Def}\,$, see Equation (2.2), and of the formula $\partial_{\boldsymbol{\nu}}^{\operatorname{Def}^{*}\,}=-\frac{1}{2}\big[\boldsymbol{\nu}\otimes\sharp+\sharp\otimes\boldsymbol{\nu}\big]\in\operatorname{Hom}(T^{*\otimes 2}M;TM)$ of Corollary 2.5(iv).
To prove (ii), we use again the definition of $\boldsymbol{D}_{\boldsymbol{\nu}}\mathrel{\,:=\,}\frac{1}{2}(\boldsymbol{\nu}\otimes\sharp+\sharp\otimes\boldsymbol{\nu})\,\operatorname{Def}\,$ and Corollary 2.5(i) to obtain
$\sigma_{1}(\boldsymbol{D}_{\boldsymbol{\nu}};\xi)X\,=\,\frac{1}{2}\big[\boldsymbol{\nu}\otimes\sharp+\sharp\otimes\boldsymbol{\nu}\big]\,\sigma_{1}(\operatorname{Def}\,;\xi)X\,=\,\frac{1}{2}\big[\boldsymbol{\nu}\otimes\sharp+\sharp\otimes\boldsymbol{\nu}\big]\,\frac{\imath}{2}\big[\xi\otimes\sharp+\sharp\otimes\xi\big]X\\ \,=\,\frac{\imath}{4}\big[\boldsymbol{\nu}\otimes\sharp+\sharp\otimes\boldsymbol{\nu}\big]\,(\xi\otimes X^{\sharp}+X^{\sharp}\otimes\xi)\\ \,=\,\frac{\imath}{4}\big(\xi(\boldsymbol{\nu})X+(\boldsymbol{\nu}\cdot X)\xi^{\sharp}+(\boldsymbol{\nu}\cdot X)\xi^{\sharp}+\xi(\boldsymbol{\nu})X\big)\\ \,=\,\frac{\imath}{2}\big(\xi(\boldsymbol{\nu})X+(\boldsymbol{\nu}\cdot X)\xi^{\sharp}\big)\,=\,\frac{\imath}{2}\big(\xi(\boldsymbol{\nu})+\xi^{\sharp}\otimes\boldsymbol{\nu}^{\sharp}\big)X\,.$
The transformation properties of the principal symbol with respect to adjoints and the formula $(v\otimes w)^{*}=w^{\sharp}\otimes v^{\sharp}$ give right away (iii). Finally, taking into account that $\sharp^{2}=id$ and that $\xi^{\sharp}(\xi)=|\xi|^{2}$ and using the multiplicativity of the principal symbol, we obtain
$\sigma_{2}(\operatorname{Def}^{*}\,\operatorname{Def}\,;\xi)$ $\,=\,\sigma_{1}(\operatorname{Def}^{*}\,;\xi)\sigma_{1}(\operatorname{Def}\,;\xi)$
$\,=\,\frac{1}{4}(\xi^{\sharp}\otimes\sharp+\sharp\otimes\xi^{\sharp})(\xi\otimes\sharp+\sharp\otimes\xi^{\sharp})\,=\,\frac{1}{2}(|\xi|^{2}+\xi^{\sharp}\otimes\xi)\,$,
as claimed in (iv). ∎
2.2. An abstract integration by parts formula
We continue to assume that $(M,g)$ is a Riemannian manifold. We now recall a general (abstract) integration by parts formula that will often be used in conjunction with the formulas of the previous subsection.
Our integration by parts formulas will be used either on the manifold $M$ or on some open subset $\Omega\subset M$ with smooth boundary $\Gamma:=\partial\Omega$.
Assumption 2.7.
Let $\Omega\subset M$ be an open subset. Recall that $\overline{\Omega}$ denotes the closure of $\Omega$ and that $\Gamma$ denotes the boundary $\partial\Omega:=\overline{\Omega}\smallsetminus\Omega$ of $\Omega$. We require in this paper that $\Omega$ be on one side of its boundary. This means that we require one of the following two equivalent conditions:
  1. (i)
    If $\Omega_{-}:=M\smallsetminus\overline{\Omega}$, then $\Gamma:=\partial\Omega=\partial\Omega_{-}$.
  2. (ii)
    $\Omega$ is the interior of $\overline{\Omega}$.
This assumption is necessary in order to define a smooth vector field $\boldsymbol{\nu}$ that, at $\Gamma$, coincides with the unit outer normal vector to $\Omega$. In the case of integration over $M$, integration by parts will be performed for compactly supported functions (or sections), so there will be no boundary terms. However, when using integration by parts on the domain $\Omega$, we will obtain boundary terms. It is the purpose of this subsection to give a general formula for these boundary terms. There is no loss of generality to assume that $\Omega$ is connected
We endow $\partial\Omega$ with the measure $dS_{g}$ induced by the metric on $M$. This allows us to introduce the inner product on the boundary $\Gamma$ of $\Omega$. Also, we let $u\cdot v$ denote the (pointwise) scalar product of two sections of our Hermitian vector bundle $E\to M$, as before. We then let
(2.6) $\begin{gathered}(u,v)\,=\,(u,v)_{\Omega}\mathrel{\,:=\,}\int_{\Omega}u(x)\cdot v(x)\,d\operatorname{vol}_{g}(x)\quad\mbox{ and}\\ (u,v)^{\prime}\mathrel{\,:=\,}\int_{\Gamma=\partial\Omega}u(x)\cdot v(x)\,dS_{g}(x)\end{gathered}$
denote the $L^{2}$-inner product of sections of $E$ over $\Omega$ and, respectively, over $\Gamma:=\partial\Omega$.
We will need some additional formulas continuing the discussion of the previous subsection. We include them in a remark.
Remark 2.8.
Recall that $\sharp:TM\to T^{*}M$ is the vector bundle isomorphism defined by the metric $g$ on $TM$. Also, if $E,F\to M$ are two Hermitian vector bundles, then recall that $\sigma_{1}(P;\xi)\in\operatorname{Hom}(E;F)$ is the value at $\xi\in T^{*}M$ of the principal symbol of an order one differential operator $P:\mathcal{C}^{\infty}(M;E)\to\mathcal{C}^{\infty}(M;F)$ and $\partial_{\boldsymbol{\nu}}^{P}:=-\imath\sigma_{1}(P;\boldsymbol{\nu}^{\sharp})$ (see Equation (2.5)). Let also $d\operatorname{vol}_{g}$ be the volume element associated to the metric $g$ on $M$. We then have the following simple results.
  1. (i)
    or any vector field $X$, we have [107, page 49]
    (2.7) $\mathcal{L}_{X}\,d\operatorname{vol}_{g}\,=\,\operatorname{div}(X)\,d\operatorname{vol}_{g}\,$,
    where $\mathcal{L}_{X}$ is the Lie derivative in the direction of $X$.
  2. (ii)
    If $P$ is a first order differential operator, then $\sigma_{1}(P^{*})=\sigma_{1}(P)^{*}$, and, hence $\partial_{\boldsymbol{\nu}}^{P^{*}}=-(\partial_{\boldsymbol{\nu}}^{P})^{*}$.
  3. (iii)
    If $X$ is a smooth vector field on $M$ and $P=X$, then $\sigma_{1}(X;\xi)=\imath\xi(X)$, and hence
    $\partial_{\boldsymbol{\nu}}^{X}\,=\,-\imath\sigma_{1}(X;\boldsymbol{\nu}^{\sharp})\,=\,-\imath\big(\imath\boldsymbol{\nu}^{\sharp}(X)\big)\,=\,\boldsymbol{\nu}\cdot X\,$,
    which, as we will see in the next proposition, is consistent with the well-known divergence formula:
    $(X(f),h)_{\Omega}\,=\,-(f,X(h)+\operatorname{div}(X)h)_{\Omega}+((X\cdot\boldsymbol{\nu})f,h)^{\prime}\,$,
    [54, 107, 138]. (See [63] for more details, where the above formula was used as the first step in the proof of Proposition 2.9.)
  4. (iv)
    If $P:=\nabla:=\sharp d:\mathcal{C}^{\infty}(M)\to\mathcal{C}^{\infty}(M;TM)$ (the gradient), then $\sigma_{1}(\nabla;\xi)=\imath\xi^{\sharp}$ and a particular case of Proposition 2.9 next is the following integration by parts formula:
    $(\nabla f,X)_{\Omega}\,=\,(f,\nabla^{*}X)_{\Omega}+(f\boldsymbol{\nu},X)_{\partial\Omega}\,=\,(f,\nabla^{*}X)_{\Omega}+(f,\boldsymbol{\nu}\cdot X)_{\partial\Omega}\,.$
    Recalling that $\nabla_{X}:=\iota_{X}\circ\nabla$, we obtain that $\sigma_{1}(\nabla;\xi)=\iota_{X}\circ\imath\xi^{\sharp}=\imath\xi(X)$.
We are ready to formulat now the needed abstract integration by parts formula, which is Proposition 9.1 from Chapter 2 of [130], see also Proposition A.3.14 from [63].
Proposition 2.9.
Let $\Omega\subset M$ be an open subset with smooth boundary $\Gamma:=\partial\Omega$ and let $P:\mathcal{C}^{\infty}(M;E)\to\mathcal{C}^{\infty}(M;F)$ be a first order differential operator. Let
$\partial_{\boldsymbol{\nu}}^{P}\,=\,-\imath\sigma_{1}(P;\boldsymbol{\nu}^{\sharp})\in\operatorname{Hom}(E;F)$
and let $P^{*}:\mathcal{C}^{\infty}(M;F)\to\mathcal{C}^{\infty}(M;E)$ be the formal adjoint of $P$. Then, for all $u\in\mathcal{C}^{\infty}_{c}(M;E)$ and $v\in\mathcal{C}^{\infty}_{c}(M;F)$, we have
$(Pu,v)_{\Omega}\,=\,(u,P^{*}v)_{\Omega}+(\partial_{\boldsymbol{\nu}}^{P}u,v)^{\prime}$
and $\partial_{\boldsymbol{\nu}}^{P}:\mathcal{C}^{\infty}(M;E)\to\mathcal{C}^{\infty}(\Gamma;F)$ is the unique operator with this property. (Recall that $(\ ,\ )^{\prime}$ is the inner product on the boundary $\partial\Omega$.)
For completeness, we sketch now a proof. See [63] for more details.
Proof.
First, if $P=a$, an order zero multiplication operator, then the formula is obviously true. Second, if $P=X$, a smooth vector field, the claimed formula is nothing but the divergence formula (Remark (2.8)(ii)), which is well known. One checks easily that if the formula of the proposition is true for $P$ and $a$ and $b$ are suitable endomorphisms, then the formula is true also for $aPb$. Since the formula is linear and all first order differential operators can be written as linear combinations of differential operators of the form $aXb$ and $a$, with $X$ a smooth vector field, the result follows. ∎
2.3. Green formulas and energy estimates for the Stokes operator
In this subsection, $(M,g)$ is a Riemannian manifold with bounded geometry, that is, we assume that it has positive injectivity radius and all covariant derivatives of the curvature are bounded. (See [48] for more on manifolds with bounded geomery, including references.) Also, $E\to M$ is a Hermitian vector bundle with connection, also assume to have bounded geometry, meaning that all the covariant derivatives of its curvature are bounded. We will let $\Omega\subset M$ be an open subset with boundary $\Gamma:=\partial\Omega$, assumed to be a submanifold with bounded geometry of $M$ (see [9, 63] definitions). In this paper, we are interested in the case when $M$ is a manifold with straight cylindrical ends (Definition 3.11), $E\to M$ a compatible Hermitian vector bundle with metric preserving connection (Definition 3.13), and $\Omega\subset M$ also compatible with the straight cylindrical ends structure of $M$ (Equation (1.2)). Then $M$, $E$, and $\Omega$ fit into the framework of this subsection (that is $M$, $E$, and $\Gamma$ have bounded geometry), so the results proved next apply to the setting we are interested in.
Our Green-type formulas are formulated in terms of the scalar product on $E$, denoted $\xi\cdot\eta$, and on the innner-product on $L^{2}(\Omega;E)$, which, we recall, is given by
$(f,g)_{\Omega}\mathrel{\,:=\,}\int_{\Omega}f(x)\cdot g(x)\,\,d\operatorname{vol}_{g}(x)\,.$
(In particular, we have $(f,g)_{\Omega}\mathrel{\,:=\,}\int_{\Omega}f(x)\overline{g(x)}\,\,d\operatorname{vol}_{g}(x)$ if $f$ and $g$ are functions.) Our convention is that our scalar products are conjugate linear in the second variable and linear in the first. Let also $\boldsymbol{u},\boldsymbol{w}\in H^{1}(\Omega;E)$. For the simplicity of the notations, we shall write
(2.8) $(\boldsymbol{u},\boldsymbol{w})^{\prime}\mathrel{\,:=\,}(\boldsymbol{u}_{+},\boldsymbol{w}_{+})^{\prime}\,.$
To avoid confusion (for instance, when $u$ is defined on both sides of $\Gamma:=\partial\Omega$), we shall sometimes write $(\boldsymbol{u}_{+},\boldsymbol{w}_{+})^{\prime}$ instead of $(\boldsymbol{u},\boldsymbol{w})^{\prime}$. Recall that $\textbf{1}_{A}$ denotes the characteristic function of the set $A$ (that is, $\textbf{1}_{A}(x)=1$ if $x\in A$ and $\textbf{1}_{A}(x)=0$ if $x\notin A$). We shall also need the following notation, which will be fixed from now on.
Notation 2.10.
We will use the following notation throughout this section:
  1. (i)
    $U\mathrel{\,:=\,}\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)\,=\,(\boldsymbol{u}\ \ p)^{\top}$ and $W\mathrel{\,:=\,}\left(\begin{array}[]{c}{\boldsymbol{w}}\\ {q}\end{array}\right)\,=\,(\boldsymbol{w}\ \ q)^{\top}$ will denote two generic sections of $H^{2}(\Omega;TM)\oplus H^{1}(\Omega)$; thus $\boldsymbol{u},\boldsymbol{w}\in H^{2}(\Omega;TM)$ are the vector components and $p,q\in H^{1}(\Omega)$ are the scalar components of $U$ and, respectively, $W$.
  2. (ii)
    The sesquilinear form
    $B_{\Omega}(U,W)\mathrel{\,:=\,}2(\operatorname{Def}\,\boldsymbol{u},\operatorname{Def}\,\boldsymbol{w})_{\Omega}+(\nabla^{*}\boldsymbol{u},q)_{\Omega}+(p,\nabla^{*}\boldsymbol{w})_{\Omega}+(V\boldsymbol{u},\boldsymbol{w})_{\Omega}-(V_{0}p,q)_{\Omega}$
    and we sometimes let $\mathfrak{v}\mathrel{\,:=\,}(V\boldsymbol{u},\boldsymbol{w})_{\Omega}-(V_{0}p,q)_{\Omega}$;
  3. (iii)
    If $w$ is the section over $\Omega$ of some vector bundle over $M$, then $\textbf{1}_{\Omega}w$ will denote its extension with zero outside $\Omega$.
Recall that, for all $\in H^{1}(\Omega;E), h+h_{+}$ denotes its trace (or restriction) at the boundary $\partial\Omega$. Also, recall the definition of distributions of the form $h\otimes\delta_{\Gamma}$,
(2.9) $\langle h\otimes\delta_{\Gamma},\phi\rangle\mathrel{\,:=\,}\int_{\Gamma}h(x)\cdot\phi(x)\,dS_{\Gamma}(x)\,.$
We then have the following two Green-type formulas formula that relate our generalise Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ with the form $B_{\Omega}$
Proposition 2.11.
Let us assume that $M$ has bounded geometry and that $\partial\Omega\subset M$ is a submanifold with bounded geometry, as before. We use the Notation 2.10, in particular, $U:=(\boldsymbol{u}\ \ p)^{\top}$ and $W:=(\boldsymbol{w}\ \ q)^{\top}$ are in $H^{2}(\Omega;TM)\oplus H^{1}(\Omega)$. Then we have the following relations:
  1. (i)
    $\big(\boldsymbol{\Xi}U,W\big)_{\Omega}\,=\,B_{\Omega}(U,W)+(\boldsymbol{T}_{\boldsymbol{\nu}}U,\boldsymbol{w})^{\prime}\,=\,B_{\Omega}(U,W)+\big(\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U,W\big)^{\prime}\,.$
  2. (ii)
    $\big(\boldsymbol{\Xi}U,W\big)_{\Omega}-\big(U,\boldsymbol{\Xi}W\big)_{\Omega}\,=\,(\boldsymbol{T}_{\boldsymbol{\nu}}U,\boldsymbol{w})^{\prime}-(\boldsymbol{u},\boldsymbol{T}_{\boldsymbol{\nu}}W)^{\prime}\,=\,(\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U,W)^{\prime}-(U,\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}W)^{\prime}\,.$
  3. (iii)
    We let $\textbf{1}_{\Omega}U$ be the section of $TM\oplus\mathbb{C}\to M$ that is equal to $U$ on $\Omega$ and equal to 0 outside $\Omega$. Then we have the following representation formula
    $\boldsymbol{\Xi}\big(\textbf{1}_{\Omega}U\big)\,=\,\textbf{1}_{\Omega}\big(\boldsymbol{\Xi}U\big)-(\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U)_{+}\otimes\delta_{\Gamma}+\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}^{*}(U_{+}\otimes\delta_{\Gamma})\,.$
If $\Omega=M$, the results are still true, but we must remove the boundary terms, that is the ones involving $(\ ,\ )^{\prime}$. More precisely, this proposition reduces to
(2.10) $\big(\boldsymbol{\Xi}U,W\big)_{M}\,=\,B_{\Omega}(U,W)\,=\,\big(\boldsymbol{\Xi}U,W\big)_{M}\,.$
Proof.
Let us assume first that $U$ and $W$ are smooth and compactly supported. Let $\mathfrak{v}:=(\boldsymbol{u},V\boldsymbol{w})_{\Omega}-(p,V_{0}q)_{\Omega}=(V\boldsymbol{u},\boldsymbol{w})_{\Omega}-(V_{0}p,q)_{\Omega}$ and $B_{\Omega}$, be as in Notation 2.10. Then, using the relation $\partial_{\boldsymbol{\nu}}^{\operatorname{Def}^{*}\,}\operatorname{Def}\,=-\boldsymbol{D}_{\boldsymbol{\nu}}$ of Corollary 2.6(i), we obtain the following relations:
$\big(\boldsymbol{\Xi}U,W\big)_{\Omega}\,=\,(2\operatorname{Def}^{*}\,\operatorname{Def}\,\boldsymbol{u},\boldsymbol{w})_{\Omega}+(\nabla^{*}\boldsymbol{u},q)_{\Omega}+(\nabla p,\boldsymbol{w})_{\Omega}+\mathfrak{v}\\ \,=\,2(\operatorname{Def}\,\boldsymbol{u},\operatorname{Def}\,\boldsymbol{w})_{\Omega}-2(\boldsymbol{D}_{\boldsymbol{\nu}}\boldsymbol{u},w)^{\prime}+(\nabla^{*}\boldsymbol{u},q)_{\Omega}+(p,\nabla^{*}\boldsymbol{w})_{\Omega}+(p\boldsymbol{\nu},\boldsymbol{w})^{\prime}+\mathfrak{v}\\ \,=\,B_{\Omega}(U,W)-2(\boldsymbol{D}_{\boldsymbol{\nu}}\boldsymbol{u},w)^{\prime}+(p\boldsymbol{\nu},\boldsymbol{w})^{\prime}\,=\,B_{\Omega}(U,W)+(\boldsymbol{T}_{\boldsymbol{\nu}}U,\boldsymbol{w})^{\prime}\\ \,=\,B_{\Omega}(U,W)+(\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U,W)^{\prime}\,.$
This proves (i) for $U$ and $W$ smooth, compactly supported. To complete the proof of (i), it suffices to notice that all the terms in (i) are continuous on $H^{2}(M;TM)\oplus H^{1}(M)$ and that $\mathcal{C}^{\infty}_{c}(M;TM\oplus\mathbb{C})$ is dense in the latter, because $M$ has bounded geometry [50, 63].
The second point is an immediate consequence of the first one using the fact that $B_{\Omega}$ is sesquilinear. Indeed,
$\big(\boldsymbol{\Xi}U,W\big)_{\Omega}-\big(U,\boldsymbol{\Xi}W\big)_{\Omega}$ $\,=\,\big(\boldsymbol{\Xi}U,W\big)_{\Omega}-\overline{\big(\boldsymbol{\Xi}W,U\big)}_{\Omega}$
$\,=\,B_{\Omega}(U,W)+(\boldsymbol{T}_{\boldsymbol{\nu}}U,\boldsymbol{w})^{\prime}-\overline{B_{\Omega}(W,U)}-\overline{(\boldsymbol{T}_{\boldsymbol{\nu}}W,\boldsymbol{u})}\,{}_{\partial\Omega}$
$\,=\,(\boldsymbol{T}_{\boldsymbol{\nu}}U,\boldsymbol{w})^{\prime}-(\boldsymbol{u},\boldsymbol{T}_{\boldsymbol{\nu}}W)^{\prime}\,.$
The function $\textbf{1}_{\Omega}U$ is potentially discontinuous on $M$, and hence $\boldsymbol{\Xi}\big(\textbf{1}_{\Omega}U\big)$ is defined in distribution sense on $M$. Its definition and the second point (already proved) finally give
$\langle\boldsymbol{\Xi}\big(\textbf{1}_{\Omega}U\big),W\rangle$ $\mathrel{\,:=\,}\big\langle\textbf{1}_{\Omega}U,\boldsymbol{\Xi}W\big\rangle\mathrel{\,:=\,}\big(\textbf{1}_{\Omega}U,\boldsymbol{\Xi}W\big)_{M}\,=\,\big(U,\boldsymbol{\Xi}W\big)_{\Omega}$
$\,=\,\big(\boldsymbol{\Xi}U,W\big)_{\Omega}-((\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U),W)^{\prime}+(U,\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}W)^{\prime}$
$\,=\,\big(\textbf{1}_{\Omega}\boldsymbol{\Xi}U,W\big)-((\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U)_{+},W)^{\prime}+(U_{+},\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}W)^{\prime}$
$\,=\,\big(\textbf{1}_{\Omega}\boldsymbol{\Xi}U,W\big)-\langle(\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U)_{+}\otimes\delta_{\Gamma},W\rangle+\langle U_{+}\otimes\delta_{\Gamma},\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}W\rangle$
$\,=\,\big\langle\textbf{1}_{\Omega}\boldsymbol{\Xi}U,W\big\rangle-\langle(\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U)_{+}\otimes\delta_{\Gamma},W\rangle+\langle\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}^{*}(U_{+}\otimes\delta_{\Gamma}),W\rangle\,.$
Because $W$ was arbitrary, this proves the relation (iii). ∎
We will need the following definition from [48].
Definition 2.12.
Let $M$ be a manifold and $E$ and $F$ be two vector bundles over $M$. If $M$ is connected, we say that a differential operator $T:L^{2}(M;E)\to\mathcal{C}^{\infty}_{c}(M;F)^{\prime}$ satisfies the $L^{2}$-unique continuation property if, given $u\in L^{2}(M;E)$ that vanishes in a non-empty open subset of $M$ and satisfies Tu= 0Tu\,=\,0, then u= 0u\,=\,0 everywhere on $M$. For general $M$, we say that $T$ satisfies the $L^{2}$-unique continuation property if it satisfies this property on any connected component of $M$.
Clearly, an operator that is injective satisfies the $L^{2}$-unique continuation property. The concept of $L^{2}$-unique continuation property just recalled allows us to obtain the following corollary. Recall that a Killing vector field $X$ is a vector field that preserves the metric, equivalently, $\operatorname{Def}\,X=0$. We shall need the following strong form of positivity for the matrix valued potential $V_{0}$ (in our previous works, we assumed that $V_{0}$ is a scalar multiple of the identity matrix, but here we allow it to be more general).
Definition 2.13.
Let $A$ be a topological space and $V:A\to M_{n}(\mathbb{C})$. We shall write $V\succ 0$ on $A$ if,
  1. (i)
    $V\geq 0$ on $A$ and,
  2. (ii)
    for every connected component $A_{0}$ of $A$, there is a point $a_{0}\in A_{0}$ such that $V(a_{0})>0$ (that is, in addition to $V(a_{0})\geq 0$, we also have that $V(a_{0})$ is invertible).
The same definition applies to sections of endomorphism bundles.
If $V$ is scalar valued or a scalar multiple of the identity $1\in M_{n}(\mathbb{C})$ (as in our previous papers on the subject), the condition $V\succ 0$ on $A$ is equivalent to $V\succ 0$ on every connected component of $A$. We have the following corollary.
Corollary 2.14.
Let $V,V_{0}\geq 0$ and $U=\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)\in H^{2}(\Omega;TM)\oplus H^{1}(\Omega)$ satisfy $\boldsymbol{\Xi}U=0$ in $\Omega$ and $(\boldsymbol{T}_{\boldsymbol{\nu}}U,\boldsymbol{u})^{\prime}=0$.
  1. (i)
    We have $\operatorname{Def}\,\boldsymbol{u}\,=\,0$, $V\boldsymbol{u}\,=\,0$, $\nabla^{*}\boldsymbol{u}\,=\,0$, $V_{0}p\,=\,0$, and $\nabla p\,=\,0$ in $\Omega$.
  2. (ii)
    If, furthermore, $V_{0}\succ 0$ in $\Omega$, then $p=0$ on $\Omega$.
  3. (iii)
    Similarly, if one of the following three conditions is satisfied:
    1. (a)
      No connected component of $\Omega$ has non-zero Killing vector fields;
    2. (b)
      $V\succ 0$ on $\Omega$; or
    3. (c)
      $\partial\Omega\neq\emptyset$ and $\boldsymbol{u}=0$ on $\partial\Omega$;
    then $\boldsymbol{u}=0$ in $\Omega$.
In particular, this corollary gives that $\boldsymbol{u}=0$ on $\operatorname{supp}(V)\cap\Omega$ and $p=0$ on $\operatorname{supp}(V_{0})\cap\Omega$. Recall that by the statement $\phi\not\equiv 0$ on AA,” we mean that there exists $a$ in the domain of $\phi$ such that $\phi(a)\not=0$. To negate this statement, we shall write $\phi=0$ in AA.”
Proof.
The real part $\operatorname{Re}(\mathfrak{w})$ of $\mathfrak{w}:=(p,\nabla^{*}\boldsymbol{u})_{\Omega}-(\nabla^{*}\boldsymbol{u},p)_{\Omega}$ vanishes. Let
$W\mathrel{\,:=\,}\left(\begin{array}[]{c}{\boldsymbol{w}}\\ {q}\end{array}\right)\,=\,\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {-p}\end{array}\right)\,=:\,U^{\prime}$
in the formula $\big(\boldsymbol{\Xi}U,W\big)_{\Omega}\,=\,B_{\Omega}(U,W)+(\boldsymbol{T}_{\boldsymbol{\nu}}U,\boldsymbol{w})^{\prime}$ of Proposition 2.11. Together with the definition of $B_{\Omega}$ in Notation 2.10 and with $\operatorname{Re}(\mathfrak{w})=0$, this gives
$0$ $\,=\,\operatorname{Re}\big[\big(\boldsymbol{\Xi}U,U^{\prime}\big)_{\Omega}-(\boldsymbol{T}_{\boldsymbol{\nu}}U,\boldsymbol{u})^{\prime}\big]\,=\,\operatorname{Re}\big[B_{\Omega}(U,U^{\prime})\big]$
$\,=\,\operatorname{Re}\big[2(\operatorname{Def}\,\boldsymbol{u},\operatorname{Def}\,\boldsymbol{u})_{\Omega}-(\nabla^{*}\boldsymbol{u},p)_{\Omega}+(p,\nabla^{*}\boldsymbol{u})_{\Omega}+(V\boldsymbol{u},\boldsymbol{u})_{\Omega}+(V_{0}p,p)_{\Omega}\big]$
= $\,=\,2(\operatorname{Def}\,\boldsymbol{u},\operatorname{Def}\,\boldsymbol{u})_{\Omega}+(V\boldsymbol{u},\boldsymbol{u})_{\Omega}+(V_{0}p,p)_{\Omega}\,.$
Because $V,V_{0}\geq 0$, all three terms in the last sum are non-negative, so each of them equals zero. Therefore $\operatorname{Def}\,\boldsymbol{u}=0$, $V\boldsymbol{u}=0$, and $V_{0}p=0$ in $\Omega$. We also have
$0\,=\,\boldsymbol{\Xi}U\,=\,\left(\begin{array}[]{c}{2\operatorname{Def}^{*}\,\operatorname{Def}\,\boldsymbol{u}+V\boldsymbol{u}+\nabla p}\\ {\nabla^{*}\boldsymbol{u}-V_{0}p}\end{array}\right)\,=\,\left(\begin{array}[]{c}{\nabla p}\\ {\nabla^{*}\boldsymbol{u}}\end{array}\right)\,$,
and hence we obtain (i). The condition $\nabla p=0$ just proved implies that $p$ is locally constant. Since, moreover, $V_{0}p=0$, this constant is zero on the connected components of $\Omega$ on which $V_{0}\not\equiv 0$, and this proves (ii). (Notice that this is exactly the $L^{2}$-unique continuation property of $\nabla$, see Definition 2.12.) Similarly, (iii) follows from the fact that $\operatorname{Def}\,$ satisfies the $L^{2}$-unique continuation property (see [48]). ∎
The above result holds true also if $\Omega=M$ (with the same proof). More precisely, we have the following consequences.
Corollary 2.15.
Let $V,V_{0}\geq 0$ and $U=\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)\in H^{2}(M;TM)\oplus H^{1}(M)$ satisfy $\boldsymbol{\Xi}U=0$ in $M$, which, we recall, is assumed to be connected. Then the following results hold:
  1. (i)
    $\operatorname{Def}\,\boldsymbol{u}\,=\,0$, $V\boldsymbol{u}\,=\,0$, $\nabla^{*}\boldsymbol{u}\,=\,0$, $V_{0}p\,=\,0$, and $\nabla p\,=\,0$ in $M$.
  2. (ii)
    If, furthermore, $V_{0}\succ 0$ in $M$, then $p=0$ on $M$.
  3. (iii)
    Similarly, if one of the following two conditions is satisfied:
    1. (a)
      $M$ has no non-zero Killing vector fields or
    2. (b)
      $V\succ 0$ on $M$
    then $\boldsymbol{u}=0$ in $M$.
We make the following simple observation.
Remark 2.16.
Let $\varepsilon$ be the diagonal matrix $1\oplus(-1)$. The last corollary and the proof of Corollary 2.14 give that the operator
$\varepsilon\boldsymbol{\Xi}\mathrel{\,:=\,}\varepsilon\boldsymbol{\Xi}_{V,V_{0}}\mathrel{\,:=\,}\left(\begin{array}[]{ccc}1&0\\ 0&-1\end{array}\right)\left(\begin{array}[]{ccc}\boldsymbol{L}+V&\nabla\\ \nabla^{*}&-V_{0}\end{array}\right)\,=\,\left(\begin{array}[]{ccc}\boldsymbol{L}+V&\nabla\\ -\nabla^{*}&V_{0}\end{array}\right)$
has positive real part, in the sense that $\operatorname{Re}(\varepsilon\boldsymbol{\Xi}U,U)_{M}\geq 0$ for all $U$ smooth with compact support.
3. General background material and manifolds with cylindrical ends
This section contains some preliminary material on pseudodifferential operators, normal lateral limits, Fourier transforms and their relation to normal lateral limits, and manifolds with straight cylindrical ends (definition and vector bundles and Sobolev spaces on these manifolds). Several related results are contained in the book [63], to which we refer for further details as well as for the concepts not defined here. See also [52, 53, 130] for related background material. Except the last subsection, where we introduce manifolds with straight cylindrical ends, in the other three subsections we work on $\mathbb{R}^{n}$.
3.1. Basic function spaces and some Fourier transforms
We now recall some basic function spaes, the Fourier transform, some other basic concepts, and compute some Fourier transforms that will be needed in what follows, especially for the limit and jump relations.
Let $p\in[1,\infty]$ and $(\mathfrak{M},\mu)$ be a measure space. If $u:\mathfrak{M}\to\mathbb{C}$ is a measurable function, then its $L^{p}$-norm is defined by
(3.1) $\|u\|_{L^{p}(\mathfrak{M},\mu)}\mathrel{\,:=\,}\,\begin{cases}\ \left(\int_{\mathfrak{M}}|u(x)|^{p}\,d\mu(x)\right)^{1/p}\,,&\quad\mbox{if }\ p<\infty\\ \ \,\operatorname{ess-sup}_{x\in X}\,|u(x)|\,,&\quad\mbox{if }\ p=\infty\,.\end{cases}$
Of course, it is possible that $\|u\|_{L^{p}(\mathfrak{M},\mu)}=\infty$. We identify functions that coincide except on a zero measure set, to obtain the Lebesgue spaces
$L^{p}(\mathfrak{M};\mu)\mathrel{\,:=\,}\{u:\mathfrak{M}\to\mathbb{C}\mid\ u\mbox{ measurable and }\|u\|_{L^{p}(\mathfrak{M})}<+\infty\}/\ker(\|\cdot\|_{L^{p}(\mathfrak{M})})\,.$
When the measure $\mu$ is clear from the context, we shall write $L^{p}(\mathfrak{M})$ instead of $L^{p}(\mathfrak{M};\mu)$. For instance, unless otherwise explicitly stated, on $\mathbb{R}^{n}$ we shall consider the Lebesque measure. Thus, $L^{p}(\mathbb{R}^{n})$ denotes the usual $L^{p}$-spaces defined using the Lebesque measure.
Our convention is that $\mathbb{N}:=\{1,2,\ldots\}$ and we let $\mathbb{Z}_{+}:=\mathbb{N}\cup\{0\}$. The elements $\alpha\in\mathbb{Z}_{+}^{n}$ will be called multi-indices. If $x=(x_{1},x_{2},\ldots,x_{n})\in\mathbb{R}^{n}$ and $\alpha\in\mathbb{Z}_{+}^{n}$ is a multi-index, we let $x^{\alpha}:=x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\ldots x_{n}^{\alpha_{n}}$. We also let $\partial_{x_{j}}\partial_{j}:=\frac{\partial}{\partial x_{j}}$ be the $j$th partial derivative and $\partial^{\alpha}:=\partial_{1}^{\alpha_{1}}\partial_{2}^{\alpha_{2}}\ldots\partial_{n}^{\alpha_{n}}$. Recall then that
(3.2) $\mathcal{S}(\mathbb{R}^{n})\mathrel{\,:=\,}\{u:\mathbb{R}^{n}\to\mathbb{C}\mid x^{\alpha}\partial^{\beta}u\in L^{2}(\mathbb{R}^{n})\,,\ \forall\,\alpha,\beta\in\mathbb{Z}_{+}^{n}\}$
denotes the space of Schwartz functions on $\mathbb{R}^{n}$ (i.e. smooth rapidly decaying functions at infinity). By $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ we denote the dual of $\mathcal{S}(\mathbb{R}^{n})$, called the space of tempered distributions on $\mathbb{R}^{n}$.
We let $\mathcal{C}^{k}(\mathbb{R}^{n})$ denote the space of $k$-times differentiable functions on $\mathbb{R}^{n}$ and $\mathcal{C}(\mathbb{R}^{n}):=\mathcal{C}^{0}(\mathbb{R}^{n})$ (simply, the space of continuous functions on $\mathbb{R}^{n}$). If $\alpha\in\mathbb{Z}_{+}^{n}$, then $|\alpha|:=\alpha_{1}+\alpha_{2}+\ldots+\alpha_{n}$. We let $\mathcal{C}_{0}^{k}(\mathbb{R}^{n})$ denote the set of functions $f:\mathbb{R}^{n}\to\mathbb{C}$ such that $\lim_{|x|\to\infty}|\partial^{\alpha}f(x)|=0$, for all $|\alpha|\leq k$. As usual, $\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n})$ denotes the set of smooth functions with compact support in $\mathbb{R}^{n}$ (this definition extends to manifolds right away).
We let $\imath^{2}=-1$ and $x\cdot\xi:=\sum_{j=1}^{n}x_{j}\xi_{j}$ denote the inner product of two vectors $x,\xi\in\mathbb{R}^{n}$. The Fourier transform of tempered distributions will be denoted $\mathcal{F}:\mathcal{S}^{\prime}(\mathbb{R}^{n})\to\mathcal{S}^{\prime}(\mathbb{R}^{n})$. More precisely, our convention is that, if $f\in L^{1}(\mathbb{R}^{n})$, then $\mathcal{F}$ and its inverse $\mathcal{F}^{-1}$ and are given by
(3.3) $\begin{gathered}\hat{f}(x)\mathrel{\,:=\,}\mathcal{F}f(x)\mathrel{\,:=\,}\int_{\mathbb{R}^{n}}e^{-\imath x\cdot\tau}f(\tau)\,d\tau\qquad\mbox{and}\\ \mathcal{F}^{-1}f(x)\,=\,\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{n}}e^{\imath x\cdot\tau}f(\tau)\,d\tau\,.\end{gathered}$
It follows that $\mathcal{F}f,\mathcal{F}^{-1}f\in\mathcal{C}_{0}(\mathbb{R}^{n}):=\mathcal{C}_{0}^{0}(\mathbb{R}^{n})$, because $f$ was assumed to be integrable.
Let us recall a few basic calculations with the Fourier transform for later use.
Remark 3.1.
First, for $f\in\mathcal{S}(\mathbb{R}^{n})$, we have
$\mathcal{F}(f^{\prime})(x)\,=\,\int_{\mathbb{R}}e^{-\imath x\tau}f^{\prime}(\tau)\,d\tau\,=\,-\int_{\mathbb{R}}\big(e^{-\imath x\tau}\big)^{\prime}f(\tau)\,d\tau\,=\,\imath x\hat{f}(x)\,.$
An equivalent way of writing this relation is $\mathcal{F}\partial_{x}f=\imath x\mathcal{F}f$, which remains valid for $f\in\mathcal{S}^{\prime}(\mathbb{R})$ by continuity using the density of $\mathcal{S}(\mathbb{R})$ in $\mathcal{S}^{\prime}(\mathbb{R})$ in the weak topology. Similarly, if $f\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$, $\partial_{j}:=\frac{\partial}{\partial x_{j}},$ and $x_{j}$ denotes also the operator of multiplication by $x_{j}$, then we have the following equalities of tempered distributions: $\mathcal{F}(\partial_{j}f)\,=\,\imath x_{j}\mathcal{F}(f)$, $\mathcal{F}(\tau_{j}f)\,=\,\imath\partial_{j}\mathcal{F}(f)$, $\mathcal{F}^{-1}(\partial_{j}f)\,=\,-\imath x_{j}\mathcal{F}^{-1}(f)$, and $\mathcal{F}^{-1}(\tau_{j}f)\,=\,-\imath\partial_{j}\mathcal{F}^{-1}(f).$
Recall that $\operatorname{p.\!\!v.}\,x^{-1}=:\operatorname{\mathfrak{V}}$ denotes the distribution
(3.4) $\langle\operatorname{p.\!\!v.}\,x^{-1},\phi\rangle\mathrel{\,:=\,}\lim_{\epsilon\searrow 0}\ \int_{|x|>\epsilon}\frac{\phi(x)}{x}\,dx\,=:\,\langle\operatorname{\mathfrak{V}},\phi\rangle\,,\ \ \phi\in\mathcal{S}(\mathbb{R})\,.$
(sometimes called Hadamard’s “partie finie” integral). (We use $\operatorname{\mathfrak{V}}:=\operatorname{p.\!\!v.}\,x^{-1}$ in order to streamline the notation.) We shall need the explicit form of the Fourier transform $\mathcal{F}\operatorname{\mathfrak{V}}:=\mathcal{F}(\operatorname{p.\!\!v.}\,x^{-1})$. This is well-known and is expressed in terms of the $\operatorname{sgn}$ function:
[…]
(the value $\operatorname{sgn}(0)$ can, in fact, be chosen arbitrarily). As usual, we let
$u(0+)\mathrel{\,:=\,}\lim_{\epsilon\searrow 0}u(\epsilon)\quad\mbox{and}\quad u(0-)\mathrel{\,:=\,}\lim_{\epsilon\searrow 0}u(-\epsilon)\,$,
when these limits exist. For example, $\operatorname{sgn}(0+)-\operatorname{sgn}(0-)=2$.
For any function $f:\mathbb{R}\to\mathbb{C}$, we let
(3.6) $\tilde{f}(x)\mathrel{\,:=\,}f(-x)\,$,
a definition that extends, by duality, to distributions. As for functions, a distribution $u\in\mathcal{C}^{\infty}_{c}(\mathbb{R})^{\prime}$ is called even (respectively, odd) if $\tilde{u}=u$ (respectively, $\tilde{u}=-u$). It is well-known (and very easy to check) that $\mathcal{F}(\tilde{u})=\widetilde{\mathcal{F}(u)}$ for all $u\in\mathcal{S}^{\prime}(\mathbb{R})$. This shows that the Fourier transform of an even (respectively, odd) tempered distribution is even (respectively, odd). This observation will be needed for the proof of the following result (which is undoubtedly well-known, see, for instance, Proposition 8.2 in Chapter 3 in [130]; it was stated without proof in [63]). For $x\in\mathbb{R}^{n}$, we let, as usual
(3.7) $\langle x\rangle\mathrel{\,:=\,}\sqrt{1+|x|^{2}}\,=\,\big(1+x_{1}^{2}+\ldots+x_{n}^{2}\big)^{1/2}\,.$
Lemma 3.2.
Let $\operatorname{\mathfrak{V}}:=\operatorname{p.\!\!v.}\,x^{-1}$ be as in (3.4). Then
  1. (i)
    $\operatorname{\mathfrak{V}}\in\mathcal{S}^{\prime}(\mathbb{R})$ and $\mathcal{F}^{-1}\operatorname{\mathfrak{V}}=\frac{\imath}{2}\operatorname{sgn}$.
  2. (ii)
    Let $\chi_{0}\in\mathcal{C}^{\infty}_{c}(\mathbb{R})$ be an even function with $\chi_{0}(x)=1$ for $x$ close to 0, let
    $Z\mathrel{\,:=\,}(1-\chi_{0})\operatorname{\mathfrak{V}}\in\mathcal{S}^{\prime}(\mathbb{R})\,,\ \mbox{ and let }\ W\mathrel{\,:=\,}\mathcal{F}^{-1}Z\mathrel{\,:=\,}\mathcal{F}^{-1}\big((1-\chi_{0})\operatorname{\mathfrak{V}})\,.$
    Then $W$ extends to a smooth function on each of the intervals $0,\infty)$ and $(-\infty,0]$ and $W(0+)=\frac{\imath}{2}=-W(0-)$.
  3. (iii)
    Also, for all $j,k\in\mathbb{Z}_{+}$, there exist $C_{j,k}>0$ such that, for all $x\neq 0$, we have
    $|\partial_{x}^{j}W(x)|\leq C_{j,k}\langle x\rangle^{-k}\,.$
    The same result holds for the restriction of $W$ to $(-\infty,0)$.
Compare with Proposition 8.2 in Chapter 3 in [130].
Proof.
Let $\delta_{0}$ be the Dirac distribution at $0\in\mathbb{R}$. The derivative $\operatorname{sgn}^{\prime}$ of the $\operatorname{sgn}$ function of Equation (3.5) is given (in distribution sense) by
$\operatorname{sgn}^{\prime}\,=\,[\operatorname{sgn}(0+)-\operatorname{sgn}(0-)]\delta_{0}\,=\,2\delta_{0}\,.$
Remark (3.1) then gives
$\imath x\mathcal{F}(\operatorname{sgn})(x)\,=\,\mathcal{F}(\operatorname{sgn}^{\prime})(x)\,=\,2\mathcal{F}(\delta_{0})(x)\,=\,2\,.$
It can be checked directly that $x\operatorname{\mathfrak{V}}:=x\big(\operatorname{pv}\,x^{-1}\big)=1$. Therefore
$x\,\Big(\frac{\imath}{2}\mathcal{F}(\operatorname{sgn})(x)-\operatorname{\mathfrak{V}}\Big)\,=\,0\in\mathcal{S}^{\prime}(\mathbb{R})\,.$
Thus the difference $\frac{\imath}{2}\mathcal{F}(\operatorname{sgn})(x)-\operatorname{\mathfrak{V}}$ is a multiple of $\delta_{0}$. Since both $\mathcal{F}(\operatorname{sgn})(x)$ and $\operatorname{\mathfrak{V}}$ are odd distributions and $\delta_{0}$ is even, it follows that $\frac{\imath}{2}\mathcal{F}(\operatorname{sgn})(x)-\operatorname{\mathfrak{V}}=0$. Applying $\mathcal{F}^{-1}$ to this equality yields the point (i).
To prove (ii), we first notice that the distribution $\chi_{0}\operatorname{\mathfrak{V}}$ is compactly supported, and hence the distribution $u:=\mathcal{F}^{-1}(\chi_{0}\operatorname{\mathfrak{V}})$ is given by a smooth function [52]. (Although we will not use this, note that $u$ is the convolution of $\mathcal{F}^{-1}\operatorname{\mathfrak{V}}$ with a function in $\mathcal{S}$ with integral one.) Since,
$W\mathrel{\,:=\,}\mathcal{F}^{-1}\big((1-\chi_{0})\operatorname{\mathfrak{V}}\big)\,=\,\mathcal{F}^{-1}\operatorname{\mathfrak{V}}-\mathcal{F}^{-1}(\chi_{0}\operatorname{\mathfrak{V}})\,=\,\frac{\imath}{2}\operatorname{sgn}-u\,$,
we obtain that $W$ is given by a locally $L^{1}$ function that is smooth outside 0 and extends to a smooth function on $0,\infty)$. Moreover, $W$ is an odd function and has jump $W(0+)-W(0-)=\frac{\imath}{2}(\operatorname{sgn}(0+)-\operatorname{sgn}(0-))=\imath$ at 0. Therefore $W$ also extends to a smooth function on $(-\infty,0]$ and $W(0+)=-W(0-)=\frac{\imath}{2}$.
It remains to prove (iii), that is, that $\partial_{x}^{j}W$ has rapid decay at $\infty$. Indeed, for all $k\geq 1$, $\big[(1-\chi_{0})\operatorname{\mathfrak{V}}\big]^{(k)}\in L^{1}(\mathbb{R})$, and hence we have
(3.8) $(-\imath x)^{k}W\,=\,(-\imath x)^{k}\mathcal{F}^{-1}\big[(1-\chi_{0})\operatorname{\mathfrak{V}}\big]\,=\,\mathcal{F}^{-1}\Big[\big[(1-\chi_{0})\operatorname{\mathfrak{V}}\big]^{(k)}\Big]\in\mathcal{C}_{0}(\mathbb{R})\,.$
The last displayed equation gives that, for all $k\geq 1$ (and, hence, for all $k$), there exists $C_{k}>0$ such that $|W(x)|\leq C_{k}|x|^{-k}$, for all $|x|\geq 1$. This proves the result for $j=0$, and hence $W$ has rapid decay at $\pm\infty$.
Recall that $u:=\mathcal{F}^{-1}(\chi_{0}\operatorname{\mathfrak{V}}):=\mathcal{F}^{-1}(\chi_{0}[\operatorname{p.\!\!v.}\,x^{-1}])$. To obtain the result for the other values of $j$ (and, hence, prove (iii)), let us use Remark (3.1) to obtain that
$\partial_{x}u\mathrel{\,:=\,}\partial_{x}\mathcal{F}^{-1}(\chi_{0}\operatorname{\mathfrak{V}})\,=\,\imath\mathcal{F}^{-1}(x\chi_{0}\operatorname{\mathfrak{V}})\,=\,\imath\mathcal{F}^{-1}(\chi_{0})\in\mathcal{S}(\mathbb{R})\,.$
By integrating to obtain $u$ from $\partial_{x}u$ (up to a constant), this gives that there exists a constant $L\in\mathbb{C}$ such that $u|_{[0,\infty)}-L\in\mathcal{S}([0,\infty))$. Since $W(x)=\frac{\imath}{2}\operatorname{sgn}(x)-u(x)$ has zero limits at infinity (see also Equation (3.8)), we obtain that $W|_{[0,\infty)}\in\mathcal{S}([0,\infty))$, as claimed. (We also obtain $L=\frac{\imath}{2}$, but that is not needed for the proof.) ∎
As a consequence, we obtain the following lemma on jump-values of certain Fourier transforms. It contains one of the most important calculations for the proof of jump relations (see Lemma A.5.5 from [63]).
Lemma 3.3.
Let $L\in\mathbb{R}$ and $u:\mathbb{R}\to\mathbb{R}$ be a continuous function such that there exist $C\geq 0$ and $\epsilon>0$ such that, for all $x\in\mathbb{R}$, we have
(3.9) $\langle x\rangle^{\epsilon}|xu(x)-L|\leq C\,.$
Let $\tilde{u}(x)=u(-x)$, as before. Then $u+\tilde{u}$ is integrable, $u$ is a tempered distribution (i.e., $u\in\mathcal{S}^{\prime}(\mathbb{R})$), and its inverse Fourier transform $\mathcal{F}^{-1}u$ is a function that is continuous everywhere, except maybe at $0\in\mathbb{R}$, with one-sided limits in 0 given by
$\begin{gathered}\mathcal{F}^{-1}u(0+)+\mathcal{F}^{-1}u(0-)\,=\,\frac{1}{2\pi}\,\int_{\mathbb{R}}\big[u(x)+u(-x)\big]\,dx\,,\\ \mathcal{F}^{-1}u(0+)-\mathcal{F}^{-1}u(0-)\,=\,\imath L\,,\quad\mbox{ and}\\ \lim_{x\to\infty}xu(x)\,=\,\lim_{x\to-\infty}xu(x)\,=\,L\,.\end{gathered}$
Proof.
As in Lemma 3.2, let $\chi_{0}\in\mathcal{C}^{\infty}_{c}(\mathbb{R})$ be an even function with $\chi_{0}=1$ in a neighborhood of 0 and let $\operatorname{\mathfrak{V}}:=\operatorname{p.\!\!v.}\,x^{-1}\in\mathcal{S}^{\prime}(\mathbb{R})$. Using these distributions, we now decompose our given function $u$ as
(3.10) $u\,=\,L\operatorname{\mathfrak{V}}+w_{1}+w_{2}\in\mathcal{S}^{\prime}(\mathbb{R})\,$,
where $w_{1}:=(u-L\operatorname{\mathfrak{V}})\chi_{0}$, $w_{2}:=(u-L\operatorname{\mathfrak{V}})(1-\chi_{0})$, and $L\in\mathbb{R}$ is as in the statement.
The distribution $w_{1}:=(u-L\operatorname{\mathfrak{V}})\chi_{0}$ is compactly supported, and hence $\mathcal{F}^{-1}w_{1}$ is a smooth function on $\mathbb{R}$ (an easy fact, see, for instance, [52]). In turn, the function $w_{2}:=(u-L\operatorname{\mathfrak{V}})(1-\chi_{0})$ is a continuous function on $\mathbb{R}$ that vanishes in a neighborhood of 0. Our assumptions then imply that there exists $C^{\prime}\geq 0$ such that, for all $x\in\mathbb{R}$, we have $|w_{2}(x)|\leq C^{\prime}\langle x\rangle^{-1-\epsilon}$. Consequently, $w_{2}\in L^{1}(\mathbb{R})$ and thus $\mathcal{F}^{-1}w_{2}$ is continuous. Combining with Equation (3.10), we obtain
(3.11) $\mathcal{F}^{-1}u-L\mathcal{F}^{-1}\operatorname{\mathfrak{V}}\,=\,\mathcal{F}^{-1}w_{1}+\mathcal{F}^{-1}w_{2}\in\mathcal{C}(\mathbb{R})\,.$
Therefore $\mathcal{F}^{-1}u$ and $L\mathcal{F}^{-1}\operatorname{\mathfrak{V}}$ have the same jump at 0. The relation $\mathcal{F}^{-1}\operatorname{\mathfrak{V}}:=\mathcal{F}^{-1}(\operatorname{p.\!\!v.}\,x^{-1}=\frac{\imath}{2}$ of Lemma 3.2(i) then gives
(3.12) $\mathcal{F}^{-1}u(0+)-\mathcal{F}^{-1}u(0-)\,=\,L\big[\mathcal{F}^{-1}\operatorname{\mathfrak{V}}(0+)-\mathcal{F}^{-1}\operatorname{\mathfrak{V}}(0-)\big]\,=\,\imath L\,.$
This proves the second displayed relation.
To prove the first relation, let $w:=u+\tilde{u}$, where $\tilde{u}(x)=u(-x)$, as before. Then the assumption gives that there exists $C^{\prime\prime}>0$ such that, for all $x\in\mathbb{R}$, we have $|w(x)|\leq C^{\prime\prime}\langle x\rangle^{-1-\epsilon}$. Consequently, $w$ is integrable and hence $\mathcal{F}^{-1}w$ is continuous on $\mathbb{R}$. Combining this fact with the relation $\mathcal{F}(\tilde{u})=\widetilde{\mathcal{F}(u)}$, we obtain
$\mathcal{F}^{-1}u(0+)+\mathcal{F}^{-1}u(0-)\,=\,\mathcal{F}^{-1}u(0+)+\widetilde{\mathcal{F}^{-1}u}(0+)\,=\,\mathcal{F}^{-1}u(0+)+\mathcal{F}^{-1}\tilde{u}(0+)\\ \,=\,\mathcal{F}^{-1}w(0+)\,=\,\mathcal{F}^{-1}w(0)\,=\,\frac{1}{2\pi}\int_{\mathbb{R}}w(x)dx\,=\,\frac{1}{2\pi}\int_{\mathbb{R}}\big[u(x)+u(-x)\big]dx\,.$
The last statement of this lemma (the limit) follows the assumption that $\langle x\rangle^{\epsilon}|xu(x)-L|\leq C$, which gives $\lim_{|x|\to\infty}(xu(x)-L)=0$. ∎
We will use the last lemma in the following form.
Corollary 3.4.
Let $u:\mathbb{R}\to\mathbb{R}$ be a function satisfying the assumptions of Lemma 3.3 (that is, satisfying the condition $\langle x\rangle^{\epsilon}|xu(x)-L|\leq C$ of Equation (3.9)). Then
  1. (i)
    $u$ defines a tempered distribution on $\mathbb{R}$;
  2. (ii)
    $\mathcal{F}^{-1}u$ is given by a locally integrable function that is continuous on $\mathbb{R}\smallsetminus\{0\}$;
  3. (iii)
    the following limit exists:
    $\operatorname{\mathfrak{L}}(u)\mathrel{\,:=\,}\lim_{|x|\to\infty}xu(x)\,;$
  4. (iv)
    $u+\tilde{u}$ is integrable on $\mathbb{R}$, and
    (3.13) $\mathcal{F}^{-1}u(0\pm)\,=\,\pm\frac{\imath\operatorname{\mathfrak{L}}(u)}{2}+\frac{1}{2\pi}\,\int_{\mathbb{R}}\frac{u(x)+u(-x)}{2}\,dx\,.$
3.2. Pseudodifferential operators on $\mathbb{R}^{n}$
Most of the operators that we will work with (if not all of them!) will be pseudodifferential operators. We thus recall now some basic definitions that are needed in the following. First, we recall that for $x\in\mathbb{R}^{n}$, we use the notation $\langle x\rangle\mathrel{\,:=\,}\sqrt{1+|x|^{2}}\,=\,\big(1+x_{1}^{2}+\ldots+x_{n}^{2}\big)^{1/2},$ see Equation (3.7).
Let $\mathbb{Z}_{+}^{n}:=\left\{\alpha=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{Z}^{n}:\alpha_{i}\geq 0\,,\ 1\leq i\leq n\right\}$, and, for $\alpha\in\mathbb{Z}_{+}^{n}$,
$\partial_{x}^{\alpha}\mathrel{\,:=\,}\partial^{\alpha}\mathrel{\,:=\,}\frac{\partial^{\alpha_{1}}}{\partial x_{1}^{\alpha_{1}}}\cdots\frac{\partial^{\alpha_{n}}}{\partial x_{n}^{\alpha_{n}}}\,$,
as before. We now recall the classical, well-known definition of Kohn-Nirenberg symbols (see also [53])
Definition 3.5.
We let $S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ denote the set of functions $a:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{C}$ such that, for every pair of multi-indices $\alpha,\beta\in\mathbb{Z}_{+}^{n}$, there exists $C_{\alpha\beta}\geq 0$ such that
$|\partial_{x}^{\alpha}\partial_{\xi}^{\beta}a(x,\xi)|\leq C_{\alpha\beta}\langle\xi\rangle^{m-|\beta|}\,.$
  1. (i)
    The function $a$ will be called a symbol.
  2. (ii)
    For $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, we define $a(x,D):\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n})\to\mathcal{C}^{\infty}(\mathbb{R}^{n})$ by the formula
    $a(x,D)u(x)\mathrel{\,:=\,}\frac{1}{(2\pi)^{n}}\,\int_{\mathbb{R}^{n}}e^{\imath x\cdot\xi}a(x,\xi)\hat{u}(\xi)\,d\xi\,.$
    The operator $a(x,D)$ is called the pseudodifferential operator with symbol $a$.
For $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, we shall denote by $a(x,\xi)$ its values, where $x,\xi\in\mathbb{R}^{n}$. The elements of $M_{k}(S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n}))$ are also called symbols. See also [54, 63, 130]. In order to clarify some later conventions, let us mention here that the second variable in $a(x,\xi)$, namely $\xi$, is really in the dual of $\mathbb{R}^{n}$, which is nevertheless identified with $\mathbb{R}^{n}$ using the standard inner product (this will be relevant for our study of the jump relations, where the last basis vector in the $\xi$ variable will be denoted $e_{n}^{\sharp}$ instead of simply $e_{n}=(0,\ldots,0,1)$).
The operator $a(x,D)$ is a particular case of a pseudodifferential operator on $\mathbb{R}^{n}$ and, locally, every differential operator on $\mathbb{R}^{n}$ is of this form [53]. For our purposes, we will not need to use pseudodifferential operators more general than $a(x,D)$ on $\mathbb{R}^{n}$. We shall denote by $\Psi^{m}(\mathbb{R}^{n})$ the set of all pseudodifferential operators on $\mathbb{R}^{n}$, it is the set of operators $T:\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n})\to\mathcal{C}^{\infty}(\mathbb{R}^{n})$ such that, for all $\phi\in\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n})$, the operator $\phi T\phi$ is of the form $a(x,D)$ for a suitable symbol $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ depending on $T$ and $\phi$. This definition extends right away to operators acting between sections of vector bundles on manifolds. Thus, if $M$ is a smooth manifold and $E,F\to M$ are smooth vector bundles, then
[…]
where, this time, $a\in M_{k,l}(S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n}))$ depending on $T$ and $\phi$, $n$ is the dimension of $M$, $E$ has rank $l$, and $F$ has rank $k$.
If $b\in S^{m}(\mathbb{R}^{n-1}\times\mathbb{R}^{n-1})$, we let $b(x^{\prime},D^{\prime})$ denote its associated operator. In general, symbols with a prime (i.e. ) will refer to objects on $\mathbb{R}^{n-1}$. The most important example is that, if $x\in\mathbb{R}^{n}$, then $x^{\prime}\in\mathbb{R}^{n-1}$ and $x_{n}\in\mathbb{R}$ are its projections, so that $x=(x^{\prime},x_{n})$.
Recall that the distribution kernel $k_{a(x,D)}$ of $a(x,D)$ is the distribution on $\mathbb{R}^{n}\times\mathbb{R}^{n}$ such that, for all $\phi,\psi\in\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n})$, we have
(3.15) $\langle a(x,D)\phi,\psi\rangle\,=\,\langle k_{a(x,D)},\psi\boxtimes\phi\rangle\,$,
where $\psi\boxtimes\phi\in\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ is given by $\psi\boxtimes\phi(x,y)=\psi(x)\phi(y)$ and the pairings $\langle\ ,\ \rangle$ are those of distributions and test functions (for the first pairing, on $\mathbb{R}^{n}$, and, for the second pairing, on $\mathbb{R}^{n}\times\mathbb{R}^{n}$). For further use, let us record the following well-known lemma [53, 54, 130].
Lemma 3.6.
The distribution kernel $k_{a(x,D)}(x,y)$ of $a(x,D)$ is smooth for $x\neq y$ and there it is given by $k_{a(x,D)}(x,y)\,=\,\mathcal{F}_{\xi}^{-1}a(x,x-y)\,.$
A smooth function $b:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{C}$ will be called essentially homogeneous of order $m$ if
(3.16) $b(x,\lambda\xi)\,=\,\lambda^{m}b(x,\lambda\xi)\,,\qquad\mbox{for all }|\xi|\geq 1\ \mbox{ and }\ \lambda|\xi|\geq 1\,.$
A symbol $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ will be called classical if there exist essentially homogeneous symbols $a_{j}\in S^{m-j}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ of order mjm-j such that, for every $N\in\mathbb{N}$, $a-\sum_{j=0}^{N}a_{j}\in S^{m-N-1}(\mathbb{R}^{n}\times\mathbb{R}^{n})$. The set of classical symbols of order $m$ will be denoted $S^{m}_{\operatorname{cl}}(\mathbb{R}^{n}\times\mathbb{R}^{n})$. We now recall the definition of the principal symbol, which will play an important role in what follows.
Definition 3.7.
If $a\in S^{m}_{\operatorname{cl}}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ and $\xi\in\mathbb{R}^{n}$, $\xi\neq 0$, we let
$\sigma_{m}(a;x,\xi)\mathrel{\,:=\,}\lim_{\lambda\to\infty}\lambda^{-m}a(x,\lambda\xi)$
denote its principal part (or symbol). We shall say that $\sigma_{m}(a(x,D)):=\sigma_{m}(a)$ is the principal symbol of the pseudodifferential operator $a(x,D)$.
Therefore, $\sigma_{m}(a):\mathbb{R}^{n}\times(\mathbb{R}^{n}\smallsetminus\{0\})\to\mathbb{C}$ satisfies $\sigma_{m}(a;x,\lambda\xi)=\lambda^{m}\sigma_{m}(a;x,\xi)$, for all $\lambda>0$. For example,
(3.17) $\sigma_{1}(\partial_{j};x,\xi)\,=\,\imath\xi_{j}\,.$
Given $a\in S_{\operatorname{cl}}^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, we can always find an essentially homogeneous function $b_{m}$ of order $m$, $b_{m}\in S_{\operatorname{cl}}^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, such that $\sigma_{m}(a;x,\xi)=b_{m}(x,\xi)$, for all $|\xi|\geq 1$ [53, 54, 131]. We shall say that $b_{m}$ is an essential homogeneous representative of $\sigma_{m}(a)$. Also, note that $a\in S^{m}_{\operatorname{cl}}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ will have a different principal symbol if regarded as an element of $S^{m+1}_{\operatorname{cl}}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, in fact, then, $\sigma_{m+1}(a)=0$.
3.3. Traces and normal lateral limits on $\mathbb{R}^{n}$
We shall need some general results on the normal lateral limits $a(x,D)u]_{\pm}$ at $\Gamma:=\{x_{n}=0\}$ of the value $a(x,D)u$ of a pseudodifferential operator $a(x,D)$ acting on suitable distribution $u=h\otimes\delta_{\Gamma}$ supported on $\Gamma$. This problem will be formulated precisely next and will be studiend in detail on Euclidean space in Section 4 and on manifolds with cylindrical ends in Section 6 following [63, 67, 68].
It will be convenient to use the following notation. Let $\Gamma_{\epsilon}\subset\mathbb{R}^{n}$ be the hyperplane
(3.18) $\Gamma_{\epsilon}\mathrel{\,:=\,}\{x=(x^{\prime},x_{n})\in\mathbb{R}^{n-1}\times\mathbb{R}\mid x_{n}=\epsilon\}\,,\qquad\Gamma:=\Gamma_{0}$
with the induced Euclidean measure. We parameterize $\Gamma_{\epsilon}$ via the diffeomorphism $\Gamma\ni(x^{\prime},\epsilon)\mapsto x^{\prime}\in\mathbb{R}^{n-1}$. Recall that by $\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n})$ we denote the set of smooth functions with compact support in $\mathbb{R}^{n}$. We let $e_{n}=(0,0,\ldots,0,1)\in\mathbb{R}^{n}$. Then $\Gamma_{\epsilon}\mathrel{\,:=\,}\Gamma+\epsilon e_{n}\mathrel{\,:=\,}\{x=(x^{\prime},x_{n})\in\mathbb{R}^{n}:x_{n}=\epsilon\}\subset\mathbb{R}^{n}$. Then $\tau_{-\epsilon}:H^{s}(\Gamma_{\epsilon})\to H^{s}(\Gamma)$ will be the natural isometries induced by translations $\tau_{\epsilon}(x):=x-\epsilon e_{n}$, $\tau_{\epsilon}:\Gamma_{\epsilon}\to\Gamma_{0}=\Gamma$. (Note that, in this setting, the role of $\Omega$ is played by $\Omega=\mathbb{R}_{+}^{n}:=\{x_{n}>0\}$, so that $\Gamma:=\partial\Omega$, as usual.)
The following lemma is often used in this paper (see also Lemma A.5.1 in [63]).
Lemma 3.8.
Let $\Gamma:=\{x_{n}=0\}\simeq\mathbb{R}^{n-1}$ be as in (3.18), $s\in\mathbb{R}$, and $h\in H^{s}(\Gamma)$, $s\in\mathbb{R}$.
  1. (i)
    Let $s\geq 0$ and let $h\otimes\delta_{\Gamma}$ be defined by
    $\langle h\otimes\delta_{\Gamma},\phi\rangle\mathrel{\,:=\,}\int_{\mathbb{R}^{n-1}}h(x^{\prime})\phi(x^{\prime},0)\,dx^{\prime}\,,\qquad\forall\,\phi\in\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n})\,.$
    Then $h\otimes\delta_{\Gamma}\in H^{s^{\prime}}(\mathbb{R}^{n};E)$ for all $s^{\prime}<-1/2$.
  2. (ii)
    Let $\langle\ ,\ \rangle_{\Gamma}$ be the pairing of distributions and test functions on $\Gamma$. If $s<0$, the formula
    $\langle h\otimes\delta_{\Gamma},\phi\rangle\mathrel{\,:=\,}\langle h,\phi|_{\Gamma}\rangle_{\Gamma}\,,\qquad\forall\,\phi\in\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n})\,$,
    defines a distribution $h\otimes\delta_{\Gamma}\in H^{s-1/2}(\mathbb{R}^{n};E)$.
Proof.
Let us assume first that $s^{\prime}<-\frac{1}{2}$, $u\in H^{-s^{\prime}}(\mathbb{R}^{n})$, and $s=0$ (so $h\in L^{2}(\Gamma)$). Then the restriction (or trace) $u|_{\Gamma}\in L^{2}(\Gamma)$ is well defined and depends continuously on $u$, by classical results [130], because $s^{\prime}>1/2$. Consequently, $\int_{\mathbb{R}^{n-1}}u(x^{\prime},0)h(x^{\prime})\,dx^{\prime}$ is also well defined and depends continuously on $u$, hence it defines an element in $H^{-s^{\prime}}(\mathbb{R}^{n})^{*}\simeq H^{s^{\prime}}(\mathbb{R}^{n})$, as claimed. Thus $h\otimes\delta_{\Gamma}\in H^{s^{\prime}}(\mathbb{R}^{n})$.
Finally, let us consider $s<0$. Let $u\in H^{-s+1/2}(\mathbb{R}^{n})$. Then $u|_{\Gamma}\in H^{-s}(\Gamma)$, because s+1/2>1/2-s+1/2>1/2 and the pairing $\langle h,u|_{\Gamma}\rangle_{\Gamma}$ is well defined and continuous with respect to $u$. Therefore $h\otimes\delta_{\Gamma}\in H^{s-1/2}(\mathbb{R}^{n})$. This completes the proof. ∎
Let $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ and $a(x,D)$ be as in Definition 3.5(ii), $m\in\mathbb{R}$. Then we define the potential operator
(3.19) $\mathcal{S}_{a}h\mathrel{\,:=\,}\mathcal{S}_{a(x,D)}h\mathrel{\,:=\,}a(x,D)(h\otimes\delta_{\Gamma})\,.$
Let s<1/2s<-1/2 and $h\in H^{s}(\Gamma)$. Then Lemma 3.8 and the mapping properties of the pseudodifferential operators [53, 67, 130] show that
$\mathcal{S}_{a}h\mathrel{\,:=\,}\mathcal{S}_{a(x,D)}h\mathrel{\,:=\,}a(x,D)(h\otimes\delta_{\Gamma})\in H^{s-m}(\mathbb{R}^{n})\,.$
Thus, if m<1m<-1, we can choose $s$ close to 1/2-1/2 such that sm>1/2s-m>1/2, and hence the trace $(\mathcal{S}_{a}h)|_{\Gamma}$ is well defined. (In the case m=1m=-1, this result does not hold. See Theorem 4.16 for a detailed discussion in this case.)
Definition 3.9.
If $u:\mathbb{R}^{n}\to\mathbb{C}$ is smooth enough, we then let
[…]
whenever these definitions make sense.
It is known that, for a suitable pseudodifferential operator $a$ and $\epsilon\neq 0$, $a_{\epsilon}(x^{\prime},D^{\prime})$ is defined and is again a pseudodifferential operator. We will prove this in the cases in which we are interested in Section 4 and we will study the properties of the family $a_{\epsilon}$, $\epsilon\neq 0$.
We let $\mathbb{R}_{\pm}^{n}:=\{x=(x^{\prime},x_{n})\in\mathbb{R}^{n-1}\times\mathbb{R}\mid\pm x_{n}>0\}$ be the two open half-spaces defined by the hyperplane $\Gamma=\{x_{n}=0\}$. Let $s>1/2$. If $u\in H_{\operatorname{loc}}^{s}(\mathbb{R}_{\pm}^{n})$, we let
[…]
using, of course, that $\Gamma=\partial\mathbb{R}_{\pm}^{n}$. If, furthermore, $u\in H_{\operatorname{loc}}^{s}(\mathbb{R}^{n})$, we shall write $u|_{\Gamma}:=\gamma_{+}(u)=\gamma_{-}(u)$. Then, we have the following useful result.
Lemma 3.10.
Let $u\in H_{\operatorname{loc}}^{s}(\mathbb{R}_{\pm}^{n})$, $s>1/2$. Then
$u_{\pm}\mathrel{\,:=\,}\lim_{t\searrow 0}\tau_{\mp t}u_{t}\,=\,\gamma_{\pm}(u)\in H_{\operatorname{loc}}^{s-1/2}(\Gamma)\,.$
In particular, if $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ for some m<1m<-1 and $h\in L^{2}(\mathbb{R}^{n})$, then
$\big[a(x,D)(h\otimes\delta_{\Gamma})\big]_{+}\,=\,\big[a(x,D)(h\otimes\delta_{\Gamma})\big]_{-}\,=\,\big[a(x,D)(h\otimes\delta_{\Gamma})\big]_{\Gamma}\,.$
Proof.
The first relation of the lemma is clearly true if $u\in\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n})$. The first relation then follows from the uniform continuity (in $t$) of the restrictions $H_{\operatorname{loc}}^{s}(\mathbb{R}_{+}^{n})\to H_{\operatorname{loc}}^{s-1/2}(\Gamma+te_{n})$ and the density of $\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n})$ in $H_{\operatorname{loc}}^{s}(\mathbb{R}_{+}^{n})$. The result on $\mathbb{R}_{-}^{n}$ is proved exactly in the same way.
The result on pseudodifferential operators follows from the mapping properties of pseudodifferential operators, from the properties of the traces on the spaces $H_{\operatorname{loc}}^{s}$, and from Lemma 3.8. ∎
Typically, in this paper, we shall work with functions (or sections of vector bundles) to which the above lemma applies, and hence we will not need to distinguish between the limits $u_{\pm}$ and the traces $\gamma_{\pm}(u)$. Of course, there exist important situations when the limits $u_{\pm}$ exist but the traces $\gamma_{\pm}(u)$ do not exist.
3.4. Manifolds with cylindrical ends, compatible vector bundles and differential operators
We continue with some standard background material, but now we turn from the case of $\mathbb{R}^{n}$ to that of a manifold with cylindrical ends $M$. We begin by recalling a definition of manifolds with straight cylindrical ends that is equivalent to the one given in the introduction (see [63, 67, 68]).
Definition 3.11.
A manifold with straight cylindrical ends $M$ is a complete non-compact Riemannian manifold that is the disjoint union
$M\,=\,M_{0}\cup\left[M^{\prime}\times(-\infty,R_{M})\right]\,$,
where $(M_{0},g_{0})$ is a compact Riemannian manifold, $M^{\prime}:=\partial M_{0}\neq\emptyset$, $R_{M}\in\mathbb{R}$, and the cylindrical end $M^{\prime}\times(-\infty,R_{M})$ is endowed with a product metric. Any Riemannian manifold isometrically diffeomorphic to one of this form will also be called a manifold with straight cylindrical ends.
There is no loss of generality to assume that $M$ is connected, so we will do that throughout the paper, in order to simplify our statements. It follows from the fact that $M$ is complete that the isometric embedding $M^{\prime}\times(-\infty,0)\hookrightarrow M$ extends to an isometry $M^{\prime}\times(-\infty,0]\hookrightarrow M$ that maps $M^{\prime}\times\{0\}$ to $M^{\prime}\simeq\partial M_{0}$ diffeomorphically. There is no loss of generality to assume that this diffeomorphism is the canonical one identifying $(x^{\prime},0)\in M^{\prime}\times\{0\}$ to $x^{\prime}\in M^{\prime}\simeq\partial M_{0}$. So, again, we will assume that this is the case throughout the paper. Also, there is no loss of generality to assume $R_{E}=0$, and, again, we will usually (but not always) do that. For instance, for a cylinder, it is convenient to take $R_{E}=-1$. The case of a cylinder is a very important one, so we discuss it more formally in the following remark.
Remark 3.12.
Recall that a closed manifold is a smooth, compact manifold without boundary. Let $\mathfrak{S}$ be a closed manifold. An important example of a manifold with straight cylindrical ends is $M=\mathfrak{S}\times\mathbb{R}$. In this case, $M_{0}=\mathfrak{S}\times[-1,1]$, $R_{M}=-1$ and $M^{\prime}:=\partial M_{0}=\mathfrak{S}\times\{-1,1\}$. Then we have a canonical isometry
$\mathfrak{S}\times\mathbb{R}\,=:\,M\simeq M_{0}\cup_{\psi}\big[\mathfrak{S}\times\{-1,1\}\times(-\infty,-1]\big]\,$,
where, on the right hand side $\psi:\mathfrak{S}\times\{-1,1\}\times\{-1\}\to\mathfrak{S}\times\{-1,1\}:=\partial M_{0}$ is given by $\psi(x^{\prime},-1)=x^{\prime}$, for all $x^{\prime}\in\mathfrak{S}\times\{-1,1\}$. (The symbol $\cup_{\psi}$ means that we identify $y$ to $\psi(y)$ in the union.)
On cylinders, it is important to notice that $M:=\mathfrak{S}\times\mathbb{R}$ carries a natural action of $\mathbb{R}$ by translations and we let $\mathcal{C}^{\infty}(M)^{\mathbb{R}}=\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R})^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(\mathfrak{S})$ denote the space of smooth functions on $\mathfrak{S}\times\mathbb{R}$ that are invariant by translations. Many objects on manifolds with cylindrical ends are defined by first defining their translation invariant versions on cylinders. If $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a general manifold with cylindrical ends, then $M^{\prime}\times(-\infty,0]$ is its end and $M^{\prime}\times\mathbb{R}$ is its associated cylinder with the induced Riemannian metric (induced from $\partial M_{0}$, which, in turn, is induced from the metric on $M$).
In the following, whenever we will write $M\,=\,M_{0}\cup\left[M^{\prime}\times(-\infty,0)\right]$, we will understand that $M$ is a manifold with straight cylindrical ends as in Definition 3.11. It is easy (and well-known) to prove (see e.g. [48] and the references therein) that all manifolds with straight cylindrical ends have bounded geometry, so the results of Subsection 2.3 apply to them.
Definition 3.13.
Let $M=M_{0}\cup\left[M^{\prime}\times(-\infty,0)\right]$ be a manifold with straight cylindrical ends (Definition 3.11) and let $E\to M$ be a Hermitian vector bundle endowed with a metric preserving connection $\nabla^{E}$. We shall say that the pair $(E,\nabla^{E})$ is compatible with the straight cylindrical ends structure on $M$ (simply, compatible) if there exists $R_{E}\leq 0$ such that $E|_{M^{\prime}\times(-\infty,R_{E})}$ is isometric to the pull-back to $M^{\prime}\times(-\infty,R_{E})$ of a Hermitian vector bundle $E^{\prime}\to M^{\prime}$ with connection $\nabla^{E^{\prime}}$ via an isomorphism that maps the connection $\nabla^{E}$ of $E$ to the pull-back connection of $\nabla^{E^{\prime}}$ on $M^{\prime}\times(-\infty,R_{E})$.
If $E=\mathbb{C}$ (the one dimensional trivial bundle), then $\nabla^{E}f:=df$, where $d$ is the de Rham differential. Let $\sharp:T^{*}M\to TM$ be the vector bundle isomorphism defined by the Riemannian metric $g$ of $M$. Note then that the gradient $\nabla f:=\sharp df$ has a similar notation to a connection, however the two should not be confused.
Let $E$ be a vector bundle on $M^{\prime}\times\mathbb{R}$ obtained from the pull-back of vector bundle on $M^{\prime}$. Thus $E\to M^{\prime}\times\mathbb{R}$ is, in, in particular, a compatible vector bunde on the manifold with straight cylindrical ends $M^{\prime}\times\mathbb{R}$. We let $\mathcal{C}^{\infty}(M^{\prime}\times\mathbb{R})^{\mathbb{R}}$ denote the set of translation invariant sections of $E\to M^{\prime}\times\mathbb{R}$. In the rest of this subsection, $M$ will be a manifold with straight cylindrical ends. (The results are true also for closed manifolds, see [69] and some of them are recalled in Subsection 12.2.) We next introduce an important class of smooth sections.
Definition 3.14.
Let $E\to M$ be a compatible vector bundle and let $u\in\mathcal{C}^{\infty}(M;E)$. We shall say that $u$ is translation invariant in a neighborhood of infinity if there exists $R_{u}\leq R_{E}$ such that $u(x^{\prime},t)=u(x^{\prime},s)$ for all $s,t\leq R_{u}$ and all $x^{\prime}\in M^{\prime}:=\partial M_{0}$. We let $\mathcal{C}^{\infty}_{\operatorname{inv}}(M;E)$ denote the set of sections of $E$ that are translation invariant in a neighborhood of infinity and let $\mathcal{C}^{\infty}(M^{\prime}\times\mathbb{R};E)^{\mathbb{R}}$ be the set of translation invariant sections on $M^{\prime}\times\mathbb{R}$. Then we let ${\mathcal{R}_{\infty}}:\mathcal{C}^{\infty}_{\operatorname{inv}}(M;E)\to\mathcal{C}^{\infty}(M^{\prime}\times\mathbb{R};E)^{\mathbb{R}}$ be defined by
${\mathcal{R}_{\infty}}(u)(x^{\prime},t)\,=\,u(x^{\prime},t)\,,\quad\mbox{ for }\ t\leq R_{u}\,.$
Remark 3.15.
If $u$ is as in Definition 3.14, then there exists $u_{0}\in\mathcal{C}^{\infty}(M^{\prime};E)$ such that $u(x^{\prime},t)=u_{0}(x^{\prime})$ for $t\leq R$. Thus ${\mathcal{R}_{\infty}}u(x^{\prime},t)\,=\,u_{0}(x^{\prime})\,=\,u(x^{\prime},t)\in E$ for $t\leq R_{u}$. Thus $u_{0}$ and ${\mathcal{R}_{\infty}}(u)$ do not coincide, but they correspond to each other under the natural isomorphism $\mathcal{C}^{\infty}(M^{\prime}\times\mathbb{R};E)^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(M^{\prime};E)$. Because of this, we will sometimes regard ${\mathcal{R}_{\infty}}(u)$ as an element of $\mathcal{C}^{\infty}(M^{\prime};E)$.
We let $\mathcal{C}^{\infty}_{\operatorname{inv}}(M):=\mathcal{C}^{\infty}_{\operatorname{inv}}(M;\mathbb{C})$. We next introduce differential operators following, for instance [66].
Definition 3.16.
Let $E,F\to M$ be compatible vector bundles on $M$ with connections $\nabla^{E}$ and $\nabla^{F}$. We let $\operatorname{Diff}_{\operatorname{inv}}^{m}(M;E,F)$ be the set of differential operators $P$ acting on sections of $E$ with values sections of $F$ with the property that there exist $a_{j}\in\mathcal{C}^{\infty}_{\operatorname{inv}}(M;\operatorname{Hom}(T^{*\otimes j}M\otimes E,F))$, $j=0,\ldots,m$, (see Definition 3.14) such that
$Pu\,=\,\sum_{j=0}^{m}a_{j}(\nabla^{E})^{j}u\,.$
For such a $P$, we let
${\mathcal{R}_{\infty}}(P)\mathrel{\,:=\,}\sum_{j=0}^{m}{\mathcal{R}_{\infty}}(a_{j})(\nabla^{E})^{j}$
be the associated differential operator on $M^{\prime}\times\mathbb{R}$ obtained by “freezing the coefficients” of $P$ near infinity. It will be called the limit at infinity operator associated to $P$.
Therefore, ${\mathcal{R}_{\infty}}(P)$ will be a translation invariant differential operator on $M^{\prime}\times\mathbb{R}$. An example is the Levi-Civita connection
$\nabla^{LC}\in\operatorname{Diff}^{1}_{\operatorname{inv}}(M;TM,T^{*}M\otimes TM)\,$,
in which case ${\mathcal{R}_{\infty}}\nabla^{LC}=\nabla^{LC}$.
For the rest of this section, $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ will be a manifold with straight cylindrical ends and $E\to M$ will be a fixed compatible Hermitian vector bundle.
Let us use the notation of Definition 3.13. Then the compatible vector bundle $E\to M$ will have a product structure on $M^{\prime}\times(-\infty,R_{E})$. Hence $\nabla^{E}=\nabla^{E^{\prime}}+dt\otimes\partial_{t}$ on $M^{\prime}\times(-\infty,R_{E})$, where $\nabla^{E^{\prime}}$ is the connection on $E^{\prime}\to M^{\prime}$. Consequently, the Bochner Laplacian of $E$ satisfies
$\Delta_{E}\mathrel{\,:=\,}(\nabla^{E})^{*}\nabla^{E}\,=\,(\nabla^{E^{\prime}})^{*}\nabla^{E^{\prime}}-\partial_{t}^{2}\,.$
Definition 3.17.
Let $\nabla^{E}$ denote the compatible connexion on the compatible vector bundle $E\to M$. Let $\Delta_{E}\mathrel{\,:=\,}(\nabla^{E})^{*}\nabla^{E}$ be the Bochner Laplacian. The Sobolev space $H^{s}(M;E)$ is defined as the domain of $(1+\Delta_{E})^{\frac{s}{2}}$ for $s\geq 0$. The norm on this space is $\|u\|_{H^{s}(M;E)}:=\|(1+\Delta_{E})^{\frac{s}{2}}u\|_{L^{2}(M;E)}^{1/2}$. For $s\leq 0$, the Sobolev space $H^{s}(M;E)$ is defined as the dual of the Sobolev space $H^{-s}(M;E)$.
See [4, 6, 7, 49, 50, 66, 67, 129].
Remark 3.18.
The space $\mathcal{C}^{\infty}_{c}(M;E)$ is dense in $H^{s}(M;E)$ for all $s\in\mathbb{R}$. This is a result that holds in the greater generality of manifolds with bounded geometry, see [4, 5, 10, 50, 119].
See [63] for a more thorough discussion of the Sobolev spaces on manifolds with cylindrical ends, including a definition using dyadic partitions of unity. See [50, 133] for a discussion of Sobolev spaces on manifolds with bounded geometry.
4. Normal lateral limits of pseudodifferential operators on $\mathbb{R}^{n}$
In this section we provide a detailed study on $\mathbb{R}^{n}$ of the normal lateral limits
$a(x,D)u]_{\pm}\in H^{s}(\Gamma)$
at $\Gamma:=\{x_{n}=0\}$ for $u=h\otimes\delta_{\Gamma}$. Here $a(x,D)$ is a pseudodifferential operator (Definition 3.5), the normal lateral limits were defined in Definition 3.9, and $u=h\otimes\delta_{\Gamma}$ was defined in Lemma 3.8. The hyperplane $\Gamma:=\{x_{n}=0\}$ separates the two half-spaces
[…]
(We remark in passing, that the open set $\Omega_{+}:=\mathbb{R}^{n}_{+}$ is on one side of its boundary, as required throughout this paper.) Recall that if $a(x,D)u$ is smooth enough on either of the half-spaces $\mathbb{R}_{\pm}^{n}$, then the corresponding normal lateral limit $a(x,D)u]_{\pm}$ at $\Gamma=\partial\Omega_{\pm}$ coincides with the trace $\gamma_{\pm}a(x,D)u$ of $a(x,D)u$ at the boundary $\Gamma$ of that half-space (see Lemma 3.10).
We have included in the previous section and in this section a few very basic definition related to pseudodifferential operators. If needed, results not recalled or proved here about pseudodifferential operators can be found in one of the following books [1, 51, 52, 116, 132, 139], as well as in may others. For instance, the book [131] contains a clear but concise account of pseudodifferential operators and applications to layer potentials. An even shorter introduction to some basic facts and definitions related to pseudodifferential operators and geared towards our applications can be found in [67].
A complete and general treatment of normal lateral limits distributions of the form $a(x,D)h\otimes\delta_{\Gamma}]_{\pm}$ for pseudodifferential operators $a(x,D)$ is contained in the book by Hsiao and Wendland [54]. Here we only deal with the results (and calculations) needed to treat suitable elliptic, second order differential operators. Our presentation in this section provide a complete, yet concise, introduction to the subject of limit and jump relations for potential operators on a half-space. These results, together with those in Section 6 and in [63] provide a short–but complete–introduction to jump relations on manifolds with cylindrical ends. (They also shed light on jump relations on general manifolds.)
The main question discussed in this section is the following.
Problem 4.1.
Given $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, to study the existence and properties of the normal lateral limits (see Definition (3.9)):
$a_{\pm}(x^{\prime},D^{\prime})h\mathrel{\,:=\,}\big[a(x,D)(h\otimes\delta_{\Gamma})\big]_{\pm}\,$,
in particular, to study their relation to the traces $\gamma_{\pm}a(x,D)(h\otimes\delta_{\Gamma})$.
Often in the literature, the non-tangential limits at the boundary are studied instead of the normal lateral limits that we study in this paper. The non-tangential limits are a stronger form of boundary limit and hence, they exist less often than the normal lateral limits, however, for functions that are smooth enough, they are the same. (In fact, they are the same for most of the applications studied in this paper. We stress, however, that proving the existence of non-tangential limits is more difficult than the proof of the existence of the normal lateral limits. In particular, the normal lateral limits may exist even if the non-tangential limits do not exist.)
A word now about the notation. First, we often parametrize $\Gamma:=\{x_{n}=0\}$ with $\mathbb{R}^{n-1}$. Thus $\Gamma\subset\mathbb{R}^{n}$, but $\Gamma\simeq\mathbb{R}^{n-1}\not\subset\mathbb{R}^{n}$. This is the reason we need the isometries $\tau_{t}$, $\tau_{t}(x):=x-te_{n}$ mapping $\Gamma_{t}:=\mathbb{R}^{n-1}\times\{t\}$ to $\Gamma$ (see Equation (3.18) and the discussion following it). We distinguish $\Gamma$ from $\mathbb{R}^{n-1}$ to make it easier to transition in the later sections to an arbitrary domain (i.e., and open connected subset) with boundary $\Gamma$. However, in this paper, there will be no situation when confusions can arise if we omit the isometries (i.e., identifications) $\tau_{t}$ from the notation, so we shall do that from now on.
The results in this section expand on those in [63]. The reader should compare to that book. Part of these results were stated also in [69], but without proofs.
4.1. General results on normal lateral limits of pseudodifferential operators
We shall use the notation of Subsection 3.2. Also, when a function $f(x,y,\ldots)$ depends on several variables, $x,y,\dots$ and we are computing its Fourier transform (only) with respect to the variable $x$, we shall write $\mathcal{F}_{x}(f)(\xi,y,\ldots)$ for the resulting function. The same comment applies to the inverse Fourier transform.
Let $a\in S_{\operatorname{cl}}^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ be a classical symbol with principal part (or symbol) $\sigma_{m}(a)$, regarded as a homogeneous function in the second variable (see Definition 3.7 and the discussion following it). We shall say that $\sigma_{m}(a)$ is odd if $\sigma_{m}(a;-\xi)=-\sigma_{m}(a;\xi)$. We then have the following result (Lemma A.5.6 from [63]), which is a consequence of Lemma 3.3. Recall that $e_{n}=(0,\ldots,0,1)\in\mathbb{R}^{n}$ and that $\tilde{u}(x):=u(-x)$, see Equation (3.6).
Lemma 4.2.
Let $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, $m\in\mathbb{R}$.
  1. (i)
    For any fixed $x\in\mathbb{R}^{n}$ and $\xi^{\prime}\in\mathbb{R}^{n-1}$, the function
    $\phi_{x,\xi^{\prime}}:\mathbb{R}\to\mathbb{C}\,,\quad\phi_{x,\xi^{\prime}}(\xi_{n}):=a(x,\xi^{\prime},\xi_{n})\,$,
    defines a tempered distribution on $\mathbb{R}$ such that $\mathcal{F}^{-1}\phi_{x,\xi^{\prime}}:=\mathcal{F}_{\xi_{n}}^{-1}\phi_{x,\xi^{\prime}}$ coincides with a smooth function outside 0.
  2. (ii)
    If m<1m<-1, then $\phi_{x,\xi^{\prime}}$ is integrable on $\mathbb{R}$, and hence $\mathcal{F}^{-1}\phi_{x,\xi^{\prime}}\in\mathcal{C}_{0}(\mathbb{R})$.
  3. (iii)
    Assume that $a$ is a classical symbol of order 1-1 and that $\sigma_{-1}(a)$ is odd. Let $e_{n}:=(0,\ldots,0,1)\in\mathbb{R}^{n}$, as before, and $\operatorname{\mathfrak{L}}(a;x):=\sigma_{-1}(a;x,e_{n})$. Then, for all $x\in\mathbb{R}^{n}$ and $\xi^{\prime}\in\mathbb{R}^{n-1}$, $\phi_{x,\xi^{\prime}}$ satisfies the assumptions of Lemma 3.3 for $\epsilon=1$ and $L:=\mathcal{L}(a;x)$. That is, for all $(x,\xi^{\prime})\in\mathbb{R}^{n}\times\mathbb{R}^{n-1}$ and $\tau\in\mathbb{R}$, there exists $C_{x,\xi^{\prime}}>0$ such that
    $|\tau\phi_{x,\xi^{\prime}}(\tau)-\operatorname{\mathfrak{L}}(a;x)|\mathrel{\,:=\,}|\tau a(x,\xi^{\prime},\tau)-\sigma_{-1}(a;x,e_{n})|\leq C_{x,\xi^{\prime}}/\langle\tau\rangle\,;$
Usually, the “jump factor” $\operatorname{\mathfrak{L}}$ will be associated rather to a pseudodifferential operator $a(x,D)$ (instead of its symbol $a$), so we will write
(4.1) $\operatorname{\mathfrak{L}}(a(x,D);x)\mathrel{\,:=\,}\operatorname{\mathfrak{L}}(a;x)\mathrel{\,:=\,}\sigma_{-1}(a;x,e_{n})\,.$
Recall that $\mathcal{C}_{0}^{k}(\mathbb{R}^{n})$ denotes the set of functions $f:\mathbb{R}^{n}\to\mathbb{C}$ such that, for all $|\alpha|\leq k$, we have $\lim_{|x|\to\infty}|\partial^{\alpha}f(x)|=0$.
Proof.
The property that $\phi_{x,\xi^{\prime}}$ defines a tempered distribution follows from the inequality $\phi_{x,\xi^{\prime}}(\tau)\leq C_{x,\xi^{\prime}}\langle\tau\rangle^{m}$. When m<1m<-1, this implies that $\phi_{x,\xi^{\prime}}\in L^{1}(\mathbb{R})$, and hence its inverse Fourier transform is in $\mathcal{C}_{0}(\mathbb{R})$. This proves (ii). To complete the proof of (i), let $n>m+k+1$. We will use $t$ as dual variable to $\xi_{n}$ for the Fourier transform in $\xi_{n}\in\mathbb{R}$. We let $t$ denote both the variable of $\mathbb{R}$ and the multiplication operator by $t$ on $\mathcal{S}(\mathbb{R})$ and on $\mathcal{S}^{\prime}(\mathbb{R})$. Then $t^{n}\mathcal{F}_{\xi_{n}}^{-1}\phi_{x,\xi^{\prime}}=\imath^{n}\mathcal{F}_{\xi_{n}}^{-1}(\partial_{\xi_{n}}^{n}\phi_{x,\xi^{\prime}})\in\mathcal{C}_{0}^{k}(\mathbb{R})$ (all derivative up to order $k$ tend to 0 at infinity) where the last relation is due to the fact that $|\partial_{\xi_{n}}^{n}\phi_{x,\xi^{\prime}}(\xi_{n})|\leq C\langle\xi_{n}\rangle^{m-n}$ and mn<k1m-n<-k-1. This shows that $\mathcal{F}_{\xi_{n}}^{-1}\phi_{x,\xi^{\prime}}$ coincides with a smooth function on $\mathbb{R}\smallsetminus\{0\}$.
Finally, to prove (iii), we will use Corollary 3.4 for each of the functions $\phi_{x,\xi^{\prime}}$, $(x,\xi^{\prime})\in\mathbb{R}^{n}\times\mathbb{R}^{n-1}$, with $\operatorname{\mathfrak{L}}_{\pm}(\phi_{x,\xi})=:\operatorname{\mathfrak{L}}_{\pm}(x)$. To that end, we need to first check that the hypotheses of Lemma 3.3 are satisfied by all functions $\phi_{x,\xi^{\prime}}$. To that end, let us write
$a\,=\,b_{-1}+r_{-2}\,,\qquad b_{-1}\in S^{-1}(\mathbb{R}^{n}\times\mathbb{R}^{n})\,,\ \mbox{ and }\ r_{-2}\in S^{-2}(\mathbb{R}^{n}\times\mathbb{R}^{n})\,$,
where $b_{-1}$ is essentially homogeneous of order 1-1. Recall that this means that $b_{-1}(x,\lambda\xi)=\lambda^{-1}b_{-1}(x,\xi)$ whenever $|\xi|,\lambda|\xi|\geq 1$. It follows that $r_{-2}$ is a classical symbol. The hypothesis that $\sigma_{-1}(a;x,\xi)$ is odd in $\xi$ and the fact that $b_{-1}(a;x,\xi)=\sigma_{-1}(a;x,\xi)$ for $|\xi|\geq 1$ imply that $b_{-1}(x,\lambda\xi)=\lambda^{-1}b_{-1}(x,\xi)$ whenever $|\xi|,|\lambda\xi|\geq 1$ (that is, we have extended the previous relation to $\lambda$ negative as well). We have that all derivatives of $b_{-1}(x,\xi)$ are bounded, and hence there exists a constant $C>0$ such that, for $|\tau|\geq 1$,
$|b_{-1}(x,\tau^{-1}\xi^{\prime},1)-b_{-1}(x,0,1)|\leq C|\tau^{-1}\xi^{\prime}|\,.$
By increasing $C$, we may assume that $|r_{-2}(x,\xi)|\leq C|\xi|^{-2}$ for all $|\xi|\geq 1$, because $r_{-2}$ is a symbol of order 2-2. Let $L=b_{-1}(x,0,1)=b_{-1}(x,e_{n}):=\sigma_{-1}(a;e_{n})$. For $|\tau|\geq 1$ and $(x,\xi^{\prime})\in\mathbb{R}^{n}\times\mathbb{R}^{n-1}$, we let $\xi:=(\xi^{\prime},\tau)$ with $|\xi|\geq|\tau|\geq 1$ and use the homogeneity relation for $\tau^{-1}\xi$ to obtain
$|\tau^{2}a(x,\xi^{\prime},\tau)-L\tau|\leq|\tau^{2}b_{-1}(x,\xi^{\prime},\tau)-L\tau|+|\tau^{2}r_{-2}(x,\xi^{\prime},\tau)|\\ \,=\,|\tau|\,\big|\,b_{-1}(x,\tau^{-1}\xi^{\prime},1)-b_{-1}(x,e_{n})\big|+|\tau^{2}r_{-2}(x,\xi^{\prime},\tau)|\leq C|\xi^{\prime}|+C\,=:\,C_{x,\xi^{\prime}}\,.$
Moreover, the function $\phi_{x,\xi^{\prime}}(\tau):=a(x,\xi^{\prime},\tau)$ is bounded for $|\tau|\leq 1$ (because $a$ is a symbol of order 1-1). This completes the proof. ∎
Recall that if $x=(x_{1},x_{2},\ldots,x_{n})\in\mathbb{R}^{n}$, we shall write $x=(x^{\prime},x_{n})$, where, $x^{\prime}:=(x_{1},x_{2},\ldots,x_{n-1})\in\mathbb{R}^{n-1}$. Similarly, if $\xi=(\xi_{1},\xi_{2},\ldots,\xi_{n})\in\mathbb{R}^{n}$, we shall write $\xi=(\xi^{\prime},\xi_{n})$, where $\xi^{\prime}\in\mathbb{R}^{n-1}$. As detailed in the Corollary 4.4 next, the last point of the last lemma in conjunction with Corollary 3.4 allows us to introduce the following definition.
Definition 4.3.
Let $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ with $m\in\mathbb{R}$ and $t\in\mathbb{R}$ and let $\phi_{x,\xi^{\prime}}(\xi_{n}):=a(x,\xi^{\prime},\xi_{n})$, $\widetilde{\phi}_{x,\xi^{\prime}}(\xi_{n}):=a(x,\xi^{\prime},-\xi_{n})$, as before. If $m\geq-1$, we assume $t\neq 0$. We then define:
  1. (i)
    $a_{x_{n},t}(x^{\prime},\xi^{\prime}):=\mathcal{F}_{\xi_{n}}^{-1}a(x^{\prime},x_{n},\xi^{\prime},t):=\big[\mathcal{F}^{-1}\phi_{x,\xi^{\prime}}\big](t).$
  2. (ii)
    If $a\in S_{\operatorname{cl}}^{-1}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ and $\sigma_{-1}(a;x,\xi)$ is odd in $\xi$, we also let
    1. (a)
      $a_{x_{n},0}(x^{\prime},\xi^{\prime})\mathrel{\,:=\,}\frac{1}{2\pi}\int_{\mathbb{R}}\frac{a(x^{\prime},x_{n},\xi^{\prime},\xi_{n})+a(x^{\prime},x_{n},\xi^{\prime},-\xi_{n})}{2}\,d\xi_{n}\,=\,\int_{\mathbb{R}}\frac{\phi_{x,\xi^{\prime}}+\tilde{\phi}_{x,\xi^{\prime}}}{4\pi}\,dt$;
    2. (b)
      $\operatorname{\mathfrak{L}}(a;x):=\sigma_{-1}(a;x,e_{n})$; and
    3. (c)
      $a_{x_{n},0\pm}(x^{\prime},\xi^{\prime})\mathrel{\,:=\,}\pm\frac{\imath\sigma_{-1}(a;x,e_{n})}{2}+a_{x_{n},0}(x^{\prime},\xi^{\prime})\mathrel{\,:=\,}\pm\frac{\imath}{2}\operatorname{\mathfrak{L}}(a;x)+a_{x_{n},0}(x^{\prime},\xi^{\prime})$.
The reason for introducing Definition 4.3 comes from Section 3.3. Of course, if m<1m<-1 in this definition, we have, for all $t\in\mathbb{R}$,
(4.2) $a_{x_{n},t}(x^{\prime},\xi^{\prime})\mathrel{\,:=\,}\mathcal{F}_{\xi_{n}}^{-1}a(x^{\prime},x_{n},\xi^{\prime},t)\mathrel{\,:=\,}\frac{1}{2\pi}\int_{\mathbb{R}}e^{\imath t\xi_{n}}a(x,\xi^{\prime},\xi_{n})\,d\xi_{n}\,.$
Recall that $\tilde{\psi}(t):=\psi(-t)$. The last point of the last lemma, in conjunction with Corollary 3.4 yields the following very useful consequence.
Corollary 4.4.
Let $a$ be a classical symbol of order 1-1, $\phi_{x,\xi^{\prime}}(\xi_{n}):=a(x,\xi^{\prime},\xi_{n})$, $\phi_{x,\xi^{\prime}}:\mathbb{R}\to\mathbb{C}$, $e_{n}:=(0,\ldots,0,1)\in\mathbb{R}^{n}$, and $\operatorname{\mathfrak{L}}(a;x):=\sigma_{-1}(a;x,e_{n})$, as before. We assume that $\sigma_{-1}(a)$ is odd in the sense that, for all $x,\xi\in\mathbb{R}^{n}$, $\xi\neq 0$, we have $\sigma_{-1}(a;x,-\xi)=-\sigma_{-1}(a;x,-\xi)$. Then, for all $x\in\mathbb{R}^{n}$ and $\xi^{\prime}\in\mathbb{R}^{n-1}$, we have the following properties:
  1. (i)
    $\phi_{x,\xi^{\prime}}$ defines a tempered distribution on $\mathbb{R}$.
  2. (ii)
    $\mathcal{F}^{-1}\phi_{x,\xi^{\prime}}\in\mathcal{S}^{\prime}(\mathbb{R})$ is a locally integrable function that is continuous on $\mathbb{R}\smallsetminus\{0\}$.
  3. (iii)
    $\operatorname{\mathfrak{L}}(a;x)\mathrel{\,:=\,}\sigma_{-1}(a;x,e_{n})\,=\,\lim_{|\tau|\to\infty}\tau a(x,\xi^{\prime},\tau)\in\mathbb{C}.$
  4. (iv)
    $\phi_{x,\xi^{\prime}}+\tilde{\phi}_{x,\xi^{\prime}}\in L^{1}(\mathbb{R})$.
  5. (v)
    The one-sided limits $\mathcal{F}^{-1}\phi_{x,\xi^{\prime}}(0\pm)$,
[…]
    of $\mathcal{F}^{-1}\phi_{x,\xi^{\prime}}\in\mathcal{S}^{\prime}(\mathbb{R})$ at 0 are given by
    $\mathcal{F}^{-1}\phi_{x,\xi^{\prime}}(0\pm)$ $\,=\,\pm\frac{\imath\sigma_{-1}(a;x,e_{n})}{2}+\frac{1}{2\pi}\,\int_{\mathbb{R}}\frac{a(x,\xi^{\prime},\tau)+a(x,\xi^{\prime},-\tau)}{2}\,d\tau$
    $\,=:\,\pm\frac{\imath\operatorname{\mathfrak{L}}(a;x)}{2}+a_{x_{n},0}(x^{\prime},\xi^{\prime})\,=:\,a_{x_{n},0\pm}(x^{\prime},\xi^{\prime})\,.$
In our book [63], the last lemma and the last corollary were proved for the case m=1m=-1 under a slightly more general assumption, as we explain next.
Remark 4.5.
Let m=1m=-1, $e_{n}=(0,\ldots,0,1)\in\mathbb{R}^{n}$ (as before), and let us assume that $a$ is classical (but not necessarily odd). Then, for all $x\in\mathbb{R}^{n}$ and $\xi^{\prime}\in\mathbb{R}^{n-1}$, the following limits exist
$\operatorname{\mathfrak{L}}_{\pm}(a;x)\mathrel{\,:=\,}\lim_{\tau\to\pm\infty}\tau a(x,\xi^{\prime},\tau)\,=\,\pm\sigma_{-1}(a;x,\pm e_{n})\in\mathbb{C}\,.$
In our book, we have proved the last lemma under the weaker assumption that $\operatorname{\mathfrak{L}}_{+}(x)=\operatorname{\mathfrak{L}}_{-}(x)$. In this paper, this relation is a consequence of the fact that $\sigma_{-1}(a)$ is assumed to be odd. This simplifies the presentation and is anyway the only case needed in this paper. From now on, we shall rather use this assumption (odd principal symbol, instead of the less restrictive assumption $\operatorname{\mathfrak{L}}_{+}(x)=\operatorname{\mathfrak{L}}_{-}(x)$ used in our book).
Before turning to the study of Problem 4.1, let us recall the following basic results.
Theorem 4.6.
Let $(\mathfrak{M},\mu)$ be a measure space, $N\in\mathbb{Z}_{+}\cup\{\infty\}$, and $U\subset\mathbb{R}^{k}$ be an open subset. Let also $f:U\times\mathfrak{M}\to\mathbb{C}$ be a measurable function satisfying, for all multi-indices $\alpha\in\mathbb{Z}_{+}^{k}$, $|\alpha|\leq N$, the following three conditions:
  1. (i)
    The derivative $\partial_{x}^{\alpha}f(x,\omega)$ is defined for all $(x,\omega)\in U\times\mathfrak{M}$.
  2. (ii)
    For all $\omega\in\mathfrak{M}$, the function $x\mapsto\partial_{x}^{\alpha}f(x,\omega)$ is continuous in $x\in U$.
  3. (iii)
    There is an integrable function $G_{\alpha}:\mathfrak{M}\to[0,\infty)$ such that, for all $x\in U$,
    $|\partial_{x}^{\alpha}f(x,\omega)|\leq G_{\alpha}(\omega)\,.$
Then the function $F:\mathbb{R}^{k}\to\mathbb{C}$, $F(x):=\int_{\mathfrak{M}}f(x,\omega)\,d\mu(\omega)$, has the property that $\partial_{x}^{\alpha}F$ is defined and continuous for all $|\alpha|\leq N$, and it is given by
$\partial_{x}^{\alpha}F(x)=\int_{\mathfrak{M}}\partial_{x}^{\alpha}f(x,\omega)\,d\mu(\omega)\,$,
and $|\partial_{x}^{\alpha}F(x)|\leq\int_{\mathfrak{M}}G_{\alpha}(\omega)\,d\mu(\omega)$.
Proof.
The result follows by applying Lebesgue’s dominated convergence theorem. See [41, Appendix E]. ∎
If $N=0$ (hence $\alpha=0$), then this theorem simply states that $F(x)$ is defined for all $x\in\mathbb{R}^{k}$, that it is continuous in $x$, and that it is bounded by $|F(x)|\leq\int_{\mathfrak{M}}G_{0}(\omega)\,d\mu(\omega).$ An immediate consequence of this theorem is the following strong operator convergence result.
To identify the limits of $\big[a(x,D)(h\otimes\delta_{\Gamma})\big]_{\epsilon}$ as $\epsilon\to 0$, we shall need the following well-known, standard result on pseudodifferential operators.
Lemma 4.7.
Let $m\in\mathbb{R}$, let $b_{\epsilon}\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, $\epsilon\in[0,1]$, and assume that the set $\{b_{\epsilon}\mid\epsilon\in[0,1]\}$ is bounded in $S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ (that is, for all $(x,\xi)\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, we have $|\partial_{x}^{\alpha}\partial_{\xi}^{\beta}b_{\epsilon}(x,\xi)|\leq C_{\alpha,\beta}\langle\xi\rangle^{m-|\beta|}$, with $C_{\alpha,\beta}$ independent of $\epsilon$). Let us also assume that, for all $(x,\xi)\in\mathbb{R}^{n}\times\mathbb{R}^{n}$ we have $\lim_{\epsilon\to 0}b_{\epsilon}(x,\xi)=b_{0}(x,\xi)$. Then, for all $h\in H^{s}(\mathbb{R}^{n})$, we have the following convergence in $H^{s-m}(\mathbb{R}^{n})$
$\lim_{\epsilon\to 0}b_{\epsilon}(x,D)h\,=\,b_{0}(x,D)h\in H^{s-m}(\mathbb{R}^{n})\,.$
Proof.
The norm of $a(x,D)$ depends on finitely many of the semi-norms defining the topology on $S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ (see [67] in addition to the standard textbooks mentioned above). The result follows by combining this property with Theorem 4.6. ∎
We now turn to the study of Problem 4.1. To that end, recall that $\Gamma:=\{(x^{\prime},x_{n})\in\mathbb{R}^{n}\mid x_{n}=0\}$ and that $h\otimes\delta_{\Gamma}$ is the distribution
$\langle h\otimes\delta_{\Gamma},\phi\rangle:=\int_{\Gamma}h(x^{\prime})\phi(x^{\prime},0)\,dx^{\prime},$
where $\phi\in\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n})\,$, see Equation (3.18) and Lemma 3.8. In particular, if $h\in L^{2}(\Gamma)$, then $h\otimes\delta_{\Gamma}\in H^{-1/2-\epsilon}(\mathbb{R}^{n})$, for all $\epsilon>0$. If $u:\mathbb{R}^{n}\to\mathbb{C}^{k}$ is continuous enough, recalling the notation introduced in Definition (3.9), we obtain that then $u_{\epsilon}:\mathbb{R}^{n-1}\to\mathbb{C}^{k}$ is given by $u_{\epsilon}(x^{\prime})\mathrel{\,:=\,}u(x^{\prime},\epsilon)$, $x^{\prime}\in\mathbb{R}^{n-1}$. Note that, we identify $\Gamma_{\epsilon}:=\Gamma+\epsilon e_{n}:=\{x=(x^{\prime},x_{n})\in\mathbb{R}^{n}:x_{n}=\epsilon\}$ with $\Gamma$ and their associated function spaces with the translation $\tau_{-\epsilon}$, $\tau_{t}(x):=x+te_{n}$, as explained in Subsection 3.3. Recall the potential operator
$\mathcal{S}_{a}(h)\mathrel{\,:=\,}a(x,D)(h\otimes\delta_{\Gamma})\in\mathcal{C}^{\infty}(\mathbb{R}^{n}\smallsetminus\Gamma)$
of Equation (3.19).
Here is our first result on $\big[a(x,D)(h\otimes\delta_{\Gamma})\big]_{\epsilon}$, regarded as a function on $\mathbb{R}^{n-1}\simeq\{x_{n}=\epsilon\}$, as explained in Subsection 3.3, see especially Definition (3.9).
Proposition 4.8.
Let $h\in L^{2}(\mathbb{R}^{n-1})$, let $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, $m\in\mathbb{R}$, and let $a_{x_{n},t}$, $t\neq 0$, be as in Definition 4.3. Then, for any $\epsilon\neq 0$, $a(x,D)(h\otimes\delta_{\Gamma})$ coincides with a smooth function on $\Gamma_{\epsilon}:=\Gamma+\epsilon e_{n}$ and its restrictions to these sets satisfy
$\big[\mathcal{S}_{a}(h)\big]_{\epsilon}\mathrel{\,:=\,}\big[a(x,D)(h\otimes\delta_{\Gamma})\big]_{\epsilon}\,=\,a_{\epsilon,\epsilon}(x^{\prime},D^{\prime})h\,.$
Proof.
Let $h\in\mathcal{C}^{\infty}_{c}(\Gamma)$ and $H:=h\otimes\delta_{\Gamma}$, see Lemma 3.8. Then $H\in H^{-1}(\mathbb{R}^{n})$ (because 1<1/2-1<-1/2), so $a(x,D)H$ is defined. Since $H$ has singularities only on $\Gamma$, $a(x,D)H$ will also have singularities only on $\Gamma$, because pseudodifferential operators preserve the singular support. Therefore, $a(x,D)H$ coincides with a smooth function outside $\Gamma$. Moreover,
$\widehat{H}(\xi)\mathrel{\,:=\,}\mathcal{F}(h\otimes\delta_{\Gamma})(\xi)\,=\,\mathcal{F}_{x^{\prime}}(h)(\xi^{\prime})\,=\,\hat{h}(\xi^{\prime})\,$,
where $\mathcal{F}_{x^{\prime}}$ is the Fourier transform on $\Gamma$. For $a$ of order <1<-1, we then obtain
$a(x,D)H(x)$ $\mathrel{\,:=\,}\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{n}}e^{\imath x\cdot\xi}a(x,\xi)\widehat{H}(\xi)\,d\xi$
$\,=\,\frac{1}{(2\pi)^{n}}\int_{\mathbb{R}^{n-1}}e^{\imath x^{\prime}\cdot\xi^{\prime}}\left(\int_{\mathbb{R}}e^{\imath x_{n}\cdot\xi_{n}}a(x,\xi)\hat{h}(\xi^{\prime})\,d\xi_{n}\right)\,d\xi^{\prime}$
$\,=\,\frac{1}{(2\pi)^{n-1}}\int_{\mathbb{R}^{n-1}}e^{\imath x^{\prime}\cdot\xi^{\prime}}a_{x_{n},x_{n}}(x^{\prime},\xi^{\prime})\hat{h}(\xi^{\prime})\,d\xi^{\prime}$
$\,=:\,a_{x_{n},x_{n}}(x^{\prime},D^{\prime})h(x^{\prime})\,$,
with all integrals being absolutely convergent, because $\hat{h}(\xi^{\prime})$ has rapid decay in $\xi^{\prime}$ and $a$ is of order 1-1, which ensures the integrability in the $\xi_{n}$ variable. (This justifies the use of Fubini’s theorem in the second equality).
For $a$ of order $m\geq-1$, let $\phi\in\mathcal{S}(\mathbb{R}^{n})$, regarded as a symbol on $\mathbb{R}^{n}$ that is constant in $x$. Let $\phi_{\epsilon}(\xi):=\phi(\epsilon\xi)$ and assume that $\phi(0)=1$. Then $b_{\epsilon}:=a\phi_{\epsilon}$ is a symbol of order -$\infty$ and, for all $u\in H^{s}(\mathbb{R}^{n})$, $b_{\epsilon}(x,D)u\to a(x,D)u$ in $H^{s-m}(\mathbb{R}^{n})$ as $\epsilon\searrow 0$. The desired result for $h\in\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n-1})$ is obtained by applying the statement already proved to $b_{\epsilon}$ and passing to the limit for $\epsilon\searrow 0$ and using Lemma 4.7. The result for general $h$ follows by continuity in $h$ and the density of the space $\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n-1})$ in $L^{2}(\mathbb{R}^{n-1})$. ∎
We have the following simple remark that simplifies our notation.
Remark 4.9.
Proposition 4.8 gives the equality $a_{\epsilon}=a_{\epsilon,\epsilon}$, which thus reconciles the definitions of $a_{\epsilon}$ (introduced in Definition 3.9) and that of $a_{\epsilon,\epsilon}$ (introduced in Definition 4.3).
We shall need the following standard lemma.
Lemma 4.10.
Let $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, $m\in\mathbb{R}$, be essentially homogeneous of order $m$ (i.e., $a(x,\lambda\xi)=\lambda^{m}a(x,\xi)$ for $\lambda|\xi|,|\xi|\geq 1$, see Equation (3.16)). Let $a_{x_{n},t}$ be as in Definition 4.3(i).
  1. (i)
    For all $\lambda|\xi^{\prime}|,|\xi^{\prime}|\geq 1$ and all $t\neq 0$, we have
    $a_{x_{n},t}(x^{\prime},\lambda\xi^{\prime})\,=\,\lambda^{m+1}a_{x_{n},\lambda t}(x^{\prime},\xi^{\prime})\,.$
  2. (ii)
    Let $\varepsilon\in\{-1,1\}$ and assume that $a(x,-\xi)=\varepsilon a(x,\xi)$ for $|\xi|\geq 1$, then, for all $\lambda|\xi^{\prime}|,|\xi^{\prime}|\geq 1$, we have
    $a_{x_{n},t}(x^{\prime},-\xi^{\prime})\,=\,\varepsilon a_{x_{n},-t}(x^{\prime},\xi^{\prime})\,.$
  3. (iii)
    If m<1m<-1, the results of the previous two points hold also for $t=0$.
Proof.
Assume first that m<1m<-1. The substitution $\xi_{n}=\lambda\tau$ yields
$a_{x_{n},t}(x^{\prime},\lambda\xi^{\prime})\mathrel{\,:=\,}\mathcal{F}_{\xi_{n}}^{-1}a(x^{\prime},x_{n},\lambda\xi^{\prime},t)\mathrel{\,:=\,}\frac{1}{2\pi}\int_{\mathbb{R}}e^{\imath t\xi_{n}}a(x^{\prime},x_{n},\lambda\xi^{\prime},\xi_{n})\,d\xi_{n}\\ \,=\,\frac{1}{2\pi}\int_{\mathbb{R}}e^{\imath t\lambda\tau}a(x^{\prime},x_{n},\lambda\xi^{\prime},\lambda\tau)\,\lambda d\tau\,=\,\frac{\lambda^{m+1}}{2\pi}\int_{\mathbb{R}}e^{\imath t\lambda\tau}a(x^{\prime},x_{n},\xi^{\prime},\tau)\,d\tau\\ \,=\,\lambda^{m+1}a_{x_{n},\lambda t}(x^{\prime},\xi^{\prime})\,.$
If $m\geq-1$, the proof is the same by pairing test functions supported away from 0. This proves (i). The proof carries over to (ii) is completely similar. If m<1m<-1, the proofs of (i) and (ii) obviously work also for $t=0$, which proves (iii). ∎
4.2. Normal lateral limits of pseudodifferential operators of orders <1<-1
Let $\Gamma:=\{x_{n}=0\}\subset\mathbb{R}^{n}$, as always in this paper. For further use, let us record the following calculation.
Lemma 4.11.
If s<1s<-1, then there exists $C_{s}>0$ such that $\int_{\mathbb{R}}\langle\xi\rangle^{s}\,d\xi_{n}\,=\,C_{s}\langle\xi^{\prime}\rangle^{s+1}.$
Proof.
The substitution $\xi_{n}=\langle\xi^{\prime}\rangle t$ gives:
$\int_{\mathbb{R}}\langle\xi\rangle^{s}\,d\xi_{n}=\int_{\mathbb{R}}\big(\langle\xi^{\prime}\rangle^{2}+\langle\xi^{\prime}\rangle^{2}t^{2}\big)^{s/2}\langle\xi^{\prime}\rangle\,dt=\langle\xi^{\prime}\rangle^{s+1}\int_{\mathbb{R}}(1+t^{2})^{s/2}\,dt=:C_{s}\langle\xi^{\prime}\rangle^{s+1},$
where the last integral is convergent because s<1s<-1. ∎
We can now prove the needed results for order m<1m<-1 operators. Recall the symbols $a_{x_{n},\epsilon}$, $\epsilon\in\mathbb{R}$, of Definition 4.3.
Proposition 4.12.
Let m<1m<-1 and let $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$.
  1. (i)
    For any $(x_{n},t)\in\mathbb{R}^{2}$, we have that the map $\mathbb{R}^{n-1}\times\mathbb{R}^{n-1}\ni(x^{\prime},\xi^{\prime})\to a_{x_{n},t}(x^{\prime},\xi^{\prime})\in\mathbb{C}$ defines a symbol in $S^{m+1}(\mathbb{R}^{n-1}\times\mathbb{R}^{n-1})$.
  2. (ii)
    The family $\{a_{x_{n},t}\,\mid\,(x_{n},t)\in\mathbb{R}^{2}\}$ is bounded in $S^{m+1}(\mathbb{R}^{n-1}\times\mathbb{R}^{n-1})$.
  3. (iii)
    The function $\mathbb{R}^{n}\times\mathbb{R}^{n}\ni(x,\xi^{\prime},t)\to a_{x_{n},t}(x^{\prime},\xi^{\prime})\in\mathbb{C}$ is continuous.
  4. (iv)
    If $h\in H^{s}(\Gamma)$, then the function $\mathbb{R}^{2}\ni(x_{n},t)\to a_{x_{n},t}(x^{\prime},D^{\prime})h\in H^{s-m-1}(\Gamma)\simeq H^{s-m-1}(\mathbb{R}^{n-1})$ is continuous.
  5. (v)
    If $a\in S_{\operatorname{cl}}^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$, then, for all $x_{n}\in\mathbb{R}$, $a_{x_{n},0}\in S_{\operatorname{cl}}^{m+1}(\mathbb{R}^{n-1}\times\mathbb{R}^{n-1})$, that is, $a_{x_{n},0}$ is also a classical symbol.
Proof.
Let $a\in S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n})$. Then, by definition, for each $\alpha,\beta\in\mathbb{Z}_{+}^{n}$, there exists $C_{\alpha,\beta}>0$ such that, for all $x,\xi\in\mathbb{R}^{n}$,
$|\partial_{x}^{\alpha}\partial_{\xi}^{\beta}a(x,\xi)|\leq C_{\alpha,\beta}\langle\xi\rangle^{m-|\beta|}\,.$
Since $\langle\xi\rangle\geq\langle\xi_{n}\rangle$ and $m-|\beta|\leq m<-1$, we have $|\partial_{x}^{\alpha}\partial_{\xi}^{\beta}a(x,\xi)|\leq C_{\alpha,\beta}\langle\xi_{n}\rangle^{m-|\beta|}$. The bound on the right hand side of the last inequality is integrable and independent of $(x,\xi^{\prime})$. The hypothesis of the theorem of derivation under the integral sign, Theorem 4.6, are thus satisfied and we obtain, for all $x=(x^{\prime},x_{n})$ that
$a_{x_{n},t}(x^{\prime},\xi^{\prime})\mathrel{\,:=\,}\mathcal{F}_{\xi_{n}}^{-1}a(x^{\prime},x_{n},\xi^{\prime},t)\mathrel{\,:=\,}\frac{1}{2\pi}\int_{\mathbb{R}}e^{\imath t\xi_{n}}a(x^{\prime},x_{n},\xi^{\prime},\xi_{n})\,d\xi_{n}\,.$
is well defined (i.e. the integral converges). We also obtain the relation
$\partial_{x}^{\alpha}\partial_{\xi^{\prime}}^{\beta}a_{x_{n},t}(x^{\prime},\xi^{\prime})\,=\,\frac{1}{2\pi}\,\int_{\mathbb{R}}\partial_{x}^{\alpha}\partial_{\xi^{\prime}}^{\beta}e^{\imath t\xi_{n}}a(x,\xi)\,d\xi_{n}$
and the estimate
(4.3) $2\pi\,|\partial_{x}^{\alpha}\partial_{\xi^{\prime}}^{\beta}a_{t}(x,\xi)|\,\leq\,C_{\alpha,\beta}\,\int_{\mathbb{R}}\langle\xi\rangle^{m-|\beta|}\,d\xi_{n}\leq C_{m-|\beta|}C_{\alpha,\beta}\langle\xi^{\prime}\rangle^{m+1-|\beta|}\,$,
by Lemma 4.11. This proves (i), (ii), and (iii) at once.
The point (iv) is a consequence of Lemma 4.7 and the points already proved. The point (v) follows from Lemma 4.10, by linearity. ∎
We are now ready to formulate our main result concerning the limit/jump values of classical matrix valued pseudodifferential operators of order <1<-1. So far, we have formulated and proved our results for scalar symbols. The proofs extend, however, immediately to the vector valued case. We thus formulate the following results for sections of a trivial $k$-dimensional vector bundle.
Theorem 4.13.
Let $a\in M_{k}(S^{m}(\mathbb{R}^{n}\times\mathbb{R}^{n}))$ for some m<1m<-1 and $k\in\mathbb{N}$. We use the relation $a_{\epsilon}=a_{\epsilon,\epsilon}$, see Remark 4.9.
  1. (i)
    The matrix valued function $a_{0}$ is an order $m+1$ symbol given by
    $a_{0}(x^{\prime},\xi^{\prime})\,=\,\frac{1}{2\pi}\int_{\mathbb{R}}a(x^{\prime},0,\xi^{\prime},\xi_{n})\,d\xi_{n}\,.$
    If $a$ is classical, then $a_{0}$ is also classical.
  2. (ii)
    For all $s\in\mathbb{R}$ and all $h\in H^{s}(\Gamma;\mathbb{C}^{k})=H^{s}(\Gamma)^{k}$ and all $\epsilon\neq 0$ small, we have $\big[a(x,D)(h\otimes\delta_{\Gamma})]_{\epsilon}=a_{\epsilon}(x^{\prime},D^{\prime})h$.
  3. (iii)
    For all $s\in\mathbb{R}$ and all $h\in H^{s}(\Gamma;\mathbb{C}^{k})=H^{s}(\Gamma)^{k}$, we have
    $\mathcal{S}_{a}h]_{\pm}\mathrel{\,:=\,}\big[a(x,D)(h\otimes\delta_{\Gamma})]_{\pm}\mathrel{\,:=\,}\lim_{\epsilon\to 0\pm}\big[a(x,D)(h\otimes\delta_{\Gamma})]_{\epsilon}\\ \,=\,\lim_{\epsilon\to 0\pm}a_{\epsilon}(x^{\prime},D^{\prime})h\,=\,a_{0}(x^{\prime},D^{\prime})h\in H^{s-m-1}(\Gamma)^{k}\,.$
    (The distribution $h\otimes\delta_{\Gamma}$ was introduced in Lemma 3.8 and $u_{\epsilon}$ is the restriction of $u$ to $\Gamma+\epsilon e_{n}\simeq\mathbb{R}^{n-1}$.)
  4. (iv)
    If $h\in L^{2}(\Gamma)^{k}$, then $\mathcal{S}_{a}h\mathrel{\,:=\,}a(x,D)(h\otimes\delta_{\Gamma})\in H^{s^{\prime}}(\mathbb{R}^{n})^{k}$ for any $s^{\prime}\in(1/2,-m-1/2)$, and hence we have the equality of traces (or restrictions)
    $\mathcal{S}_{a}h]_{+}\,=\,[\mathcal{S}_{a}h]_{-}\,=\,[\mathcal{S}_{a}h]|_{\Gamma}\,=\,a_{0}(x^{\prime},D^{\prime})h\in H^{s^{\prime}-1/2}(\Gamma)^{k}\,.$
  5. (v)
    Let $k_{a_{0}(x^{\prime},D^{\prime})}$ be the distribution kernel of $a_{0}(x^{\prime},D^{\prime})$ and $k_{a(x,D)}$ be the distribution kernel of $a(x,D)$. Then
    $k_{a_{0}(x^{\prime},D^{\prime})}(x^{\prime},y^{\prime})\,=\,k_{a(x,D)}(x^{\prime},0,y^{\prime},0)\,,\quad x^{\prime},y^{\prime}\in\mathbb{R}^{n-1}\,,\ x^{\prime}\neq y^{\prime}\,.$
Proof.
We can assume that $k=1$. The general case is completely similar.
The fact that $a_{0}$ is a symbol of order $m+1$ in (i) was proved already in Proposition 4.12. (The fact that $a_{0}$ is classical if $a$ is will be proved in a slightly greater generality in Theorem 4.16.) Finally, the stated equality in (i) follows from the fact that $\mathbb{R}\ni\xi_{n}\to a(x,\xi^{\prime},\xi_{n})\in\mathbb{C}$ is integrable on $\mathbb{R}$ and the definition of $a_{0}=a_{0,0}$ (see also Equation (4.2)).
The point (ii) is the content of Proposition 4.8. The first two equalities in (iii) are definitions. The third equality in (iii) is the consequence of (ii) just discussed, and, finally, the last equality in (iii) follows from Proposition 4.12(iv) for $x_{n}=t=\epsilon$. The results and limits are in the stated space because $a_{0}$ is a symbol of order $m+1$, as proved in (i).
The first two equalities in (iv) follow from Lemma 3.10 and the last equality is a consequence of (iii) already proved.
Finally, let us prove (v). Let $\mathcal{F}_{\xi}$ is the Fourier transform in the second variable (in $\xi$). Since $a$ defines a tempered distribution, $\mathcal{F}_{\xi}^{-1}a$ is also a tempered distribution. We next use Lemma 3.6 and the formula for $a_{0}=a_{0,0}$ in (i) to obtain
$k_{a_{0}(x^{\prime},D^{\prime})}(x^{\prime},y^{\prime})\,=\,\mathcal{F}_{\xi^{\prime}}^{-1}a_{0}(x^{\prime},x^{\prime}-y^{\prime})\,=\,\mathcal{F}_{\xi^{\prime}}^{-1}\mathcal{F}_{\xi_{n}}^{-1}a(x^{\prime},0,x^{\prime}-y^{\prime},0)\\ \,=\,\mathcal{F}_{\xi}^{-1}a(x^{\prime},0,x^{\prime}-y^{\prime},0)\,=\,k_{a(x,D)}(x^{\prime},0,y^{\prime},0)\,$,
as claimed. ∎
Remark 4.14.
It is known that the distribution kernel of a negative order operator cannot be supported on the diagonal (because the pseudodifferential operators supported on the diagonal are local, and hence they are multiplication operators, by Peetre’s theorem [54, 52, 53], that is, order zero operators). Therefore, the restriction of the distribution kernel of a negative order operator outside the diagonal completely determines the operator (this is relevant in view of Theorems 4.13(v) and 4.16(i) and (ii), for whose proof, a similar argument was used). Thus, if m<1m<-1, the knowledge of the distribution kernel $k_{a(x,D)}$ of a pseudodifferential operator $a(x,D)$ of order $m$ completely determines the distribution kernel $k_{a_{0}(x^{\prime},D^{\prime})}$ and, hence, it completely determines $a_{0}$ and $a_{0}(x^{\prime},D^{\prime})$.
4.3. Normal lateral limits of pseudodifferential operators of orders 1-1
We now study normal lateral limits of pseudodifferential operators of order m=1m=-1. For simplicity, we consider only classical pseudodifferential operators. Several points of the next proposition have already been proved, but the most significant calculation (the point (iii) below) has not yet been performed. We again formulate our result for matrices of symbols.
Proposition 4.15.
Let $a\in M_{k}(S^{-1}_{\operatorname{cl}}(\mathbb{R}^{n}\times\mathbb{R}^{n})$. We let $a_{x_{n},t}(x^{\prime},\xi^{\prime})$ be as in Definition 4.3 (we apply that definition to each entry of the matrix). We assume that $\sigma_{-1}(a;x,\xi)$ is odd in $\xi\in\mathbb{R}^{n}$.
  1. (i)
    For any multi-indices $\alpha\in\mathbb{Z}_{+}^{n}$ and $\beta\in\mathbb{Z}_{+}^{n-1}$, there exists $C_{\alpha,\beta}>0$ such that, for all $(x,\xi^{\prime},t)\in\mathbb{R}^{n}\times\mathbb{R}^{n-1}\times\mathbb{R}$, we have
    (4.4) $|\partial_{x}^{\alpha}\partial_{\xi^{\prime}}^{\beta}a_{x_{n},t}(x^{\prime},\xi^{\prime})|\leq C_{\alpha,\beta}\langle\xi^{\prime}\rangle^{-|\beta|}\,.$
  2. (ii)
    The set $\{a_{x_{n},t},a_{x_{n},0\pm}\mid x_{n},t\in\mathbb{R}\}$ is a bounded subset of $S^{0}(\mathbb{R}^{n-1}\times\mathbb{R}^{n-1})$.
  3. (iii)
    The function $\mathbb{R}^{n}\times\mathbb{R}^{n-1}\times\mathbb{R}\ni(x,\xi^{\prime},t)\mapsto a_{x_{n},t}(x^{\prime},\xi^{\prime})\in M_{k}(\mathbb{C})$ is continuous except at $t=0$, where it has lateral limits $a_{x_{n},0\pm}(x^{\prime},\xi^{\prime})$.
Proof.
We proceed as in the proof of Lemma 4.2, that is, we write
$a\,=\,b_{-1}+r_{-2}\,,\qquad b_{-1}\in S^{-1}(\mathbb{R}^{n}\times\mathbb{R}^{n})\ \mbox{ and }\ r_{-2}\in S^{-2}(\mathbb{R}^{n}\times\mathbb{R}^{n})\,$,
where $b_{-1}$ is essentially homogeneous of order 1-1. It follows that $r_{-2}$ is a classical symbol and hence it satisfies all the statements of our proposition, by Proposition 4.12 (in the same order of points). Since all the statements of our proposition depend linearly on $a$, it is enough to assume that $a=b_{-1}$, that is, that $a$ is is eventually homogeneous of order 1-1.
To prove (i), we shall consider separately the cases $t\neq 0$ and $t=0$.
Case 1: $t\neq 0$; Step 1.1: The main calculations The idea is to substract from $a$ the “main singularity” using the function $Z(x):=x^{-1}(1-\chi_{0}(x))$ of Lemma 3.2, whose notation we will use in the rest of the proof. This will allow us to express the desired $a_{x_{n},t}(x^{\prime},\xi^{\prime})$ in terms of the inverse Fourier transform of an integrable function and of $W:=\mathcal{F}^{-1}Z$. Recall from Lemma 3.2 that $\chi_{0}\in\mathcal{C}^{\infty}_{c}(\mathbb{R})$ is an even function that is equal to 1 in a neighborhood of 0. In addition to the assumptions of that lemma, let us assume that $\chi_{0}$ has support in (1,1)(-1,1), because then $Z$ will also be essentially homogeneous of order 1-1. We shall thus study the function
(4.5) $u(x,\xi^{\prime},\tau)\mathrel{\,:=\,}a(x,\xi^{\prime},\tau)-\operatorname{\mathfrak{L}}(a;x)Z(\tau)\,=\,a(x,\xi^{\prime},\tau)-a(x,e_{n})\tau^{-1}(1-\chi_{0}(\tau))\,$,
which is essentially homogeneous of order 1-1 (because the support of $\chi_{0}$ is contained in (1,1)(-1,1)). The function $u$ will play an important role in what follows. It will be written, for short, as $u:=a-\operatorname{\mathfrak{L}}Z$. For any multi-index $\alpha\in\mathbb{Z}_{+}^{n}$, the fact that $a$ and $u$ are essential homogeneous of order 1-1 with respect to $\tau$ gives (by dividing by $|\tau|\geq 1$)
$\tau^{2}|\partial_{x}^{\alpha}u(x,\xi^{\prime},\tau)|$ $\mathrel{\,:=\,}|\tau|\,\big|\partial_{x}^{\alpha}a(x,|\tau|^{-1}\xi^{\prime},\operatorname{sgn}(\tau))-\partial_{x}^{\alpha}a(x,e_{n})\operatorname{sgn}(\tau)\big|$
$\,=\,|\tau|\,\big|\partial_{x}^{\alpha}a(x,|\tau|^{-1}\xi^{\prime},\operatorname{sgn}(\tau))-\partial_{x}^{\alpha}a(x,0,\operatorname{sgn}(\tau))\big|\leq C|\xi^{\prime}|\,$,
where $C=\sup|\nabla_{\xi}\partial_{x}^{\alpha}a(x,\xi)|$, which is finite, because $a$ is an order 1-1 symbol and where we have used the hypothesis that $\sigma_{-1}(a;x,\xi)$ is odd in $\xi$. For $|\tau|\geq 1$, this shows that
(4.6) $|\partial_{x}^{\alpha}u(x,\xi^{\prime},\tau)|\mathrel{\,:=\,}|\partial_{x}^{\alpha}a(x,\xi^{\prime},\tau)-\partial_{x}^{\alpha}a(x,e_{n})Z(\tau)|\leq G_{\alpha}\langle\xi^{\prime}\rangle\langle\tau\rangle^{-2}\,$,
for some $G_{\alpha}>0$ independent of $(x,\xi^{\prime})$. The left hand side $|\partial_{x}^{\alpha}u(x,\xi^{\prime},\tau)|$ of the last displayed relation is bounded, so the inequality (4.6) actually holds for all $\tau$ (with possibly different constants $G_{\alpha}$). Consequently, $\langle\xi^{\prime}\rangle^{-1}\partial_{x}^{\alpha}u(x,\xi^{\prime},\tau)$ is integrable with respect to $\tau$ uniformly in $(x,\xi^{\prime})$. That is, the assumptions of the theorem giving the derivability under the integral sign (Theorem 4.6) are satisfied and we obtain that, for all $(x,\xi^{\prime})\in\mathbb{R}^{n}\times\mathbb{R}^{n-1}$ and $t\neq 0$,
$|\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}u(x,\xi^{\prime},t)|\leq G_{\alpha}^{\prime}\langle\xi^{\prime}\rangle\,.$
Let $W:=\mathcal{F}^{-1}Z$, again as in Lemma 3.2, and recall, from that lemma, that $W\in L^{\infty}(\mathbb{R})$, more precisely, we shall use $|W|\leq 1$. By applying the inverse Fourier transform, we then obtain that
(4.7) $|\partial_{x}^{\alpha}a_{x_{n},t}(x^{\prime},\xi^{\prime})|\mathrel{\,:=\,}\big|\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}a(x,\xi^{\prime},t)\big|\,=\,\big|\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}u(x,\xi^{\prime},t)+\partial_{x}^{\alpha}a(x,e_{n})\mathcal{F}_{\xi_{n}}^{-1}Z(t)\big|\\ \leq\big|\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}u(x,\xi^{\prime},t)\big|+\big|\partial_{x}^{\alpha}a(x,e_{n})W(t)\big|\leq G_{\alpha}^{\prime}\langle\xi^{\prime}\rangle+\sup|\partial_{x}^{\alpha}a(x,e_{n})|\leq C_{\alpha}\langle\xi^{\prime}\rangle\,.$
For $|\beta|>0$, the derivatives $\partial_{\xi_{j}}$ kill the term $a(x,e_{n})Z(\tau)$, in $u:=a-\operatorname{\mathfrak{L}}Z$ (Equation (4.5)) which is independent of $\xi$ and hence, in analogy with the estimate (4.6), we have
(4.8) $|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}u(x,\xi^{\prime},\tau)|\mathrel{\,:=\,}|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a(x,\xi^{\prime},\tau)-\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a(x,e_{n})Z(\tau)|\\ \,=\,|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a(x,\xi^{\prime},\tau)|\leq G_{\alpha,\beta}\langle\xi\rangle^{-1-|\beta|}\leq G_{\alpha,\beta}\langle\tau\rangle^{-2}\langle\xi^{\prime}\rangle^{1-|\beta|}\,$,
for some $G_{\alpha,\beta}>0$ independent of $(x,\xi^{\prime})$, where we have used that $\langle\tau\rangle\leq\langle(\xi^{\prime},\tau)\rangle$ and $\langle\xi^{\prime}\rangle\leq\langle(\xi^{\prime},\tau)\rangle$.
In view of our results above, all the $\xi^{\prime}$-derivatives $\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}u(x,\xi^{\prime},\tau)$ of the function $\partial_{x}^{\alpha}u(x,\xi^{\prime},\tau)$ are integrable in $\tau$ uniformly in $(x,\xi^{\prime})$. Therefore
$|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}u(x,\xi^{\prime},\tau)|\leq G_{\alpha,\beta}^{\prime}\langle\xi^{\prime}\rangle^{1-|\beta|}\,.$
We then proceed as in Step 1.1, but using that $\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}\big[a(x,e_{n})\mathcal{F}_{\xi_{n}}^{-1}Z(t)\big]=0$ (which we have already noticed above). We obtain
(4.9) $|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a_{x_{n},t}(x^{\prime},\xi^{\prime})|\mathrel{\,:=\,}\big|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}a(x,\xi^{\prime},t)\big|\\ \,=\,\big|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}u(x,\xi^{\prime},t)+\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a(x,e_{n})\mathcal{F}_{\xi_{n}}^{-1}Z(t)\big|\,=\,\big|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}u(x,\xi^{\prime},t)\big|\\ \leq G_{\alpha,\beta}^{\prime}\langle\xi^{\prime}\rangle^{1-|\beta|}\leq G_{\alpha,\beta}^{\prime}\,.$
Step 1.2: Proof of (i) when $t\neq 0$ and $|\xi^{\prime}|\leq 1$. In this case, the estimates (4.4) of point (i) follow from Equation (4.9) (the last displayed equation) and from Equation (4.7).
Step 1.3: Proof of (i) when $t\neq 0$ and $|\xi^{\prime}|\geq 1$. Again, we notice that we have proved that $\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a_{x_{n},t}(x^{\prime},\xi^{\prime})$ exists for all $(x,\xi^{\prime})$. We shall now remove the constraint $|\xi^{\prime}|\leq 1$ using the essential homogeneity of order 1-1 of $a$ and the homogeneity of the Fourier transform (this easy fact is proved in most basic textbooks, see, for instance, [52, 130]). More precisely, Lemma 4.10 states that $a_{x_{n},t}(x^{\prime},\lambda\xi^{\prime})\,=\,a_{x_{n},\lambda t}(x^{\prime},\xi^{\prime})$ (we use the case m=1m=-1). By replacing $\xi^{\prime}$ with $\lambda^{-1}\xi^{\prime}$ and by taking the product of derivatives $\partial_{x}^{\alpha}$, we obtain $\partial_{x}^{\alpha}a_{x_{n},t}(x^{\prime},\xi^{\prime})\,=\,\partial_{x}^{\alpha}a_{x_{n},\lambda t}(x^{\prime},\lambda^{-1}\xi^{\prime})$. This relation, for $\lambda:=|\xi^{\prime}|$, gives
$|\partial_{x}^{\alpha}a_{x_{n},t}(x^{\prime},\xi^{\prime})|\,=\,|\partial_{x}^{\alpha}a_{x_{n},\lambda t}(x^{\prime},\lambda^{-1}\xi^{\prime})|\leq\sqrt{2}C_{\alpha}\,$,
by Equation (4.7) for $t\neq 0$ replaced with $\lambda^{-1}t\neq 0$ and $\xi^{\prime}$ replaced with $\lambda^{-1}\xi^{\prime}$, the later having length =1=1 (so $\langle\lambda^{-1}\xi^{\prime}\rangle=\sqrt{2}$). Further taking the $\partial_{\xi^{\prime}}^{\beta}$ derivative, this further gives
$|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a_{x_{n},t}(x^{\prime},\xi^{\prime})|\,=\,\big|\partial_{\xi^{\prime}}^{\beta}\big[\partial_{x}^{\alpha}a_{x_{n},\lambda t}(x^{\prime},\lambda^{-1}\xi^{\prime})\big]\big|\\ \,=\,\lambda^{-|\beta|}\big|\big[\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a_{x_{n},\lambda t}\big](x^{\prime},\lambda^{-1}\xi^{\prime})\big|\leq G_{\alpha,\beta}^{\prime}\langle\xi^{\prime}\rangle^{-|\beta|}\,$,
where the last step is by Equation (4.9). The relation (4.4) is thus also proved in this case.
Case 2: $t=0$. This is similar to the case $t\neq 0$. In fact, the only place where we have used $t\neq 0$ was when we estimated the function $u:=a-\operatorname{\mathfrak{L}}Z$ of Equation (4.5) and its Fourier transform employing the properties of the Fourier transform function $W:=\mathcal{F}^{-1}Z$. For $t=0$, we replace $u$ with the symmetrization $u_{0}(x,\xi^{\prime},\tau):=\frac{1}{2}\big(a(x,\xi^{\prime},\tau)+a(x,\xi^{\prime},-\tau)\big)$ and the function $W$ becomes unnecessary. We include now the calculation, but we do not repeat the arguments that are easy and very similar to the ones used in the previous case. Also, we do not separate the proof formally in substeps. The function $u$ will be replaced with a function $u_{0}$ that has similar decay properties and is moreover even in $\tau$. We now proceed according to this plan.
First, for any multi-index $\alpha\in\mathbb{Z}_{+}^{n}$, the essential homogeneity of $a$ with respect to $\tau\geq 1$ gives successively the following:
$2\tau|\partial_{x}^{\alpha}u_{0}(x,\xi^{\prime},\tau)|\mathrel{\,:=\,}\big|\partial_{x}^{\alpha}a(x,\tau^{-1}\xi^{\prime},1)+\partial_{x}^{\alpha}a(x,\tau^{-1}\xi^{\prime},-1)\big|\\ \leq\big|\partial_{x}^{\alpha}a(x,\tau^{-1}\xi^{\prime},1)-\partial_{x}^{\alpha}a(x,0,1)\big|+\big|\partial_{x}^{\alpha}a(x,\tau^{-1}\xi^{\prime},-1)-\partial_{x}^{\alpha}a(x,0,-1)\big|\leq C|\xi^{\prime}|/\tau\,$,
where $C=\sup|\nabla_{\xi}\partial_{x}^{\alpha}a(x,\xi)|$ and where we have used the hypothesis $\sigma_{-1}(a;x,\xi)$ is odd in $\xi$. This shows that
$|\partial_{x}^{\alpha}u_{0}(x,\xi^{\prime},\tau)|\leq G_{\alpha}\langle\xi^{\prime}\rangle\langle\tau\rangle^{-2}\,$,
for some $G_{\alpha}>0$ independent of $(x,\xi^{\prime})$, that is, $\langle\xi^{\prime}\rangle^{-1}\partial_{x}^{\alpha}u_{0}(x,\xi^{\prime},\tau)$ is uniformly integrable in $(x,\xi^{\prime})$, with $|\xi^{\prime}|\leq 1$. We can therefore use the derivability under the integral sign (Theorem 4.6) to obtain that $|\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}u_{0}(x,\xi^{\prime},\tau)|\leq G_{\alpha}^{\prime}\langle\xi^{\prime}\rangle$. By applying the inverse Fourier transform, we then obtain for $|\xi^{\prime}|\leq 1$ that
$|\partial_{x}^{\alpha}a_{x_{n},0}(x^{\prime},\xi^{\prime})|\mathrel{\,:=\,}\big|\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}u_{0}(x,\xi^{\prime},0)\big|\leq G_{\alpha}^{\prime}\langle\xi^{\prime}\rangle\leq G_{\alpha}^{\prime}$
is uniformly bounded in $(x,\xi^{\prime})$. This proves the estimate (4.4) for $t=0$, $|\beta|=0$, and $|\xi^{\prime}|\leq 1$.
Let us assume now that $t=0$, $|\beta|>0$, and $|\xi^{\prime}|\leq 1$. This follows more or less the same calculation as in the previous case. For $|\beta|>0$, in analogy with the estimate (4.6), we have
(4.10) $2|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}u_{0}(x,\xi^{\prime},\tau)|\mathrel{\,:=\,}|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}\big[a(x,\xi^{\prime},\tau)+a(x,\xi^{\prime},-\tau)\big]|\leq G_{\alpha,\beta}\langle\tau\rangle^{-2}\langle\xi^{\prime}\rangle^{-|\beta|}\,.$
In Equation (4.10), all the derivatives $\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}u_{0}(x,\xi^{\prime},\tau)$ of the function $\partial_{x}^{\alpha}u_{0}(x,\xi^{\prime},\tau)$ are uniformly integrable in $(x,\xi^{\prime})$. Therefore $|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}u_{0}(x,\xi^{\prime},\tau)|\leq G_{\alpha,\beta}^{\prime}\langle\xi^{\prime}\rangle^{-|\beta|}$ and
$|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a_{x_{n},0}(x^{\prime},\xi^{\prime})|\mathrel{\,:=\,}\big|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}\mathcal{F}_{\xi_{n}}^{-1}u_{0}(x,\xi^{\prime},0)\big|\leq G_{\alpha,\beta}^{\prime}\langle\xi^{\prime}\rangle^{-|\beta|}\leq G_{\alpha,\beta}^{\prime}\,.$
This gives the estimate (4.4) with $C_{\alpha,\beta}=G_{\alpha,\beta}^{\prime}$ when $|\xi^{\prime}|\leq 1$.
To complete the proof of (i), let now $t=0$ and $|\xi^{\prime}|\geq 1$. Again, we notice that we have proved that $\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a_{x_{n},0}(x^{\prime},\xi^{\prime})$ exists for all $(x,\xi^{\prime})$ and that it is obtained by commuting differentiation with integration. We shall now remove the constraint $|\xi^{\prime}|\leq 1$ using the essential homogeneity of order 1-1 of $a$ and of $u_{0}$ and the homogeneity of the Fourier transform [52, 130]. We can nolonger use Lemma 4.10, but we can proceed as in its proof, since we are dealing with integrable functions. Again for $\lambda=|\xi^{\prime}|\geq 1$ and using in the last step the substitution $\tau=\lambda s$, $s\in\mathbb{R}$, we have
$\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a_{x_{n},0}(x^{\prime},\xi^{\prime})\,=\,\frac{1}{4\pi}\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}\int_{\mathbb{R}}\big(a(x,\xi^{\prime},\tau)+a(x,\xi^{\prime},-\tau))\,d\tau\\ \,=\,\frac{1}{4\pi\lambda}\int_{\mathbb{R}}\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}\big(a(x,\lambda^{-1}\xi^{\prime},\lambda^{-1}\tau)+a(x,\lambda^{-1}\xi^{\prime},-\lambda^{-1}\tau))\,d\tau\\ \,=\,\frac{\lambda^{-1-|\beta|}}{4\pi}\int_{\mathbb{R}}\big((\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a)(x,\lambda^{-1}\xi^{\prime},\lambda^{-1}\tau)+(\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a)(x,\lambda^{-1}\xi^{\prime},-\lambda^{-1}\tau))\,d\tau\\ \,=\,\frac{\lambda^{-|\beta|}}{4\pi}\int_{\mathbb{R}}\big((\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a)(x,\lambda^{-1}\xi^{\prime},s)+(\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a)(x,\lambda^{-1}\xi^{\prime},-s))\,ds\,.$
Equation (4.10) then gives (note that, in that equation, the term $\langle\xi^{\prime}\rangle^{-|\beta|}\leq 1$)
$|\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a_{x_{n},0}(x^{\prime},\xi^{\prime})|\leq\frac{\lambda^{-|\beta|}}{4\pi}\int_{\mathbb{R}}\big|(\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a)(x,\lambda^{-1}\xi^{\prime},s)+(\partial_{\xi^{\prime}}^{\beta}\partial_{x}^{\alpha}a)(x,\lambda^{-1}\xi^{\prime},-s)|\,ds\\ \leq G_{\alpha,\beta}\lambda^{-|\beta|}\int_{\mathbb{R}}\langle s\rangle^{-2}ds\,=\,C_{\alpha,\beta}\langle\xi^{\prime}\rangle^{-|\beta|}\,.$
This completes the proof of (i).
The point (ii) now follows easily. The point (i) gives by definition that $\{a_{x_{n},t}\mid x_{n},t\in\mathbb{R}\}$ is a bounded subset of $S^{0}(\mathbb{R}^{n}\times\mathbb{R}^{n})$. Then, we notice that $\operatorname{\mathfrak{L}}(a;x):=\sigma_{-1}(a;x,e_{n})\in S_{\operatorname{cl}}^{0}(\mathbb{R}^{n-1}\times\mathbb{R}^{n-1})$ and it is a bounded family of order 0 classical symbols on $\mathbb{R}^{n-1}$ parametrized by $x_{n}\in\mathbb{R}$ (all these symbols are multiplications operators, since they are independent of $\xi^{\prime}$). The formula $a_{x_{n},0\pm}(x,\xi^{\prime})\,=\,\pm\frac{\imath}{2}\operatorname{\mathfrak{L}}(a;x)+a_{x_{n},0}(x,\xi^{\prime})$ then gives (ii).
Finally, (iii) follows from Corollary 4.4 using the notation of Definition 4.3. ∎
To formulate our main result on lateral (or boundary) limits of pseudodifferential operators of order =1=-1 (Theorem 4.16 next), recall the following notation and results:
  1. (i)
    $u_{\epsilon}:=u|{\Gamma_{\epsilon}}$ and $u_{\pm}:=\lim_{\epsilon\to 0\pm}u_{\epsilon}$ were introduced in Definition (3.9).
  2. (ii)
    The distribution $\langle h\otimes\delta_{\Gamma},\phi\rangle:=\int_{\Gamma}h\phi dx^{\prime}$ was introduced in Lemma 3.8.
  3. (iii)
    We know that $a_{t}=a_{t,t}$, $t\in\mathbb{R}$, and $a_{0\pm}:=a_{0,0\pm}$, see Definition (4.3), Proposition 4.8, and the comment following its proof.
We again formulate the following theorem for sections of a $k$-dimensional trivial vector bundle on $\mathbb{R}^{n}$.
Theorem 4.16.
We use the notation in Definition 4.3 and the relation $a_{\epsilon}=a_{\epsilon,\epsilon}$, see Remark 4.9. Let $a\in S^{-1}_{\operatorname{cl}}(\mathbb{R}^{n}\times\mathbb{R}^{n};M_{k}(\mathbb{C}))$ and assume that $\sigma_{-1}(a)$ is odd (in the sense that $\sigma_{-1}(a;x,-\xi)=-\sigma_{-1}(a;x,\xi)$ for all $\xi\in\mathbb{R}^{n}$).
  1. (i)
    Let $k_{a_{0}(x^{\prime},D^{\prime})}$ be the distribution kernel of $a_{0}(x^{\prime},D^{\prime})$ and $k_{a(x,D)}$ be the distribution kernel of $a(x,D)$. Then both $k_{a_{0}(x^{\prime},D^{\prime})}(x^{\prime},y^{\prime})$ and $k_{a(x,D)}(x,y)$ are smooth for $x^{\prime},y^{\prime}\in\mathbb{R}^{n-1}$, $x^{\prime}\neq y^{\prime}$, and
    $k_{a_{0}(x^{\prime},D^{\prime})}(x^{\prime},y^{\prime})\,=\,k_{a(x,D)}(x^{\prime},0,y^{\prime},0)\,.$
  2. (ii)
    $\sigma_{0}(a_{0})$ is also odd, and, for all $h\in\mathcal{C}^{\infty}_{c}(\mathbb{R}^{n-1})$,
    $a_{0}(x^{\prime},D^{\prime})h(x)\,=\,\operatorname{p.\!\!v.}\,\int_{\mathbb{R}^{n-1}}k_{a(x,D)}(x^{\prime},y^{\prime})h(y^{\prime})\,dy^{\prime}\,.$
  3. (iii)
    The three symbols $a_{0}:=a_{0,0}$ and $a_{0\pm}:=a_{0,0\pm}$ are order zero classical, with principal parts (or symbols)
    $\sigma_{0}(a_{0};x^{\prime},\xi^{\prime})\,=\,\frac{1}{2\pi}\int_{\mathbb{R}}\frac{\sigma_{-1}(a;x^{\prime},0,\xi^{\prime},\xi_{n})+\sigma_{-1}(a;x^{\prime},0,\xi^{\prime},-\xi_{n})}{2}\,d\xi_{n}$
    and $\sigma_{0}(a_{0\pm};x^{\prime},\xi^{\prime})=\pm\frac{\imath}{2}\sigma_{-1}(a;x^{\prime},e_{n})+\sigma_{0}(a_{0};x^{\prime},\xi^{\prime})$
  4. (iv)
    For all $s\in\mathbb{R}$ and all $h\in H^{s}(\Gamma;\mathbb{C}^{k})=H^{s}(\Gamma)^{k}$ and all $\epsilon\neq 0$ small, we have $\big[a(x,D)(h\otimes\delta_{\Gamma})]_{\epsilon}=a_{\epsilon}(x^{\prime},D^{\prime})h$.
  5. (v)
    For all $s\in\mathbb{R}$ and for all $h\in H^{s}(\mathbb{R}^{n-1};\mathbb{C}^{k})=H^{s}(\mathbb{R}^{n-1})^{k}$, we have
    $\mathcal{S}_{a}h]_{\pm}\mathrel{\,:=\,}\lim_{\epsilon\to\pm 0}\big[a(x,D)(h\otimes\delta_{\Gamma})]_{\epsilon}\,=\,\lim_{\epsilon\to\pm 0}a_{\epsilon}(x^{\prime},D^{\prime})h\,=\,a_{0\pm}(x^{\prime},D^{\prime})h\in H^{s}(\mathbb{R}^{n-1})^{k}\,.$
Proof.
We again assume $k=1$. Let us prove (i). As is well known the distribution kernel $k_{a(x,D)}(x,y)$ of $a(x,D)$ is smooth for $x\neq y$ and there it is given by $k_{a(x,D)}(x,y)\,=\,\mathcal{F}_{\xi}^{-1}a(x,x-y)\,$, where $\mathcal{F}_{\xi}^{-1}$ is the inverse Fourier transform in the second variable (i.e., in $\xi\in\mathbb{R}^{n}$; see Lemma 3.6). Let $\tilde{a}(x,\xi):=a(x,\xi^{\prime},-\xi_{n})$. Using the formula defining $a_{0}=a_{0,0}$ (see Definition 4.3), we obtain
$k_{a_{0}(x^{\prime},D^{\prime})}(x^{\prime},y^{\prime})$ $\,=\,\mathcal{F}_{\xi^{\prime}}^{-1}a_{0}(x^{\prime},x^{\prime}-y^{\prime})$
$\,=\,\frac{1}{2}\mathcal{F}_{\xi^{\prime}}^{-1}\mathcal{F}_{\xi_{n}}^{-1}(a+\tilde{a})(x^{\prime},0,x^{\prime}-y^{\prime},0)$
$\,=\,\mathcal{F}_{\xi}^{-1}a(x^{\prime},0,x^{\prime}-y^{\prime},0)$
$\,=\,k_{a}(x^{\prime},0,x^{\prime}-y^{\prime},0)\,$,
because $\mathcal{F}_{\xi}^{-1}\tilde{a}(x,y)=\mathcal{F}_{\xi}^{-1}a(x,y^{\prime},-y_{n})$ and $y_{n}=0$ in our equation. This proves the first point.
Property (ii) is a general result on pseudodifferential operators of order 0 with odd symbol [54]. Indeed, (ii) was already proved for operators of order <1<-1, see Theorem 4.13. We can thus assume that $a$ is odd and essentially homogeneous of order 1-1. Then it follows from Lemma 4.10 that $a_{0}$ is also odd and homogeneous of order 0. Let $Bu(x^{\prime}):=\operatorname{p.\!\!v.}\,\int_{\mathbb{R}^{n-1}}k_{a_{0}(x,D)}(x^{\prime},y^{\prime})u(y^{\prime})\,dy^{\prime}$. In view of (i), it is enough to prove that $f:=a_{0}(x^{\prime},D^{\prime})-B=0$. The distribution kernel of the operator $f$ is hence supported on the diagonal. This means that $f$ preserves supports. It is then known (Peetre’s theorem) that $f:=a_{0}(x^{\prime},D^{\prime})-B$ is a multiplication operator by some function $f_{0}$, see, for example, Theorem 8.1.8 (ii) in [54] (on pages 508 and 509). Thus the distribution kernel of $f$ is of the form
$k_{f}(x^{\prime},y^{\prime})\,=\,f_{0}(x^{\prime})\delta(x^{\prime}-y^{\prime})\,=\,f_{0}(x^{\prime})\delta(y^{\prime}-x^{\prime})\,=\,k_{f}(x^{\prime},2x^{\prime}-y^{\prime})\,.$
But $k_{a_{0}(x,D)}(x^{\prime},y^{\prime})=-k_{a_{0}(x,D)}(x^{\prime},2x^{\prime}-y^{\prime})$, since $a_{0}$ is an odd symbol (we have used a similar calculation to the one just performed, the formula for the distribution kernel in Lemma 3.6, and the fact that the Fourier transform of an odd function is an odd function). The same property is satisfied by the distribution kernel of $B$. Therefore $k_{f}(x^{\prime},y^{\prime})=-k_{f}(x^{\prime},2x^{\prime}-y^{\prime})=-k_{f}(x^{\prime},y^{\prime})$, and, hence $f=0$.
Let us prove the third point. As in the proof of Proposition 4.15, we may assume that $a$ is essentially homogeneous of order 1-1. Then $\sigma_{-1}(a;x,\xi)=a(x,\xi)$ for $|\xi|\geq 1$. Lemma 4.10 then gives that $a_{x_{n},0}$ is also essentially homogeneous of order 0, and hence $\sigma_{0}(a_{x_{n},0};x^{\prime},\xi^{\prime})=a_{x_{n},0}(x^{\prime},\xi^{\prime})$ for $|\xi|\geq 1$. We then have for $|\xi^{\prime}|\geq 1$
$\sigma_{0}(a_{x_{n},0};x^{\prime},\xi^{\prime})\,=\,a_{x_{n},0}(x^{\prime},\xi^{\prime})\,=\,\int_{\mathbb{R}}\frac{a(x,\xi^{\prime},\tau)+a(x,\xi^{\prime},-\tau)}{4\pi}\,d\tau\\ \,=\,\int_{\mathbb{R}}\frac{\sigma_{-1}(a;x,\xi^{\prime},\tau)+\sigma_{-1}(a;x,\xi^{\prime},-\tau)}{4\pi}\,d\tau\,.$
Setting $x_{n}=0$ and using the definitions of $a_{0,0\pm}:=a_{0,0}\pm\frac{\imath}{2}\operatorname{\mathfrak{L}}(a)$ (Definition 4.3) we obtain (iii).
The point (iv) is the content of Proposition 4.8 (see also Remark 4.9).
Recall that, for every inner product vector space $V$, the map $\sharp:V\to V^{*}$ is defined by $v^{\sharp}(w):=(w,v)$. Finally, to prove (v), let us consider the matrix unit $e_{ij}:=e_{i}\otimes e_{j}^{\sharp}\in M_{k}(\mathbb{C})$. By splitting our symbol as $a=\sum a_{ij}e_{ij}$, we can assume that $k=1$. The first equality of (v) is the definition of $\mathcal{S}_{a}$. The second equality of (v) is the content of (iv) just discussed. Then, Proposition 4.15 shows that the assumptions of Lemma 4.7 are satisfied by the family $a_{\epsilon,\epsilon}(x^{\prime},D^{\prime})$ as $\epsilon\to 0$ with either positive or negative values, which gives the last equality in (v) and allows us to conclude (iv) and completes the proof of the theorem. ∎
Recall that we have $a_{\epsilon}=a_{\epsilon,\epsilon}$ and that we write $a_{0,0\pm}=:a_{0\pm}$, for simplicity. Most of the relations of Theorem 4.16 have been written in a compact form. The expanded form of these relations amounts to the following five relations:
$\begin{gathered}\big[a(x,D)(h\otimes\delta_{\Gamma})]_{+}\mathrel{\,:=\,}\lim_{\epsilon\searrow 0}\big[a(x,D)(h\otimes\delta_{\Gamma})]_{\epsilon}\,=\,\lim_{\epsilon\searrow 0}a_{\epsilon}(x^{\prime},D^{\prime})h\,=\,a_{0+}(x^{\prime},D^{\prime})h\\ \big[a(x,D)(h\otimes\delta_{\Gamma})]_{-}\mathrel{\,:=\,}\lim_{\epsilon\nearrow 0}\big[a(x,D)(h\otimes\delta_{\Gamma})]_{\epsilon}\,=\,\lim_{\epsilon\nearrow 0}a_{\epsilon}(x^{\prime},D^{\prime})h\,=\,a_{0-}(x^{\prime},D^{\prime})h\\ \sigma_{0}(a_{x_{n},0};x^{\prime},\xi^{\prime})\,=\,\frac{1}{2\pi}\int_{\mathbb{R}}\frac{\sigma_{-1}(a;x^{\prime},x_{n},\xi^{\prime},\xi_{n})+\sigma_{-1}(a;x^{\prime},x_{n},\xi^{\prime},-\xi_{n})}{2}\,d\xi_{n}\\ \sigma_{0}(a_{x_{n},0+};x^{\prime},\xi^{\prime})\,=\,\frac{\imath}{2}\sigma_{-1}(a;\lambda e_{n})+\sigma_{0}(a_{x_{n},0};x^{\prime},\xi^{\prime})\qquad\mbox{and}\\ \sigma_{0}(a_{x_{n},0-};x^{\prime},\xi^{\prime})\,=\,-\frac{\imath}{2}\sigma_{-1}(a;\lambda e_{n})+\sigma_{0}(a_{x_{n},0};x^{\prime},\xi^{\prime})\,.\end{gathered}$
Example 4.17.
Let us look at the example of the double layer potential associated to $\Delta=\sum_{j=1}^{n}\partial_{j}^{2}$ on the half-space $\mathbb{R}_{+}^{n}$ for $n\geq 3$. The outer unit normal vector field is $e_{n}$. Hence the pseudodifferential operator $a(x,D)$ giving the double layer potential has distribution kernel $\partial_{y_{n}}N(x-y)$, where $N(x)=\gamma_{n}|x|^{2-n}$, for some constant $\gamma_{n}$, whose value is irrelevant for the current considerations. Let $C_{N}$ be the operator with kernel N(xy)N(x-y). Then $a(x,D)=C_{N}(-\partial_{n})^{*}=C_{N}\partial_{n}$. The symbol of the operator $C_{N}$ with kernel N(xy)N(x-y) is $|\xi|^{-2}$ (that is, the symbol of $\Delta^{-1}$, because $\sigma_{2}(\Delta)=-|\xi|^{2}$). The symbol of $\partial_{n}$ is $\sigma_{1}(\partial_{n})=\imath\xi_{n}$. Therefore, the symbol of $a(x,D)$ is
$a(x,\xi)\,=\,-\imath\xi_{n}|\xi|^{-2}\,$,
which is essentially homogeneous of order 1-1. This gives $a_{x_{n},0}=0$. Consequently
$a(x,D)h]_{\pm}\,=\,\pm\frac{\imath}{2}\operatorname{\mathfrak{L}}(a)h\,=\,\pm\frac{1}{2}\,h\,.$
5. Pseudodifferential operators on manifolds with straight cylindrical ends
From now on, $M$ will be a manifold with straight cylindrical ends, $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$. Also, in the following, $E\to M$ will be a fixed compatible Hermitian vector bundle. We now define two needed classes of pseudodifferential operators on $M$, the $\operatorname{inv}$-calculus” and the $\operatorname{ess}$-calculus.” For $M$ closed, these definitions are trivial (they recover the usual pseudodifferential operators, so the give nothing new), so the only interesting case is the one when $M$ is not compact. Including the case $M$ compact is useful, however, in order to formulate our results in a more general setting, as they are true for both closed manifolds and for manifolds with straight cylindrical ends.
The resulting classes of pseudodifferential operators are very closely related to the $b$- and $c$- calculi of Melrose [93, 84]. Related calculi have also been considered by Schulze and Schrohe [121, 122, 124]. See [67] for more references and for a discussion of the relation between our calculi and the ones of Mazzeo, Melrose, Schrohe, and Schulze. The results of this section are close to those in [63, 67]. The presentation is, however, different and, occasionally, it provides more general results. For instance, in [63], we work only with classical pseudodifferential operators, whereas here we allow the Kohn-Nirenberg classes of symbols recalled in Definition 3.5.
5.1. Translation invariance in a neighborhood at infinity: inv-calculus
Let $M=M_{0}\cup\left[M^{\prime}\times(-\infty,0)\right]$ be a manifold with straight cylindrical ends, such that $M_{0}$ is a compact manifold and $M^{\prime}:=\partial M_{0}\neq\emptyset$. Let $\Psi^{m}(M;E,F)$ denote the set of all order $m\in\mathbb{R}$ pseudodifferential operators $P:\mathcal{C}^{\infty}_{c}(M;E)\to\mathcal{C}^{\infty}(M;F)$. For any continuous, linear operator $P:\mathcal{C}^{\infty}_{c}(M;E)\to\mathcal{C}^{\infty}_{c}(M;F)^{\prime}$, we let $k_{P}$ denote its distribution kernel.
For all $R,s\geq 0$, we let
(5.1) $\Phi_{s}:M^{\prime}\times(-\infty,R]\to M^{\prime}\times(-\infty,R-s]\,,\quad\Phi_{s}(x,t)=(x,t-s)\,$,
which is a bijective isometry (given by the translation with s-s in the $t$-direction). If $s<0$, then $\Phi_{s}:=\Phi_{-s}$. We fix $R\ll 0$ (i.e., much less than 0) so the maps $\Phi_{s}$ extend to sections of our given compatible vector bundles by
[…]
We let $\operatorname{dist}_{g}$ denote the distance function on $M$ defined by the given complete metric $g$ (which, we recall, was assumed to be a product on the cylindrical end $M^{\prime}\times(-\infty,0]$). Then, for $\varepsilon>0$, we define the the $\varepsilon$–neighborhood of the diagonal to be the set
(5.2) $\{(x,y)\in M\times M:\operatorname{dist}_{g}(x,y)<\varepsilon\}\,.$
Recall the definition of pseudodifferential operators $\Psi^{m}(M;E,F)$, Equation (3.14).
Definition 5.1.
A continuous linear operator $P:\mathcal{C}^{\infty}_{c}(M,E)\to\mathcal{C}^{\infty}(M,F)^{\prime}$ is called translation invariant in a neighborhood of infinity if the following two conditions are satisfied:
  1. (i)
    There exists $\varepsilon_{P}>0$ such that its distribution kernel $k_{P}$ is supported in the $\varepsilon_{P}$–neighborhood of the diagonal, see Equations (3.15) and (5.2).
  2. (ii)
    There exists $R_{P}<R_{E}-\varepsilon_{P}$ such that, for all $u\in\mathcal{C}^{\infty}_{c}\big(M^{\prime}\times(-\infty,R_{P}),E\big)$ and all $s>0$, we have $P\Phi_{s}(u)=\Phi_{s}(Pu).$
We let $\Psi_{\operatorname{inv}}^{m}(M;E,F)$ denote the space of all classical pseudodifferential operators $P:\mathcal{C}^{\infty}_{c}(M,E)\to\mathcal{C}^{\infty}(M,F)$ of order $m\in\mathbb{R}\cup\{\pm\infty\}$ that are translation invariant in a neighborhood of infinity.
Thus $\Psi_{\operatorname{inv}}^{m}(M;E,F)\subset\Psi^{m}(M;E,F)$. We also let $\Psi_{\operatorname{inv}}^{m}(M;E):=\Psi_{\operatorname{inv}}^{m}(M;E,E)$, $\Psi_{\operatorname{inv}}^{m}(M):=\Psi_{\operatorname{inv}}^{m}(M;\mathbb{C}):=\Psi_{\operatorname{inv}}^{m}(M;\mathbb{C},\mathbb{C})$, $\Psi_{\operatorname{inv}}^{-\infty}(M;E,F):=\bigcap_{m\in{\mathbb{Z}}}\Psi_{\operatorname{inv}}^{m}(M;E,F)$, and $\Psi_{\operatorname{inv}}^{\infty}(M;E,F):=\bigcup_{m\in{\mathbb{Z}}}\Psi_{\operatorname{inv}}^{m}(M;E,F)$. We let $\Psi_{\operatorname{inv}}^{m}(M;E,F):=\Psi^{m}(M;E,F)$ if $M$ is closed, and similarly for the other notation.
Remark 5.2.
Let us notice that, if $u$ has support in $M^{\prime}\times(-\infty,R)$, $R<0$ small, then $\Phi_{s}(u)$ has support in $M^{\prime}\times(-\infty,R-s)$ and $Pu$ has support in $M^{\prime}\times(-\infty,R+\varepsilon_{P})$. Consequently, if $u$ has support in $M^{\prime}\times(-\infty,R)$ and $s<0$, then both $P\Phi_{s}(u)$ and $\Phi_{s}(Pu)$ are defined, because $R_{P}<R_{E}-\varepsilon_{P}$. Consequently, the condition $P\Phi_{s}(u)=\Phi_{s}(Pu)$ of Definition 5.1 makes sense.
The calculus $\Psi_{\operatorname{inv}}^{m}(M;E,F)$ of Definition 5.1 will be called the $\operatorname{inv}$-calculus [67]. Our $\operatorname{inv}$-calculus is smaller than the “$b$-calculus” as follows from the following remark (see also [67]).
Remark 5.3.
Let $\overline{M}$ denote the compactification of $M$ using the function $r:=e^{t}$ (the Kondratiev transform), so that $M=\{r>0\}$ and $\partial\overline{M}=\{r=0\}$. Then
$\Psi_{\operatorname{inv}}^{m}(M)\subset\Psi_{b}^{m}(\overline{M})\,$,
where $\Psi_{b}^{m}(\overline{M})$ is the “$b$-calculus” of Melrose [93, 94] and Schulze [122, 123]. (Schulze calls it the “cone-calculus,” see an important paper by Lauter and Seiler [80] comparing the two approaches. See also [34, 47, 67, 83, 84, 124, 126] for further results on the $b$-calculus.)
Recall that $\operatorname{Hom}(E;F)\to M$ denotes the vector bundle of all vector bundle homomorphisms $E\to F$. Then $\operatorname{Hom}(E;F)\to M$ is also a compatible vector bundle over $M$ (because $E$ and $F$ are compatible). The next remark provides some examples of operators in $\Psi_{\operatorname{inv}}^{m}(M;E,F)$.
Remark 5.4.
We have
$\mathcal{C}^{\infty}_{\operatorname{inv}}(M;\operatorname{Hom}(E;F))$ $\,=\,\mathcal{C}^{\infty}(M;\operatorname{Hom}(E;F))\cap\Psi_{\operatorname{inv}}^{m}(M;E,F)\quad\mbox{and}$
$\operatorname{Diff}_{\operatorname{inv}}^{m}(M;E,F)$ $\,=\,\{P\mbox{ a differential operator on }M\}\cap\Psi_{\operatorname{inv}}^{m}(M;E,F)\,.$
see Definitions 3.14 and 3.16.
Recall that $k_{P}$ denotes the distribution kernel of a continuous, linear operator $P:\mathcal{C}^{\infty}_{c}(M;E)\to\mathcal{C}^{\infty}_{c}(M;F)^{\prime}$.
Definition 5.5.
Let $\Psi_{\operatorname{comp}}^{m}(M;E,F)$ denote the subspace of pseudodifferential operators $P\in\Psi^{m}(M;E,F)$ with compactly supported distribution kernel $k_{P}$.
The following result follows from the definition of the $\operatorname{inv}$-calculus.
Proposition 5.6.
Let $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ be a manifold with cylindrical ends. Then $\Psi_{\operatorname{comp}}^{m}(M;E,F)\subset\Psi_{\operatorname{inv}}^{m}(M;E,F)$.
5.2. Principal symbol
The principal symbol plays a crucial role in the study of the $\operatorname{inv}$-calculus. Let $S^{*}M$ be the set of unit vectors of $T^{*}M$ and let
$\sigma_{m}:\Psi^{m}(M;E,F)\to\mathcal{C}^{\infty}(S^{*}M;\operatorname{Hom}(E,F))$
be the principal symbol map of classical pseudodifferential operators of order $\leq m$ (see Definition 3.7). We have the usual notion of ellipticity.
Definition 5.7.
An operator $P\in\Psi_{\operatorname{inv}}^{m}(M;E,F)$ is called elliptic if its principal symbol $\sigma_{m}(P)\in\mathcal{C}^{\infty}(S^{*}M;\operatorname{Hom}(E,F))$ is invertible on $S^{*}M\subset T^{*}M$.
In the case of $\operatorname{inv}$-calculus, we have the additional property that the principal symbol of a pseudodifferential operator that is translation invariant in a neighborhood of infinity is also translation invariant at infinity [67].
Proposition 5.8.
The manifold $S^{*}M:=\{\xi\in T^{*}M:\|\xi\|=1\}$ is also a manifold with straight cylindrical ends. The restriction of the principal symbol $\sigma_{m}:\Psi^{m}(M;E,F)\to\mathcal{C}^{\infty}(S^{*}M;\operatorname{Hom}(E;F))$ to $\Psi_{\operatorname{inv}}^{m}(M;E,F)$ maps it to $\mathcal{C}^{\infty}_{\operatorname{inv}}(S^{*}M;\operatorname{Hom}(E,F))$, and hence we obtain a map $\sigma_{m}:\Psi_{\operatorname{inv}}^{m}(M;E,F)\to\mathcal{C}^{\infty}_{\operatorname{inv}}(S^{*}M;\operatorname{Hom}(E,F))\,.$
5.3. Algebra properties of the $\rm{inv}$}-calculus
We summarize now some of the algebra properties of the $\operatorname{inv}$-calculus $\Psi_{\operatorname{inv}}^{m}$. All of them are direct analogues of the corresponding properties for the usual pseudodifferential calculus on compact manifolds or for other calculi on manifolds with cylindrical ends. For simplicity, we shall mention our results only in the case $E=F$. The general case is obtained easily by replacing $E$ with $E\oplus F$. See also [98, 63, 67, 68].
Theorem 5.9.
Let $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ be a manifold with straight cylindrical ends and let $E,F,G\to M$ be compatible vector bundles on $M$. Then the following properties hold.
  1. (i)
    If $P\in\Psi_{\operatorname{inv}}^{m}(M;E,F)$, then its adjoint $P^{*}\in\Psi_{\operatorname{inv}}^{m}(M;F,E)$, and, for $m,m^{\prime}\in\mathbb{R}$,
    $\Psi_{\operatorname{inv}}^{m^{\prime}}(M;F,G)\Psi_{\operatorname{inv}}^{m}(M;E,F)\subset\Psi_{\operatorname{inv}}^{m^{\prime}+m}(M;E,G)\,.$
  2. (ii)
    The principal symbol $\sigma_{m}:\Psi_{\operatorname{inv}}^{m}(M;E,F)\to\mathcal{C}^{\infty}_{\operatorname{inv}}(S^{*}M;\operatorname{Hom}(E;F))$ of Propositioin 5.8 is onto and its kernel is $\Psi_{\operatorname{inv}}^{m-1}(M;E,F)$. It is *-multiplicative, in the sense that
    $\sigma_{m^{\prime}+m}(QP)\,=\,\sigma_{m^{\prime}}(Q)\sigma_{m}(P)\ \mbox{ and }\ \sigma_{m}(P^{*})\,=\,\sigma_{m}(P)^{*}\,$,
    if $Q\in\Psi_{\operatorname{inv}}^{m^{\prime}}(M;F,G)$ and $P\in\Psi_{\operatorname{inv}}^{m}(M;E,F)$.
  3. (iii)
    An operator $P\in\Psi_{\operatorname{inv}}^{m}(M;E,F)$ is elliptic if, and only if, there exists $Q\in\Psi_{\operatorname{inv}}^{-m}(M;F,E)$ such that $QP-1\in\Psi_{\operatorname{inv}}^{-\infty}(M;E)$ and $PQ-1\in\Psi_{\operatorname{inv}}^{-\infty}(M;F)$.
5.4. Translation invariant operators on a cylinder and their kernels
We now look at the special case $M=\mathfrak{S}\times\mathbb{R}$, where $\mathfrak{S}$ is a given Riemannian manifold. This special example of a manifold with cylindrical ends is useful in studying the general case. Of course, the cylinder $M=\mathfrak{S}\times\mathbb{R}$ is a special case of manifolds with cylindrical ends (with $M_{0}=\mathfrak{S}\times[-1,1]$ and $M^{\prime}:=\{-1,1\}$, see Definition 3.11 and Remark 3.12.)
To simplify the notation, we shall set as in Melrose’s work
(5.3) $\psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F)\mathrel{\,:=\,}\Psi_{\operatorname{inv}}^{m}(\mathfrak{S}\times\mathbb{R};E,F)^{\mathbb{R}}\,.$
We begin with some simple observations on the supports of our operators that will allow us to define the normal (or limit) operator associated to an operator that is translation invariant in a neighborhood of infinity. Let $E,F\to\mathfrak{S}$ be two Hermitian vector bundles that we lift to compatible vector bundles on $\mathfrak{S}\times\mathbb{R}$ denoted with the same symbols. We have the following useful characterization of the space $\Psi_{\operatorname{inv}}^{m}(\mathfrak{S}\times\mathbb{R};E,F)^{\mathbb{R}}$ [67].
Lemma 5.10.
The space $\psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F):=\Psi_{\operatorname{inv}}^{m}(\mathfrak{S}\times\mathbb{R};E,F)^{\mathbb{R}}$ consists of the set of order $m$, translation invariant, properly supported, pseudodifferential operators on $\mathfrak{S}\times\mathbb{R}$ (invariant with respect to the translations by $t\in\mathbb{R}$), that is
$\psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F)\,=\,\{P\in\Psi^{m}(\mathfrak{S}\times\mathbb{R};E,F)^{\mathbb{R}}:\ P\mbox{ is properly supported}\,\}.$
5.5. Limit operators
An important concept associated to the operators in the $b$-calculus is that of “normal operator” (in Melrose’s terminology) or “conormal principal symbol” (in Schulze’s terminology, who uses the term “cone-calculus” instead of the $b$-calculus). This concept is specific to non-compact manifolds and is easy to be defined for the $\rm{inv}$}-calculus. We shall call these operators “limit at infinity operators,” (or “limit operators,” for short). None of the results of this section have analogues for closed manifolds.
The translation invariant operators appear in the following lemma that will provide the definition of limit operators. The proof of this lemma can be found in [67] (see also [63], [68]).
Lemma 5.11.
Let $M=M_{0}\cup\left[M^{\prime}\times(-\infty,0)\right]$ be a manifold with straight cylindrical ends. Let $P\in\Psi_{\operatorname{inv}}^{m}(M;E,F)$ and $u\in\mathcal{C}^{\infty}_{c}(M^{\prime}\times\mathbb{R};E)$. Then, there exists $R_{P,u}>0$ such that $\Phi_{-s}P\Phi_{s}(u)$ is defined and independent of $s>R_{P,u}$. Then, for $s>R_{P,u}$, we let
${\mathcal{R}_{\infty}}(u)\mathrel{\,:=\,}\Phi_{-s}P\Phi_{s}(u)\,.$
  1. (i)
    If $a\in\mathcal{C}^{\infty}_{\operatorname{inv}}(M;\operatorname{Hom}(E;F))\subset\Psi_{\operatorname{inv}}^{m}(M;E,F)$, then ${\mathcal{R}_{\infty}}(a)$ is the same as the one in Definition 3.14.
  2. (ii)
    In general, ${\mathcal{R}_{\infty}}(P)\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F):=\Psi_{\operatorname{inv}}^{m}(M^{\prime}\times\mathbb{R};E,F)^{\mathbb{R}}$.
  3. (iii)
    If $M=M^{\prime}\times\mathbb{R}$ and $P$ is translation invariant (that is, $P\in\Psi_{\operatorname{inv}}^{m}(M^{\prime}\times\mathbb{R};E,F)^{\mathbb{R}}$), then ${\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}=P$.
Definition 5.12.
Let $P\in\Psi_{\operatorname{inv}}^{m}(M;E,F)$ and
${\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)\mathrel{\,:=\,}\Psi^{m}(M\times\mathbb{R};E,F)^{\mathbb{R}}$
be as in Lemma 5.11. Then $\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)$} is called the limit operator associated to $P$. The resulting map
${\mathcal{R}_{\infty}}:\Psi_{\operatorname{inv}}^{m}(M;E,F)\ni P\longmapsto{\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)$
will be called the limit operator map.
The following proposition provides a good intuition about the limit operator ${\mathcal{R}_{\infty}}P$ [67].
Proposition 5.13.
Let $P\in\Psi_{\operatorname{inv}}^{m}(M;E,F)$. Then, for all $(x,t)\neq(y,s)\in M^{\prime}\times(-\infty,0)$ and all $\lambda\in\mathbb{R}$ large enough
$k_{{\mathcal{R}_{\infty}}(P)}(x,t,y,s)\,=\,k_{P}(x,t-\lambda,y,s-\lambda)\in(F\boxtimes E)_{(x,t,y,s)}\mathrel{\,:=\,}F_{x}\otimes E_{y}$
5.6. The structure of the $\operatorname{inv}$-calculus
We shall make repeated us of the following cut-off function $\eta$ that will be fixed throughout this paper.
Notation 5.14.
We fix a smooth “cut-off” function $\eta:M^{\prime}\times\mathbb{R}\to[0,1]$
$\eta(x,t)\,=\,\begin{cases}\ 1\ \mbox{ for }\ t\leq-2\\ \ 0\ \mbox{ for }\ t\geq-1\,.\end{cases}$
We extend $\eta$ to a function on $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ by setting it to be equal to 0 on $M_{0}$. Then, for $T\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F):=\Psi_{\operatorname{inv}}^{m}(M^{\prime}\times\mathbb{R};E,F)^{\mathbb{R}}$, we let
$s_{0}(T)\mathrel{\,:=\,}\eta T\eta:\mathcal{C}^{\infty}_{c}(M;E)\to\mathcal{C}^{\infty}_{c}(M;F)\,.$
The following lemma from [67], proves, in particular, that the limit operator $\mathcal{R}_{\infty}$} of Definition 5.12 is surjective. Recall that $\Psi_{\operatorname{comp}}^{m}(M;E,F)\subset\Psi^{m}(M;E,F)$ denote the subspace of pseudodifferential operators with compactly supported distribution kernel, see Definition 5.5. Some of these properties follow from the corresponding properties of the $b$-calculus.
Proposition 5.15.
Let $T\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)$ and let $s_{0}(T):=\eta T\eta$ be as in 5.14. Then the following properties hold.
  1. (i)
    $s_{0}(T)\in\Psi_{\operatorname{inv}}^{m}(M;E,F)$ satisfies ${\mathcal{R}_{\infty}}(s_{0}(T))=T.$
  2. (ii)
    The map ${\mathcal{R}_{\infty}}:\Psi_{\operatorname{inv}}^{m}(M;E,F)\to\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)$ is multiplicative and surjective and $\ker({\mathcal{R}_{\infty}}):={\mathcal{R}_{\infty}}^{-1}(0)=\Psi_{\operatorname{comp}}^{m}(M;E,F)$.
  3. (iii)
    The map $P\mapsto\big(s_{0}({\mathcal{R}_{\infty}}(P)),P-s_{0}({\mathcal{R}_{\infty}}(P))\big)$ yields the following direct sum decomposition
    $\Psi_{\operatorname{inv}}^{m}(M;E,F)\,=\,s_{0}(\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F))+\Psi_{\operatorname{comp}}^{m}(M;E,F)\,.$
  4. (iv)
    For any $T\in\Psi_{\operatorname{inv}}^{m}(M;E,F)$, we have
    $\sigma_{m}({\mathcal{R}_{\infty}}(T))\,=\,{\mathcal{R}_{\infty}}(\sigma_{m}(T))\in\mathcal{C}^{\infty}(S^{*}(M^{\prime}\times\mathbb{R});\operatorname{Hom}(E;F))^{\mathbb{R}}\,.$
We also have the following characterization of Sobolev spaces [67], [63].
Corollary 5.16.
Let $u\in H^{t}(M;E)$, for some $t\in\mathbb{R}$. We have $u\in H^{s}(M;E)$ if, and only if, $Pu\in L^{2}(M;E)$ for all $P\in\Psi_{\operatorname{inv}}^{s}(M;E)$.
5.7. The $\operatorname{ess}$-calculus of essentially translation invariant operators
We continue to assume that $M=M_{0}\cup\left[M^{\prime}\times(-\infty,0)\right]$ is a manifold with straight cylindrical ends. The algebra $\Psi_{\operatorname{inv}}^{\infty}(M)$ (the $\operatorname{inv}$-calculus, Definition 5.1) is not stable under inversion in the sense that if $T\in\Psi_{\operatorname{inv}}^{\infty}(M)$ is an invertible operator on the space $L^{2}(M)$, then its inverse does not necessarily belong to $\Psi_{\operatorname{inv}}^{\infty}(M)$. In this section we introduce the space $\Psi_{\operatorname{ess}}^{m}(M;E,F)$ of essentially translation invariant (classical) pseudodifferential operators $P:\mathcal{C}^{\infty}_{c}(M;E)\to\mathcal{C}^{\infty}(M;F)$, such that $\Psi_{\operatorname{inv}}^{\infty}(M)\subset\Psi_{\operatorname{ess}}^{\infty}(M)$ and the latter is stable under inversion. The operators in $\Psi_{\operatorname{ess}}^{\infty}(M)$ are called essentially translation invariant [67] (in [98], they were called almost invariant). The presentation in [63] is very detailed and complete and we refer the reader to that book for more details.
We shall continue to use the cut-off function $\eta:M\to[0,1]$ of Notation 5.14. (Recall that $\eta\in\mathcal{C}^{\infty}(M)$ has support in $M^{\prime}\times(-\infty,-1]$ and $\eta=1$ on $M^{\prime}\times(-\infty,-2]$.) The function $\eta$ allowed us to associate to any $T\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E):=\Psi_{\operatorname{inv}}^{m}(M^{\prime}\times\mathbb{R};E)^{\mathbb{R}}$ the operator
$s_{0}(T)\mathrel{\,:=\,}\eta T\eta\in\Psi_{\operatorname{inv}}^{m}(M;E)\,$,
which is well-defined (see Notation 5.14 and Proposition 5.15(i)).
Definition 5.17.
Let $E\to M$ be a compatible Hermitian vector bundle over the manifold with straight cylindrical ends $M$.
  1. (i)
    Let $\Psi_{\operatorname{sus}}^{-\infty}(M;E)$ consist of the linear operators $T\in\Psi^{-\infty}(M\times\mathbb{R};E)^{\mathbb{R}}$ with convolution kernel $C_{T}\in\mathcal{S}(M\times M\times\mathbb{R};E\boxtimes E)$.
  2. (ii)
    We define $\Psi_{\mathcal{S}}^{-\infty}(M;E)$ to be the set of pseudodifferential operators $P\in\Psi^{-\infty}(M;E)$ with distribution kernel $k_{P}\in\mathcal{S}(M\times M;E\boxtimes E)$.
  3. (iii)
    Let $\Psi_{\operatorname{sus}}^{-\infty}(M^{\prime};E)\subset\Psi^{-\infty}(M^{\prime}\times\mathbb{R};E)^{\mathbb{R}}$. We then define the space $\Psi_{\operatorname{ess}}^{-\infty}(M;E)$ of essentially translation invariant, order -$\infty$ pseudodifferential operators on $M$ by
    $\Psi_{\operatorname{ess}}^{-\infty}(M;E)\mathrel{\,:=\,}s_{0}(\Psi_{\operatorname{sus}}^{-\infty}(M^{\prime};E))+\Psi_{\mathcal{S}}^{-\infty}(M;E)\,.$
  4. (iv)
    Finally, for $m\in\mathbb{R}$, we define the space $\Psi_{\operatorname{ess}}^{m}(M;E)$ of essentially translation invariant, order $m$ pseudodifferential operators on $M$ by
    $\Psi_{\operatorname{ess}}^{m}(M;E)\mathrel{\,:=\,}\Psi_{\operatorname{inv}}^{m}(M;E)+\Psi_{\operatorname{ess}}^{-\infty}(M;E)\,.$
It then follows easily from the last definition that $\Psi_{\operatorname{sus}}^{-\infty}(M;E)=\Psi_{\operatorname{ess}}^{-\infty}(M\times\mathbb{R};E)^{\mathbb{R}}$.
Remark 5.18.
The last two definitions provide the following simple inclusions. First,
$\Psi_{\operatorname{comp}}^{-\infty}(M;E)\subset\Psi_{\mathcal{S}}^{-\infty}(M;E)\,$,
because the space of Schwartz sections contains the space of compactly supported sections. Let again $s_{0}$ be as in 5.14. This inclusion implies that
$\Psi_{\operatorname{inv}}^{-\infty}(M;E)\,=\,s_{0}(\psi_{\operatorname{sus}}^{-\infty}(M^{\prime};E))+\Psi_{\operatorname{comp}}^{-\infty}(M;E)\\ \subset s_{0}(\Psi_{\operatorname{sus}}^{-\infty}(M^{\prime};E))+\Psi_{\mathcal{S}}^{-\infty}(M;E)\,=:\,\Psi_{\operatorname{ess}}^{-\infty}(M;E)\,.$
Finally, for all $m\in\mathbb{Z}\cup\{\pm\infty\}$, we also obtain the inclusion
$\Psi_{\operatorname{inv}}^{m}(M;E)\subset\Psi_{\operatorname{inv}}^{m}(M;E)+\Psi_{\operatorname{ess}}^{-\infty}(M;E)\,=:\,\Psi_{\operatorname{ess}}^{m}(M;E)\,.$
5.8. Properties of essentially translation invariant operators
We continue to assume that $M$ is a manifold with straight cylindrical ends and that $E\to M$ is a compatible vector bundle. The next theorem gives the main properties of the calculus
$\Psi_{\operatorname{ess}}^{m}(M;E)\mathrel{\,:=\,}\Psi_{\operatorname{inv}}^{m}(M;E)+\Psi_{\operatorname{ess}}^{-\infty}(M;E)$
introduced in Definition 5.17 (see [63], [67] for the proof).
Theorem 5.19.
Let $m,m^{\prime},s,s^{\prime}\in\mathbb{R}$, let $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ be a manifold with straight cylindrical ends, and let $E\to M$ a compatible vector bundle, as before. Then we have the following properties.
  1. (i)
    $\Psi_{\operatorname{ess}}^{m}(M;E)\Psi_{\operatorname{ess}}^{m^{\prime}}(M;E)\subset\Psi_{\operatorname{ess}}^{m+m^{\prime}}(M;E)$ and $\Psi_{\operatorname{ess}}^{m}(M;E)^{*}\subset\Psi_{\operatorname{ess}}^{m}(M;E)$.
  2. (ii)
    If $P\in\Psi_{\operatorname{ess}}^{m}(M;E)$, then $P:H^{s}(M;E)\to H^{s-m}(M;E)$ is well-defined and continuous.
  3. (iii)
    The morphism $\sigma_{m}:\Psi_{\operatorname{ess}}^{m}(M;E)/\Psi_{\operatorname{ess}}^{m-1}(M;E)\to\mathcal{C}^{\infty}_{\operatorname{inv}}(S^{*}M;\operatorname{End}(E))$ defined by the principal symbol is an isomorphism.
  4. (iv)
    Let $P\in\Psi_{\operatorname{ess}}^{m}(M;E)$. Then $P$ is elliptic if, and only if, there exists $Q\in\Psi_{\operatorname{inv}}^{-m}(M;E)$ such that $PQ-1,QP-1\in\Psi_{\operatorname{ess}}^{-\infty}(M;E)$.
5.9. Translation invariance, the Fourier transform, and indicial operators
The translation invariance of the limit operators allows us to consider the one-dimensional Fourier transform to study them. To do that, in this subsection, we shall restrict ourselves to a cylinder. Let $\mathfrak{S}$ be a generic closed manifold (i.e., a smooth, compact, boundaryless manifold). Thus, in this subsection, we will work on $\mathfrak{S}\times\mathbb{R}$ and $E,F$ will be vector bundles on $\mathfrak{S}\times\mathbb{R}$ that are lifts of bundles from $\mathfrak{S}$. We then consider the one dimensional Fourier transform
(5.4) $\mathcal{F}:L^{2}\big(\mathfrak{S}\times\mathbb{R};E\big)\to L^{2}\big(\mathfrak{S}\times\mathbb{R};E\big),\quad\mathcal{F}(f)(y,\tau)\mathrel{\,:=\,}\int_{\mathbb{R}}e^{-\imath\tau x}f(y,x)dx\,$,
with $\imath^{2}=-1$ (see also Equation (3.3)). We no longer assume that $F=E.$ We next introduce the space $\Psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F)$ in general.
Definition 5.20.
We let $\Psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F):=\Psi_{\operatorname{ess}}^{m}(\mathfrak{S}\times\mathbb{R};E,F)^{\mathbb{R}}$.
The definition of $s_{0}$ in Notation 5.14 then extends to $T\in\Psi_{\operatorname{sus}}^{m}(M^{\prime};E)$ and we obtain $s_{0}(\Psi_{\operatorname{sus}}^{m}(M^{\prime};E))\subset\Psi^{m}(M;E)$. We obtain the following remark.
Remark 5.21.
Let $m\in\mathbb{R}$ and $Q\in\Psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F):=\Psi_{\operatorname{ess}}^{m}(\mathfrak{S}\times\mathbb{R};E,F)^{\mathbb{R}}$. Let $u_{0}\in\mathcal{C}^{\infty}(\mathfrak{S};E)$ and $\tau\in\mathbb{R}$ and set $u(x^{\prime},t):=e^{\imath t\tau}u_{0}(x^{\prime})$. Recall that $\Phi_{s}$ denotes the translation by $s$, see Equation (5.1). Then
$\Phi_{s}(u)(x^{\prime},t)\mathrel{\,:=\,}u(x,t+s)\,=\,e^{\imath(t+s)\tau}u_{0}(x^{\prime})\,=\,e^{\imath s\tau}u(x^{\prime},t)\,.$
Thus, $\Phi_{s}(u)=e^{\imath\tau s}u$. If $Q\in\psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F)$, then $Qu\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R};F)$ is defined, because $Q$ is properly supported. In general, $Qu\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R};F)$ is defined by continuity. This gives that the function $v(x^{\prime},t):=e^{-\imath t\tau}[Qu](x^{\prime},t)$ satisfies
$\Phi_{s}(v)](x^{\prime},t)\mathrel{\,:=\,}v(x^{\prime},t+s)\mathrel{\,:=\,}e^{-\imath(t+s)\tau}[Qu](x^{\prime},t+s)\,=\,e^{-\imath(t+s)\tau}\Phi_{s}[Qu](x^{\prime},t)\\ \,=\,e^{-\imath(t+s)\tau}[Q\Phi_{s}(u)](x^{\prime},t)\,=\,e^{-\imath(t+s)\tau}e^{\imath s\tau}[Q(u)](x^{\prime},t)\,=\,v(x^{\prime},t)\,$,
because $\Phi_{s}$ and $Q$ commute. That is $\Phi_{s}(v)=v$, and hence $v$ depends only on $x^{\prime}$ and can be identified with a function on $M^{\prime}$ denoted $\widehat{Q}(\tau)u_{0}$.
This allows us to extend the Fourier transform to operators as follows.
Definition 5.22.
Let $m\in\mathbb{R}$, let $Q\in\Psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F):=\Psi_{\operatorname{ess}}^{m}(\mathfrak{S}\times\mathbb{R};E,F)^{\mathbb{R}}$, and let $u_{0}\in\mathcal{C}^{\infty}(\mathfrak{S};E)$. We write $u:=e^{\imath t\tau}u_{0}$ for the section $u\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R};E)$ satisfying $u(x^{\prime},t):=e^{\imath t\tau}u_{0}(x^{\prime})$. In view of Remark 5.21, we can define $\widehat{Q}(\tau)u_{0}\in\mathcal{C}^{\infty}(\mathfrak{S};F)$ by the formula
$\widehat{Q}(\tau)u_{0}](x^{\prime})\,=\,e^{-\imath t\tau}[Q(u)](x^{\prime},t)\mathrel{\,:=\,}e^{-\imath t\tau}[Q(e^{\imath t\tau}u_{0})](x^{\prime},t)\,.$
It will be convenient to write the formula in Definition 5.22 as
(5.5) $Q(e^{\imath t\tau}u_{0})\,=\,e^{\imath t\tau}[\widehat{Q}(\tau)u_{0}]\,$,
where $u_{0}\in\mathcal{C}^{\infty}(\mathfrak{S};E)$, $Q\in\Psi_{\operatorname{ess}}^{m}(\mathfrak{S};E,F)$, and $\widehat{Q}(\tau)\in\Psi^{m}(\mathfrak{S};E,F)$. This allows us to obtain the following proposition whose proof can be found in [63] and [67]. Recall that, when defined, $k_{P}$ denotes the distribution kernel of an operator $P$ and that $C_{P}$ denotes its convolution kernel. In particular, if $Q$ is a multiplication operator only with respect to $x^{\prime}$, that is, if $Qu(x^{\prime},t)=a(x^{\prime})u(x^{\prime},t)$, then Definition 5.22 gives
(5.6) $\widehat{Q}(\tau)u_{0}](x^{\prime})\,=\,e^{-\imath t\tau}[Q(u)](x^{\prime},t)\mathrel{\,:=\,}e^{-\imath t\tau}[e^{\imath t\tau}au_{0}](x^{\prime},t)\,=\,a(x^{\prime})u_{0}(x^{\prime})\,$,
and, hence, $\widehat{Q}(\tau)=a$ (for this particular choice of $Q$ given as a multiplication operator with $a$, which is a function of $x^{\prime}$ alone).
Proposition 5.23.
Let $Q\in\Psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F):=\Psi_{\operatorname{ess}}^{m}(\mathfrak{S}\times\mathbb{R};E,F)^{\mathbb{R}}$, $P\in\Psi_{\operatorname{ess}}^{m^{\prime}}(\mathfrak{S}\times\mathbb{R};F,G)$. We then have the following properties:
  1. (i)
    $\widehat{PQ}(\tau)\,=\,\widehat{P}(\tau)\widehat{Q}(\tau)$, for all $\tau\in\mathbb{R}$.
  2. (ii)
    The distribution kernel $k_{\widehat{P}(\tau)}\in\mathcal{C}^{\infty}(\mathfrak{S}^{2};G\boxtimes F)$ of $\widehat{P}(\tau)$ and the convolution kernel $C_{P}\in\mathcal{C}^{\infty}(\mathfrak{S}^{2}\times\mathbb{R};G\boxtimes F)$ of $P$ are related by the formula $k_{\widehat{P}(\tau)}(x^{\prime},y^{\prime})=\mathcal{F}C_{P}(x^{\prime},y^{\prime},\tau),$ where $x^{\prime},y^{\prime}\in\mathfrak{S}$.
  3. (iii)
    If $u\in\mathcal{C}^{\infty}_{c}(\mathfrak{S}\times\mathbb{R};F)$, then $\widehat{Pu}(\tau)\,=\,\widehat{P}(\tau)\hat{u}(\tau)$.
The operators $\widehat{Q}(\tau)$ are useful for checking the invertibility of $Q\in\Psi_{\operatorname{sus}}^{m}(M;E,F)$, as follows from the following well-known result [84, 93, 123].
Theorem 5.24.
Let $\mathfrak{S}$ be a smooth, closed manifold and $Q\in\Psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F)$. The operator $Q:H^{s}(\mathfrak{S}\times\mathbb{R};E)\to H^{s-m}(\mathfrak{S}\times\mathbb{R};F)$ is invertible if, and only if, all the operators $\widehat{Q}(\tau):H^{s}(\mathfrak{S};E)\to H^{s-m}(\mathfrak{S};F)$, $\tau\in{\mathbb{R}}$, are invertible.
We let $t$ denote the second coordinate function on the half-infinite cylinder $M^{\prime}\times(-\infty,0)$ and we extend it to a smooth function (denoted also by $t$) on $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ such that $t\geq 0$ on $M_{0}$ (recall that $M^{\prime}:=\partial M_{0}$). This allows us to write $M^{\prime}\times(-\infty,0)=\{t\in\mathbb{R}:t<0\}$. Let $\Phi_{s}(x,t)=(x,t-s)$ be as in Equation (5.1). Then we obtain the following norm convergence result [63], [67].
Lemma 5.25.
If $P\in\Psi_{\operatorname{ess}}^{m}(M;E,F)$ and $u\in\mathcal{C}^{\infty}_{c}(M^{\prime}\times\mathbb{R};E)$, then, for $s$ large enough, $P\Phi_{s}(u)$ is defined and we have the following norm convergence in $L^{2}(M^{\prime}\times\mathbb{R};F)$,
${\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}(u)\mathrel{\,:=\,}\lim_{s\to\infty}\Phi_{-s}\big[P\Phi_{s}(u)|_{\{t<0\}}\big]\,.$
In addition, the resulting operator satisfies ${\mathcal{R}_{\infty}}(P)\in\Psi_{\operatorname{sus}}^{m}(M^{\prime};E)$ and ${\mathcal{R}_{\infty}}\circ s_{0}(P_{0})=P_{0}$ for $P_{0}\in\Psi_{\operatorname{sus}}^{m}(M^{\prime};E)$.
Using the Fourier transform for $Q={\mathcal{R}_{\infty}}(P)$ in Definition 5.22, we obtain Melrose’s indicial operators.
Definition 5.26.
Let $P\in\Psi_{\operatorname{ess}}^{m}(M;E)$ and $\widehat{P}(\tau):=\widehat{{\mathcal{R}_{\infty}}(P)}(\tau)$ be as in Remark 5.21. Then the family
$\{\widehat{P}(\tau)\}_{\tau\in\mathbb{R}}\mathrel{\,:=\,}\{\widehat{{\mathcal{R}_{\infty}}(P)}(\tau)\}_{\tau\in\mathbb{R}}\mathrel{\,:=\,}\{\mathcal{F}{\mathcal{R}_{\infty}}(P)\mathcal{F}^{-1}(\tau)\}_{\tau\in\mathbb{R}}$
is called the indicial family of $P$. The operator $\widehat{P}(\tau)$ is called the indicial operator of $P$.
As for the limit operators, the indicial operators are multiplicative [63, 67, 68].
Theorem 5.27.
Let $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ be a manifold with cylindrical ends and $E,F,G\to M$ be compatible Hermitian vector bundles. Let $m,m^{\prime}\in\mathbb{R}$.
  1. (i)
    If $P\in\Psi_{\operatorname{ess}}^{m}(M;E,F)$ and $Q\in\Psi_{\operatorname{ess}}^{m^{\prime}}(M;F,G)$, then ${\mathcal{R}_{\infty}}(PQ)\,=\,{\mathcal{R}_{\infty}}(P){\mathcal{R}_{\infty}}(Q)$.
  2. (ii)
    For $P$ and $Q$ as in (i), their indicial families (Definition 5.26) satisfy $\widehat{QP}(\tau)\,=\,\widehat{Q}(\tau)\widehat{P}(\tau)$, $\tau\in\mathbb{R}\,.$
  3. (iii)
    We have ${\mathcal{R}_{\infty}}\circ s_{0}=id$ on $\Psi_{\operatorname{sus}}^{m}(M^{\prime};E)$ and the map ${\mathcal{R}_{\infty}}:\Psi_{\operatorname{ess}}^{m}(M;E,F)\to\Psi_{\operatorname{sus}}^{m}(M^{\prime};E)$ is surjective with kernel $\Psi_{\mathcal{S}}^{-\infty}(M;E,F)+\Psi_{\operatorname{comp}}^{m}(M;E,F)$.
  4. (iv)
    We have the following exact sequence of algebras
    $0\longrightarrow\Psi_{\mathcal{S}}^{-\infty}(M;E)\longrightarrow\Psi_{\operatorname{ess}}^{-\infty}(M;E)\stackrel{{\scriptstyle{\mathcal{R}_{\infty}}}}{{-\!\!\!\longrightarrow}}\Psi_{\operatorname{sus}}^{-\infty}(M^{\prime};E)\longrightarrow 0\,.$
  5. (v)
    For any $T\in\Psi_{\operatorname{inv}}^{m}(M;E,F)$, we have $\sigma_{m}({\mathcal{R}_{\infty}}(T))\,=\,{\mathcal{R}_{\infty}}(\sigma_{m}(T))\,.$
The indicial operator of a matrix of operators is defined componentwise.
Remark 5.28.
Let us also notice that, given a translation invariant operator
$T\in\Psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F)\mathrel{\,:=\,}\Psi_{\operatorname{ess}}^{m}(\mathfrak{S}\times\mathbb{R};E,F)^{\mathbb{R}}\subset\Psi_{\operatorname{ess}}^{m}(\mathfrak{S}\times\mathbb{R};E,F)\,$,
we can define its indicial operators $\widehat{T}(\tau)$ in two ways, first as an operator in $\Psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F)$ and then as an operator in $\Psi_{\operatorname{ess}}^{m}(\mathfrak{S}\times\mathbb{R};E,F)$. The definition in the second case gives us “twice” each of the operators $\widehat{T}(\tau)$ defined in the first case, so the corresponding families will have the same properties. Unless stated otherwise, our definition of $\widehat{T}(\tau)$ will be given by regarding the operator $T$ as an operator in $\Psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F)$ (i.e., as an invariant operator, which gives us “fewer” indicial operators $\widehat{T}(\tau)$).
6. Normal lateral limits on manifolds with straight cylindrical ends
In this section, we adapt the results from Section 4 on normal lateral limits on half-spaces to smooth, open domains in manifolds with cylindrical ends. We thus continue to assume that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a smooth Riemannian manifold with straight cylindrical ends (Definition 3.11) and that $E,F\to M$ are two compatible vector bundle (see Definition 3.13). Most importantly, we let $\Omega\subset M$ be an open subset with smooth boundary $\Gamma:=\partial\Omega\neq\emptyset$ (so the assumption is that $\Gamma$ is also a smooth manifold) that is compatible with the structure of cylindrical ends of $M$, as in the introduction, Equation (1.2). That is:
$\Omega\cap\big[M^{\prime}\times(-\infty,R_{\Omega}]\big]\,=\,\Omega^{\prime}\times(-\infty,R_{\Omega}]\,$,
for an open subset $\Omega^{\prime}\subset M^{\prime}$. Note that $\Gamma$ is also a smooth manifold with straight cylindrical ends if we write
(6.1) $\Gamma\cap\big[M^{\prime}\times(-\infty,R_{\Omega}]\big]\,=\,\Gamma^{\prime}\times(-\infty,R_{\Omega}]\,$,
where $\Gamma^{\prime}:=\partial\Omega^{\prime}$. Recall that in this paper we assume that $\Omega$ is on one side of its boundary, see Assumption 2.7 (in particular, $\Gamma$ is also the boundary of $\Omega_{-}:=M\smallsetminus\overline{\Omega}$). It is no loss of generality to assume that $\Omega$ is connected; by contrast, we cannot assume that $\Omega^{\prime}$ is connected. However, $\Omega^{\prime}$ also must be on one side of its boundary. See Figure 2:
Figure 2. A manifold with boundary and straight cylindrical ends and a boundaryless manifold with straight cylindrical ends containing it as in Equation (1.2).
6.1. Normal tubular neighborhoods
Recall that the (global) spaces $H^{s}(M)$ are defined using the metric (see Definition 3.17 and Subsection 3.4 in general). Their variants with values in smooth vector bundles are defined similarly. We shall denote by $\cdot$ the inner product on the compatible vector bundle $E$. Distributions of the form $h\otimes\delta_{\Gamma}$ (see Equation (2.9)), that is,
$\langle h\otimes\delta_{\Gamma},\phi\rangle\mathrel{\,:=\,}\int_{\Gamma}h(x)\cdot\phi(x)\,dS_{\Gamma}(x)\,$,
will play an important role in what follows. We have the following analog of Lemma 3.8.
Lemma 6.1.
Let $E$ be a compatible, smooth Hermitian vector bundle on $M$ with inner product denoted $\cdot$, let $h\otimes\delta_{\Gamma}$ be as in Equation (2.9), and let $s\in\mathbb{R}$. Let $s^{\prime}\in\mathbb{R}$ be such that
$\begin{cases}\ s^{\prime}\mathrel{\,:=\,}s-1/2&\mbox{ if }s<0\\ \ s^{\prime}<-1/2&\mbox{ arbitrary, if }\ s\geq 0\,.\end{cases}$
Then, for all $h\in H^{s}(\Gamma;E)$, we have that $h\otimes\delta_{\Gamma}\in H^{s^{\prime}}(M;E)$.
Proof.
The proof is almost identically to that of Lemma 3.8, except that we use that the maps $H^{r}(M;E)\to H^{r-1/2}(\Gamma;E)$, $r>1/2$ are continuous [50, 63]. ∎
Let $P$ be an order $m$ pseudodifferential operator acting on the sections of $E$, that is $P\in\Psi^{m}(M;E)$. Recall that we are assuming that $\Omega$ is on one side of its boundary. We are interested in studying
(6.2) $\mathcal{S}_{P}h\mathrel{\,:=\,}P(h\otimes\delta_{\Gamma})\,$,
provided that the latter is defined. For the convenience of the notation, we also let
[…]
where the last relation is equivalent to saying that $\Omega$ is on one side of its boundary. We are especially interested in the following two restrictions and their traces
[…]
Notation 6.2.
We let $\boldsymbol{\nu}$ be the outer unit normal vector to $\Gamma:=\partial\Omega$. We extend this vector field to a global (smooth) vector field on $M$ (not necessarily unit everywhere), still denoted $\boldsymbol{\nu}$. Also, let $\sharp:TM\to T^{*}M$ be the isomorphism defined by the metric of $M$, as before. We shall write $\boldsymbol{\nu}^{\sharp}:=\sharp\boldsymbol{\nu}$.
If $v\in TM$ is a tangent vector to $M$ in $x$, we let $\exp(tv)$ denote the image of $tv$ under the exponential map, which is defined for $|t|$ small (depending on $v$).
Definition 6.3.
If $\epsilon>0$ is such that the normal exponential map
$\exp^{\perp}:\Gamma\times(-\epsilon,\epsilon)\ni(x,t)\,\mapsto\,\exp(t\boldsymbol{\nu}(x))\in M$
($\Gamma=\partial\Omega$) is well defined and is a diffeomorphism onto its image, then we shall say that $\Gamma$ has an $\epsilon$-normal tubular neighborhood.
It is well-known that if $\Gamma$ is compact or if $M$ has straight cylindrical ends and $\Omega$ is compatible with this structure, then $\Gamma$ will have an $\epsilon$-normal tubular neighborhood, for some $\epsilon>0$ small enough, see [107, Corollary 5.5.3]. The curves $t\mapsto\exp(t\boldsymbol{\nu}(x))$, $x\in\partial\Omega$, will be called the normal geodesics to $\partial\Omega$. If $u$ is a section of $E$ over $M$, $\Gamma$ has an $\epsilon$-normal tubular neighborhood, and $t\in(-\epsilon,\epsilon)$, we let
(6.5) $u_{t}\mathrel{\,:=\,}u|_{\exp^{\perp}(\Gamma\times\{t\})}\in\mathcal{C}^{\infty}(\Gamma\times\{t\};E)\simeq\mathcal{C}^{\infty}(\Gamma;E)\,$,
where the last isomorphism is obtained via parallel transport along the normal geodesics $(-\epsilon,\epsilon)\ni t\to\exp(t\boldsymbol{\nu}(x))\in M$, $x\in\partial\Omega$. (These isomorphisms are the analogues of the ones induced by $\tau_{\epsilon}(x):=x-x_{n}e_{n}$, which was introduced after Equation (3.18).)
It will be important for us to study the limits $u_{\pm}:=\lim_{t\to\pm 0}u_{t}$ in some function space on $\Gamma:=\partial\Omega$, for suitable $u$. When they exist, we call these limits, the normal lateral limits of $u$. In case $u$ is smooth enough on $\Omega_{+}:=\Omega$ and on $\Omega_{-}:=M\smallsetminus\overline{\Omega}$, then $u_{+}$ is the trace of $u|_{\Omega_{+}}:=u|_{\Omega}$ at the boundary and, similarly, $u_{-}$ is the trace of $u|_{\Omega_{-}}:=u|_{M\smallsetminus\overline{\Omega}}$ at the boundary, see Lemma 3.10.
6.2. Normal lateral limits for manifolds with cylindrical ends
We now turn to the study of normal lateral limits of pseudodifferential operators on general Riemannian manifolds with straight cylindrical ends. Note that, unlike in the other chapters, the case $M$ compact is not trivial, but it can be handled exactly as the case of manifolds with cylindrical ends. (This has been done in [69], so we will not review it in detail here.)
As usual, the case of pseudodifferential operators of order m<1m<-1 is easier. We begin with the case of operators with compactly supported distribution kernels in $M\times M$, which will then be used to deal with the general case. Let $F\to M$ be a second hermitian vector bundle (in addition to $E$). We have the following simple calculation that will be used repeatedly, so we formulate it as a lemma.
Lemma 6.4.
Let $\Omega$, $\Gamma:=\partial\Omega$, and $M$ have compatible straight cylindrical ends, see Equations (1.2) and (6.1), with $\Omega$ on one side of its boundary $\Gamma$, as before. Also, we assume that $E,F\to M$ are two compatible hermitian vector bundles, again, as before. We also let
  1. (i)
    $P\in\Psi_{\operatorname{ess}}^{m}(M;E,F)$, m<1m<-1;
  2. (ii)
    $s^{\prime}\in(1/2,-m-1/2)\not$$\emptyset$; and
  3. (iii)
    $h\in L^{2}(\Gamma;E)$;
then $\mathcal{S}_{P}h:=P(h\otimes\delta_{\Gamma})\in H^{s^{\prime}}(M;F)$, and hence,
$\mathcal{S}_{P}h]_{+}\,=\,[\mathcal{S}_{P}h]_{-}\,=\,[\mathcal{S}_{P}h]|_{\Gamma}\in H^{s^{\prime}-1/2}(\Gamma;F)\,.$
In particular, the trace (or restriction) $\mathcal{S}_{P}h|_{\Gamma}$ of $\mathcal{S}_{P}h:=P(h\otimes\delta_{\Gamma})$ at $\Gamma$ is defined and it coincides with the traces $\mathcal{S}_{P}h]_{\pm}$ associated to the domains $\Omega=:\Omega_{+}$ and $\Omega_{-}:=M\smallsetminus\overline{\Omega}$ with boundary $\Gamma$.
Proof.
The condition m1/2>1/2-m-1/2>1/2 shows that the set (1/2,m1/2)(1/2,-m-1/2) is non-empty, so we can choose $s^{\prime}\in(1/2,-m-1/2)$. Because $s^{\prime}+m<-1/2$, Lemma 6.1(i) shows that $h\otimes\delta_{\Gamma}\in H^{s^{\prime}+m}(M;E)$, and therefore $\mathcal{S}_{P}h:=P(h\otimes\delta_{\Gamma})\in H^{s^{\prime}}(M;E)$. Since $s^{\prime}>1/2$, the trace $\mathcal{S}_{P}h\in H^{s^{\prime}-1/2}(\Gamma;E)$ is defined and it coincides with the traces from the two open subsets with boundary $\Gamma$, by Lemma 3.10. ∎
Because the trace of $\mathcal{S}_{P}h:=P(h\otimes\delta_{\Gamma})$ at $\Gamma$ is defined and it coincide with the traces associated to the domains $\Omega_{\pm}$ with boundary $\Gamma$, we shall concentrate on the restriction (or trace) $\mathcal{S}_{P}h|_{\Gamma}$ of $\mathcal{S}_{P}h$ to $\Gamma$. The behavior of this restriction is the content of the following theorem.
For the rest of this section, we keep the assumptions of Lemma 6.4, that is, we assume that $\Omega$, $\Gamma:=\partial\Omega$, and $M$ have compatible straight cylindrical ends (see Equations (1.2) and (6.1)) with $\Omega$ on one side of its boundary $\Gamma$ and that let $E,F\to M$ are two compatible hermitian vector bundles.
Theorem 6.5.
We keep the assumptions of Lemma 6.4, as explained above. Let m<1m<-1 and $s^{\prime}\in(1/2,-m-1/2)$. Then, for $|\epsilon|$ small and any $P\in\Psi_{\operatorname{ess}}^{m}(M;E,F)$, there exists a unique $P_{\epsilon}\in\Psi_{\operatorname{ess}}^{m+1}(\Gamma;E,F)$ with the following properties:
  1. (i)
    For any $h\in L^{2}(\Gamma;E)$, $\mathcal{S}_{P}h:=P(h\otimes\delta_{\Gamma})\in H^{s^{\prime}}(M;F)$, and hence the traces of $\mathcal{S}_{P}h:=P(h\otimes\delta_{\Gamma})$ at the two sides of $\Gamma$ are defined and they satisfy
    $\mathcal{S}_{P}h]_{+}\,=\,[\mathcal{S}_{P}h]_{-}\,=\,[\mathcal{S}_{P}h]|_{\Gamma}\,=\,P_{0}h\in H^{s^{\prime}-1/2}(\Gamma;F)\,.$
  2. (ii)
    For any $x\in\Gamma:=\partial\Omega$ and $\xi^{\prime}\in T_{x}^{*}\Gamma$, let $\xi\in T_{x}^{*}M$ be a lift of $\xi^{\prime}$. The principal symbol of $P_{0}$ is then given by
    $\sigma_{m+1}(P_{0};\xi^{\prime})\,=\,\frac{1}{2\pi}\int_{\mathbb{R}}\sigma_{m}(P;\xi+t\boldsymbol{\nu}^{\sharp}_{x})\,dt\,.$
  3. (iii)
    The distribution kernel of the operator $P_{0}$ satisfies $k_{P_{0}}(x^{\prime},y^{\prime})=k_{P}(x^{\prime},y^{\prime})$ for all $x^{\prime}\neq y^{\prime}$ in $\Gamma$, and hence $(\phi P\psi)_{0}=\phi P_{0}\psi$, for all $\phi,\psi\in C_{0}^{\infty}(M)$.
  4. (iv)
    Let $s\in\mathbb{R}$ and $h\in H^{s}(\Gamma;E)$, then $\mathcal{S}_{P}h]_{\epsilon}\mathrel{\,:=\,}\big[P(h\otimes\delta_{\Gamma})]_{\epsilon}\,=\,P_{\epsilon}h$ and
    $\mathcal{S}_{P}h]_{\pm}\mathrel{\,:=\,}\lim_{\epsilon\to\pm 0}\big[P(h\otimes\delta_{\Gamma})]_{\epsilon}\,=\,\lim_{\epsilon\to\pm 0}P_{\epsilon}h\,=\,P_{0}h\in H^{s-m-1}(\Gamma;F)\,$,
    where the limit is in the topology of $H^{s-m-1}(\Gamma;F)$.
The operator $P_{0}$ will be called the restriction at $\Gamma$ operator associated to $P$.
Proof.
Let us notice that the relations $\mathcal{S}_{P}h]_{+}\,=\,[\mathcal{S}_{P}h]_{-}\,=\,[\mathcal{S}_{P}h]|_{\Gamma}\in H^{s^{\prime}-1/2}(\Gamma;F)$ have already been proved (see Lemma 6.4). Also, the last equality in (iii) is an immediate consequence of the equality of kernels (because $k_{\phi P\psi}=\phi k_{P}\psi)$. Thus we do not have to prove these properties anymore, so we will ignore these questions from now on.
Let us assume that the distribution kernel $k_{P}$ of $P$ is compactly supported (in $M\times M$) and prove our theorem in this case. We may also assume that $E$ and $F$ are trivial, one dimensional. Since we have assumed that the support $\operatorname{supp}k_{P}\subset M\times M$ of the distribution kernel of $P$ is compactly supported, its two projections $K_{1}:=p_{1}\operatorname{supp}k_{P}\subset M$ and $K_{2}:=p_{2}\operatorname{supp}k_{P}\subset M$ are also compact. Hence $K:=K_{1}\cup K_{2}\cup\operatorname{supp}h$ is also compact.
For each $x\in\Gamma$, we choose local coordinates $y$ in a neighborhood $V_{x}$ of $x$ that straighten out the boundary to the hyperplane by mapping it to $\{x_{n}=0\}\subset\mathbb{R}^{n}$. We can choose these coordinates such that they map $\exp(t\boldsymbol{\nu})$ to $(y^{\prime},t)\in\mathbb{R}^{n-1}\times(-\epsilon,\epsilon)$. Let us cover $\Gamma\cap K$ with finitely many such neighborhoods $V_{j}:=V_{x_{j}}$, which is possible since $K$ is compact. Let us then choose a smooth partition of unity $\phi_{0},\phi_{1},\ldots,\phi_{N}$ on $M$ subordinated to $M\smallsetminus K,V_{1},\ldots,V_{N}$. We can assume that $\phi_{0}$ vanishes in a neighborhood of $\Gamma\cap K$. By refining the covering $\{V_{j}\}$ of $\Gamma$, we can assume that the support of each $\phi_{i}P\phi_{j}$, $1\leq i,j\leq N$, is completely contained in a set of the form $V_{x}$. Then we use Theorem 4.13 for each of the operators $\phi_{i}P\phi_{j}$ on the coordinate neighborhood $V_{x}$ to obtain the limit operator $P_{0ij}\in\Psi^{m+1}(\Gamma)$. We define $P_{0}:=\sum_{i,j=1}^{N}P_{0ij}$. Then, for each of these operators, we have
(6.6) $\begin{gathered}\phi_{i}\big[\mathcal{S}_{P}(\phi_{j}h)\big]|_{\Gamma}\,=\,P_{0ij}h\,,\\ \sigma_{m+1}(P_{0ij};\xi^{\prime})\,=\,\frac{\phi_{i}}{2\pi}\left(\int_{\mathbb{R}}\sigma_{m}(P;\xi+t\boldsymbol{\nu}^{\sharp}_{x})\,dt\,\right)\phi_{j}\,,\quad\mbox{and}\\ k_{P_{0ij}}(x^{\prime},y^{\prime})\,=\,k_{\phi_{i}P\phi_{j}}(x^{\prime},y^{\prime})\,=\,\phi_{i}(x^{\prime})k_{P}(x^{\prime},y^{\prime})\phi_{j}(y^{\prime})\,.\end{gathered}$
by Theorem 4.13. Adding up all the corresponding relations for $i,j=1,\ldots,N$, and noticing that $\sum_{i=1}^{N}\phi_{i}=1$ on $\Gamma\cap K$ (recall that $\phi_{0}$ vanishes in a neighborhood of $\Gamma\cap K$), we obtain (ii) and
$k_{P_{0}}(x^{\prime},y^{\prime})\mathrel{\,:=\,}\sum_{i,j=1}^{N}k_{P_{0ij}}(x^{\prime},y^{\prime})\,=\,\sum_{i,j=1}^{N}\phi_{i}(x^{\prime})k_{P}(x^{\prime},y^{\prime})\phi_{j}(y^{\prime})\,=\,k_{P}(x^{\prime},y^{\prime})$
for all $x^{\prime},y^{\prime}\in\Gamma$, $x^{\prime}\neq y^{\prime}$. We have thus proved also (iii). To complete (i), let $h\in L^{2}(\Gamma;E)$. Then $\mathcal{S}_{P}(\phi_{j}h):=P\big[(\phi_{j}h)\otimes\delta_{\Gamma}\big]\in H^{s^{\prime}}(M;E)$, by Lemma 6.4, because $s^{\prime}\in(1/2,-m-1/2)$. Therefore, $\mathcal{S}_{P}h\in H^{s^{\prime}}(M;E)$ as well, by linearity. This gives
$\big[\mathcal{S}_{P}h\big]|_{\Gamma}\,=\,\sum_{i,j=1}^{N}\phi_{i}\big[\mathcal{S}_{P}(\phi_{j}h)\big]|_{\Gamma}\,=\,\sum_{i,j=1}^{N}P_{0ij}h\,=:\,P_{0}h\,$,
where the second equality is from Equation (6.6) (a consequence of Theorem 4.13). This gives the last equality of (i) and hence completes the proof of (i) as well. (We have already noticed that the first two equalities in (i) are the standard properties of Sobolev spaces discussed in Lemma 6.4.)
We have thus proved the first three points of our theorem under the additional hypothesis that $k_{P}$ is compactly supported. The general case follows immediately from this one by using the results already proved for operators of the form $\phi P\psi$, where $\phi,\psi:M\to\mathbb{C}$ are smooth and compactly supported. (Operators of this form will have compactly supported distribution kernels.) Indeed, there is no loss of generality to assume that $h$ has compact support. Let $\psi\in\mathcal{C}^{\infty}_{c}(M)$ be equal to 1 on the support of $h$. Let $x\in\Gamma$ arbitrary and $U$ a relatively compact neighborhood of $x$ in $U$. Let $\phi\in\mathcal{C}^{\infty}_{c}(M)$ be equal to $1$ on $U$. We first define $P_{0}h|_{\Gamma\cap U}:=(\phi P\psi)_{0}h|_{\Gamma\cap U}$. This definition is independent of $\phi$ and $\psi$ by (iii) for compactly supported distribution kernels already proved. We then have
$\mathcal{S}_{P}h]|_{\Gamma\cap U}\,=\,[\mathcal{S}_{\phi P\psi}h]|_{\Gamma\cap U}\\ \,=\,(\phi P\psi)_{0}h|_{\Gamma\cap U}\,=:\,P_{0}h|_{\Gamma\cap U}\,.$
Since $x$ was arbitrary, we obtain that $\mathcal{S}_{P}h]|_{\Gamma}$ and $P_{0}h$ coincide in the neighborhood of every point, and hence they are equal.
To prove the last point, let us assume first that $m=$-$\infty$. Then the result follows from the theorem on derivation under the integral sign (Theorem 4.6). Since the result depends linearly on $P$, we may therefore assume then that $P\in\Psi_{\operatorname{inv}}^{m}(\Gamma;E,F)$ and hence that $P$ is properly supported. If this is the case and $h$ has compact support, then it follows from the point (ii) already proved that $P_{0}\in\Psi^{m+1}(\Gamma;E,F)$. The fact that $P_{0}\in\Psi_{\operatorname{inv}}^{m+1}(\Gamma;E,F)$ then follows from the form of the distribution kernels (our constructions commute with translations). The result for a general $h\in H^{s}(\Gamma;E)$ (not necessarily compactly supported) is obtained using the result already proved, the translation invariance at infinity and a dyadic partition of unity on the cylindrical ends of $M$, as in [63]. ∎
We shall need mapping properties of the layer potentials, for which our main reference is [54], where symbols of rational type are discussed in detail and where further references are given. Recall [54] that a symbol is of rational type if in every fiber it is a quotient of polynomial functions. In particular, a symbol of rational type is classical.
Theorem 6.6.
We keep the assumptions of Lemma 6.4. Let $P\in\Psi_{\operatorname{ess}}^{m}(M;E,F)$ with symbol of rational type (a quotient of polynomial functions). Then, for any $s\in\mathbb{R}$, there exists $C_{s}\geq 0$ such that
$\|\mathcal{S}_{P}h|_{\Omega_{\pm}}\|_{H^{s-m-\frac{1}{2}}(\Omega_{\pm};F)}\mathrel{\,:=\,}\|P(h\otimes\delta_{\Gamma})|_{\Omega_{\pm}}\|_{H^{s-m-\frac{1}{2}}(\Omega_{\pm};F)}\leq C_{s}\|h\|_{H^{s}(\Gamma;E)}\,.$
Proof.
If the distribution kernel $k_{P}$ of $P$ has compact support, the result follows from Theorem 9.4.7 on page 584 of [54]. If $P$ is properly supported, the result follows from this case and a translation invariant partition of unity on the cylindrical ends of $M$. We have $h\otimes\delta_{\Gamma}\in H^{s^{\prime}}(M;E)$. If $P\in\Psi_{\operatorname{ess}}^{-\infty}(M;E,F)$, the mapping properties of the operators in the essentially translation invariant calculus gives that $\mathcal{S}_{P}(h):=P(h\otimes\delta_{\Gamma})\in H^{\infty}(M;F)$, so the result is true in this case as well. Since every $P\in\Psi_{\operatorname{ess}}^{m}(M;E,F)$ is the sum of an operator in $\Psi_{\operatorname{inv}}^{m}(M;E,F)$ (which is properly supported) and an operator in $\Psi_{\operatorname{ess}}^{-\infty}(M;E,F)$, the result follows from the particular cases we have already proved. ∎
Recall from Notation 6.2 that $\boldsymbol{\nu}$ is a fixed vector field on $M$ that is the outer unit normal vector to $\Gamma:=\partial\Omega$. Also, recall that $\sharp:TM\to T^{*}M$ is the isomorphism defined by the metric. We now turn to the study of the normal lateral limits of operators of order m=1m=-1. We first need to establish some more notation.
Notation 6.7.
For $P\in\Psi_{\operatorname{ess}}^{-1}(M;E,F)$, we let $\sigma_{-1}(P)\in\mathcal{C}^{\infty}(T^{*}M\smallsetminus\{0\};\operatorname{Hom}(E;F))$ denote its principal symbol. We shall write $\sigma_{-1}(P;\xi)\in\operatorname{Hom}(E_{x},F_{x})$ for its value at $\xi\in T_{x}^{*}M\smallsetminus\{0\}$ and we assume that $\sigma_{-1}(P;\xi)$ is odd in $\xi\in T^{*}M$. We then let $\operatorname{\mathfrak{L}}(P;x)\in\operatorname{Hom}(E_{x},F_{x})$ be given by
$\operatorname{\mathfrak{L}}(P;x)\mathrel{\,:=\,}\sigma_{-1}(P;-\boldsymbol{\nu}_{x}^{\sharp})\,.$
We also let, for all $0\neq\xi^{\prime}\in T^{*}\Gamma$, $\xi\in T_{x}^{*}M$ be such that it projects onto $\xi^{\prime}$ and is orthogonal to $\boldsymbol{\nu}_{x}^{\sharp}$ and
$b_{0}(\xi^{\prime})\mathrel{\,:=\,}\frac{1}{4\pi}\,\int_{\mathbb{R}}\big[\sigma_{-1}(P;\xi+\tau\boldsymbol{\nu}_{x}^{\sharp})+\sigma_{-1}(P;\xi-\tau\boldsymbol{\nu}_{x}^{\sharp})\big]\,d\tau\in\operatorname{Hom}(E_{x},F_{x})\,.$
The form of the definition of $\operatorname{\mathfrak{L}}$ is due to the fact that $e_{n}=-\boldsymbol{\nu}$ in the Euclidean case (see Section 4) if $\Omega=\Omega_{+}=\mathbb{R}_{+}^{n}$. We let $\mathcal{S}_{P}h:=P(h\otimes\delta_{\Gamma})$, as usual.
Theorem 6.8.
Let $P\in\Psi_{\operatorname{ess}}^{-1}(M;E,F)$. We continue to assume the conditions of Lemma 6.4, as before. Let $\operatorname{\mathfrak{L}}(P)$ and $b_{0}$ be as in Notation 6.7 and assume that $\sigma_{-1}(P;\xi)$ is odd in $\xi\in T^{*}M$. Then, for $|\epsilon|$ small, there exist pseudodifferential operators $P_{\epsilon}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;E,F)$, such that, if we let $P_{0\pm}\mathrel{\,:=\,}\pm\frac{\imath}{2}\operatorname{\mathfrak{L}}(P)+P_{0}$, then we have the following properties.
  1. (i)
    $\operatorname{\mathfrak{L}}_{\pm}(P)\in C_{\operatorname{inv}}^{\infty}(M;\operatorname{Hom}(E;F))$.
  2. (ii)
    $P_{0\pm}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;E,F)$.
  3. (iii)
    $\sigma_{0}(P_{0})=b_{0}$.
  4. (iv)
    $k_{P_{0}}(x^{\prime},y^{\prime})=k_{P}(x^{\prime},y^{\prime})$ for all $x^{\prime}\neq y^{\prime}$ in $\Gamma$.
  5. (v)
    for all $s\in\mathbb{R}$ and all $h\in H^{s}(\Gamma;E)$, we have $\big[\mathcal{S}_{P}h\big]_{\epsilon}:=\big[P(h\otimes\delta_{\Gamma})\big]_{\epsilon}=P_{\epsilon}h$.
  6. (vi)
    $\big[\mathcal{S}_{P}h\big]_{\pm}\mathrel{\,:=\,}\lim_{\epsilon\to\pm 0}\big[P(h\otimes\delta_{\Gamma})\big]_{\epsilon}\,=\,\lim_{\epsilon\to\pm 0}P_{\epsilon}h\,=\,P_{0\pm}h\in H^{s}(\Gamma;F)\,.$
  7. (vii)
    Let us assume that $P$ is of rational type, that $s>0$, and that $h\in H^{s}(M;E)$. Then $\mathcal{S}_{P}(h)\in H^{s+1/2}(\Omega_{\pm})$, and hence the traces satisfy
    $\gamma_{\pm}(\mathcal{S}_{P}(h))\,=\,\mathcal{S}_{P}(h)_{\pm}\,=\,P_{0\pm}h\in H^{s}(\Gamma;F)\,.$
Proof.
We have $\operatorname{\mathfrak{L}}(P)\in C_{\operatorname{inv}}^{\infty}(M;\operatorname{Hom}(E;F))\subset\Psi_{\operatorname{inv}}^{0}(M;E,F)$ by the definitions of $\operatorname{\mathfrak{L}}_{\pm}(P)$ and the properties of the principal symbol. Except for the first and last points, the proof is to a large extent, word-for-word the same as the one of Theorem 6.5, whose notations we use here. We sketch the argument for the benefit of the reader.
We agaoin begin with the case when $P$ has compactly supported distribution kernel. Then, the first two relations in the crucial Equation (6.6) are replaced with
[…]
where $\xi\perp\boldsymbol{\nu}_{x}$. (The last relation of that equation does not change.)
Let us now turn to points (i) and (vii), which are slightly different from the proof of Theorem 6.5. First, the only thing that we need to add to complete the proofs of (i) is that $\operatorname{\mathfrak{L}}_{\pm}(P)=\sum_{ij=1}^{N}\operatorname{\mathfrak{L}}_{\pm}(\phi_{i}P\phi_{j})$. Finally, the last point follows from Lemma 3.10 and Theorem 6.6. ∎
If in the previous theorem $P$ is the Laplacian, $P=\Delta:=-d^{*}d$ or the classical Stokes operator $\boldsymbol{\Xi}_{0,0}$, then $b_{0}=0$ and hence $P_{0}$ is of order 1-1, but that is not true in general. For instance, as we will see below, it is not true if $P=\boldsymbol{\Xi}_{V,V_{0}}$ and $V_{0}$ does not vanish identically on $\Gamma=\partial\Omega$.
Corollary 6.9.
Let $P\in\Psi_{\operatorname{ess}}^{m}(M;E,F)$ be as in Theorem 6.5 (i.e. m<1m<-1) or as in Theorem 6.8 (i.e. m=1m=-1 and $P$ is classical). Then $(P^{*})_{0}=(P_{0})^{*}$ and, when $P$ is classical, $\operatorname{\mathfrak{L}}(P^{*})=\operatorname{\mathfrak{L}}(P)^{*}$.
Proof.
Let $x^{\prime},y^{\prime}\in\Gamma$, $x^{\prime}\neq y^{\prime}$. Theorems 6.5 and 6.8 give the second and the last of the following equalities:
$k_{P_{0}^{*}}(x^{\prime},y^{\prime})\,=\,k_{P_{0}}(y^{\prime},x^{\prime})^{*}\,=\,k_{P}(y^{\prime},x^{\prime})^{*}\,=\,k_{P^{*}}(x^{\prime},y^{\prime})\,=\,k_{(P^{*})_{0}}(x^{\prime},y^{\prime})\,.$
As we have already noticed, the operators $P_{0}^{*}$ and $(P^{*})_{0}$ are determined by the values of their distribution kernels outside the diagonal (see Remark 4.14 and Theorem 4.16(ii)). Since the corresponding distribution kernels of $(P^{*})_{0}$ and $(P_{0})^{*}$ coincide, we have $P_{0}^{*}=(P^{*})_{0}$, as claimed. The last relation of the Corollary follows from the sequence of equalities
$\operatorname{\mathfrak{L}}(P^{*})\,=\,\sigma_{-1}(P^{*};-\boldsymbol{\nu}^{\sharp})\,=\,\sigma_{-1}(P;-\boldsymbol{\nu}^{\sharp})^{*}\,=\,\operatorname{\mathfrak{L}}(P)^{*}\,.$
This completes the proof. ∎
Without ambiguity, we shall therefore write $P_{0}^{*}=(P^{*})_{0}=(P_{0})^{*}$ from now on.
As noticed in [69, 63], the results we have obtained so far in this section remain valid in the case $\Omega$ bounded. We conclude this section with a property that is specific to manifolds with cylindrical ends. Recall the indicial operator ${\mathcal{R}_{\infty}}(P)$ and the indicial family $\widehat{P}(\tau)$, obtained by taking the Fourier transform of ${\mathcal{R}_{\infty}}(P)$.
Theorem 6.10.
Let $P\in\Psi_{\operatorname{ess}}^{m}(M;E,F)$, $m\leq-1$. Then
${\mathcal{R}_{\infty}}(P)]_{0}\,=\,{\mathcal{R}_{\infty}}(P_{0})\quad\mbox{and}\quad(\widehat{P}(\tau))_{0}\,=\,\widehat{(P_{0})}(\tau)\,.$
If m=1m=-1, we also have ${\mathcal{R}_{\infty}}[\operatorname{\mathfrak{L}}_{\pm}(P)]\,=\,\operatorname{\mathfrak{L}}_{\pm}[{\mathcal{R}_{\infty}}(P)]$ and, consequently, $\widehat{[\operatorname{\mathfrak{L}}_{\pm}(P)]}(\tau)\,=\,\operatorname{\mathfrak{L}}_{\pm}[\widehat{P}(\tau)]$ and $(\widehat{P}(\tau))_{0\pm}\,=\,\widehat{(P_{0\pm})}(\tau)$.
Proof.
The first relation is follows from the fact that limits at infinity and restrictions to $\Gamma$ commute. Both of them commute also with translations, and hence with the Fourier transform. This gives the second relation. The last relation follows from the fact that $\sigma_{-1}$ and $\mathcal{R}_{\infty}$} commute, see Proposition 5.15(iv). ∎
6.3. Douglis-Nirenberg operators: jumps, ellipticity, and Fredholmness
The generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (Equation 1.3) is not elliptic in the usual sense, but is elliptic in Douglis-Nirenberg sense. As with the usual ellipticity property, the ellipticity in Douglis-Nirenberg sense (together with the invertibility of the limit operators) is enough to establish the Fredholm property of our generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$, but under different Sobolev spaces. This is the content of Theorem 6.15. The definition of DouglisNirenberg{Douglis-Nirenberg}-operators makes sense for general manifolds $M$, so, in the beginning of this subsection, we allow $M$ to be a general smooth manifold.
Definition 6.11.
Let $M$ be a smooth manifold, let $E_{i}\to M$ smooth vector bundles, $1\leq i\leq k$, and let $s_{i},t_{j}\in\mathbb{R}$, $1\leq i,j\leq k$. We set $\boldsymbol{s}:=(s_{1},\ldots,s_{k})$, $\boldsymbol{t}:=(t_{1},\ldots,t_{k})$, and $E:=E_{1}\oplus\ldots\oplus E_{k}$. Then
$\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E)\mathrel{\,:=\,}\{P=[P_{ij}]\mid\,P_{ij}\in\Psi^{s_{i}+t_{j}}(M;E_{j},E_{i}),\ 1\leq i,j\leq k\}\,.$
An operator $P=[P_{ij}]\in\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M;TM\oplus\mathbb{C})$ is said to be of Douglis-Nirenberg-order $\leq[\boldsymbol{s}+\boldsymbol{t}]$. For $P=[P_{ij}]\in\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M;TM\oplus\mathbb{C})$ let
${\operatorname{Symb}}_{\boldsymbol{s},\boldsymbol{t}}(P)\mathrel{\,:=\,}[\sigma_{s_{i}+t_{j}}(P_{ij})]$
be its $(\boldsymbol{s},\boldsymbol{t})$–principal symbol, which is a suitable section of the lift of the endomorphism bundle $\operatorname{End}(E)$ to $T^{*}M\smallsetminus 0$. The operator $P$ is said to be $(\boldsymbol{s},\boldsymbol{t})$–Douglis-Nirenberg elliptic if its $(\boldsymbol{s},\boldsymbol{t})$–principal symbol matrix ${\operatorname{Symb}}_{\boldsymbol{s},\boldsymbol{t}}(P)$ is invertible outside the zero section (i.e., on $T^{*}M\smallsetminus 0$).
The reason for introducing the spaces $\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E)$ is that our generalized Stokes operator fits into this framework.
Example 6.12.
Let $k=2$, let $\boldsymbol{s}=(s_{1},s_{2})=\boldsymbol{t}=(t_{1},t_{2})=(1,0)$, let ${E_{1}}=TM$, and let $E_{2}$$\mathbb{C}$. Then
$\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}\in\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M;TM\oplus\mathbb{C})\mathrel{\,:=\,}\{P=[P_{ij}]\mid\,P_{ij}\in\Psi^{s_{i}+t_{j}}(M;E_{j},E_{i})\}\,.$
More importantly, $\boldsymbol{\Xi}$ is not elliptic in the usual sense, but we will see in Proposition 6.13 that $\boldsymbol{\Xi}$ is $(\boldsymbol{s},\boldsymbol{t})$-Douglis-Nirenberg-elliptic.
Recall that $\sharp:T^{*}M\to TM$ denotes the isomorphism defined by the Riemannian metric $g$ of $M$ (it is sometimes called the “musical isomorphism”). For instance, using the notation introduced in Notation (2.3), if $0\neq\eta\in T_{x}M$, then $\frac{1}{\|\eta\|^{2}}\eta\otimes\eta^{\sharp}$ is the projection onto $\eta\neq 0$, regarded as a linear map $T_{x}M\to T_{x}M$.
Proposition 6.13.
Let $(\boldsymbol{s},\boldsymbol{t})=(1,0)$. Then the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (1.3) belongs to $\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M;TM\oplus\mathbb{C})$ and is $(\boldsymbol{s},\boldsymbol{t})$–Douglis-Nirenberg elliptic (Definition 6.11). The $(\boldsymbol{s},\boldsymbol{t})$–principal symbol of $\boldsymbol{\Xi}$ is
${\operatorname{Symb}}_{\boldsymbol{s},\boldsymbol{t}}\left(\boldsymbol{\Xi}\right)(\xi)=\left(\begin{array}[]{cc}|\xi|^{2}+\xi^{\sharp}\otimes\xi&\imath\xi^{\sharp}\\ -\imath\xi&-V_{0}\end{array}\right)\in\operatorname{End}(TM\oplus\mathbb{C})\,$,
which is invertible with inverse
$\left(\begin{array}[]{cc}\frac{1}{|\xi|^{2}}-\frac{V_{0}+1}{2V_{0}+1}\frac{1}{|\xi|^{4}}\xi^{\sharp}\otimes\xi&\frac{\imath}{(2V_{0}+1)|\xi|^{2}}\xi^{\sharp}\\ -\frac{\imath}{(2V_{0}+1)|\xi|^{2}}\xi&-\frac{2}{2V_{0}+1}\end{array}\right)\in\operatorname{End}(TM\oplus\mathbb{C})\,.$
Proof.
The given formula for principal symbol of $\boldsymbol{\Xi}$ is a consequence of [130, Chapter 5 §12, (12.20)] (but see also Corollary 2.6 for the principal symbol of $\operatorname{Def}^{*}\,\operatorname{Def}\,$ and Remark 2.8(iii) for the principal symbol of $\nabla$). The result then follows by verifying that the product of these two matrices is the identity. ∎
See also [68, Theorem 5.11 and Remark 5.12] and [36]. We notice that the formula for the inverse of the principal symbol was obtained using Schur’s complement formula.
We now return to the case when $M$ is a manifold with straight cylindrical ends (Definition 3.11) and the vector bundles are compatible (Definition 3.13). We begin by adapting Definition 6.11 to the setting cylindrical ends.
Remark 6.14.
Unless explicitly stated otherwise, we use the notation of the last definition, Definition 6.11. Let $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ be a manifold with straight cylindrical ends and $E_{i}\to M$ be compatible vector bundles. Recall the definitions of $\Psi_{\operatorname{ess}}^{m}(M;E)$ and $\Psi_{\operatorname{sus}}^{m}(M^{\prime};E)$ from Definitions 5.17 and 5.20. We now want to introduce the analogous spaces for Douglis-Nirenberg-operators. First, we let
$\Psi_{\operatorname{ess}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E)\mathrel{\,:=\,}\{P=[P_{ij}]\mid\,P_{ij}\in\Psi_{\operatorname{ess}}^{s_{i}+t_{j}}(M;E_{j},E_{i}),\ 1\leq i,j\leq k\}\,.$
Let also $m=\max\{s_{i}+t_{j}\}$. Then $\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E)\subset\Psi^{m}(M;E)$ and we have
$\Psi_{\operatorname{ess}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E)\,=\,\Psi_{\operatorname{ess}}^{m}(M;E)\cap\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E)\,.$
Second, by considering $M^{\prime}\times\mathbb{R}$ instead of $M$, we can define similarly,
$\Psi_{\operatorname{sus}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M^{\prime};E)\,=\,\Psi_{\operatorname{sus}}^{m}(M^{\prime};E)\cap\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M^{\prime}\times\mathbb{R};E)\\ \mathrel{\,:=\,}\Psi_{\operatorname{ess}}^{m}(M^{\prime}\times\mathbb{R};E)^{\mathbb{R}}\cap\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M^{\prime}\times\mathbb{R};E)\,.$
The limit and indicial operators ${\mathcal{R}_{\infty}}(P)$ and $\widehat{P}(\tau)$ of $P\in\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E)$ are defined by restriction from $\Psi_{\operatorname{ess}}^{m}(M;E)$. In particular, ${\mathcal{R}_{\infty}}(P)\in\Psi_{\operatorname{sus}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M^{\prime};TM\oplus\mathbb{C})$ (See Lemma 5.25) and their indicial operators (of Fourier transforms) $\widehat{P}(\tau)\in\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M^{\prime};TM\oplus\mathbb{C})$ (see Definition 5.26).
We will use the Douglis-Nirenberg-ellipticity of $\boldsymbol{\Xi}$ to establish (in the next subsection, under certain positivity assumptions on the potentials $V$ and $V_{0}$) its Fredholm property. To that end, we shall use the following characterization of the Fredholm property in the essentially translation invariant calculus. (This result was discussed and proved in great detail in [63, Theorem 15.4.19] and [67, Theorem 7.12], so we will not include a proof here. It builds on earlier results of many people, including Kondratiev, Melrose, and Schulze, see the above quoted works for references.) Recall the notation of the last remark, Remark 6.14. Then we have the following characterization of Fredholm property for a Douglis-Nirenberg-elliptic operator on a manifold with straight cylindrical ends (see [68] and [63]).
Theorem 6.15.
Let $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ be a manifold with straight cylindrical ends and $s_{i},t_{j}\in\mathbb{R}$, $1\leq i,j\leq k$, as before. Let $P=[P_{ij}]\in\Psi_{\operatorname{ess}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E)$ and $m\in\mathbb{R}$. Then the map
$P:H^{m+[\boldsymbol{t}]}(M;E)\mathrel{\,:=\,}\oplus_{j=1}^{k}H^{m+t_{1}}(M;E_{j})\to H^{m-[\boldsymbol{s}]}(M;E)\mathrel{\,:=\,}\oplus_{i=1}^{k}H^{m-s_{1}}(M;E_{i})$
is Fredholm if, and only if,
  1. (i)
    $P$ is Douglis-Nirenberg-elliptic and
  2. (ii)
    the limit operator ${\mathcal{R}_{\infty}}(P)\in\Psi_{\operatorname{sus}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M^{\prime};E)$ is invertible.
(If $M$ is closed, we replace $\Psi_{\operatorname{ess}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E)$ with $\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E)$ and we remove the condition (ii).)
The case $M$ closed is, of course, very classical, due to Seeley, see [63, 53, 54, 125, 138]. The results of the previous subsections of this section on normal lateral limites then extend immediately to the components $P_{ij}$ of the operators $P=[P_{ij}]\in\Psi^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E)$ that satisfy $s_{i}+t_{j}\leq-1$. (See also [29, 44, 56, 68, 63, 71, 95, 121] and the references therein.)
7. Green formulas and invertibility of $\boldsymbol{\Xi}$ on cylinders
We now extend the Green formulas and the results in Section 2 to the case of a cylinder. We then use these results to prove the invertibility of our generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (Equation (1.3)) on a cylinder.
We thus assume in this section that $M:=\mathfrak{S}\times\mathbb{R}$, where $\mathfrak{S}$ is a closed manifold (i.e., smooth, compact, without boundary). Moreover, all other objects used in this section will be translation invariant, thus $\Omega=\Omega^{\prime}\times\mathbb{R}$ and $V$ and $V_{0}$ descend to functions on $\mathfrak{S}$, denoted, respectively, $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ (see Equation (5.6) for a justification of this notation). Explicitly, we have $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$ for $x^{\prime}\in\mathfrak{S}$ and $t\in\mathbb{R}$. The results of this section will be used for $\mathfrak{S}$ possibly disconnected, so we will not assume that $\mathfrak{S}$ is connected, although the proofs of the results in this section can be reduced immediately to the connected case. Note also that we are continuing to assume that $M$ is a manifold with cylindrical ends, albeit of a very particular kind. Recall that the scalar product on $L^{2}(M;E)$ is
$(f,g)_{\Omega}\mathrel{\,:=\,}\int_{\Omega}f(x)\cdot g(x)\,\,d\operatorname{vol}_{g}(x)\,.$
(We have $(f,g)_{\Omega}\mathrel{\,:=\,}\int_{\Omega}f(x)\overline{g(x)}\,\,d\operatorname{vol}_{g}(x)$ if $f$ and $g$ are functions.) Our convention is that our scalar products are conjugate linear in the second variable and linear in the first.
7.1. Differential operators on cylinders and their indicial operators
We now make explicit the differential operators studied in the previous subsection for our manifold $M=\mathfrak{S}\times\mathbb{R}$ (a cylinder). We then use these explicit formulas to study the associated indicial operators and prove that they have the $L^{2}$-unique continuation property.
Let $p_{i}$, $i=1,2$ be the projections of $M=\mathfrak{S}\times\mathbb{R}$ onto $\mathfrak{S}$ and, respectively, $\mathbb{R}$. Let $p_{1}^{*}(T\mathfrak{S})$ be the pull-back of the tangent vector bundle to $\mathfrak{S}$ to $M$. It coincides with the vertical tangent bundle to the fibers of $\mathfrak{S}\times\mathbb{R}\to\mathbb{R}$. The pull-back $p_{2}^{*}(T\mathbb{R})$ is defined similarly. We will use the following decomposition:
(7.1) $TM=p_{1}^{*}(T\mathfrak{S})\oplus p_{2}^{*}(T\mathbb{R})=p_{1}^{*}(T\mathfrak{S})\times\mathbb{R}\,.$
Accordingly, any section $X$ of $TM$ will decompose as $X=(X_{1},f_{X})$, where $X_{1}$ is a section of $p_{1}^{*}(T\mathfrak{S})\to M:=\mathfrak{S}\times\mathbb{R}$ and $f_{X}\in\mathcal{C}^{\infty}(M)$. Thus $X_{1}$ is a smooth family of vertical vector fields on $M$ (i.e., tangent to the submanifolds $\mathfrak{S}\times\{t\}$, $t\in\mathbb{R}$). Similarly, $f_{X}$ corresponds to $f_{X}\partial_{t}$, where $t$ is the variable on $\mathbb{R}$.
We shall denote with a ‘prime’ (i.e., with ) the “vertical” objects associated to the fibers $\mathfrak{S}\times\{t\}$, $t\in\mathbb{R}$, of the projection $M\to\mathbb{R}$. We shall sometimes write simply $T\mathfrak{S}$ instead of its pull-back $p_{1}^{*}T\mathfrak{S}$. We shall do that especially when considering sections of the latter bundle, thus
$\mathcal{C}^{\infty}(M;T\mathfrak{S})\mathrel{\,:=\,}\mathcal{C}^{\infty}(M;p_{1}^{*}T\mathfrak{S})\,.$
Let $\nabla f:=(df)^{\sharp}$ be the usual gradient. If $f\in\mathcal{C}^{\infty}(M)$ and $x=(x^{\prime},t)\in M=\mathfrak{S}\times\mathbb{R}$, then $\nabla^{\prime}f(x^{\prime},t)$ is the gradient at $x^{\prime}$ of the restriction $f|_{\mathfrak{S}\times\{t\}}$, $t\in\mathbb{R}$, called the vertical gradient. The vertical gradient $\nabla^{\prime}f$ is thus a section of $p_{1}^{*}T\mathfrak{S}$. Similarly, $d^{\prime}$ is the vertical differential and $\operatorname{Def}\,^{\prime}$ is the vertical deformation tensor. In particular, the vertical gradient $\nabla^{\prime}:\mathcal{C}^{\infty}(M)\to\mathcal{C}^{\infty}(M;T\mathfrak{S})$ is given by $\nabla^{\prime}(f)=(d^{\prime}f)^{\sharp}$.
Proposition 7.1.
Let $M=\mathfrak{S}\times\mathbb{R}$ and let $\nabla^{\prime}$ be the vertical part of the gradient. In terms of the decomposition of Equation (7.1), we obtain that the gradient of $f\in\mathcal{C}^{\infty}(M)$, $M=\mathfrak{S}\times\mathbb{R}$, is $\nabla f=(\nabla^{\prime}f,\partial_{t}f)$. In matrix notation, we have
(7.2) $\nabla\,=\,\left(\begin{array}[]{c}{\nabla^{\prime}}\\ {\partial_{t}}\end{array}\right):\mathcal{C}^{\infty}(M)\to\begin{array}[]{c}\mathcal{C}^{\infty}(M;T\mathfrak{S})\\ \oplus\\ \mathcal{C}^{\infty}(M)\end{array}$
This follows by noticing that $df=d^{\prime}f+\partial_{t}fdt$, where $t\in\mathbb{R}$ is the canonical variable on the cylinder $\mathfrak{S}\times\mathbb{R}$.
Proposition 7.2.
Let $M=\mathfrak{S}\times\mathbb{R}$ and let $\operatorname{Def}\,^{\prime}$ be the vertical deformation operator (acting just on the $\mathfrak{S}$-directions) and similarly $d^{\prime}$ is the vertical differentiation. Using the decomposition (7.1), the deformation operator $\operatorname{Def}\,$ becomes
(7.3) $\operatorname{Def}\,\,=\,\left(\begin{array}[]{cc}\operatorname{Def}\,^{\prime}&0\\ \frac{1}{\sqrt{2}}\sharp\partial_{t}&\frac{1}{\sqrt{2}}d^{\prime}\\ 0&\partial_{t}\end{array}\right)\ :\ \begin{array}[]{c}\mathcal{C}^{\infty}(M;T\mathfrak{S})\\ \oplus\\ \mathcal{C}^{\infty}(M)\end{array}\ \rightarrow\ \begin{array}[]{c}\mathcal{C}^{\infty}(M;T^{*\otimes 2}\mathfrak{S})\\ \oplus\\ \mathcal{C}^{\infty}(M;T^{*}\mathfrak{S})\\ \oplus\\ \mathcal{C}^{\infty}(M)\,.\end{array}$
Proof.
Let $X,Y,Z$ be three smooth vector fields on $M=\mathfrak{S}\times\mathbb{R}$. The decomposition $TM=p_{1}^{*}(T\mathfrak{S})\oplus p_{2}^{*}(T\mathbb{R})=p_{1}^{*}(T\mathfrak{S})\times\mathbb{R}$ of Equation (7.1) gives $X=(X_{1},f_{X})$, where $X_{1}$ is a section of $p_{1}^{*}(T\mathfrak{S})\to M:=\mathfrak{S}\times\mathbb{R}$ and $f_{X}\in\mathcal{C}^{\infty}(M)$, as explained above. We decompose similarly $Y=(Y_{1},f_{Y})$ and $Z=(Z_{1},f_{Z})$. The Levi-Civita connection $\nabla_{X}^{LC,M}$ on $M$ is then, by unicity,
(7.4) $\nabla_{X}^{LC,M}Y\,=\,(\nabla_{X_{1}}^{LC,\mathfrak{S}}Y_{1}+f_{X}\partial_{t}Y_{1},X_{1}(f_{Y})+f_{X}\partial_{t}f_{Y})\,.$
The formula (7.1) gives that the symmetric tensor product $S^{2}T^{*}M\subset T^{*\otimes 2}M$ can be written as
(7.5) $S^{2}T^{*}M\,=\,S^{2}p_{1}^{*}T^{*}\mathfrak{S}\oplus p_{1}^{*}T^{*}\mathfrak{S}\oplus\mathbb{C}\,.$
(The factors involving $\sqrt{2}$ appear in order to make the above identification isometric.) With these identifications, we have
(7.6) $\big(\nabla_{Y}^{LC,M}X\big)\cdot Z\,=\,\big(\nabla_{Y_{1}}^{LC,\mathfrak{S}}X_{1}\big)\cdot Z_{1}+f_{Y}(\partial_{t}X_{1})\cdot Z_{1}+Y_{1}(f_{X})f_{Z}+f_{Y}f_{Z}\partial_{t}f_{X}\,.$
Formula 2.1 then yields the result. ∎
Recall that if $P\in\Psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F):=\Psi_{\operatorname{ess}}^{m}(\mathfrak{S}\times\mathbb{R};E,F)^{\mathbb{R}}$, then the Fourier transform yields the indicial operators $\widehat{P}(\tau)\in\Psi^{m}(\mathfrak{S};E,F)$, $\tau\in\mathbb{R}$, (see Definition 5.26). We can now prove the following result.
Proposition 7.3.
Let $M=\mathfrak{S}\times\mathbb{R}$ and let $\nabla^{\prime}$ be the vertical gradient. In terms of the decomposition of Equation (7.1), we obtain that the indicial operator $\widehat{\nabla}(\tau)$ is given by
$\widehat{\nabla}(\tau)\,=\,\left(\begin{array}[]{c}{\nabla^{\prime}}\\ {\imath\tau}\end{array}\right):\mathcal{C}^{\infty}(\mathfrak{S})\to\mathcal{C}^{\infty}(\mathfrak{S};TM)\ \simeq\begin{array}[]{c}\mathcal{C}^{\infty}(\mathfrak{S};T\mathfrak{S})\\ \oplus\\ \mathcal{C}^{\infty}(\mathfrak{S})\,.\end{array}$
Let $\Omega^{\prime}\subset\mathfrak{S}$ be an open subset. We have that $\widehat{\nabla}(\tau)$ is injective on $L^{2}(\Omega^{\prime})$ for $\tau\neq 0$ and $\widehat{\nabla}(\tau)$ satisfies the $L^{2}$–unique continuation on $L^{2}(\Omega^{\prime})$ property for all $\tau$. The kernel of $\widehat{\nabla}(0)$ on $L^{2}(\Omega^{\prime})$ consists of locally constant functions.
Proof.
The operator $\nabla^{\prime}$ acts only in the vertical directions, and hence $\widehat{\nabla}^{\prime}(\tau)=\nabla^{\prime}$ (it is independent of $\tau$). On the other hand, $\widehat{(\partial_{t})}(\tau)=\imath\tau$. The formula for $\widehat{\nabla}(\tau)$ follows then from these formulas and from Equation (7.2). For the rest of the proposition, it suffices to assume that $\mathfrak{S}$ is connected. Let $f\in\mathcal{C}^{\infty}(\mathfrak{S})$ be such that $\widehat{\nabla}(\tau)f=0$. This is equivalent to $\nabla^{\prime}f=0$ and $\imath\tau f=0$. If $\tau\neq 0$, then, of course, $f=0$, and hence $\widehat{\nabla}(\tau)$, besides being injective, satisfies also the $L^{2}$–unique continuation property on $\Omega^{\prime}$. Let us assume then that $\tau=0$. Then $f$ is constant on $\Omega^{\prime}$. If, moreover, $f$ vanishes on some non-empty subset of some connected component $\Omega_{0}^{\prime}$ of $\Omega^{\prime}$, then $f=0$ in $\Omega_{0}^{\prime}$. (This is the $L^{2}$–unique continuation property of $\nabla^{\prime}$.) ∎
This result and its proof extend to the $\operatorname{Def}\,$ operator.
Proposition 7.4.
Let $M=\mathfrak{S}\times\mathbb{R}$ and let $\operatorname{Def}\,^{\prime}$ be the vertical deformation operator and $d^{\prime}$ be the vertical exterior derivative. In terms of the decomposition of Equation (7.1), the indicial operator $\widehat{\operatorname{Def}}\,(\tau)$ (Definition 5.26) of the deformation operator $\operatorname{Def}\,$ on $M$ is given by
(7.7) $\widehat{\operatorname{Def}\,}(\tau)\,=\,\left(\begin{array}[]{cc}\operatorname{Def}\,^{\prime}&0\\ \frac{\imath}{\sqrt{2}}\sharp\tau&\frac{1}{\sqrt{2}}d^{\prime}\\ 0&\imath\tau\end{array}\right)\ :\ \begin{array}[]{c}\mathcal{C}^{\infty}(\mathfrak{S};T\mathfrak{S})\\ \oplus\\ \mathcal{C}^{\infty}(\mathfrak{S})\end{array}\ \rightarrow\ \begin{array}[]{c}\mathcal{C}^{\infty}(\mathfrak{S};T^{*\otimes 2}\mathfrak{S})\\ \oplus\\ \mathcal{C}^{\infty}(\mathfrak{S};T^{*}\mathfrak{S})\\ \oplus\\ \mathcal{C}^{\infty}(\mathfrak{S})\,.\end{array}$
Let $\Omega^{\prime}\subset\mathfrak{S}$ be an open subset. We have that the operator $\widehat{\operatorname{Def}}\,(\tau)$ is injective on $L^{2}(\Omega^{\prime};TM)\simeq L^{2}(\Omega^{\prime};T\mathfrak{S})\oplus L^{2}(\Omega^{\prime})$ for $\tau\neq 0$. For $\tau=0$ we have
$\ker\widehat{\operatorname{Def}}\,(0)\,=\,\{(\boldsymbol{u},c)\in\mathcal{C}^{\infty}(\Omega^{\prime};T\mathfrak{S})\oplus\mathcal{C}^{\infty}(\Omega^{\prime}):\operatorname{Def}\,^{\prime}\boldsymbol{u}=0\,,\ d^{\prime}c=0\}\,.$
In particular, $\widehat{\operatorname{Def}}\,(\tau)$ satisfies the $L^{2}$–unique continuation property on $\Omega^{\prime}$ for all $\tau\in\mathbb{R}$.
Proof.
The proof of the formula for $\widehat{\operatorname{Def}}\,(\tau)$ is similar to the proof of the formula for $\widehat{\nabla}(\tau)$ obtained in the proof of Proposition 7.3. Indeed, we notice first that the operator $\operatorname{Def}\,^{\prime}$ acts only in the vertical directions, and hence $\widehat{\operatorname{Def}}\,^{\prime}(\tau)=\operatorname{Def}\,^{\prime}$ (it is again independent of $\tau$). This formula, together with the formula $\widehat{(\partial_{t})}(\tau)=\imath\tau$ already discussed (see the proof of Proposition 7.3) and Equation (7.3) then yield the claimed formula for $\widehat{\operatorname{Def}}\,(\tau)$.
Let $U:=\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)\in\mathcal{C}^{\infty}(\Omega^{\prime};TM)\oplus\mathcal{C}^{\infty}(\Omega^{\prime})$ be such that $\widehat{\operatorname{Def}}\,(\tau)U=0$. Our assumption implies, in particular, that $\imath\tau p=0$. Let us assume first that $\tau\neq 0$. Then $p=0$. The second equation then becomes $\sharp\boldsymbol{u}=0$, and hence $\boldsymbol{u}=0$ also. Thus $\widehat{\operatorname{Def}}\,(\tau)$ is injective for $\tau\neq 0$. In particular, it satisfies the $L^{2}$-unique continuation property, as well.
It remains to prove that $\widehat{\operatorname{Def}}\,(0)$ satisfies the $L^{2}$-unique continuation property on $\Omega^{\prime}$ and to identify its kernel. Again, for the rest of the proof, it suffices to assume that $\Omega^{\prime}$ is connected. Let $U=\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)$, as before. The relation $\widehat{\operatorname{Def}}\,(0)U=0$ is equivalent to $\operatorname{Def}\,^{\prime}\boldsymbol{u}=0$ (from the first equation) and $d^{\prime}p=0$ (from the second equation, the third one being automatically satisfied). This gives the stated form for the kernel of $\widehat{\operatorname{Def}}\,(0)$. Therefore, if $\boldsymbol{u}$ and $p$ vanish on some non-empty open subsets of $\Omega^{\prime}$, then $\boldsymbol{u}=0$, because $\operatorname{Def}\,^{\prime}$ satisfies the $L^{2}$-unique continuation property on $\Omega^{\prime}$, and $p=0$, because $d^{\prime}$ satisfies the $L^{2}$-unique continuation property on $\Omega^{\prime}$. Thus $\widehat{\operatorname{Def}}\,(0)$ satisfies the $L^{2}$-unique continuation property on $\Omega^{\prime}$ and this completes the proof. ∎
7.2. Green formulas and energy estimates for indicial families
In view of further applications to manifolds with cylindrical ends, we want to extend the Green and representation formulas and the energy estimates of Section 2 (especially, Subsection 2.3) to the case of the indicial family $\widehat{\boldsymbol{\Xi}}$ of $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (see Definition 5.26). Thus, in this subsection, we continue to consider the cylinder $M=\mathfrak{S}\times\mathbb{R}$, where $\mathfrak{S}$ is a closed manifold. The results of this section will be used for $\mathfrak{S}$ not necessarily connected, so we will not assume that $\mathfrak{S}$ is connected, although the general case can immediately be reduced to the connected case. Also, we let $\color[rgb]{1,0,1}\widehat{V}(x^{\prime})={\mathcal{R}_{\infty}}V(x^{\prime},t)=V(x^{\prime},t)\geq 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})={\mathcal{R}_{\infty}}V_{0}(x^{\prime},t)=V_{0}(x^{\prime},t)\geq 0$ be smooth, translation invariant functions (this notation is justified by Equation (5.6)). Again, this choice of $V$ and $V_{0}$ satisfies all our previous assumptions. (Note that in some of our previous papers, the notation $\tilde{V}$ and $\tilde{V}_{0}$ was used instead of $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$. However, in this paper, the notation $\tilde{u}$ means $\tilde{u}(x):=u(-x)$.)
Recall the Fourier transform (in the variable $t\in\mathbb{R}$) $\mathcal{F}:L^{2}(\mathfrak{S}\times\mathbb{R})\to L^{2}(\mathfrak{S}\times\mathbb{R})$, $\hat{u}=\mathcal{F}u$, which satisfies
(7.8) $2\pi(u,v)_{M}\,=\,(\mathcal{F}u,\mathcal{F}v)_{M}\,=\,\int_{\mathbb{R}}(\hat{u}(\tau),\hat{v}(\tau))_{\mathfrak{S}}\,d\tau\,.$
Note that the first equality in Equation (7.8) is a consequence of the Parceval theorem for the Fourier transform, and we use the simplified notation $\hat{u}(\tau)$ instead of $\hat{u}(x,\tau)$ for $(x,\tau)\in\mathfrak{S}\times\mathbb{R}=M$. We shall use the notations $\hat{u}(\tau)$ and $\mathcal{F}u(\tau)$ interchangeably.
For any $P\in\Psi_{\operatorname{ess}}^{m}(M;E)^{\mathbb{R}}$ we have
(7.9) $\widehat{Pu}(\tau)\mathrel{\,:=\,}(\mathcal{F}Pu)(\tau)\,=\,\widehat{P}(\tau)\hat{u}(\tau)\,.$
This gives
(7.10) $2\pi(Pu,v)_{M}\,=\,(\mathcal{F}Pu,\mathcal{F}v)_{M}\,=\,\int_{\mathbb{R}}(\widehat{P}(\tau)\hat{u}(\tau),\hat{v}(\tau))_{\mathfrak{S}}\,d\tau\,.$
For $\phi\in L^{\infty}(\mathbb{R})$, we shall need the Fourier multiplier $\mathbb{M}_{\phi}$ defined by $\mathcal{F}\mathbb{M}_{\phi}=\phi\mathcal{F}$, or, explicitely,
(7.11) $\mathcal{F}\mathbb{M}_{\phi}u](\tau)\mathrel{\,:=\,}\phi(\tau)\hat{u}(\tau)\,.$
All of the above definitions and relations extend to vector bundles that are lifted from vector bundles on $\mathfrak{S}$ (and hence they come with a natural action of $\mathbb{R}$ by translations). We then have that the Fourier multiplier operators commute with the operators in $\Psi_{\operatorname{sus}}^{m}(M;E,F):=\Psi_{\operatorname{ess}}^{m}(M;E,F)^{\mathbb{R}}$ (recall that $M=\mathfrak{S}\times\mathbb{R}$).
Also, in this subsection, we let $\Omega^{\prime}\subset\mathfrak{S}$ be a smooth open subset with boundary $\Gamma^{\prime}:=\partial\Omega^{\prime}$ (a smooth manifold). Our results are formulated in general, but there is no loss of generality to assume that $\Omega^{\prime}$ is connected (that is, a smooth domain). Also, recall that, in this section, $\color[rgb]{1,0,1}\widehat{V}(x^{\prime})=V(x^{\prime},t)$ and $\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})=V_{0}(x^{\prime},t)$, as motivated by Equation (5.6). In particular, $V=\color[rgb]{1,0,1}\widehat{V}$ and $V_{0}=\color[rgb]{1,0,1}\widehat{V}_{0}$ are translation invariant. The main equation of Notation 2.10 becomes.
(7.12) $\begin{gathered}U\mathrel{\,:=\,}(\boldsymbol{u}\ \ p)^{\top}\,,\ W\mathrel{\,:=\,}(\boldsymbol{w}\ \ q)^{\top}\in H^{1}(\Omega^{\prime};TM)\oplus L^{2}(\Omega^{\prime})\,,\\ \qquad\mathfrak{v}\mathrel{\,:=\,}({\color[rgb]{1,0,1}\widehat{V}}\boldsymbol{u},\boldsymbol{w})_{\Omega^{\prime}}-({\color[rgb]{1,0,1}\widehat{V}_{0}}p,q)_{\Omega^{\prime}}\,,\quad\mbox{and}\\ B_{\Omega^{\prime},\tau}(U,W)\mathrel{\,:=\,}2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{w})_{\Omega^{\prime}}+(\widehat{\nabla}^{*}(\tau)\boldsymbol{u},q)_{\Omega^{\prime}}+(p,\widehat{\nabla}^{*}(\tau)\boldsymbol{w})_{\Omega^{\prime}}+\mathfrak{v}\,.\end{gathered}$
Also, in this subsection, we let $(\ ,\ )^{\prime}:=(\ ,\ )_{\partial\Omega^{\prime}},$ the scalar product on the boundary $\Gamma^{\prime}=\partial\Omega^{\prime}$ of $\Omega^{\prime}$.
Recall that $M=\mathfrak{S}\times\mathbb{R}$ and $\Omega=\Omega^{\prime}\times\mathbb{R}$ and all the operators of the form $\widehat{P}(\tau)$ are obtained from the Fourier transform with respect to the action of $\mathbb{R}$ by translation, see, for example, Equation (7.9). Proposition 2.11 becomes:
Proposition 7.5.
Let $U^{\top}:=(\boldsymbol{u}\ \ p)$ and $W^{\top}:=(\boldsymbol{w}\ \ q)$ belong to $H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$ and let $(\ ,\ )^{\prime}:=(\ ,\ )_{\partial\Omega^{\prime}}$. Then we have the following formulas:
  1. (i)
    $(\widehat{\boldsymbol{\Xi}}(\tau)U,W\big)_{\Omega^{\prime}}\,=\,B_{\Omega^{\prime},\tau}(U,W)+\big(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)U,\boldsymbol{w}\big)^{\prime}.$
  2. (ii)
    $\big(\widehat{\boldsymbol{\Xi}}(\tau)U,W\big)_{\Omega^{\prime}}-\big(U,\widehat{\boldsymbol{\Xi}}(\tau)W\big)_{\Omega^{\prime}}\,=\,(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)U,\boldsymbol{w})^{\prime}-\big(\boldsymbol{u},\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)W\big)^{\prime}\,.$
If $\partial\Omega^{\prime}$ is empty, then we drop the the inner products $(\ ,\ )^{\prime}$ on $\partial\Omega^{\prime}$.
Proof.
The proof is similar to that of Proposition 2.11 (sometimes it follows from that proposition using the Fourier transform).s Let then $U_{0}:=(\,\boldsymbol{u}_{0}\ \ p_{0}\,)^{\top},W_{0}:=(\,\boldsymbol{w}_{0}\ \ q_{0}\,)^{\top}\in H^{2}(\Omega^{\prime}\times\mathbb{R};TM)\oplus H^{1}(\Omega^{\prime}\times\mathbb{R})$, as in Notation (2.10) with $\Omega$ replaced with $\Omega^{\prime}\times\mathbb{R}$. The first part of Proposition 2.11 and Equation (7.10) give then that
$\frac{1}{2\pi}\int_{\mathbb{R}}(\widehat{\boldsymbol{\Xi}}(\tau)\widehat{U}_{0}(\tau),\widehat{W}_{0}(\tau)\big)_{\Omega^{\prime}}\,d\tau\,=\,(\boldsymbol{\Xi}U_{0},W_{0})_{\Omega}\,=\,B_{\Omega}(U_{0},W_{0})+(\boldsymbol{T}_{\boldsymbol{\nu}}U_{0},\boldsymbol{w}_{0})^{\prime}\\ \,=\,\frac{1}{2\pi}\int_{\mathbb{R}}\big[B_{\Omega^{\prime},\tau}(\widehat{U}_{0}(\tau),\widehat{W}_{0}(\tau))+(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)\widehat{U}_{0}(\tau),\hat{\boldsymbol{w}}_{0}(\tau))^{\prime}\big]\,d\tau\,.$
By replacing $U_{0}$ with its image $\mathbb{M}_{\phi}U_{0}$ through the Fourier multiplier $\mathbb{M}_{\phi}$, we obtain
$\int_{\mathbb{R}}\phi(\tau)(\widehat{\boldsymbol{\Xi}}(\tau)\widehat{U}_{0}(\tau),\widehat{W}_{0}(\tau)\big)_{\Omega^{\prime}}\,d\tau\\ \,=\,\int_{\mathbb{R}}\phi(\tau)\big[B_{\Omega^{\prime},\tau}(\widehat{U}_{0}(\tau),\widehat{W}_{0}(\tau))+(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)\widehat{U}_{0}(\tau),\hat{\boldsymbol{w}}_{0}(\tau))^{\prime}\big]\,d\tau\,.$
Since $\phi\in L^{\infty}(\mathbb{R})$ is arbitrary, we obtain, for every $\tau\in\mathbb{R}$, that
$(\widehat{\boldsymbol{\Xi}}(\tau)\widehat{U}_{0}(\tau),\widehat{W}_{0}(\tau)\big)_{\Omega^{\prime}}\,=\,B_{\Omega^{\prime},\tau}(\widehat{U}_{0}(\tau),\widehat{W}_{0}(\tau))+(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)\widehat{U}_{0}(\tau),\hat{\boldsymbol{w}}_{0}(\tau))^{\prime}\,.$
The point (i) (for any fixed $\tau$) then follows by choosing $U_{0}$ and $W_{0}$ such that $\widehat{U}_{0}(\tau)=U$ and $\widehat{W}_{0}(\tau)=W$.
The second point follows immediately from (i) using that $B_{\Omega^{\prime},\tau}$ is sesquilinear (following the same method of proof as that for Proposition 2.11). ∎
It follows from the definition of indicial operators (Definition 5.26) that
(7.13) $\widehat{\boldsymbol{\Xi}}(\tau)\,=\,\left(\begin{array}[]{cc}2\widehat{\operatorname{Def}}^{*}\,(\tau)\widehat{\operatorname{Def}}\,(\tau)+\color[rgb]{1,0,1}\widehat{V}&\widehat{\nabla}(\tau)\\ \widehat{\nabla}^{*}(\tau)&-\color[rgb]{1,0,1}\widehat{V}_{0}\end{array}\right)\,$,
where, we recall, we are identifying $\color[rgb]{1,0,1}\widehat{V}=\widehat{V}(\tau)$ and $\color[rgb]{1,0,1}\widehat{V}_{0}=\widehat{V}_{0}(\tau)$. (The formula (7.13) is valid for any manifold with straight cylindrical ends, not just for our manifold $M:=\mathfrak{S}\times\mathbb{R}$.) The last proposition gives then the following lemma analogous to Corollary 2.14. Recall that, we are assuming that $V,V_{0}\geq 0$. Also, recall that $\Omega^{\prime}\subset\mathfrak{S}$ is a smooth domain.
Lemma 7.6.
Let $\tau\in\mathbb{R}$, let $\color[rgb]{1,0,1}\widehat{V}\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R};\operatorname{End}(TM))^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(\mathfrak{S};TM)$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R})^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(\mathfrak{S})$ be non-negative (i.e., $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$), and let
$U=\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$
satisfy $\widehat{\boldsymbol{\Xi}}(\tau)U=0$ in $\Omega^{\prime}$ and $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)U,\boldsymbol{u})^{\prime}=0$. Then the following vanishing relations hold in $\Omega^{\prime}$:
$\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}\,=\,0\,,\quad\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u}\,=\,0\,,\quad\widehat{\nabla}^{*}(\tau)\boldsymbol{u}\,=\,0\,,\quad\widehat{\nabla}(\tau)p\,=\,0\,,\quad\mbox{and}\quad\color[rgb]{1,0,1}\widehat{V}_{0}p\,=\,0\,.$
Proof.
Let us take $W=U^{\prime}:=\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {-p}\end{array}\right)$ in the formula
$\big(\widehat{\boldsymbol{\Xi}}(\tau)U,W\big)_{\Omega^{\prime}}\,=\,B_{\Omega^{\prime},\tau}(U,W)+(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)U,\boldsymbol{w})^{\prime}$
of Proposition 7.5. Let $\operatorname{Re}(z)$ denote the real part of $z\in\mathbb{C}$. Then $\operatorname{Re}\big[-(\widehat{\nabla}^{*}(\tau)\boldsymbol{u},p)_{\Omega^{\prime}}+(p,\widehat{\nabla}^{*}(\tau)\boldsymbol{u})_{\Omega^{\prime}}\big]=0$. Let $\mathfrak{v}:=(\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u},\boldsymbol{u})_{\Omega^{\prime}}+(\color[rgb]{1,0,1}\widehat{V}_{0}p,p)_{\Omega^{\prime}}$, as before. Together with the definition of $B_{\Omega^{\prime},\tau}$ in Equation (7.12), this gives
$0$ $\,=\,\operatorname{Re}\big[\big(\widehat{\boldsymbol{\Xi}}(\tau)U,U^{\prime}\big)_{\Omega^{\prime}}-(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)U,\boldsymbol{u})^{\prime}\big]\,=\,\operatorname{Re}\big[B_{\Omega^{\prime},\tau}(U,U^{\prime})\big]$
$\,=\,\operatorname{Re}\big[2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u})_{\Omega^{\prime}}-(\widehat{\nabla}^{*}(\tau)\boldsymbol{u},p)_{\Omega^{\prime}}+(p,\widehat{\nabla}^{*}(\tau)\boldsymbol{u})_{\Omega^{\prime}}+\mathfrak{v}\big]$
= $\,=\,2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u})_{\Omega^{\prime}}+\mathfrak{v}$
= $\,=\,2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u})_{\Omega^{\prime}}+(\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u},\boldsymbol{u})_{\Omega^{\prime}}+(\color[rgb]{1,0,1}\widehat{V}_{0}p,p)_{\Omega^{\prime}}\,.$
All three terms in the last sum are non-negative, so each of them equals zero. Therefore $\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}=0$, $\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u}=0$, and $\color[rgb]{1,0,1}\widehat{V}_{0}p=0$. We also have
$0\,=\,\widehat{\boldsymbol{\Xi}}(\tau)U\,=\,\left(\begin{array}[]{c}{2\widehat{\operatorname{Def}}^{*}\,(\tau)\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}+\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u}+\widehat{\nabla}(\tau)p}\\ {\widehat{\nabla}^{*}(\tau)\boldsymbol{u}-\color[rgb]{1,0,1}\widehat{V}_{0}p}\end{array}\right)\,=\,\left(\begin{array}[]{c}{\widehat{\nabla}(\tau)p}\\ {\widehat{\nabla}^{*}(\tau)u}\end{array}\right)\,.$
Hence $\widehat{\nabla}^{*}(\tau)\boldsymbol{u}=0$ and $\widehat{\nabla}(\tau)p=0$, as claimed. This completes the proof. ∎
Recall that $V\succ 0$ on $\Omega$ means that $V\geq 0$ and that $V$ is (strictly) positive definite at at least one point in every connected component of $\Omega$ (see Definition 2.13). We now take $\tau=0$ in the above lemma. The $L^{2}$-unique continuation properties of $\widehat{\operatorname{Def}}\,(0)$ and $\widehat{\nabla}(0)$ then give the following result.
Corollary 7.7.
Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$ and let $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$ satisfy $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega^{\prime}$ and $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(0)U,\boldsymbol{u})_{\Gamma^{\prime}}=0$ (as in Lemma 7.6, but for $\tau=0$).
  1. (i)
    If $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$ (Definition 2.13), then $p=0$ on $\Omega^{\prime}$.
  2. (ii)
    Similarly, if one of the following conditions is satisfied:
    1. (a)
      $\boldsymbol{u}=0$ on $\Gamma^{\prime}$ or
    2. (b)
      $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega^{\prime}$,
    then $\boldsymbol{u}=0$ in $\Omega^{\prime}$.
Proof.
Lemma 7.6 gives $\widehat{\nabla}(0)p=0$ in $\Omega^{\prime}$. The assumption that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ in $\Omega$ together with Proposition 7.3 (stating the $L^{2}$-unique continuation property of $\widehat{\nabla}(0)$ in $\Omega^{\prime}$) then give $p=0$. Similarly, Lemma 7.6 gives $\widehat{\operatorname{Def}}\,(0)\boldsymbol{u}=0$. Our assumptions allow us to use Proposition 7.4 (the $L^{2}$-unique continuation property of $\widehat{\operatorname{Def}}\,(0)$) to conclude that $\boldsymbol{u}=0$. This completes the proof. ∎
As before, we obtain the following corollaries.
Corollary 7.8.
Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$ and let $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\mathfrak{S};TM)\oplus H^{1}(\mathfrak{S})$ satisfy $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\mathfrak{S}$.
  1. (i)
    If $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$, then $p=0$ on $\mathfrak{S}$.
  2. (ii)
    Similarly, if $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\mathfrak{S}$, then $\boldsymbol{u}=0$ in $\mathfrak{S}$.
The case $\tau\neq 0$ is very similar, in fact, even easier than the case $\tau=0$, because of the injectivity of the relevant operators for $\tau\neq 0$.
Corollary 7.9.
Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$, $\tau\neq 0$, and $U^{\top}=(\boldsymbol{u}\ \ p)\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$ satisfy $\widehat{\boldsymbol{\Xi}}(\tau)U=0$ in $\Omega^{\prime}$ and $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)U,\boldsymbol{u})^{\prime}=0$ (the same assumptions as in Lemma 7.6, except that $\tau\neq 0$ in this corollary). Then $U=0$.
Proof.
Lemma 7.6 gives $\widehat{\nabla}(\tau)p=0$ and $\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}=0$. Because both operators $\widehat{\nabla}(\tau)$ and $\widehat{\operatorname{Def}}\,(\tau)$ are injective for $\tau\neq 0$ (by Propositions 7.3 and 7.4), we obtain $\boldsymbol{u}=0$ and $p=0$. ∎
In the last Corollary, if $\Omega^{\prime}=\mathfrak{S}$, then the condition $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)U,\boldsymbol{u})^{\prime}=0$ is void. It is automatically satisfied, so, in particular, it is unnecessary. This yields the following result.
Corollary 7.10.
Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$, $\tau\neq 0$, and $U^{\top}=(\boldsymbol{u}\ \ p)\in H^{2}(\mathfrak{S};TM)\oplus H^{1}(\mathfrak{S})$ satisfy $\widehat{\boldsymbol{\Xi}}(\tau)U=0$. Then $U=0$.
7.3. The invertibility of $\boldsymbol{\Xi}$ on a cylinder
Recall from Theorem 6.15 that the Fredholm property of an operator depends on the invertibility of its limit operator, which acts on the cylinder associated to $M$. The limit operator of $\boldsymbol{\Xi}_{V,V_{0}}$ is $\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$. It is the purpose of this subsection to establish the invertibility of $\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$ acting on the associated cylinder under the assumption that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
We thus continue to assume, in this subsection, that $M$ is the cylinder $M=\mathfrak{S}\times\mathbb{R}$ with $\mathfrak{S}$ a closed manifold and with a product metric. We also assume that $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and that $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$, and hence $V$ and $V_{0}$ are thus translation invariant. We also assume that they are both $\succ 0$ on $M$ (see Definition 2.13). This condition is equivalent to $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
Let us notice that, since $M:=\mathfrak{S}\times\mathbb{R}$ is endowed with the product metric, this metric is translation invariant, and hence $\boldsymbol{\Xi}$ is also translation invariant. We can then define its Fourier transform (or indicial family) $\widehat{\boldsymbol{\Xi}}(\tau)$ as in Section 5.9, especially Definition 5.22 (and the discussion surrounding it). In particular, we choose the definition that avoids duplications, see Remark also 5.28.
Proposition 7.11.
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$. Then $\widehat{\boldsymbol{\Xi}}(0):=\widehat{\boldsymbol{\Xi}}_{V,V_{0}}(0)$ is invertible and $\widehat{\boldsymbol{\Xi}}(0)^{-1}\in\Psi^{[-\boldsymbol{s}-\boldsymbol{t}]}(\mathfrak{S};TM\oplus\mathbb{C})$.
Proof.
The generalized Stokes operator $\widehat{\boldsymbol{\Xi}}(0)$ is $(\boldsymbol{s},\boldsymbol{t})$–Douglis-Nirenberg elliptic because $\boldsymbol{\Xi}$ is $(\boldsymbol{s},\boldsymbol{t})$–Douglis-Nirenberg elliptic (Proposition 6.13) and because ellipticity is preserved by taking indicial families. Because $\mathfrak{S}$ is compact, Theorem 6.15 then gives that $\widehat{\boldsymbol{\Xi}}(0)$ is Fredholm as an operator
$\widehat{\boldsymbol{\Xi}}(0):H^{1}(\mathfrak{S};TM)\oplus L^{2}(\mathfrak{S})\to H^{-1}(\mathfrak{S};TM)\oplus L^{2}(\mathfrak{S})\,.$
The second Green formula of Proposition 7.5 shows that $\widehat{\boldsymbol{\Xi}}(0)$ is self-adjoint, and hence it has index zero. Therefore, in order to show that $\widehat{\boldsymbol{\Xi}}(0)$ induces a bijection between the above spaces, it suffices to show that it is injective. Let then $U=\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)=(\,\boldsymbol{u}\ \ p\,)^{\top}\in L^{2}(\mathfrak{S};TM\oplus\mathbb{C})$ be such that $\widehat{\boldsymbol{\Xi}}(0)U=0$. The ellipticity of $\widehat{\boldsymbol{\Xi}}(0)$ guarantees then that $U\in H^{\infty}(\mathfrak{S};TM\oplus\mathbb{C})$, by elliptic regularity (see, for example, Theorem 15.3.30 in [63]). Assumption 8.3 allow us to use Corollary 7.8 to conclude that $U=0$. Therefore $\widehat{\boldsymbol{\Xi}}(0)$ is invertible as an operator between the indicated spaces. The property that its inverse belongs to the space $\Psi^{[-\boldsymbol{s}-\boldsymbol{t}]}(\mathfrak{S};TM\oplus\mathbb{C})$ follows from the work of Beals and others (see Theorem 7.4 in [67] for more details). ∎
The case $\tau\neq 0$ is completely similar, but only requires the weaker assumptions $\color[rgb]{1,0,1}\widehat{V}:={\mathcal{R}_{\infty}}(V)\geq 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}:={\mathcal{R}_{\infty}}(V_{0})\geq 0$.
Proposition 7.12.
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $\tau\neq 0$, and that $\color[rgb]{1,0,1}\widehat{V}_{0},\color[rgb]{1,0,1}\widehat{V}\geq 0$. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$. Then $\widehat{\boldsymbol{\Xi}}(\tau)$ is invertible and $\widehat{\boldsymbol{\Xi}}(\tau)^{-1}\in\Psi^{[-\boldsymbol{s}-\boldsymbol{t}]}(\mathfrak{S};TM\oplus\mathbb{C})$.
Proof.
The proof is almost the same as that of Proposition 7.11 (but easier). The only difference is that, instead of Corollary 7.8, we use Corollary 7.10, which only requires that $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ be non-negative. ∎
Standard properties of pseudodifferential operators yield the following result, which will, however, not be used in this paper. It does specify, however, in what sense we are talking about the invertibility of our operators $\widehat{\boldsymbol{\Xi}}(\tau)$.
Remark 7.13.
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ with $\mathfrak{S}$ compact and that $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ are non-negative. Then, for any $\tau,r\in\mathbb{R}$, $\tau\neq 0$, we have an isomorphism
$\widehat{\boldsymbol{\Xi}}(\tau)\mathrel{\,:=\,}\widehat{\boldsymbol{\Xi}}_{V,V_{0}}(\tau):H^{r+1}(\mathfrak{S};TM)\oplus H^{r}(\mathfrak{S})\to H^{r-1}(\mathfrak{S};TM)\oplus H^{r}(\mathfrak{S})\,.$
The result is true also for $\tau=0$ if we further assume that $M$, $V$, and $V_{0}$ satisfy the non-vanishing Assumption 8.3. Under the same conditions, we have $\widehat{\boldsymbol{\Xi}}(\tau)^{-1}\in\Psi^{[-\boldsymbol{s}-\boldsymbol{t}]}(\mathfrak{S};TM\oplus\mathbb{C})$.
The last two propositions then give the invertibility of $\boldsymbol{\Xi}$ on cylinders (provided that the positivity at infinity assumption 8.1 is satisfied).
Theorem 7.14.
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$, that $V$ and $V_{0}$ are translation invariant and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. Then the operator
$\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}:H^{1}(M;TM)\oplus L^{2}(M)\to H^{-1}(M;TM)\oplus L^{2}(M)$
is invertible. If $\boldsymbol{s}=\boldsymbol{t}=(1,0)$, its inverse satisfies $\boldsymbol{\Xi}^{-1}\in\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM\oplus\mathbb{C})$ and maps $H^{r+1}(M;TM)\oplus H^{r}(M)\to H^{r-1}(M;TM)\oplus H^{r}(M)$ isomorphically for all $r\in\mathbb{R}$.
Proof.
We know from Propositions 7.11 and 7.12 that $\widehat{\boldsymbol{\Xi}}(\tau)$ is invertible for all $\tau\in\mathbb{R}$. It is well-known from the works of Melrose, Schulze and others (see also Theorem 5.11 in [68]) that then $\boldsymbol{\Xi}$ is invertible as stated. Once we have proved that $\boldsymbol{\Xi}$ is invertible, the fact that $\boldsymbol{\Xi}^{-1}\in\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM\oplus\mathbb{C})$ follows from the work of Beals and others (see Theorem 7.4 in [67] for further details). The claimed isomorphisms are obtained from the mapping properties of translation invariant operators. ∎
8. Invertibility of $\boldsymbol{\Xi}$, layer potentials, and jump relations on manifolds with cylindrical ends
We now use the results of the previous sections and prove the invertibility of several Stokes-type operators (under suitable positivity conditions on the potentials $V$ and $V_{0}$) and use these results to define our layer potentials. We then prove the main jump formulas for these layer potentials.
We assume from now on that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a manifold with straight cylindrical ends, so $M^{\prime}$ is a closed manifold (i.e., compact, boundaryless). The case of closed manifolds was treated in [69] and will be briefly reminded in Subsection 12.2. Also, from now on, we shall assume that $V$ and $V_{0}$ are translation invariant at infinity and that they are are smooth. (Some more precise assumption will be formulated later, see Assumptions 8.1 and 8.3.)
8.1. The Fredholm and invertibility properties of $\boldsymbol{\Xi}$
Using the results of the previous subsection, we now recall the proof of the Fredholm property, respectively, of the invertibility property of $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ under suitable positivity assumptions on $V$ and $V_{0}$ (Assumptions 8.1 and, respectively, 8.3). These results were first established in [68]. The proof relies on the “Mitrea-Taylor” trick [99] and on the results and methods from [67]. To establish the Fredholm property, we shall need the following assumption (slightly stronger than Assumption ).
Assumption 8.1.
Let $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ be a smooth, connected manifold with straight cylindrical ends with $M^{\prime}:=\partial M_{0}$ closed, non-empty and let $V\in\mathcal{C}^{\infty}_{\operatorname{inv}}(M;TM)$ and $V_{0}\in\mathcal{C}^{\infty}_{\operatorname{inv}}(M)$ be our potentials. We shall say that $V$ and $V_{0}$ satisfy the positivity assumption at infinity on $M$ if
$\color[rgb]{1,0,1}\widehat{V}\mathrel{\,:=\,}{\mathcal{R}_{\infty}}V\succ 0\quad\mbox{and}\quad\color[rgb]{1,0,1}\widehat{V}_{0}\mathrel{\,:=\,}{\mathcal{R}_{\infty}}V_{0}\succ 0\quad\mbox{on }M\,.$
Notice that we are not assuming that, globally, $V$ and $V_{0}$ are non-negative. This will be done in Assumption 8.3, which is, hence slightly stronger than Assumption 8.1. We also recall that, if $a$ is a multiplication operator, then we use also the notation $\color[rgb]{1,0,1}\widehat{a}$ for ${\mathcal{R}_{\infty}}(a)$ (see Equation (5.6) for a justification of this notation). We also recall (see Definition 2.13) that the explicit forms of the last two assumptions amount to the following:
  1. (i)
    $\color[rgb]{1,0,1}\widehat{V}\geq 0$ and, in every connected component of $M^{\prime}$, there is a point $x$ (and hence a whole open set containing it) where $\color[rgb]{1,0,1}\widehat{V}(x)$ is invertible (i.e., $V_{0}(x)>0$ in $M_{n}(\mathbb{C})$).
  2. (ii)
    $\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$ and, in every connected component of $M^{\prime}$, there is a point $x$ (and hence a whole open set containing it) where $\color[rgb]{1,0,1}\widehat{V}_{0}(x)$ is invertible (i.e., $V_{0}(x)\in(0,\infty)$).
Recall that by the statement $\phi\not\equiv 0$ on AA,” we mean that there exists $a$ in the domain of $\phi$ such that $\phi(a)\not=0$. To negate this statement, we shall write $\phi=0$ in AA.” The role of this assumption (Assumption 8.1) is to be able to use the results of the last subsection (which relied on the second part of Corollary 7.7 to obtain the vanishing of $p$ and $\boldsymbol{u}$ in $\Omega^{\prime}$, if $\Omega=M$ we used Corollary 7.7 instead). The assumption that $M$ is connected is just to simplify some of our statements. For instance, it ensures that $V_{0}\not\equiv 0$ on all connected components of $M$ (since there is only one such connected component!). The general case follows from the case $M$ connected easily. By contrast, we cannot assume $M^{\prime}$ to be connected, as it may be disconnected even if $M$ is connected.
Let us study first the Fredholm property of $\boldsymbol{\Xi}$ on our manifold $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ with straight cylindrical ends. Recall that $M^{\prime}:=\partial M_{0}$ is always assumed to be compact. We first notice that the invertibility of the limit operator ${\mathcal{R}_{\infty}}\boldsymbol{\Xi}$ follows from the results of the previous section.
8.2. The Moore-Penrose pseudoinverse of $\boldsymbol{\Xi}$
Let $\mathcal{N}$ be the kernel of $\boldsymbol{\Xi}:H^{1}(M;TM)\oplus L^{2}(M)\to H^{-1}(M;TM)\oplus L^{2}(M)$. The elliptic regularity of $\boldsymbol{\Xi}$ implies that $\mathcal{N}$ will consist of smooth sections. Assume that $\mathcal{N}\subset L^{2}(M;TM\oplus\mathbb{C})$ and that it is finite dimensional and let $p_{\mathcal{N}}$ be the $L^{2}$–orthogonal projection onto $\mathcal{N}$. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$ in Definition 6.11. Assume also that there exists a pseudodifferential operator
(8.1) $\boldsymbol{\Xi}^{(-1)}\in\Psi_{\operatorname{cl}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM\oplus\mathbb{C})\mathrel{\,:=\,}\left(\begin{array}[]{cc}\Psi^{-2}(M;TM)&\Psi^{-1}(M;TM,\mathbb{C})\\ \Psi^{-1}(M;\mathbb{C},TM)&\Psi^{0}(M)\end{array}\right)$
such that
(8.2) $\begin{gathered}\boldsymbol{\Xi}^{(-1)}(1-p_{\mathcal{N}})\,=\,(1-p_{\mathcal{N}})\boldsymbol{\Xi}^{(-1)}\,=\,\boldsymbol{\Xi}^{(-1)}\ \mbox{ and }\\ \boldsymbol{\Xi}\boldsymbol{\Xi}^{(-1)}\,=\,\boldsymbol{\Xi}^{(-1)}\boldsymbol{\Xi}=1-p_{\mathcal{N}}\end{gathered}$
(i.e., we assume that $\boldsymbol{\Xi}$ is invertible on the orthogonal complement of its kernel $\mathcal{N}$). The operator $\boldsymbol{\Xi}^{(-1)}$ is uniquely determined by the relations (8.2) and will be called the Moore-Penrose pseudoinverse of $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$.
We also obtain the following result:
Theorem 8.2.
Let $M$ be a manifold with straight cylindrical ends and assume that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ (Assumption 8.1). Then the following properties hold.
  1. (i)
    The limit operator ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})\in\Psi_{\operatorname{sus}}^{m}(M^{\prime};TM,\mathbb{C})$ is invertible (see Lemma 5.11).
  2. (ii)
    The generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ is Fredholm.
  3. (iii)
    Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$. Consequently, the Moore-Penrose pseudoinverse of $\boldsymbol{\Xi}$ satisfies $\boldsymbol{\Xi}^{(-1)}\in\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM\oplus\mathbb{C})$.
Proof.
Assumption 8.1 implies that ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})=\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$ satisfies the assumptions of Theorem 7.14, which then yields right away the first point. The generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ is elliptic by Proposition 6.13. Since ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ is invertible, Theorem 6.15 implies that $\boldsymbol{\Xi}$ is Fredholm. The last point follows from the Beals-type result proved in [67, Theorem 7.4] (this result can also be found in [63]). ∎
We now use this result to prove the invertibility of $\boldsymbol{\Xi}$ under the following assumption.
Assumption 8.3.
We say that $V$ and $V_{0}$ satisfy the positivity assumption on $M$ if they satisfy the positivity assumption at infinity on $M$ (Assumption 8.1) and $V\geq 0$ and $V_{0}\geq 0$.
It follows from this formulation that Assumption 8.3 implies Assumption 8.1. We can now prove our main invertibility result (Theorem 5.11 from [68]).
Theorem 8.4.
Let $M$ be a manifold with straight cylindrical ends and assume that $V$ and $V_{0}$ satisfy the positivity assumption (Assumption 8.3). Then the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ is invertible. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$. Consequently, its inverse satisfies $\boldsymbol{\Xi}^{-1}\in\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM\oplus\mathbb{C})$ and
$\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}:H^{r+1}(M;TM)\oplus H^{r}(M)\to H^{r-1}(M;TM)\oplus H^{r}(M)$
is an isomorphism for all $r\in\mathbb{R}$.
Proof.
The induced operator
$\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}:H^{1}(M;TM)\oplus L^{2}(M)\to H^{-1}(M;TM)\oplus L^{2}(M)$
is Fredholm by Theorem thm2.main.Fredholm just proved. It is also self-adjoint, and hence it is of index zero (between the indicated spaces). To prove that $\boldsymbol{\Xi}$ is invertible, it hence suffices to prove that it is injective. We proceed as in the proof of the last two propositions, using energy estimates.
Let $U\in H^{1}(M;TM)\oplus L^{2}(M)$ be such that $\boldsymbol{\Xi}U=0$ on $M$. Then $U\in H^{2}(M;TM)\oplus H^{1}(M)$, by elliptic regularity. Let us notice that the fact that $V$ and $V_{0}$ satisfy the non-vanishing assumption on $M$ (Assumption 8.3) and that $M$ is connected implies that $V_{0}\not\equiv 0$ and $V\succ 0$ on $M$. The assumption $\boldsymbol{\Xi}U=0$ then allows us to use Corollary 2.15 to conclude that $U=0$. Thus $\boldsymbol{\Xi}$ is invertible between the indicated spaces, and hence also as a pseudodifferential operator. The fact that its inverse belongs to the set $\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM\oplus\mathbb{C})$ follows from a “Beals-type theorem” (Theorem 7.4 in [67]). The last stated isomorphism follows from the mapping properties of operators in the spaces $\Psi_{\operatorname{ess}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M;TM)$ and in $\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM)$, as in the last corollaries. This completes the proof. ∎
The invertibility of $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ in the case $M$ closed is obtained when both $V$ and $V_{0}$ do not vanish identically on $M$. Moreover, the condition that $V$ does not vanish identically on $M$ can be replaced with the condition that $M$ does not have non-vanishing Killing vector fields. These results were proved in [69, Theorem 6.1(4)]. In relation to this result, the reader might have noticed that we are not considering the case when $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ has straight cylindrical ends but does not have non-vanishing Killing vector fields. In fact, this condition is not very useful for manifolds with straight cylindrical ends because, even if $M$ does not have non-vanishing Killing vector fields, its “end” $M^{\prime}\times\mathbb{R}$ will always have non-vanishing Killing vector fields, for instance $\partial_{t}$ (the infinitesimal generator of the translations along $\mathbb{R}$). (See Subsection 12.2 for more on the case $M$ closed.)
8.3. Definition of layer potentials on $M$
The Fredholm result of the previous section allow us to define and study several layer potential type operators. We continue to assume that $M$ has straight cylindrical ends. We also assume that $V$ and $V_{0}$ satisfy the positivity at infinity Assumption 8.1, in order to be able to use the Fredholmness results of the previous section. (The case $M$ closed was treated similarly in great detail in [69].) Recall that the Moore-Penrose pseudoinverse $\boldsymbol{\Xi}^{(-1)}$ of $\boldsymbol{\Xi}$ (which was proved to be Fredholm in Theorem 7.14) is defined as the inverse of the operator $\boldsymbol{\Xi}$ acting from the orthogonal of $\ker\boldsymbol{\Xi}$ to its image (see [69] for details and references). As in [69], we also obtain the following result:
Proposition 8.5.
Let $M$ be a manifold with straight cylindrical ends. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$, as in Proposition 6.13. We assume that $V$ and $V_{0}$ satisfy the non-vanishing Assumption 8.3. Then
  1. (i)
    The pseudoinverse $\boldsymbol{\Xi}^{(-1)}$ has the form
    $\boldsymbol{\Xi}^{(-1)}\,=:\,\left(\begin{array}[]{cc}\mathcal{A}&\mathcal{B}\\ \mathcal{C}&\mathcal{D}\end{array}\right)\in\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM\oplus\mathbb{C})\,$,
    where $\mathcal{A}\in\Psi_{\operatorname{ess}}^{-2}(M;TM)$, $\mathcal{C}=\mathcal{B}^{*}\in\Psi_{\operatorname{ess}}^{-1}(M;TM,\mathbb{C})$, and $\mathcal{D}\in\Psi_{\operatorname{ess}}^{0}(M)$.
  2. (ii)
    Consequently, we have $\sigma_{-2}(\mathcal{A})(x,\xi)\,=\,\frac{1}{|\xi|^{2}}-\frac{V_{0}+1}{2V_{0}+1}\frac{1}{|\xi|^{4}}\xi^{\sharp}\otimes\xi\,$, $\sigma_{-1}(\mathcal{B})(x,\xi)\,=\,\sigma_{-1}(\mathcal{C})(x,\xi)^{*}\,=\,\frac{\imath}{(2V_{0}+1)|\xi|^{2}}\xi^{\sharp}\,$, and $\sigma_{0}(\mathcal{D})(x,\xi)\,=\,-\frac{2}{2V_{0}+1}.$
Proof.
The point (i) makes explicit (using the definition of Douglis-Nirenberg-operators) the fact that $\boldsymbol{\Xi}^{-1}\in\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM\oplus\mathbb{C})$, a result that was already proved in Theorem 8.4. Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$ be as in the statement. The multiplicativity of the principal symbol $\operatorname{Symb}$} gives that
$\quad{\operatorname{Symb}}_{\boldsymbol{s},\boldsymbol{t}}(\boldsymbol{\Xi}){\operatorname{Symb}}_{\bf-t,-s}(\boldsymbol{\Xi}^{(-1)})\,=\,{\operatorname{Symb}}_{\boldsymbol{0},0}(1)\,=\,1\,.$
Therefore, the $({-\boldsymbol{s}},{-\boldsymbol{t}})$-principal symbol of the inverse $\boldsymbol{\Xi}^{(-1)}$ of $\boldsymbol{\Xi}$ is the inverse of the $(\boldsymbol{s},\boldsymbol{t})$-principal symbol of $\boldsymbol{\Xi}$, which is given by Proposition 6.13. Thus, the principal symbols of the operators $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$, and $\mathcal{D}$ (the entries of $\boldsymbol{\Xi}^{(-1)}$) are as stated (as given by Proposition 6.13). ∎
The existence of the inverse $\boldsymbol{\Xi}^{(-1)}$ of $\boldsymbol{\Xi}$ allows us to use classical methods to define the single and double layer potential operators for the Stokes operator (see also [39, 76, 99, 134]). We let $\Omega\subset M$ be a subdomain with straight cylindrical ends, that is
$\Omega\cap\big[M^{\prime}\times(-\infty,0)\big]\,=\,\Omega^{\prime}\times(-\infty,0)\,$,
as before (recall that such an open subset is called a compatible open subset of $M$, see Equation (1.2)). Let $\Gamma:=\partial\Omega$, as usual in this paper. Thus $\Gamma$ is also a manifold with straight cylindrical ends. For the following definition, recall the Stokes operator $\boldsymbol{\Xi}=\boldsymbol{\Xi}_{V,V_{0}}$ (1.3). Also, recall the distribution $\boldsymbol{h}\otimes\delta_{\Gamma}$ of Lemma 3.8 and the operator $\boldsymbol{T}_{\boldsymbol{\nu}}^{*}$ of Lemma 2.1. We now introduce the layer potential operators associated to our generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$.
Definition 8.6.
Let $\boldsymbol{h}\in H^{s}(\Gamma;TM)$, $s\in\mathbb{R}$. The single-layer potential $\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$ for $\boldsymbol{\Xi}$ is then
$\mathcal{S}_{\rm{ST}}(\boldsymbol{h})\mathrel{\,:=\,}\boldsymbol{\Xi}^{(-1)}\,\left[\left(\begin{array}[]{c}{\boldsymbol{h}}\\ {0}\end{array}\right)\otimes\delta_{\Gamma}\right]\,.$
Similarly, the double-layer potential $\mathcal{D}_{\rm{ST}}(\boldsymbol{h})$ for $\boldsymbol{\Xi}$ is given by:
$\mathcal{D}_{\rm{ST}}(\boldsymbol{h})\mathrel{\,:=\,}\boldsymbol{\Xi}^{(-1)}\,\big[\boldsymbol{T}_{\boldsymbol{\nu}}^{*}\left(\boldsymbol{h}\otimes\delta_{\Gamma}\right)\big]\,.$
It is useful sometimes to separate the components of $\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$ and $\mathcal{D}_{\rm{ST}}(\boldsymbol{h})$ into velocity and pressure components.
Definition 8.7.
We continue to assume that $\boldsymbol{h}\in H^{s}(\Gamma;TM)$, $s\in\mathbb{R}$. The components of these definitions yield the single-layer velocity potential $\mathcal{V}_{\rm{ST}}(\boldsymbol{h})$ and the single-layer pressure potential $\mathcal{P}_{\rm{ST}}(\boldsymbol{h})$ for $\boldsymbol{\Xi}$:
$\left(\begin{array}[]{c}\mathcal{V}_{\rm{ST}}(\boldsymbol{h})\\ \mathcal{P}_{\rm{ST}}(\boldsymbol{h})\end{array}\right)\mathrel{\,:=\,}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})\mathrel{\,:=\,}\boldsymbol{\Xi}^{(-1)}\,\left[\left(\begin{array}[]{c}{\boldsymbol{h}}\\ {0}\end{array}\right)\otimes\delta_{\Gamma}\right]\,.$
Similarly, we obtain the double-layer velocity potential $\mathcal{W}_{\rm{ST}}(\boldsymbol{h})$ and the double-layer pressure potential $\mathcal{Q}_{\rm{ST}}(\boldsymbol{h})$ for $\boldsymbol{\Xi}$
$\left(\begin{array}[]{c}\mathcal{W}_{\rm{ST}}(\boldsymbol{h})\\ \mathcal{Q}_{\rm{ST}}(\boldsymbol{h})\end{array}\right)\mathrel{\,:=\,}\mathcal{D}_{\rm{ST}}(\boldsymbol{h})\mathrel{\,:=\,}\boldsymbol{\Xi}^{(-1)}\,\big[\boldsymbol{T}_{\boldsymbol{\nu}}^{*}\left(\boldsymbol{h}\otimes\delta_{\Gamma}\right)\big]\,.$
These definitions can be made more explicitly as follows.
Remark 8.8.
We have
$\mathcal{W}_{\rm{ST}}(\boldsymbol{h})\,=\,\left(\begin{array}[]{cc}\mathcal{A}&\mathcal{B}\end{array}\right)\left(\begin{array}[]{c}-2\boldsymbol{D}_{\boldsymbol{\nu}}^{*}\\ \boldsymbol{\nu}^{\sharp}\end{array}\right)(\boldsymbol{h}\otimes\delta_{\Gamma})\,=\,(-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp})\,(\boldsymbol{h}\otimes\delta_{\Gamma})$
and
$\mathcal{Q}_{\rm{ST}}(\boldsymbol{h})\,=\,\left(\begin{array}[]{cc}\mathcal{C}&\mathcal{D}\end{array}\right)\left(\begin{array}[]{c}-2\boldsymbol{D}_{\boldsymbol{\nu}}^{*}\\ \boldsymbol{\nu}^{\sharp}\end{array}\right)(\boldsymbol{h}\otimes\delta_{\Gamma})\\ \ \,=\,(-2\mathcal{C}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{D}\boldsymbol{\nu}^{\sharp})\,(\boldsymbol{h}\otimes\delta_{\Gamma})\,.$
Similarly, $\mathcal{S}_{\rm{ST}}(\boldsymbol{h})\,=\,\big(\,\mathcal{A}(\boldsymbol{h}\otimes\delta_{\Gamma})\ \ \ \mathcal{C}(\boldsymbol{h}\otimes\delta_{\Gamma})\,\big)^{\top}$.
Theorem 6.6 gives the following result.
Proposition 8.9.
We continue to assume that $M$ is a manifold with straight cylindrical ends and that $\Omega\subset M$ is a compatible open subset (Equation (1.2)). Let $\Omega_{+}:=\Omega$, $\Omega_{-}:=M\smallsetminus\overline{\Omega}$, and $\Gamma:=\partial\Omega$ be as before. Let $s\in\mathbb{R}$ and $\boldsymbol{h}\in H^{s}(\Gamma;TM)$. Then
$\mathcal{V}_{\rm{ST}}(\boldsymbol{h})|_{\Omega_{\pm}}\in H^{s+\frac{3}{2}}(\Omega_{\pm};TM)\ \mbox{ and }\ \mathcal{P}_{\rm{ST}}(\boldsymbol{h})|_{\Omega_{\pm}}\in H^{s+\frac{1}{2}}(\Omega_{\pm})\,.$
Similarly, let $s\in\mathbb{R}$ and $\boldsymbol{h}\in H^{s}(\Gamma;TM)$. Then
$\mathcal{W}_{\rm{ST}}(\boldsymbol{h})|_{\Omega_{\pm}}\in H^{s+\frac{1}{2}}(\Omega_{\pm};TM)\ \mbox{ and }\ \mathcal{Q}_{\rm{ST}}(\boldsymbol{h})|_{\Omega_{\pm}}\in H^{s-\frac{1}{2}}(\Omega_{\pm})\,.$
Proof.
Indeed, this follows from Theorem 6.6 and Remark 8.8 because $\mathcal{A}$ has order 2-2, $\mathcal{C}$ and $\boldsymbol{P}:=-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp}$ have order 1-1, and $2\mathcal{C}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{D}\boldsymbol{\nu}^{\sharp}$ has order zero. Moreover, all these four operators have rational type symbols.
We obtain in this way the eight stated relations (including the choices of +/+/-). Just to make things clearer, let us show in detail how the first relation is proved. Theorem 6.6 for m=2m=-2, $s\in\mathbb{R}$, $P=\mathcal{A}$, gives
$\mathcal{V}_{\rm{ST}}(\boldsymbol{h})|_{\Omega_{+}}\mathrel{\,:=\,}\mathcal{A}(\boldsymbol{h}\otimes\delta_{\Gamma})|_{\Omega_{+}}\in H^{s-(-2)-\frac{1}{2}}(\Omega_{+})\,=\,H^{s+\frac{3}{2}}(\Omega_{+})\,.$
This proves the first of the eight mentioned relations. The other seven are proved in exactly the same way. ∎
The following result is a consequence of the definition of the single and double layer potentials. Recall that $\mathcal{N}$ is the kernel of $\boldsymbol{\Xi}$ and $p_{\mathcal{N}}$ is the $L^{2}$-orthogonal projection onto $\mathcal{N}$.
Proposition 8.10.
Let us assume that $V$ and $V_{0}$ satisfy the positivity assumption on $M$ (Assumption 8.3, and hence that $\boldsymbol{\Xi}$ is invertible). Let $\boldsymbol{h}\in H^{s}(\Gamma;TM)$, $s\in\mathbb{R}$. Then $\boldsymbol{\Xi}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$ and $\boldsymbol{\Xi}\mathcal{D}_{\rm{ST}}(\boldsymbol{h})$ vanish on $M\smallsetminus\Gamma$.
Proof.
We have
$\boldsymbol{\Xi}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})\,=\,\boldsymbol{\Xi}\boldsymbol{\Xi}^{-1}(\boldsymbol{h}\otimes\delta_{\Gamma})\,=\,\boldsymbol{h}\otimes\delta_{\Gamma}\,$
which vanishes outside $\Gamma$, as claimed. ∎
We notice that in [69] we have considered also the case when $\boldsymbol{\Xi}$ is only Fredholm, when some additional compatibility relations need to be added to obtain the vanishing outside $\Gamma$ (respectively, $\Gamma^{\prime}$) in the above result. We shall need the following consequences of the representation formula in Proposition 2.11, the first one of which we will call Pompeiu’s formula. Recall that $f_{+}$ denotes the trace of $f\in H^{1}(\Omega)$ at $\Gamma:=\partial\Omega$ (thus the trace from the interior).
Proposition 8.11.
Let us assume that $V$ and $V_{0}$ satisfy the positivity assumption on $M$ (Assumption 8.3, and hence that $\boldsymbol{\Xi}$ is invertible). Let $U$ be such that $U|_{\Omega}\in H^{2}(\Omega;TM)\oplus H^{1}(\Omega)$. We extend $U$ to be 0 outside $\Omega$, and call the extension ${\bf 1}_{\Omega}U$, where ${\bf 1}_{\Omega}$ be the characteristic function of $\Omega$ (as before).
  1. (i)
    ${\bf 1}_{\Omega}U\,=\,\boldsymbol{\Xi}^{-1}\left({\bf 1}_{\Omega}\big(\boldsymbol{\Xi}U\big)\right)-\mathcal{S}_{\rm{ST}}\big[(\boldsymbol{T}_{\boldsymbol{\nu}}U)_{+}\big]+\mathcal{D}_{\rm{ST}}(\boldsymbol{u}_{+})\,.$
  2. (ii)
    Consequently, if we also have $\boldsymbol{\Xi}U=0$ in $\Omega$, then
    $\mathcal{D}_{\rm{ST}}(\boldsymbol{u}_{+})(x)-\mathcal{S}_{\rm{ST}}\big[(\boldsymbol{T}_{\boldsymbol{\nu}}U)_{+}\big](x)\,=\,\begin{cases}\ U(x)&\mbox{ if }x\in\Omega\\ \ \ \,0&\mbox{ if }x\in M\smallsetminus\overline{\Omega}\,.\end{cases}$
Proof.
Recall the representation formulas of Proposition 2.11(iii)
$\boldsymbol{\Xi}\big(\textbf{1}_{\Omega}U\big)$ $\,=\,\textbf{1}_{\Omega}\big(\boldsymbol{\Xi}U\big)-(\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U)_{+}\otimes\delta_{\partial\Omega}+\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}^{*}(U_{+}\otimes\delta_{\partial\Omega})$
$\,=\,\textbf{1}_{\Omega}\big(\boldsymbol{\Xi}U\big)-\left(\begin{array}[]{c}{(\boldsymbol{T}_{\boldsymbol{\nu}}U)_{+}}\\ {0}\end{array}\right)\otimes\delta_{\partial\Omega}+\boldsymbol{T}_{\boldsymbol{\nu}}^{*}(\boldsymbol{u}_{+}\otimes\delta_{\partial\Omega})\,.$
By applying $\boldsymbol{\Xi}^{-1}$ to this formula, we obtain Pompeiu’s formula:
(8.5) $\textbf{1}_{\Omega}U$ $\,=\,\boldsymbol{\Xi}^{-1}\left(\textbf{1}_{\Omega}\big(\boldsymbol{\Xi}U\big)-\left(\begin{array}[]{c}{(\boldsymbol{T}_{\boldsymbol{\nu}}U)_{+}}\\ {0}\end{array}\right)\otimes\delta_{\Gamma}+\boldsymbol{T}_{\boldsymbol{\nu}}^{*}\,(\boldsymbol{u}_{+}\otimes\delta_{\Gamma})\right)$
$\,=\,\boldsymbol{\Xi}^{-1}\left(\textbf{1}_{\Omega}\big(\boldsymbol{\Xi}U\big)\right)-\mathcal{S}_{\rm{ST}}({(\boldsymbol{T}_{\boldsymbol{\nu}}U)_{+}})+\mathcal{D}_{\rm{ST}}(\boldsymbol{u}_{+})\,.$
The second part follows immediately from Pompeiu’s formula 8.5. ∎
8.4. Jump relations
As usual, the limit and jump relations of our layer potentials will play a crucial role. They are a delicate and difficult topic, especially for the Stokes operator. We now establish the needed limit and jump relations following Subsection 2.1 and, especially, Section 6 to establish some needed jump relations for the needed components of our potential operators $\mathcal{S}_{\rm{ST}}$ and $\mathcal{D}_{\rm{ST}}$ (see Remark 8.8). We continue to assume that $M$ is with straight cylindrical ends and that $V$ and $V_{0}$ satisfy the positivity assumption at infinity (Assumption 8.1), which allows us to conclude that $\boldsymbol{\Xi}$ is Fredholm and hence that the layer potentials are defined. Close versions of the relations established in this section are contained also in [63] and [69]. They were also proved by Mitrea and Taylor [99] and Dindos̆ and Mitrea [39] with a different approach.
Proposition 8.12.
Let $\boldsymbol{P}:=-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp}$ and denote $\mathfrak{f}=(V_{0}+1)/(2V_{0}+1)$ and $\mathfrak{g}=1/(2V_{0}+1)$, as before. Then
$\sigma_{-1}(\boldsymbol{P};\xi)\,=:\,a(\xi)\,=\,\frac{\imath}{|\xi|^{2}}\left(\xi(\boldsymbol{\nu})+\boldsymbol{\nu}\otimes\xi-\frac{2\mathfrak{f}\xi(\boldsymbol{\nu})}{|\xi|^{2}}\xi^{\sharp}\otimes\xi+\mathfrak{g}\xi^{\sharp}\otimes\boldsymbol{\nu}^{\sharp}\right)\,.$
Consequently, $\operatorname{\mathfrak{L}}=-\imath$ for this operator.
Proof.
The calculations are local, so we can assume that $\Omega=\mathbb{R}_{+}^{n}$. Then $\boldsymbol{\nu}=-e_{n}=(0,\ldots,0,-1)$. We shall write, $\xi=\xi^{\prime}+t\boldsymbol{\nu}$. Substituting $\xi=\xi^{\prime}+te_{n}$, Proposition 8.5 gives
$\begin{gathered}\sigma_{-2}(\mathcal{A};\xi)\,=\,\frac{1}{|\xi|^{2}}-\frac{\mathfrak{f}}{|\xi|^{4}}\xi^{\sharp}\otimes\xi\,,\qquad\sigma_{-1}(\mathcal{B})(x,\xi)\,=\,\frac{\imath\mathfrak{g}}{|\xi|^{2}}\xi^{\sharp}\,,\\ \mbox{and }\quad\sigma_{1}(\boldsymbol{D}_{\boldsymbol{\nu}}^{*};\xi)\,=\,-\frac{\imath}{2}\big[\xi(\boldsymbol{\nu})+\boldsymbol{\nu}\otimes\xi\big]\,=\,\frac{\imath}{2}\big[t+e_{n}\otimes\xi\big]\,.\end{gathered}$
Using that $\boldsymbol{\nu}^{\sharp}=-e_{n}^{\sharp}$, we can compute
$\sigma_{-1}(\boldsymbol{P};\xi)$ $\,=\,-2\sigma_{-2}(\mathcal{A};\xi)\sigma_{1}(\boldsymbol{D}_{\boldsymbol{\nu}}^{*};\xi)+\sigma_{-1}(\mathcal{B};\xi)\boldsymbol{\nu}^{\sharp}$
$\,=\,-2\left(\frac{1}{|\xi|^{2}}-\frac{\mathfrak{f}}{|\xi|^{4}}\xi^{\sharp}\otimes\xi\right)\frac{\imath}{2}(t+e_{n}\otimes\xi)-\frac{\imath\mathfrak{g}}{|\xi|^{2}}\xi^{\sharp}\otimes e_{n}^{\sharp}$
$\,=\,-\frac{\imath}{|\xi|^{2}}\big(t+e_{n}\otimes\xi-\frac{2\mathfrak{f}t}{|\xi|^{2}}\xi^{\sharp}\otimes\xi+\mathfrak{g}\xi^{\sharp}\otimes e_{n}^{\sharp}\big)\,.$
To obtain the “jump” coefficient $\operatorname{\mathfrak{L}}$, we expand the formula for the principal symbol from the last equation according to the highest powers of $t$, using $\xi=\xi^{\prime}+te_{n}^{\sharp}$, with $\xi^{\prime}=(\xi_{1},\ldots,\xi_{n-1},0)$, to obtain
$\lim_{t\to\pm\infty}t\sigma_{-1}(\boldsymbol{P};\xi)$ $\,=\,-\lim_{t\to\pm\infty}\frac{\imath t}{|\xi|^{2}}\big(t+e_{n}\otimes\xi-\frac{2\mathfrak{f}t}{|\xi|^{2}}\xi^{\sharp}\otimes\xi+\mathfrak{g}\xi^{\sharp}\otimes{e_{n}^{\sharp}}\big)$
$\,=\,-\lim_{t\to\pm\infty}\frac{\imath t^{2}}{|\xi|^{2}}\big(1+e_{n}\otimes e_{n}^{\sharp}-\frac{2\mathfrak{f}t^{2}}{|\xi|^{2}}e_{n}\otimes e_{n}^{\sharp}+\mathfrak{g}e_{n}\otimes e_{n}^{\sharp}\big)$
$\,=\,-\imath\big(1+e_{n}\otimes e_{n}^{\sharp}-2\mathfrak{f}e_{n}\otimes e_{n}^{\sharp}+\mathfrak{g}e_{n}\otimes e_{n}^{\sharp}\big)$
$\,=\,-\imath\big[1+(1-2\mathfrak{f}+\mathfrak{g})e_{n}\otimes e_{n}^{\sharp}\big]\,=\,-\imath\,$,
because $1-2\mathfrak{f}+\mathfrak{g}=0$. ∎
The following integrals can be obtained by using the Residue Theorem (the details can be found in the proof of Lemma 16.3.2 of [63], see also [60]).
Lemma 8.13.
Let $a>0$. We have
$\int_{\mathbb{R}}\frac{x^{2}dx}{(a^{2}+x^{2})^{2}}\,=\,\displaystyle\frac{\pi}{2a}\,,\quad\displaystyle\int_{\mathbb{R}}\frac{dx}{(a^{2}+x^{2})^{2}}\,=\,\displaystyle\frac{\pi}{2a^{3}}\,,\quad\mbox{and}\quad\displaystyle\int_{\mathbb{R}}\frac{dx}{a^{2}+x^{2}}\,=\,\displaystyle\frac{\pi}{a}\,.$
For the rest of the paper, we shall let $\boldsymbol{P}:=-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp}$ be the pseudodifferential operator defining the vector part $\mathcal{W}_{\rm{ST}}$ of the double layer potential (Definition 8.6) and let
(8.6) $\boldsymbol{K}\mathrel{\,:=\,}\boldsymbol{P}_{0}\mathrel{\,:=\,}\big(-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp})_{0}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$
be the restriction of $\boldsymbol{P}:=-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp}$ to $\Gamma:=\partial\Omega$ (see Theorem 6.8).
Theorem 8.14.
Let $\boldsymbol{K}:=\boldsymbol{P}_{0}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ be as in Equation (8.6). Then
$\mathcal{W}_{\rm{ST}}(\boldsymbol{h})_{\pm}\mathrel{\,:=\,}\big[\boldsymbol{P}(\boldsymbol{h}\otimes\delta_{\Gamma})\big]_{\pm}\,=\,\left[\pm\frac{1}{2}+\boldsymbol{K}\right]\boldsymbol{h}\,$,
where $\sigma_{0}(\boldsymbol{K};\xi^{\prime})=\frac{\imath V_{0}}{2(2V_{0}+1)|\xi^{\prime}|}\big(\boldsymbol{\nu}\otimes\xi^{\prime}-\xi^{{}^{\prime}\sharp}\otimes\boldsymbol{\nu}^{\sharp}\big)$. In particular, the two operators $\pm\frac{1}{2}+\boldsymbol{K}$ are elliptic and have self-adjoint principal symbols.
Proof.
Let $\mathfrak{f}=(V_{0}+1)/(2V_{0}+1)$ and $\mathfrak{g}=1/(2V_{0}+1)$ be as in Proposition 8.12. As in the proof of Theorem 6.8, we use local coordinates such that $\boldsymbol{\nu}=-e_{n}$. We need only to identify $\sigma_{0}(\boldsymbol{K};\xi^{\prime})=\sigma_{0}(\boldsymbol{P}_{0};\xi^{\prime})$ (the rest follows from Theorem 6.8 and Proposition 8.12). To that end, we separate the terms that are even in $t$ in $\sigma_{-1}(\boldsymbol{P};\xi)$. For instance, recalling that $\xi=\xi^{\prime}+te_{n}$, the even part of $\xi^{\sharp}\otimes\xi$ is $\xi^{{}^{\prime}\sharp}\otimes\xi^{\prime}+t^{2}e_{n}\otimes e_{n}^{\sharp}$, whereas its odd part is $t(e_{n}\otimes\xi^{\prime}+\xi^{{}^{\prime}\sharp}\otimes e_{n}^{\sharp})$. This gives
$b(\xi^{\prime},t)\mathrel{\,:=\,}\sigma_{-1}(\boldsymbol{P};\xi)+\sigma_{-1}(\boldsymbol{P};\xi^{\prime},-t)\\ \,=\,-\frac{2\imath}{|\xi|^{2}}\Big[e_{n}\otimes\xi^{\prime}+\mathfrak{g}\xi^{{}^{\prime}\sharp}\otimes e_{n}^{\sharp}-\frac{2\mathfrak{f}t^{2}}{|\xi|^{2}}(e_{n}\otimes\xi^{\prime}+\xi^{{}^{\prime}\sharp}\otimes e_{n}^{\sharp})\Big]\\ \,=\,-2\imath\Big[\left(\frac{1}{|\xi|^{2}}-\frac{2\mathfrak{f}t^{2}}{|\xi|^{4}}\right)e_{n}\otimes\xi^{\prime}+\left(\frac{\mathfrak{g}}{|\xi|^{2}}-\frac{2\mathfrak{f}t^{2}}{|\xi|^{4}}\right)\xi^{{}^{\prime}\sharp}\otimes e_{n}^{\sharp}\Big]\,.$
We next use Theorem 6.8 and the relations
$\int_{\mathbb{R}}\frac{1}{|\xi|^{2}}\,dt\,=\,\frac{\pi}{|\xi^{\prime}|}\ \mbox{ and }\ \int_{\mathbb{R}}\frac{t^{2}}{|\xi|^{4}}\,dt\,=\,\frac{\pi}{2|\xi^{\prime}|}\,,\quad\xi^{\prime}\neq 0\,$,
(see Lemma 8.13) to obtain
$\sigma_{0}(\boldsymbol{K};\xi^{\prime})$ $\,=\,\frac{1}{4\pi}\int_{\mathbb{R}}b(\xi^{\prime},t)\,dt$
$\,=\,-\frac{\imath}{2\pi}\int_{\mathbb{R}}\Big[\left(\frac{1}{|\xi|^{2}}-\frac{2\mathfrak{f}t^{2}}{|\xi|^{4}}\right)e_{n}\otimes\xi^{\prime}+\left(\frac{\mathfrak{g}}{|\xi|^{2}}-\frac{2\mathfrak{f}t^{2}}{|\xi|^{4}}\right)\xi^{{}^{\prime}\sharp}\otimes e_{n}^{\sharp}\Big]\,dt$
$\,=\,\frac{\imath V_{0}}{2(2V_{0}+1)|\xi^{\prime}|}\big(\xi^{{}^{\prime}\sharp}\otimes e_{n}^{\sharp}-e_{n}\otimes\xi^{\prime}\big)$
$\,=\,\frac{\imath V_{0}}{2(2V_{0}+1)|\xi^{\prime}|}\big(\boldsymbol{\nu}\otimes\xi^{\prime}-\xi^{{}^{\prime}\sharp}\otimes\boldsymbol{\nu}^{\sharp}\big)\,.$
This explicit formula gives that $\sigma_{0}(\boldsymbol{K})^{*}=\sigma_{0}(\boldsymbol{K})$. To complete the proof, we notice that the non-zero eigenvalues $\lambda=\pm\frac{V_{0}}{2(2V_{0}+1)}$ of $\sigma_{0}(\boldsymbol{K};\xi^{\prime})$ satisfy $|\lambda|<1/4$, and hence the operators $\pm\frac{1}{2}+\boldsymbol{K}$ are elliptic. ∎
Remark 8.15.
Let us notice that $\boldsymbol{K}=\boldsymbol{P}_{0}\in\Psi^{0}(\Gamma;TM)$ is of order 1-1 if, and only if, $V_{0}=0$ on $\Gamma$ (see Equation (8.6) and Theorem 8.14 above).
Let $\mathfrak{f}=(V_{0}+1)/(2V_{0}+1)$ and $\mathfrak{g}=1/(2V_{0}+1)$, as in Proposition 8.12. Also, recall the pseudodifferential operators $\mathcal{A}$ and $\mathcal{C}$ introduced in Proposition 8.5 as two of the matrix components of $\boldsymbol{\Xi}^{(-1)}$. (These operators were used then in the definition of the single layer potential $\mathcal{S}_{\rm{ST}}$.) We obtain the following result.
Theorem 8.16.
Let $\boldsymbol{h}\in L^{2}(\Gamma;TM)$, $\Gamma:=\partial\Omega$, $\mathfrak{f}:=\frac{V_{0}+1}{2V_{0}+1}$ and $\mathfrak{g}=\frac{1}{2V_{0}+1}$. Then
  1. (i)
    $\mathcal{V}_{\rm{ST}}(\boldsymbol{h})_{\pm}\,=\,\mathcal{A}_{0}(\boldsymbol{h})$ and $\sigma_{-1}(\mathcal{A}_{0};\xi^{\prime})\,=\,\frac{1}{4|\xi^{\prime}|}(2-\mathfrak{f}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp}-\mathfrak{f}\eta^{\sharp}\otimes\eta)\,$, where $\eta:=|\xi^{\prime}|^{-1}\xi^{\prime}$. If $V_{0}\geq 0$, then $\mathcal{A}_{0}$ is elliptic with self-adjoint symbol.
  2. (ii)
    $\mathcal{P}_{\rm{ST}}(\boldsymbol{h})_{\pm}=\left(\mp\displaystyle\frac{\mathfrak{g}}{2}\boldsymbol{\nu}^{\sharp}+\mathcal{C}_{0}\right)\boldsymbol{h}$, where $\sigma_{0}(\mathcal{C}_{0};\xi^{\prime})=-\displaystyle\frac{\imath\mathfrak{g}}{2|\xi^{\prime}|}\xi^{\prime}.$
  3. (iii)
    $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})]_{\pm}\,=\,\left(\mp\frac{1}{2}+{\boldsymbol{K}}^{*}\right)\boldsymbol{h}\,$, where $\boldsymbol{K}=\boldsymbol{P}_{0}:=(-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp})_{0}$ is defined as in Theorem 8.14.
Proof.
Recall that, for $\xi\in T_{x}^{*}M$, the linear map $\xi^{\sharp}\otimes\xi\in\operatorname{End}(T_{x}M)$ is defined by $(\xi^{\sharp}\otimes\xi)(v):=\xi(v)\xi^{\sharp}$. In particular, $\sigma_{-2}(\mathcal{A},\xi)=\frac{1}{|\xi|^{4}}\big(|\xi|^{2}-\mathfrak{f}\xi^{\sharp}\otimes\xi\big)$ (recall that $\mathfrak{f}:=\frac{V_{0}+1}{2V_{0}+1}$). For $\xi\in T^{*}M$, let us write $\xi=\xi^{\prime}+t\boldsymbol{\nu}^{\sharp}$, with $\xi^{\prime}(\boldsymbol{\nu})=0$ and we identify $\xi^{\prime}$ with an element of $T^{*}\Gamma$, when $\xi^{\prime}\in T_{x}^{*}M$ with $x\in\Gamma$. To prove the first equality, we use Theorem 6.5(ii) (see also Theorem 4.13) and then Proposition 8.5 to obtain the following relation
$\sigma_{-1}(\mathcal{A}_{0};\xi^{\prime})$ $\,=\,\frac{1}{2\pi}\int_{\mathbb{R}}\sigma_{-2}(\mathcal{A};\xi)\,dt$
$\,=\,\frac{1}{2\pi}\int_{\mathbb{R}}\frac{1}{|\xi|^{4}}\big(|\xi|^{2}-\mathfrak{f}\xi^{\sharp}\otimes\xi\big)\,dt$
$\,=\,\frac{1}{2\pi}\int_{\mathbb{R}}\frac{1}{|\xi|^{4}}\big[|\xi|^{2}-\mathfrak{f}(\xi^{\prime\sharp}\otimes\xi^{\prime}+t\boldsymbol{\nu}\otimes\xi^{\prime}+t\xi^{\prime\sharp}\otimes\boldsymbol{\nu}^{\sharp}+t^{2}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp})\big]\,dt$
$\,=\,\frac{1}{2\pi}\int_{\mathbb{R}}\frac{1}{|\xi|^{4}}\big[|\xi|^{2}-\mathfrak{f}(\xi^{\prime\sharp}\otimes\xi^{\prime}+t^{2}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp})\big]\,dt$
$\,=\,\frac{1}{2\pi}\int_{\mathbb{R}}\left(\frac{1}{|\xi^{\prime}|^{2}+t^{2}}-\frac{\mathfrak{f}}{(|\xi^{\prime}|^{2}+t^{2})^{2}}\xi^{\prime\sharp}\otimes\xi^{\prime}-\frac{\mathfrak{f}t^{2}}{(|\xi^{\prime}|^{2}+t^{2})^{2}}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp}\right)\,dt$
$\,=\,\frac{1}{2\pi}\left(\frac{\pi}{|\xi^{\prime}|}-\frac{\mathfrak{f}\pi}{2|\xi^{\prime}|^{3}}\xi^{\prime\sharp}\otimes\xi^{\prime}-\frac{\mathfrak{f}\pi}{2|\xi^{\prime}|}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp}\right)$
$\,=\,\frac{1}{2|\xi^{\prime}|}-\frac{\mathfrak{f}}{4|\xi^{\prime}|^{3}}\xi^{\prime\sharp}\otimes\xi^{\prime}-\frac{\mathfrak{f}}{4|\xi^{\prime}|}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp}$
$\,=\,\frac{1}{4|\xi^{\prime}|}\left(2-\mathfrak{f}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp}-\mathfrak{f}\eta^{\sharp}\otimes\eta\right)\,.$
This proves (i). The operator $\mathcal{A}_{0}$ is elliptic because $\mathfrak{f}<1$, and hence $\sigma_{-1}(\mathcal{A}_{0})$ is invertible.
For the relation (ii), we use the identity $\sigma_{-1}(\mathcal{C};\xi)=-\frac{g\imath}{|\xi|^{2}}\xi$ (see Proposition 8.5) and then Theorem 6.8 (see also Theorem 4.16). We also write $\xi=\xi^{\prime}+t\boldsymbol{\nu}^{\sharp}$ with $\xi^{\prime}\perp\boldsymbol{\nu}$, as in the proof of the previous point. Then we notice that the even part of $\sigma_{-1}(\mathcal{C};\xi)$ (in $\tau$) is $\frac{g\imath}{|\xi|^{2}}\xi^{\prime}$. Therefore
$\sigma_{0}(\mathcal{C}_{0};\xi^{\prime})\,=\,-\frac{1}{2\pi}\int_{\mathbb{R}}\frac{g\imath}{|\xi|^{2}}\xi^{\prime}\,dt\,=\,-\frac{g\imath}{2\pi}\left(\int_{\mathbb{R}}\frac{1}{|\xi|^{2}}\,dt\right)\xi^{\prime}\,=\,-\frac{g\imath}{2|\xi^{\prime}|}\xi^{\prime}\,.$
The “jump part” is also obtained from the principal symbol of $\mathcal{C}$, namely, it is $\mp\frac{\imath}{2}\sigma_{-1}(\mathcal{C};\boldsymbol{\nu}^{\sharp})=\mp\frac{g}{2}\boldsymbol{\nu}^{\sharp}$.
Let us now prove the third relation. We have $\boldsymbol{\Xi}^{*}=\boldsymbol{\Xi}$, and hence $\big(\boldsymbol{\Xi}^{(-1)}\big)^{*}=\boldsymbol{\Xi}^{(-1)}$. Theorem 6.8 and Proposition 8.12 then give $\operatorname{\mathfrak{L}}(\boldsymbol{T}_{\boldsymbol{\nu}}\boldsymbol{\Xi}^{(-1)})=\operatorname{\mathfrak{L}}(\boldsymbol{\Xi}^{(-1)}\boldsymbol{T}_{\boldsymbol{\nu}}^{*})^{*}=\overline{(-\imath)}=\imath$. Moreover, $(\boldsymbol{T}_{\boldsymbol{\nu}}\boldsymbol{\Xi}^{(-1)})_{0}=(\boldsymbol{\Xi}^{(-1)}\boldsymbol{T}_{\boldsymbol{\nu}}^{*})_{0}^{*}=\boldsymbol{K}^{*}$. This then gives the following relation:
$\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})]_{\pm}\,=\,[\boldsymbol{T}_{\boldsymbol{\nu}}\boldsymbol{\Xi}^{(-1)}(\boldsymbol{h}\otimes\delta_{\Gamma})]_{\pm}\,=\,\left(\pm\frac{\imath}{2}\operatorname{\mathfrak{L}}(\boldsymbol{T}_{\boldsymbol{\nu}}\boldsymbol{\Xi}^{(-1)})+(\boldsymbol{T}_{\boldsymbol{\nu}}\boldsymbol{\Xi}^{(-1)})_{0}\right)\boldsymbol{h}\\ \,=\,\left(\pm\frac{\imath}{2}\operatorname{\mathfrak{L}}(\boldsymbol{\Xi}^{(-1)}\boldsymbol{T}_{\boldsymbol{\nu}}^{*})^{*}+(\boldsymbol{\Xi}^{(-1)}\boldsymbol{T}_{\boldsymbol{\nu}}^{*}\big)_{0}^{*}\right)\boldsymbol{h}\,=\,\left(\mp\frac{1}{2}+\boldsymbol{K}^{*}\right)\boldsymbol{h}\,.$
The proof is complete. ∎
For the Stokes operator $\boldsymbol{\Xi}_{0,0}$, see also[39], [62, Lemma 3.1], [70, (6.1), (6.2)], [99, Theorem 3.1], [100], [134, Lemma 1.3]. Note that $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ and $\frac{1}{2}+\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$. Therefore $\Big(\frac{1}{2}+\boldsymbol{K}\Big)\boldsymbol{S}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ and $\boldsymbol{S}\Big(\frac{1}{2}+\boldsymbol{K}^{*}\Big)\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$. In addition, we obtain the following formula.
Proposition 8.17.
We have the equality
$\Big(\frac{1}{2}+\boldsymbol{K}\Big)\boldsymbol{S}\,=\,\boldsymbol{S}\Big(\frac{1}{2}+\boldsymbol{K}^{*}\Big)\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)\,.$
Proof.
Let $\boldsymbol{h}\in H^{1/2}(\Gamma;T^{*}M)$. Theorem 6.6 then gives that
$\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)\mathrel{\,:=\,}U\mathrel{\,:=\,}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})|_{\Omega}\in H^{2}(\Omega;TM)\oplus H^{1}(\Omega)\,.$
Since $\boldsymbol{\Xi}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})=0$ in $M\smallsetminus\Gamma$ (Proposition 8.10), we can use the Green representation formula (Proposition 8.11) applied to $\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right):=U$, and this gives:
$\mathcal{D}_{\rm{ST}}(\boldsymbol{u}_{+})(x)-\mathcal{S}_{\rm{ST}}\big[(\boldsymbol{T}_{\boldsymbol{\nu}}U)_{+}\big](x)\,\,=\,\,U(x)\ \mbox{ if }\ x\in\Omega\,.$
The formula $U:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$ means that $\boldsymbol{u}:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})|_{\Omega}$ and $p:=\mathcal{P}_{\rm{ST}}(\boldsymbol{h})|_{\Omega}$ are the restrictions to $\Omega$ of the single-layer velocity and pressure potentials. By taking the traces on $\Gamma:=\partial\Omega$ from $\Omega$, we obtain
$\mathcal{D}_{\rm{ST}}\big(\big(\mathcal{V}_{\rm{ST}}(\boldsymbol{h})_{+}\big)_{+}-\mathcal{S}_{\rm{ST}}\big(\big(\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})\big)_{+}\big)_{+}\,=\,\mathcal{S}_{\rm{ST}}(\boldsymbol{h})_{+}\in H^{1}(\Gamma;TM)\oplus L^{2}(\Gamma)\,.$
The vector part of this relation is
$\mathcal{W}_{\rm{ST}}\big(\mathcal{V}_{\rm{ST}}(\boldsymbol{h})_{+}\big)_{+}-\mathcal{V}_{\rm{ST}}\big(\big(\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})\big)_{+}\big)_{+}\,=\,\mathcal{V}_{\rm{ST}}(\boldsymbol{h})_{+}\in H^{1}(\Gamma;TM)\,.$
Using the limit relations in Theorems 8.14 and 8.16, we obtain the formula
$\Big(\frac{1}{2}+\boldsymbol{K}\Big)\boldsymbol{S}\boldsymbol{h}-\boldsymbol{S}\Big(-\frac{1}{2}+\boldsymbol{K}^{*}\Big)\boldsymbol{h}\,=\,\boldsymbol{S}\boldsymbol{h}\,$,
which gives the desired result:
$\Big(\frac{1}{2}+\boldsymbol{K}\Big)\boldsymbol{S}\boldsymbol{h}-\boldsymbol{S}\Big(\frac{1}{2}+\boldsymbol{K}^{*}\Big)\boldsymbol{h}\,=\,0\,.$
Since $\boldsymbol{h}\in H^{1/2}(\Gamma;T^{*}M)$ is arbitrary, the previous formula leads to the desired property. This completes the proof. ∎
9. Calculations on a cylinder
We shall need definitions and results for the limit operators ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ that are analogous to those in the previous two sections. To this end, in this section we consider the particular case when our manifold $M$ with straight cylindrical ends is actually a cylinder $M=\mathfrak{S}\times\mathbb{R}$ with a product metric. We do not assume $\mathfrak{S}$ to be connected, though. Then, as in Section 7, we let $\Omega=\Omega^{\prime}\times\mathbb{R}\subset M$ be a product type open subset, and hence with straight cylindrical ends, where $\Omega^{\prime}\subset\mathfrak{S}$. Let $\Gamma^{\prime}:=\partial\Omega^{\prime}$. In this section, in addition to $V$ and $V\geq 0$ smooth, we assume that they are translation invariant, and hence $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$ identify to sections on $\mathfrak{S}$. Also, in this section, we use Assumption 8.1, that is that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$ (that is, they have points where they are strictly positive in every connected component of the space of definition, see Definition 2.13). Also, recall that $\Omega$ is on one side of its boundary (Assumption 2.7), and hence $\Omega^{\prime}$ is also on one side of its boundary.
We follow the usual convention and we let
(9.1) $\Omega_{-}^{\prime}\mathrel{\,:=\,}\mathfrak{S}\smallsetminus\overline{\Omega}^{\prime}\quad\mbox{and}\quad\Omega_{+}^{\prime}\mathrel{\,:=\,}\Omega^{\prime}\,.$
We also let $\Gamma^{\prime}:=\partial\Omega^{\prime}$, the boundary of $\Omega^{\prime}$, a smooth manifold, because the boundary $\Gamma$ of $\Omega$ is a smooth manifold (throughout this paper) and
(9.2) $\Gamma\,=\,\Gamma^{\prime}\times\mathbb{R}$
in this section. The assumption that $\Omega^{\prime}$ be on one side of its boundary is then equivalent to
(9.3) $\partial\Omega_{-}^{\prime}\,=\,\Gamma^{\prime}\mathrel{\,:=\,}\partial\Omega_{+}^{\prime}\mathrel{\,:=\,}\partial\Omega^{\prime}$
(see also Assumption 2.7). We also note that, for translation invariant potentials $V$ and $V_{0}$ the assumptions 8.1 and 8.3 are equivalent.
9.1. Layer potentials and jump relations for the indicial operators
With these hypotheses, we know from Theorem 7.14 that $\boldsymbol{\Xi}$ is invertible on $M=\mathfrak{S}\times\mathbb{R}$ and hence we can define the layer potential operators. These layer potential operators are translation invariant, so we can consider their indicial families (their Fourier transforms, see Definition 5.26). We have $\widehat{\boldsymbol{\Xi}}(\tau)]^{-1}=\widehat{(\boldsymbol{\Xi}^{-1})}(\tau)$, so $\widehat{\boldsymbol{\Xi}}(\tau)]^{-1}$ has a similar structure to $\boldsymbol{\Xi}^{-1}$ of Proposition 8.5. That is, we have the following remark.
Remark 9.1.
The inverse $\widehat{\boldsymbol{\Xi}}(\tau)]^{-1}$ has the form
$\widehat{\boldsymbol{\Xi}}(\tau)]^{-1}\,=:\,\left(\begin{array}[]{cc}\widehat{\mathcal{A}}(\tau)&\widehat{\mathcal{B}}(\tau)\\ \widehat{\mathcal{C}}(\tau)&\widehat{\mathcal{D}}(\tau)\end{array}\right)\in\Psi_{\operatorname{ess}}^{[-\boldsymbol{s}-\boldsymbol{t}]}(M;TM\oplus\mathbb{C})\,$,
where $\widehat{\mathcal{A}}(\tau)\in\Psi_{\operatorname{ess}}^{-2}(M;TM)$, $\widehat{\mathcal{C}}(\tau)=\widehat{\mathcal{B}}^{*}(\tau)\in\Psi_{\operatorname{ess}}^{-1}(M;TM,\mathbb{C})$, and $\widehat{\mathcal{D}}(\tau)\in\Psi_{\operatorname{ess}}^{0}(M)$.
Definition 9.2.
Assume that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$ (Assumption 8.1). Let $\boldsymbol{h}\in H^{s}(\Gamma^{\prime};TM)$, $s\in\mathbb{R}$. The indicial layer potentials operators associated to $\widehat{\boldsymbol{\Xi}}(\tau)$ are defined by
$\mathcal{S}_{\tau}(\boldsymbol{h})\mathrel{\,:=\,}\left(\begin{array}[]{c}\mathcal{V}_{\tau}(\boldsymbol{h})\\ \mathcal{P}_{\tau}(\boldsymbol{h})\end{array}\right)\mathrel{\,:=\,}\widehat{\boldsymbol{\Xi}}(\tau)^{-1}\,\left[\left(\begin{array}[]{c}{\boldsymbol{h}}\\ {0}\end{array}\right)\otimes\delta_{\Gamma}\right]\,$,
and by
$\mathcal{D}_{\tau}(\boldsymbol{h})\mathrel{\,:=\,}\left(\begin{array}[]{c}\mathcal{W}_{\tau}(\boldsymbol{h})\\ \mathcal{Q}_{\tau}(\boldsymbol{h})\end{array}\right)\mathrel{\,:=\,}\widehat{\boldsymbol{\Xi}}(\tau)^{-1}\,\big[\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}^{*}(\tau)\left(\boldsymbol{h}\otimes\delta_{\Gamma}\right)\big]\,.$
We have the following analogue of Remark 8.8.
Remark 9.3.
We have
$\mathcal{W}_{\tau}(\boldsymbol{h})\,=\,\left(\begin{array}[]{cc}\widehat{\mathcal{A}}(\tau)&\widehat{\mathcal{B}}(\tau)\end{array}\right)\left(\begin{array}[]{c}-2\boldsymbol{D}_{\boldsymbol{\nu}}^{*}\\ \boldsymbol{\nu}^{\sharp}\end{array}\right)(\boldsymbol{h}\otimes\delta_{\Gamma})\,=\,(-2\widehat{\mathcal{A}}(\tau)\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\widehat{\mathcal{B}}(\tau)\boldsymbol{\nu}^{\sharp})\,(\boldsymbol{h}\otimes\delta_{\Gamma})$
and
$\mathcal{Q}_{\tau}(\boldsymbol{h})\,=\,\left(\begin{array}[]{cc}\widehat{\mathcal{C}}(\tau)&\widehat{\mathcal{D}}(\tau)\end{array}\right)\left(\begin{array}[]{c}-2\boldsymbol{D}_{\boldsymbol{\nu}}^{*}\\ \boldsymbol{\nu}^{\sharp}\end{array}\right)(\boldsymbol{h}\otimes\delta_{\Gamma})\\ \ \,=\,(-2\widehat{\mathcal{C}}(\tau)\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\widehat{\mathcal{D}}(\tau)\boldsymbol{\nu}^{\sharp})\,(\boldsymbol{h}\otimes\delta_{\Gamma})\,.$
Similarly, $\mathcal{S}_{\tau}(\boldsymbol{h})\,=\,\big(\,\widehat{\mathcal{A}}(\tau)(\boldsymbol{h}\otimes\delta_{\Gamma})\ \ \ \widehat{\mathcal{C}}(\tau)(\boldsymbol{h}\otimes\delta_{\Gamma})\,\big)^{\top}$.
The following result follows from the definition of the indicial layer potentials and it has the same proof as Proposition 8.10
Proposition 9.4.
Let $\tau\in\mathbb{R}$ and $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$. Then $\widehat{\boldsymbol{\Xi}}(\tau)\mathcal{S}_{\tau}(\boldsymbol{h}^{\prime})$ and $\widehat{\boldsymbol{\Xi}}(\tau)\mathcal{D}_{\tau}(\boldsymbol{h}^{\prime})$ vanish on $\mathfrak{S}\smallsetminus{\Gamma^{\prime}}$.
The following result result for the layer potential operators associated to $\widehat{\boldsymbol{\Xi}}(\tau)$ and its proof are similar to those of Proposition 8.9.
Proposition 9.5.
Let $\Omega_{+}^{\prime}:=\Omega^{\prime}\subset\mathfrak{S}$, $\Omega_{-}^{\prime}:=\mathfrak{S}\smallsetminus\overline{\Omega}^{\prime}$, and $\Gamma^{\prime}:=\partial\Omega^{\prime}$, as before. Let $s\in\mathbb{R}$ and $\boldsymbol{h}^{\prime}\in H^{s}(\Gamma^{\prime};TM)$. Then
$\mathcal{V}_{\tau}(\boldsymbol{h}^{\prime})|_{\Omega_{\pm}^{\prime}}\in H^{s+\frac{3}{2}}(\Omega_{\pm}^{\prime};TM)\ \mbox{ and }\ \mathcal{P}_{\tau}(\boldsymbol{h}^{\prime})|_{\Omega_{\pm}^{\prime}}\in H^{s+\frac{1}{2}}(\Omega_{\pm}^{\prime})\,.$
Similarly, let $\boldsymbol{h}\in H^{s}(\Gamma^{\prime};TM)$. Then
$\mathcal{W}_{\tau}(\boldsymbol{h}^{\prime})|_{\Omega_{\pm}^{\prime}}\in H^{s+\frac{1}{2}}(\Omega_{\pm}^{\prime};TM)\ \mbox{ and }\ \mathcal{Q}_{\tau}(\boldsymbol{h}^{\prime})|_{\Omega_{\pm}^{\prime}}\in H^{s-\frac{1}{2}}(\Omega_{\pm}^{\prime})\,.$
Recall the indicial layer potentials of Definition 9.2. Then we have the following result.
Proposition 9.6.
Let $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$, $\Gamma^{\prime}:=\partial\Omega^{\prime}$, $\color[rgb]{1,0,1}\widehat{\mathfrak{f}}:=\frac{\color[rgb]{1,0,1}\widehat{V}_{0}+1}{2\color[rgb]{1,0,1}\widehat{V}_{0}+1}$ $\color[rgb]{1,0,1}\widehat{\mathfrak{g}}:=\frac{1}{2\color[rgb]{1,0,1}\widehat{V}_{0}+1}$. Let $\boldsymbol{K}:=\boldsymbol{P}_{0}\in\Psi^{0}(\Gamma^{\prime};TM)$ be as in Equation (8.6). Then
  1. (i)
    $\mathcal{V}_{\tau}(\boldsymbol{h}^{\prime})_{\pm}\,=\,\widehat{\mathcal{A}}_{0}(\tau)\boldsymbol{h}^{\prime}$, and $\widehat{\mathcal{A}}_{0}(\tau)$ is elliptic with self-adjoint symbol $\sigma_{-1}(\mathcal{A}_{0}(\tau);\xi^{\prime})\,=\,\frac{1}{4|\xi^{\prime}|}(2-\color[rgb]{1,0,1}\widehat{\mathfrak{f}}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp}-\color[rgb]{1,0,1}\widehat{\mathfrak{f}}\eta^{\sharp}\otimes\eta)\,$, where $\eta:=|\xi^{\prime}|^{-1}\xi^{\prime}$.
  2. (ii)
    $\mathcal{P}_{\tau}(\boldsymbol{h}^{\prime})_{\pm}=\big(\mp\frac{\color[rgb]{1,0,1}\widehat{\mathfrak{g}}}{2}\boldsymbol{\nu}^{\sharp}+\widehat{\mathcal{C}}_{0}(\tau)\big)\boldsymbol{h}^{\prime}$.
  3. (iii)
    $\mathcal{W}_{\tau}(\boldsymbol{h}^{\prime})_{\pm}\mathrel{\,:=\,}\big[\widehat{\boldsymbol{P}}(\tau)(\boldsymbol{h}\otimes\delta_{\Gamma})\big]_{\pm}\,=\,\left[\pm\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)\right]\boldsymbol{h}^{\prime}$ and the two operators $\pm\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)$ are elliptic and have self-adjoint principal symbols.
  4. (iv)
    $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\tau}(\boldsymbol{h}^{\prime})]_{\pm}\,=\,\left(\mp\frac{1}{2}+\widehat{\boldsymbol{K}}{}^{*}(\tau)\right)\boldsymbol{h}^{\prime},$ where $\boldsymbol{K}=\boldsymbol{P}_{0}:=(-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp})_{0}$, as in Theorem 8.14.
Proof.
The result follows from the “jump relations” established in Theorems 8.14 and 8.16 and the compatibility of the normal limits with the indicial operators, Theorem 6.10. ∎
9.2. Invertibility of $\boldsymbol{S}$ on a cylinder
We keep the assumptions of the previous subsection. That is,
  • $M=\mathfrak{S}\times\mathbb{R}$,
  • $V$ and $V\geq 0$ are smooth, translation invariant, and hence $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$ identify to sections on $\mathfrak{S}$.
  • $V$ and $V_{0}$ satisfy the positivity assumption at infinity (Assumption 8.1, that is, $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$).
(Recall that $\succ$ was introduced in Definition 2.13.) We also keep the notation and assumptions of equations (9.2), (9.1), and (9.3).
From now on, all layer potential operators $\boldsymbol{S}$, $\boldsymbol{K}$, … (possibly decorated with indices) are associated to the boundaries $\Gamma$ and $\Gamma^{\prime}$ of the open sets $\Omega=:\Omega_{+}$ and $\Omega^{\prime}=:\Omega_{+}^{\prime}$, see Definitions 8.7, 8.7, and 9.2.
We now prove the invertibility of the single layer potential operator $\boldsymbol{S}:=\mathcal{A}_{0}$, see Theorem 8.16 on our cylinder $M=\mathfrak{S}\times\mathbb{R}$, if $V$ and $V_{0}$ satisfy the positivity assumption at infinity (Assumption 8.1). In order to do that, we need to first study its indicial operators $\widehat{\boldsymbol{S}\,}(\tau)$, $\tau\in\mathbb{R}$. We begin with the case $\tau=0$, which requires more assumptions than the case $\tau\neq 0$.
Proposition 9.7.
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ (is a cylinder), that $V$ and $V_{0}$ are smooth, translation invariant at infinity, that $\Omega=\Omega^{\prime}\times\mathbb{R}$, and that $\Omega^{\prime}\subset\mathfrak{S}$ is on one side of its boundary $\Gamma^{\prime}:=\partial\Omega^{\prime}$ (i.e., $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}^{\prime}\subset\mathfrak{S}$). We also assume that
  1. (i)
    $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\mathfrak{S}$ and
  2. (ii)
    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}\smallsetminus\Gamma^{\prime}$.
Let $\widehat{\boldsymbol{S}\,}(0)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16. Then $\widehat{\boldsymbol{S}\,}(0)$ is invertible and $\widehat{\boldsymbol{S}\,}(0)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$.
Proof.
Our assumptions imply that $V$ and $V_{0}$ satisfy the non-vanishing conditions of Assumption 8.3 on $M=\mathfrak{S}\times\mathbb{R}$, and hence the layer potential operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ is defined (see Proposition 8.5). We know from Proposition 9.6 that $\widehat{\boldsymbol{S}\,}(0)$ is elliptic with self-adjoint principal symbol. Because $\mathfrak{S}$ is compact, $\widehat{\boldsymbol{S}\,}(0)$ is then Fredholm of index zero, by classical results (the “classical case” of a closed manifold in Theorem 6.15).
Let us show now that $\ker\widehat{\boldsymbol{S}\,}(0)=0$. Let then $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$ be such that $\widehat{\boldsymbol{S}\,}(0)\boldsymbol{h}^{\prime}=0$. Our assumptions imply that $\widehat{\boldsymbol{\Xi}}(0)$ is invertible (Proposition 7.11), and hence the indicial single layer potential $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{0}(\boldsymbol{h}^{\prime})$ is defined (see Definition 9.2). Since $\widehat{\boldsymbol{S}\,}(0)$ is elliptic and $\Gamma^{\prime}$ is smooth and compact, we have that $\boldsymbol{h}^{\prime}\in H^{s}(\Gamma^{\prime};TM)$ for all $s\in\mathbb{R}$, by elliptic regularity. Proposition 9.5 for $s=1/2$ then gives that the restrictions of $U$ to $\Omega_{+}^{\prime}:=\Omega^{\prime}$ and to $\Omega_{-}^{\prime}:=\mathfrak{S}\smallsetminus\overline{\Omega}{\,}^{\prime}$ satisfy
(9.4) $U|_{\Omega_{\pm}^{\prime}}\mathrel{\,:=\,}(\,\boldsymbol{u}\ \ p\,)^{\top}|_{\Omega_{\pm}^{\prime}}\mathrel{\,:=\,}\mathcal{S}_{0}(\boldsymbol{h}^{\prime})|_{\Omega_{\pm}^{\prime}}\in H^{2}(\Omega_{\pm}^{\prime};TM)\oplus H^{1}(\Omega_{\pm}^{\prime})\,.$
We also know that $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega_{\pm}^{\prime}$ by Proposition 8.10. Proposition 9.6 gives then that $\boldsymbol{u}_{+}=\boldsymbol{u}_{-}=\widehat{\boldsymbol{S}\,}(0)\boldsymbol{h}^{\prime}=0$. Therefore $\boldsymbol{u}=0$ in $\Omega_{-}^{\prime}$ and in $\Omega_{+}^{\prime}$, by Corollary 7.7. The same corollary gives that $p$ is constant on each connected component of $\mathfrak{S}\smallsetminus\Gamma^{\prime}$. Because $\color[rgb]{1,0,1}\widehat{V}_{0}$ is not identically equal to zero on any connected component of $\mathfrak{S}\smallsetminus\Gamma^{\prime}$, we obtain that, in fact, $p=0$. We have thus obtained that $U=0$ and hence $\widehat{\boldsymbol{S}\,}(0)$ is injective.
Finally, the fact that $\widehat{\boldsymbol{S}\,}(0)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$ follows from Beals’ theorem (see, for instance, Theorem 7.4 in [67]). ∎
The same result holds for $\tau\neq 0$ under the weaker hypothesis that $V\succ 0$ and $V_{0}\succ 0$ on $M$ (equivalently, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$).
Proposition 9.8.
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder and that $V$ and $V_{0}$ are smooth, translation invariant, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Let $\widehat{\boldsymbol{S}\,}(\tau)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16. If $\tau\neq 0$, then $\widehat{\boldsymbol{S}\,}(\tau)$ is invertible and $\widehat{\boldsymbol{S}\,}(\tau)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$.
Proof.
The proof is exactly the same as that of Proposition 9.7, except that we are using Corollary 7.9 (which only requires $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ to be non-negative) instead of Corollary 7.7. ∎
We can now prove the invertibility of the single layer potential $\boldsymbol{S}$ on $\Gamma=\Gamma^{\prime}\times\mathbb{R}$ in the case of cylinders and under Assumption 8.1.
Theorem 9.9.
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\mathfrak{S}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}\smallsetminus\Gamma^{\prime}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Let $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ be as in Theorem 8.16. Then $\boldsymbol{S}$ is invertible, $\boldsymbol{S}^{-1}\in\Psi_{\operatorname{ess}}^{1}(\Gamma;TM)$, and hence $\boldsymbol{S}:H^{s}(\Gamma;TM)\to H^{s+1}(\Gamma;TM)$ is an isomorphism for all $s\in\mathbb{R}$.
Proof.
Propositions 9.7 and 9.8 show that $\widehat{\boldsymbol{S}}(\tau)$ is invertible for all $\tau\in\mathbb{R}$. It is well-known from the work of Melrose, Schulze and others (see also Theorem 5.11 in [68]) that $\boldsymbol{S}$ is then invertible. ∎
9.3. Invertibility of $\frac{1}{2}+\boldsymbol{K}$ on a cylinder
In this subsection, we keep the assumptions of the previous two subsections. In particular, we assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$.
In the following, we will need the operators $\boldsymbol{P}:=-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp}$ and $\boldsymbol{K}\mathrel{\,:=\,}\boldsymbol{P}_{0}$. The first one is needed in the definition of the double layer potential operator $\mathcal{W}_{\rm{ST}}$ and the second one was introduced in Equation (8.6) and studied in Theorem 8.14. Each of these operators depends on the potentials $V$ and $V_{0}$. (The correspondence $\boldsymbol{P}\mapsto\boldsymbol{P}_{0}$ is the basic correspondence studied, for example, in Theorems 4.16 and 6.8.)
The proof of the invertibility of $\frac{1}{2}+\boldsymbol{K}$ (under suitable conditions on $V$ and $V_{0}$) requires the invertibility of its indicial operators. As for the operator $\boldsymbol{S}$, we first consider the case $\tau=0$.
Proposition 9.10.
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega_{+}^{\prime}:=\Omega^{\prime}$. We furthermore assume that $\Omega^{\prime}$ is on one side of its boundary, as usual (i.e., $\Omega^{\prime}$ is the interior of $\overline{\Omega}^{\prime}$, see Assumption 2.7). Let $\widehat{\boldsymbol{K}}(0)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ given in Theorem 8.14 (see also Equation (6.1)). Then $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
Recall that by the statement $\phi\not\equiv 0$ on AA,” we mean that there exists $a$ in the domain of $\phi$ such that $\phi(a)\not=0$. To negate this statement, we shall write $\phi=0$ in AA.” Of course, in view of our conventions, the condition $\color[rgb]{1,0,1}\widehat{V}_{0}\not\equiv 0$ on $\Omega^{\prime}$ implies the condition $V_{0}\not\equiv 0$ on $\Omega$ (since $V$ and $V_{0}$ are translation invariant at infinity).
Proof.
We know that that the operator $\boldsymbol{K}\mathrel{\,:=\,}\boldsymbol{P}_{0}\mathrel{\,:=\,}\big(-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp})_{0}$ satisfies $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$, by Theorem 6.8. Therefore $\widehat{\boldsymbol{K}}(0)\in\Psi^{0}(\Gamma^{\prime};TM)$, by the definition of the indicial family. We also know from Theorem 8.14 that $\frac{1}{2}+\boldsymbol{K}$ is elliptic with self-adjoint principal symbol. Consequently, $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)$ is also elliptic with self-adjoint principal symbol. Because $\Gamma^{\prime}$ is compact, $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)$ is then Fredholm of index zero by classical results [53, 54]. Therefore, in order to prove that $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)$ is invertible, it suffices to prove that its adjoint operator $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)^{*}$ is injective.
Let then $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$ be such that $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)^{*}\big]\boldsymbol{h}^{\prime}=0$. Since $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)^{*}$ is elliptic and $\Gamma^{\prime}$ is smooth and compact, we have that $\boldsymbol{h}^{\prime}\in H^{s}(\Gamma^{\prime};TM)$ for all $s\in\mathbb{R}$, by elliptic regularity. Let us consider the indicial single layer potential $U:=\mathcal{S}_{0}\boldsymbol{h}^{\prime}$ (see Definition 9.2 with $\tau=0$). Proposition 9.5 then gives that the restrictions of $U$ to $\Omega_{+}^{\prime}:=\Omega^{\prime}$ and to $\Omega_{-}^{\prime}:=\mathfrak{S}\smallsetminus\overline{\Omega}^{\prime}$ satisfy
$U|_{\Omega_{\pm}^{\prime}}\mathrel{\,:=\,}(\,\boldsymbol{u}\ \ p\,)^{\top}|_{\Omega_{\pm}^{\prime}}\mathrel{\,:=\,}\mathcal{S}_{0}(\boldsymbol{h}^{\prime})|_{\Omega_{\pm}^{\prime}}\in H^{2}(\Omega_{\pm}^{\prime};TM)\oplus H^{1}(\Omega_{\pm}^{\prime})\,.$
We also have that $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega_{\pm}^{\prime}$ by Proposition 8.10. Proposition 9.6 then gives that $\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(0)U]_{-}=\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)^{*}\big]\boldsymbol{h}^{\prime}=0$. (Recall that $\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(0)U]_{-}$ is the trace at $\Gamma^{\prime}$ of the restriction of $\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(0)U$ to the domain $\Omega_{-}^{\prime}$.)
Let next $\Omega_{0}$ be a connected component of $\Omega_{-}^{\prime}$. Our assumptions give that $\color[rgb]{1,0,1}\widehat{V}\not\equiv 0$ on $\Omega_{0}$. Hence we obtain that $\boldsymbol{u}=0$ on $\Omega_{0}$, by Corollary 7.7. The same corollary gives that $p$ is constant in $\Omega_{0}$. Let us now show that this constant is actually 0. Indeed, the relation
$0\,=\,[\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(0)U]_{-}\mathrel{\,:=\,}[-2\widehat{\boldsymbol{D}}_{\boldsymbol{\nu}}(0)\boldsymbol{u}+p\boldsymbol{\nu}]_{-}\,=\,p_{-}\boldsymbol{\nu}$
on $\Gamma$ implies that $p_{-}=0$, and hence $p=0$ on $\Omega_{0}$. Since $\Omega_{0}$ was an arbitrary connected component of $\Omega_{-}^{\prime}$, we obtain that $U=0$ on $\Omega_{-}^{\prime}$.
We have already observed that $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega_{+}^{\prime}$. Proposition 9.6 then gives the “no-jump relation” $\boldsymbol{u}_{+}=\boldsymbol{u}_{-}=0$ at the boundary $\Gamma^{\prime}:=\partial\Omega^{\prime}$ of $\Omega^{\prime}$ (interior and exterior traces). Corollary 7.7 then gives that $\boldsymbol{u}=0$ on $\Omega^{\prime}$. Because $\color[rgb]{1,0,1}\widehat{V}_{0}\not\equiv 0$ on every connected component of $\Omega^{\prime}$, the same corollary gives that $p=0$ on $\Omega^{\prime}$. We have thus obtained that $U=0$ on both $\Omega_{\pm}^{\prime}$, and hence $U=0$ on $\mathfrak{S}$.
We next use relation (iv) of Proposition 9.6 to conclude that
$\boldsymbol{h}^{\prime}$ $\,=\,\left(\frac{1}{2}+\widehat{\boldsymbol{K}}{}^{*}(\tau)\right)\boldsymbol{h}^{\prime}-\left(-\frac{1}{2}+\widehat{\boldsymbol{K}}{}^{*}(\tau)\right)\boldsymbol{h}^{\prime}$
$\,=\,[\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\tau}(\boldsymbol{h}^{\prime})]_{-}-[\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\tau}(\boldsymbol{h}^{\prime})]_{+}$
$\,=\,[\boldsymbol{T}_{\boldsymbol{\nu}}U]_{-}-[\boldsymbol{T}_{\boldsymbol{\nu}}U]_{+}\,=\,0\,.$
Therefore $\widehat{\boldsymbol{S}}(0)$ is injective, and hence it is invertible.
The fact that $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$ is also a classical result on pseudodifferential operators [16] (a proof can also be found in [63], Theorem 15.4.11, and in [67], Theorem 7.4). This completes the proof. ∎
As for the single layer potential, the case $\tau\neq 0$ requires weaker assumptions on the potentials.
Proposition 9.11.
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder $V$ and $V_{0}$ are smooth, translation invariant, and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Let $\widehat{\boldsymbol{K}}(\tau)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ of Theorem 8.14 (see also Equation (6.1)). If $\tau\neq 0$, then $\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
Proof.
The proof follows the same arguments as those in the proof of Proposition 10.5, except that, instead of Corollary 7.7, we use Corollary 7.9. ∎
We are ready to prove the invertibility of the double layer potential operator $\frac{1}{2}+\boldsymbol{K}$ on the cylinder $\Gamma:=\Gamma^{\prime}\times\mathbb{R}$.
Theorem 9.12.
Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega_{+}^{\prime}:=\Omega^{\prime}$. We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$, as usual. Let $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ be as in Theorem 8.14 (see also Equation (6.1)). Then $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi^{0}(\Gamma;TM)$.
Proof.
The proof is the same as that of Theorem 9.9, but using Propositions 9.10 and 9.11 instead of Propositions 9.7 and 9.8. ∎
10. Fredholmness and invertibility of layer potential operators on manifolds with straight cylindrical ends
We now come back to the case of a general manifold with straight cylindrical ends $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ (Definition 3.11), where $M^{\prime}:=\partial M_{0}$ is compact, as usual. We will also assume that $V$ and $V_{0}$ satisfy the corresponding non-vanishing assumptions (Assumption 8.3). Also, we assume that we are given a compatible open subset $\Omega\subset M\,=\,M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$, as in Equation (1.2). Recall that this means that $\Omega$ also has straight cylindrical end at infinity, which are assumed to “match” that of $M$, in the sense that there exists $R_{\Omega}\leq 0$ and an open subset $\Omega^{\prime}\subset M$ such that
(10.1) $\Omega\cap\big[M^{\prime}\times(-\infty,R_{\Omega})\big]\,=\,\Omega^{\prime}\times(-\infty,R_{\Omega})\,.$
Then $\Gamma:=\partial\Omega$ is also a manifold with straight cylindrical ends. Let $\Gamma^{\prime}:=\partial\Omega^{\prime}\subset M^{\prime}$, we thus have also the compatibility relation
(10.2) $\Gamma\cap\big[M^{\prime}\times(-\infty,R_{\Omega})\big]\,=\,\Gamma^{\prime}\times(-\infty,R_{\Omega})\,$,
(compare to Equation (6.1)). We can assume, in fact, that $R_{\Omega}=0$ in the last equations. We set $\Omega_{-}:=M\smallsetminus\overline{\Omega}$, $\Omega_{-}^{\prime}:=M^{\prime}\smallsetminus\overline{\Omega}{\,}^{\prime}$, $\Omega_{+}:=\Omega$, and $\Omega_{+}^{\prime}:=\Omega^{\prime}$. In the results of this section, we can split our generalized Stokes operator as a direct sum according to the connected components of $\Omega$, so there is no loss of generality to assume that $\Omega$ is connected. (We have assumed that $M$ is connected for the same reason.) The general case follows immediately from this particular case.
10.1. Invertibility of $\boldsymbol{S}$
We first prove the invertibility of the single layer potential operator $\boldsymbol{S}:=\mathcal{A}_{0}$, see Theorem 8.16. In order to do that, we need to first study its indicial operators $\widehat{\boldsymbol{S}\,}(\tau)$, $\tau\in\mathbb{R}$. We begin with the case $\tau=0$, which requires more assumptions than the case $\tau\neq 0$. Recall the definition of $\succ$ (Definition 2.13). We notice that Assumption 8.3 implies Assumption 8.1 on $M^{\prime}\times\mathbb{R}$.
Proposition 10.1.
Let us assume that $M$ is a manifold with straight cylindrical ends and that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$. Let $\widehat{\boldsymbol{S}\,}(0)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16. Then $\widehat{\boldsymbol{S}\,}(0)$ is invertible and $\widehat{\boldsymbol{S}\,}(0)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$.
Proof.
The hypotesis imply that the operator $\boldsymbol{\Xi}$ is Fredholm, and hence the layer potentials are defined. The result then follows from Proposition 9.7 applied to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ acting on $M^{\prime}\times\mathbb{R}$ and the fact that layer potentials commute with limit operators, Theorem 6.10. ∎
The same result holds for $\tau\neq 0$ under the weaker hypothesis that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$.
Proposition 10.2.
Let us assume that $M$ is a manifold with straight cylindrical ends and that $V$ and $V_{0}$ are smooth, asymptotically translation invariant at infinity, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}$. Let $\widehat{\boldsymbol{S}\,}(\tau)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16. If $\tau\neq 0$, then $\widehat{\boldsymbol{S}\,}(\tau)$ is invertible and $\widehat{\boldsymbol{S}\,}(\tau)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$.
Proof.
The proof is the same as that of Proposition 10.1, but using Proposition 9.8 instead of Proposition 9.7. (As in that proof, the assumptions $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}$ are needed for $\boldsymbol{\Xi}$ to be Fredholm, so that the layer potentials are defined.) ∎
The above propositions lead to the following useful corollary.
Corollary 10.3.
Let us assume that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a manifold with straight cylindrical ends, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\Omega^{\prime}$ is the interior of $\overline{\Omega}^{\prime}$. We also assume that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$. Then $\boldsymbol{S}$ is an unbounded Fredholm operator on $L^{2}(\Gamma;TM)$. Equivalently, it is a (bounded) Fredholm operator $\boldsymbol{S}:H^{-1/2}(\Gamma;TM)\to H^{1/2}(\Gamma;TM)$.
Proof.
The operator ${\mathcal{R}_{\infty}}(\boldsymbol{S})$ is the single layer potential operator associated to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi}_{V,V_{0}})=\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$, by Theorem 6.10. The assumptions of the corollary give that the assumptions of Theorem 9.9 are therefore satisfied by ${\mathcal{R}_{\infty}}(\boldsymbol{S})$, which is hence invertible. Because $\boldsymbol{S}$ is elliptic, Theorem 6.15 gives that $\boldsymbol{S}$ is Fredholm. ∎
We can now prove the invertibility of the single layer potential $\boldsymbol{S}$ on $\Gamma$.
Theorem 10.4.
Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions 8.3 (hence, in particular, $M$ is with straight cylindrical ends and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$). Let $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ be as in Theorem 8.16. We furthermore assume the following conditions.
  1. (i)
    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$.
  2. (ii)
    $V_{0}\succ 0$ on $M\smallsetminus\Gamma$.
  3. (iii)
    $\Omega$ is on one side of its boundary $\Gamma$ (i.e., $\Omega$ is the interior of $\overline{\Omega}$).
Then $\boldsymbol{S}$ is invertible, $\boldsymbol{S}^{-1}\in\Psi_{\operatorname{ess}}^{1}(\Gamma;TM)$, and hence $\boldsymbol{S}:H^{s}(\Gamma;TM)\to H^{s+1}(\Gamma;TM)$ is an isomorphism for all $s\in\mathbb{R}$.
Proof.
Because $V$ and $V_{0}$ satisfy the non-vanishing assumptions 8.3 on $M$, Theorem 8.4 implies that the operator $\boldsymbol{\Xi}$ is invertible, and hence the layer potentials are defined. In particular, the single layer potential operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ is defined. In addition, Theorem 8.16 shows that $\boldsymbol{S}$ is elliptic and self-adjoint from $H^{-1/2}(\Gamma;TM)$ to $H^{1/2}(\Gamma;TM)$ with respect to the $L^{2}$-inner product. We know from Corollary 10.3 just proved that $\boldsymbol{S}:H^{-1/2}(\Gamma;TM)\to H^{1/2}(\Gamma;TM)$ is a Fredholm operator. The self-adjoint property of $\boldsymbol{S}$ shows that its index is zero. Therefore, if we show that this operator is injective, then it will be invertible.
To this end, let $\boldsymbol{h}\in H^{-1/2}(\Gamma;TM)$ be such that $\boldsymbol{S}\boldsymbol{h}=0$ and let $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$ be the corresponding single layer potential. The ellipticity of $\boldsymbol{S}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ and the elliptic regularity property of Theorem 8.16 assure that $\boldsymbol{h}\in H^{s}(\Gamma;TM)$ for all $s\in\mathbb{R}$. Then, by Proposition 8.9, the restrictions of $U$ to $\Omega_{+}:=\Omega$ and to $\Omega_{-}$ satisfy the property
(10.3) $U|_{\Omega_{\pm}}\mathrel{\,:=\,}(\,\boldsymbol{u}\ \ p\,)^{\top}|_{\Omega_{\pm}}\mathrel{\,:=\,}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})|_{\Omega_{\pm}}\in H^{2}(\Omega_{\pm};TM)\oplus H^{1}(\Omega_{\pm})\,.$
We also know that $\boldsymbol{\Xi}U=0$ in $\Omega_{\pm}$ by Proposition 8.10. Theorem 8.16 gives then that $\boldsymbol{u}_{+}=\boldsymbol{u}_{-}=\boldsymbol{S}\boldsymbol{h}=0$, and by Corollary 2.14, $\boldsymbol{u}=0$ in $\Omega_{-}$ and in $\Omega_{+}$. The Stokes equation and again Corollary 2.14 give that $p$ is constant on each connected component of $M\smallsetminus\Gamma$, but this constant is 0, because $V_{0}$ is not identically equal to zero on any connected component of $M\smallsetminus\Gamma$. Thus, $U=0$. This implies that $\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U]_{\pm}=0$ on $\Gamma$, and Theorem 8.16(iii) gives
$\boldsymbol{h}\,=\,\Big(\frac{1}{2}+K^{*}\Big)\boldsymbol{h}-\Big(-\frac{1}{2}+K^{*}\Big)\boldsymbol{h}\,=\,[\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U]_{-}-[\tilde{\boldsymbol{T}}_{\boldsymbol{\nu}}U]_{+}\,=\,0\,.$
Consequently, $\boldsymbol{S}$ is injective and, in view of the above arguments, it is invertible. Finally, the property that $\boldsymbol{S}^{-1}\in\Psi_{\operatorname{ess}}^{1}(\Gamma;TM)$ follows from Theorem 7.4 in [67] (see also Theorem 15.4.11 in [63]). ∎
10.2. Invertibility of $\frac{1}{2}+\boldsymbol{K}$
In the following, we will use the operators $\boldsymbol{P}:=-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp}$ and $\boldsymbol{K}\mathrel{\,:=\,}\boldsymbol{P}_{0}$. The first one is needed in the definition of the double layer potential operator $\mathcal{W}_{\rm{ST}}$ and the second one was introduced in Equation (8.6) and studied in Theorem 8.14. Each of these operators depends on the potentials $V$ and $V_{0}$. (The correspondence $\boldsymbol{P}\mapsto\boldsymbol{P}_{0}$ is the basic correspondence studied, for example, in Theorems 4.16 and 6.8.)
The proof of the invertibility of $\frac{1}{2}+\boldsymbol{K}$ (under suitable conditions on $V$ and $V_{0}$) requires the invertibility of its indicial operators. We first consider the case $\tau=0$. This case needs additional assumptions and arguments. Recall that we have $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime},t)$ for $(x^{\prime},t)\in M^{\prime}\times(-\infty,0)$ and $t\ll 0$ and that $\color[rgb]{1,0,1}\widehat{V}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ is independent of $t$. The function $\color[rgb]{1,0,1}\widehat{V}_{0}$ on $M^{\prime}$ is defined similarly. The functions $\color[rgb]{1,0,1}\widehat{V}_{0}$ and $\color[rgb]{1,0,1}\widehat{V}$ correspond to the limit operators of the operators defined by $V_{0}$ and $V$.
Proposition 10.5.
Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions (Assumption 8.3) and let $\widehat{\boldsymbol{K}}(0)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ given in Theorem 8.14 (see also Equation (10.2)). We furthermore assume that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and that $\Omega^{\prime}$ on one side of its boundary $\Gamma^{\prime}:=\overline{\Omega^{\prime}}$. Then $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
Of course, in view of our conventions, the condition $\color[rgb]{1,0,1}\widehat{V}_{0}\not\equiv 0$ on $\Omega^{\prime}$ implies the condition $V_{0}\not\equiv 0$ on $\Omega$ (since $V$ and $V_{0}$ are translation invariant at infinity).
Proof.
The proof is very similar to that of Proposition 10.1, but using Proposition 9.10 instead of 9.7. Indeed, first, the hypotesis imply that the operator $\boldsymbol{\Xi}$ is Fredholm, and hence the layer potentials are defined. The result then follows from Proposition 9.10 applied to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ acting on $M^{\prime}\times\mathbb{R}$ and the fact that layer potentials commute with limit operators, Theorem 6.10. ∎
Proposition 10.6.
Let us assume that $V,V_{0}\geq 0$ satisfy Assumption 8.3 (so $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}$)and let $\widehat{\boldsymbol{K}}(\tau)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ of Theorem 8.14 (see also Equation (10.2)). If $\tau\neq 0$, then $\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
Proof.
The proof is the same as that of Proposition 10.5, but using Proposition 9.11 instead of Proposition 9.10. ∎
We above propositions lead to the following useful corollary.
Corollary 10.7.
Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions 8.3 and let $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ be as in Theorem 8.14 (as before, see also Equation (10.2)). We furthermore assume that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Then $\frac{1}{2}+\boldsymbol{K}$ is a Fredholm operator on $L^{2}(\Gamma;TM)$.
Proof.
The proof is similar to that of Corollary 10.3, but using Theorem 9.12 instead of Theorem 9.9. Indeed, the operator ${\mathcal{R}_{\infty}}(\frac{1}{2}+\boldsymbol{K})$ is the double layer potential operator associated to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi}_{V,V_{0}})=\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$, by Theorem 6.10. The assumptions of the corollary give that the assumptions of Theorem 9.12 are therefore satisfied by ${\mathcal{R}_{\infty}}(\frac{1}{2}+\boldsymbol{K})$, which is hence invertible. Because $\frac{1}{2}+\boldsymbol{K}$ is elliptic, Theorem 6.15 gives that $\frac{1}{2}+\boldsymbol{K}$ is Fredholm. ∎
We are ready to prove the invertibility of the double layer potential operator $\frac{1}{2}+\boldsymbol{K}$ on $\Gamma$.
Theorem 10.8.
Let us assume that $M,V$, and $V_{0}$ satisfy Assumption 8.3 (so, in particular $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$), that $\Omega$ is connected, and that $\Omega$ is the interior of $\overline{\Omega}$. We furthermore assume the following:
  1. (i)
    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$,
  2. (ii)
    $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and
  3. (iii)
    $V\succ 0$ on $\Omega_{-}$.
Let $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ be as in Theorem 8.14 (see also Equation (10.2)). Then $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi^{0}(\Gamma;TM)$.
Proof.
We split our proof into five steps.
Step 1 ($\frac{1}{2}+\boldsymbol{K}$ has index zero). Corollary 10.7 implies that $\frac{1}{2}+\boldsymbol{K}$ is a Fredholm operator. Proposition 8.17 then gives that $\frac{1}{2}+\boldsymbol{K}$ has index zero. Indeed, for a Fredholm operator $T:X\to Y$, let $\operatorname{ind}_{X}(T)$ be its index. Proposition 8.17 implies the relation $\big(\frac{1}{2}+\boldsymbol{K}\big)\boldsymbol{S}\,=\,\boldsymbol{S}\big(\frac{1}{2}+\boldsymbol{K}^{*}\big):L^{2}(M;TM)\to H^{1}(M;TM)$ and the properties of the index give that
$\operatorname{ind}_{H^{1}}\Big(\frac{1}{2}+\boldsymbol{K}\Big)+\operatorname{ind}_{L^{2}}(\boldsymbol{S})\,=\,\operatorname{ind}_{L^{2}}(\boldsymbol{S})+\operatorname{ind}_{L^{2}}\Big(\frac{1}{2}+\boldsymbol{K}^{*}\Big)\,$,
and hence $\operatorname{ind}_{H^{1}}\big(\frac{1}{2}+\boldsymbol{K}\big)=\operatorname{ind}_{L^{2}}\big(\frac{1}{2}+\boldsymbol{K}^{*}\big)$. On the other hand, Corollary 15.4.18 in [63] implies that the indices above do not depend on the corresponding spaces. Thus, we have $\operatorname{ind}_{L^{2}}\big(\frac{1}{2}+\boldsymbol{K}^{*}\big)=-\operatorname{ind}_{L^{2}}\big(\frac{1}{2}+\boldsymbol{K}\big)=-\operatorname{ind}_{H^{1}}\big(\frac{1}{2}+\boldsymbol{K}\big)$, and hence $\operatorname{ind}_{H^{1}}\big(\frac{1}{2}+\boldsymbol{K}\big)=\operatorname{ind}_{L^{2}}\big(\frac{1}{2}+\boldsymbol{K}^{*}\big)=-\operatorname{ind}_{H^{1}}\big(\frac{1}{2}+\boldsymbol{K}\big)$. Therefore $\operatorname{ind}_{H^{1}}\big(\frac{1}{2}+\boldsymbol{K}\big)=\operatorname{ind}_{L^{2}}\big(\frac{1}{2}+\boldsymbol{K}\big)=0$, as claimed.
To complete the proof of our result, it is enough to determine the image of $\frac{1}{2}+\boldsymbol{K}$. This property will be obtained in the next four steps.
Step 2 (Single layer $U:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$). To prove that $\frac{1}{2}+\boldsymbol{K}$ is onto, it suffices to show that $\ker(\frac{1}{2}+\boldsymbol{K}^{*})=0$. To this end, let $\boldsymbol{h}\in L^{2}(\Gamma;TM)$ be such that $(\frac{1}{2}+\boldsymbol{K}^{*})\boldsymbol{h}=0$. The ellipticity of $\frac{1}{2}+\boldsymbol{K}^{*}$, the fact that $\Gamma$ has straight cylindrical ends, and the elliptic regularity property ([67, Proposition 6.5]) imply that $\boldsymbol{h}\in H^{s}(\Gamma;TM)$ for all $s\in\mathbb{R}$. Let now the single layer potential $\mathcal{S}_{\rm{ST}}\boldsymbol{h}$. In view of Proposition 8.9 we obtain that the restrictions of $U$ to $\Omega_{+}:=\Omega$ and to $\Omega_{-}$ satisfy
(10.4) $U\mathrel{\,:=\,}(\,\boldsymbol{u}\ \ p\,)^{\top}\mathrel{\,:=\,}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})\in H^{2}(\Omega_{\pm};TM)\oplus H^{1}(\Omega_{\pm})\,.$
We next study the restrictions of $U$ on both domains $\Omega_{+}$ and $\Omega_{-}$.
Step 3 (Study of $U$ on $\Omega_{-}$). Let $U:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$ be as in the previous step. Proposition 8.10 implies that $\boldsymbol{\Xi}U=0$ in $\Omega_{-}$, and Theorem 8.16 gives that $\boldsymbol{T}_{\boldsymbol{\nu}}U]_{-}=(\frac{1}{2}+\boldsymbol{K}^{*})\boldsymbol{h}=0$. (Recall that $\boldsymbol{T}_{\boldsymbol{\nu}}U]_{-}$ is the trace at $\Gamma$ of $\boldsymbol{T}_{\boldsymbol{\nu}}U$ from the domain $\Omega_{-}$). Because $V$ does not vanish identically in any connected component of $\Omega_{-}$, we obtain $\boldsymbol{u}=0$ in $\Omega_{-}$, by Corollary 2.14. The same corollary gives that $p$ is constant in all connected components of $\Omega_{-}$. Moreover, the relation
$0\,=\,[\boldsymbol{T}_{\boldsymbol{\nu}}U]_{-}\mathrel{\,:=\,}[-2\boldsymbol{D}_{\boldsymbol{\nu}}\boldsymbol{u}+p\boldsymbol{\nu}]_{-}\,=\,p_{-}\boldsymbol{\nu}$
on $\Gamma$ gives that $p_{-}=0$, and hence $p=0$ on $\Omega_{-}$, since $p$ is constant in all connected components of $\Omega_{-}$.
Step 4 (Study of $U$ on $\Omega_{+}$). Let $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$, as in the previous two steps. Recall that $\boldsymbol{\Xi}U=0$ in $M\smallsetminus\Gamma$. Theorem 8.16 gives the “no-jump relation” $\boldsymbol{u}_{+}=\boldsymbol{u}_{-}=0$ at the boundary $\Gamma:=\partial\Omega$ of $\Omega$ (interior and exterior traces). The fact that $\boldsymbol{u}_{+}=0$ at $\Gamma$ allows us to use Corollary 2.14 that gives then that $\boldsymbol{u}=0$ in $\Omega$. Because $\Omega$ is connected and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, we also have $V_{0}\succ 0$ on $\Omega$. Therefore, the same corollary implies that $p$ is constant on $\Omega$. Because $V_{0}$ does not vanish identically on $\Omega$ (as we have just noticed) we even obtain that $p=0$ in $\Omega$ and $p_{+}=0$. Recalling that we have already proved that $\boldsymbol{u}=0$ in $\Omega_{-}$, we see that $\boldsymbol{u}=0$ in $M\smallsetminus\Gamma$ and hence $\boldsymbol{D}_{\boldsymbol{\nu}}\boldsymbol{u}=0$ in $M\smallsetminus\Gamma$.
Step 5 (Injectivity). The definitions of $\boldsymbol{T}_{\boldsymbol{\nu}}$ and $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$ then give $\boldsymbol{T}_{\boldsymbol{\nu}}U:=-2\boldsymbol{D}_{\boldsymbol{\nu}}\boldsymbol{u}+p\boldsymbol{\nu}\,=\,p\boldsymbol{\nu}$ on $M\smallsetminus\Gamma$, and
(10.5) $\boldsymbol{h}\,=\,[\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})]_{-}-[\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})]_{+}\,=\,(p_{-}-p_{+})\boldsymbol{\nu}\,=\,0\,.$
Therefore the operator $\frac{1}{2}+\boldsymbol{K}^{*}:L^{2}(\Gamma;TM)\to L^{2}(\Gamma;TM)$ is injective implying that the Fredholm operator of index zero $\frac{1}{2}+\boldsymbol{K}:L^{2}(\Gamma;TM)\to L^{2}(\Gamma;TM)$ is onto and hence an isomorphism, as asserted.
The fact that $(\frac{1}{2}+\boldsymbol{K})^{-1}\in\Psi^{0}(\Gamma;TM)$ follows from Theorem 5.4.11 in [63] (see also Theorem 7.4 in [67]). This completes the proof. ∎
In all the results above, the assumptions that $M$ and $\Omega$ are connected do not really decrease the generality. Note, however, that we are not assuming $\Omega_{-}:=M\smallsetminus\overline{\Omega}$, $\Omega^{\prime}$, or $M^{\prime}$ to be connected.
11. The Dirichlet problem for the generalized Stokes system via layer potentials
In this section, we present two applications of the results of the previous section. First, we obtain the well-posedness result of the Dirichlet problem for the generalized Stokes system on a smooth domain $\Omega$ with straight cylindrical ends. Then, we use this well-posedness result to show the no-jump property of the co-normal derivative $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{h})$ across the boundary of the domain $\Omega$.
11.1. The homogeneous Stokes system
The invertibility of each of the operators $\frac{1}{2}+\boldsymbol{K}$ and $\boldsymbol{S}$ gives the following well-posedness result for the modified Stokes system with Dirichlet boundary condition, problem (1.4). This result is known for the usual Stokes operator (i.e. when $V=0$ and $V_{0}=0$, see [100, Proposition 10.5.1, Theorem 10.6.2] in the case of a bounded Lipschitz domain in $\mathbb{R}^{n}$, $n\geq 2$, and [39, Theorem 5.1] in the case of a $C^{1}$ domain on a compact manifold, see also [63] for further related applications). We consider the homogeneous Dirichlet problem
(11.1) $\boldsymbol{\Xi}U\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}U\,=\,0\ \mbox{ in }\ \Omega\quad\mbox{and}\quad\boldsymbol{u}\,=\,\boldsymbol{f}\ \mbox{ on }\ \Gamma\mathrel{\,:=\,}\partial\Omega\,.$
The following result gives the well-posedness of this problem and the representability of its solutions under global assumptions on $M$ for $V$ and $V_{0}$. Recall Assumption 8.3.
Proposition 11.1.
Let $M$, $V$, and $V_{0}$ satisfy Assumption 8.3 (so, in particular, $M$ is has straight cylindrical ends) and let $\Omega\subset M$ be a compatible smooth domain (so $\Omega$ also has straight cylindrical ends, see Equation (1.2)). We assume that $V_{0}\succ 0$ on $\Omega$, that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, and that $\Omega$ is on one side of its boundary $\Gamma:=\partial{\Omega}$. We further assume one of the following conditions:
  1. (a)
    $V\succ 0$ on $\Omega_{-}$ and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ or
  2. (b)
    $V_{0}\succ 0$ on $\Omega_{-}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega_{-}^{\prime}$.
Then, for every $m\in\mathbb{Z}_{+}$ and any $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$, the homogeneous Diriechlet problem (11.1) has a unique solution $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$. Moreover, there exists a constant $C_{m}\geq 0$ such that, for any $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$, the unique solution $U$ of this problem satisfies
$\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}+\|p\|_{H^{m}(\Omega)}\leq C_{m}\|\boldsymbol{f}\|_{H^{m+1/2}(\Gamma;TM)}\,.$
Furthermore, according to which assumption is satisfied ((a) or (b)), the solution $U$ can be represented by at least one of the formulas
  • $U\mathrel{\,:=\,}\mathcal{D}_{\rm{ST}}\Big(\Big(\frac{1}{2}+\boldsymbol{K}\Big)^{-1}\boldsymbol{f}\Big)$, if (a) is satisfied, respectively
  • $U\mathrel{\,:=\,}\mathcal{S}_{\rm{ST}}\left(\boldsymbol{S}^{-1}\boldsymbol{f}\right)$, if (b) is satisfied.
Proof.
Let us check that the solution $U$ is unique. To this end, let $U_{0}=(\boldsymbol{u}_{0}\ \ p_{0})^{\top}$ satisfies the homogeneous Stokes system in $\Omega$ (i.e., $\boldsymbol{\Xi}U_{0}=0$) and the homogeneous Dirichlet condition on $\Gamma:=\partial\Omega$ (i.e., $\boldsymbol{u}_{0}=0$ on $\Gamma$). Because $V_{0}\not\equiv 0$ on $\Omega$ and $\boldsymbol{u}_{0}=0$ on $\partial\Omega$, Corollary 2.14 then ensures that $U_{0}=0$ in $\Omega$, which implies the desired uniqueness.
Our assumption (a) implies that the operator $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi^{0}(\Gamma;TM)$, by Theorem 10.8. Moreover, the elliptic regularity of the operator $\frac{1}{2}+\boldsymbol{K}$ (Theorem 8.14) implies that $\boldsymbol{h}:=(\frac{1}{2}+\boldsymbol{K})^{-1}\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$. Then $U:=\mathcal{D}_{\rm{ST}}(\boldsymbol{h})$ satisfies the required properties, including its boundedness with respect to the given datum $\boldsymbol{f}$, with a constant $C_{m}$ given by $\|\mathcal{D}_{\rm{ST}}(\frac{1}{2}+\boldsymbol{K})^{-1}\|$.
If, on the other hand, condition (b) is satisfied, then the operator $\boldsymbol{S}$ is invertible on $L^{2}(\Gamma;TM)$ and $\boldsymbol{S}^{-1}\in\Psi^{1}(\Gamma;TM)$, by Theorem 10.4. Moreover, the elliptic regularity of the operator $\boldsymbol{S}:=\mathcal{A}_{0}$ (Theorem 8.16) implies that $\boldsymbol{h}:=\boldsymbol{S}^{-1}\boldsymbol{f}\in H^{m-1/2}(\Gamma;TM)$. Theorem 6.6 then gives
$U\mathrel{\,:=\,}\mathcal{S}_{\rm{ST}}(\boldsymbol{h})\mathrel{\,:=\,}\mathcal{S}_{\rm{ST}}(\boldsymbol{S}^{-1}\boldsymbol{f})\in H^{m+1}(M;TM)\oplus H^{m}(M)\,$,
and the estimate
$\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}+\|p\|_{H^{m}(\Omega)}\leq\|\mathcal{S}_{\rm{ST}}\boldsymbol{S}^{-1}\|\|\boldsymbol{f}\|_{H^{m+1/2}(\Gamma;TM)}\,.$
Moreover, Proposition 8.10 gives that $\boldsymbol{\Xi}U=0$ in $\Omega$, and
$\boldsymbol{u}|_{\Gamma}\mathrel{\,:=\,}\big[\mathcal{V}_{\rm{ST}}(\boldsymbol{h})\big]_{+}\,=\,\boldsymbol{S}\boldsymbol{h}\mathrel{\,:=\,}\boldsymbol{S}\boldsymbol{S}^{-1}\boldsymbol{f}\,=\,\boldsymbol{f}\,.$
Hence $U$ satisfies all required properties. This completes the proof of the theorem. ∎
This proposition then allows us to prove the main result of this paper.
Proof of Theorem 1.1.
Let us first prove the uniqueness result stated in Theorem 1.1. To this end, let $U_{0}=(\boldsymbol{u}_{0}\ \ p_{0})^{\top}$ be the difference of two solutions of our Dirichlet problem, i.e. $U_{0}$ satisfies the homogeneous Stokes system in $\Omega$ and the homogeneous Dirichlet condition on $\partial\Omega$. Because $V_{0}\succ 0$ on $\Omega$ and $\boldsymbol{u}_{0}=0$ on $\partial\Omega$, Corollary 2.14 gives that $U_{0}=0$ in $\Omega$, which implies the desired uniqueness.
Next we prove the existence part of Theorem 1.1. Given the potential $V$ satisfying the assumptions of this theorem, we can modify it on $\Omega_{-}$ such that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $V\succ 0$ on $\Omega_{-}$. Then our assumptions imply that the operator $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi^{0}(\Gamma;TM)$, by Theorem 10.8. Moreover, the elliptic regularity of the operator $\frac{1}{2}+\boldsymbol{K}$ (Theorem 8.14) implies that $\boldsymbol{h}:=(\frac{1}{2}+\boldsymbol{K})^{-1}\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$. Then $U_{1}:=\mathcal{D}_{\rm{ST}}(\boldsymbol{h})$ satisfies the required properties, including its boundedness with respect to the given datum $\boldsymbol{f}$, with a constant $C_{m}$ given by $\|\mathcal{D}_{\rm{ST}}(\frac{1}{2}+\boldsymbol{K})^{-1}\|$. This completes the proof. ∎
The analogous result for the Laplace operator was proved in [98], but using a different method.
Remark 11.2.
We notice that the solution $U$ of the above theorem can be represented either by a single or as a double layer potential, depending on how we modify $V$ and $V_{0}$ on $\Omega_{-}$. Indeed, modifying $V$ on $\Omega_{-}$ such that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $V\succ 0$ on $\Omega_{-}$, then the assumptions of Theorem 10.8 are satisfied and the solution $U$ can be represented as a double layer potential
$U\mathrel{\,:=\,}\mathcal{D}_{\rm{ST}}\Big(\Big(\frac{1}{2}+\boldsymbol{K}\Big)^{-1}\boldsymbol{f}\Big)\,.$
On the other hand, modifying $V$ and $V_{0}$ on $\Omega_{-}$ such that the conditions of Theorem 10.4 be satisfied, the solution $U$ can be represented as a single layer potential
$U\mathrel{\,:=\,}\mathcal{S}_{\rm{ST}}\left(\boldsymbol{S}^{-1}\boldsymbol{f}\right)\,.$
Remark 11.3.
The constant $\mathcal{C}_{m}$ in Theorem 1.1 is independent of the given datum $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$. Indeed, it depends only on the norms of the operators $\mathcal{D}_{\rm{ST}}$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}$ and can be chosen the product of them.
11.2. The non-homogeneous Stokes system
Let us now prove the well-posedness result for the non-homogeneous problem using Theorems 1.1 and 8.4.
Theorem 11.4.
Let the assumptions of Theorem 1.1 hold. Then, for every $m\in\mathbb{Z}_{+}$, there exists a constant ${\mathcal{C}}_{m}\in(0,\infty)$ such that, for all $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$, $r\in H^{m-1}(\Omega)$, and $\boldsymbol{h}\in H^{m-1}(\Omega;TM)$, the non-homogeneous Dirichlet problem
(11.2) $\boldsymbol{\Xi}U\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}U\,=\,\left(\begin{array}[]{c}{\boldsymbol{h}}\\ {r}\end{array}\right)\ \mbox{ in }\Omega\,,\ \mbox{ and }\ \boldsymbol{u}|_{\Gamma}\,=\,\boldsymbol{f}\,$,
has a unique solution $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ and this solution satisfies
$\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}+\|p\|_{H^{m}(\Omega)}\leq{\mathcal{C}}_{m}\left(\|\boldsymbol{h}\|_{H^{m-1}(\Omega;TM)}+\|r\|_{H^{m-1}(\Omega)}+\|\boldsymbol{f}\|_{H^{m+1/2}(\Gamma;TM)}\right)\,.$
Proof.
There exists $C_{1,m}\geq 1$ such that, for all $\boldsymbol{h}\in H^{m-1}(\Omega;TM)$ and $\boldsymbol{h}\in H^{m}(\Omega)$ there are ${\boldsymbol{H}}\in H^{m-1}(M;TM)$ and $R\in H^{m}(M)$ such that $\boldsymbol{H}|_{\Omega}\,=\,\boldsymbol{h}$, $R|_{\Omega}=r$, and
$\|\boldsymbol{H}\|_{H^{m-1}(M)}+\|R\|_{H^{m}(M)}\leq C_{1,m}\big(\|\boldsymbol{h}\|_{H^{m-1}(\Omega;TM)}+\|r\|_{H^{m}(\Omega)}\big).$
Then the invertibility of the generalized Stokes operator $\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}$, Theorem 8.4, implies that there exists a unique $U_{1}=(\boldsymbol{u}_{1}\ \ p_{1})^{\top}\in H^{m+1}(M;TM)\oplus H^{m}(M)$ such that $\boldsymbol{\Xi}U_{1}=(\boldsymbol{H}\ \ R)^{\top}$ in $M$ and there exists a constant $C_{0;m}>0$ (which does not depend on $\boldsymbol{h}$ and $r$) such that
$\|U_{1}\|_{H^{m+1}(M;TM)\oplus H^{m}(M)}\leq C_{0;m}\big(\|\boldsymbol{H}\|_{H^{m}(M;TM)}+\|R\|_{H^{m}(M)}\big).$
Consequently,
$\boldsymbol{\Xi}U_{1}\,=\,\left(\begin{array}[]{c}{\boldsymbol{h}}\\ {r}\end{array}\right)\ \mbox{ in }\Omega\,.$
Now let $\boldsymbol{f}_{1}:=\boldsymbol{f}-\boldsymbol{u}_{1}|_{\Gamma}$. Hence $\boldsymbol{f}_{1}\in H^{m+\frac{1}{2}}(\Gamma;TM)$ and Theorem 1.1 implies that the Dirichlet problem
$\boldsymbol{\Xi}U_{2}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}U_{2}\,=\,0\ \mbox{ in }\Omega\ \mbox{ and }\ \boldsymbol{u}_{2}|_{\Gamma}\,=\,\boldsymbol{f}_{1}$
has a unique solution $U_{2}=(\boldsymbol{u}_{2}\ \ p_{2})^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ and
$\|\boldsymbol{u}_{2}\|_{H^{m+1}(\Omega;TM)}$ $+\|p_{2}\|_{H^{m}(\Omega)}\leq C_{m}\|\boldsymbol{f}_{1}\|_{H^{m+1/2}(\Gamma;TM)}$
$\leq C_{2;m}\left(\|\boldsymbol{f}\|_{H^{m+1/2}(\Gamma;TM)}+\|\boldsymbol{h}\|_{H^{m-1}(\Omega;TM)}+\|r\|_{H^{m-1}(\Omega)}\right)\,$,
with some constants $C_{m},C_{2;m}>0$ independent of $\boldsymbol{f}$, $r$ and $\boldsymbol{h}$.
Finally, let $U:=U_{1}+U_{2}$, $U=(\boldsymbol{u}\ \ p)^{\top}$. The above properties of $U_{1}$ and $U_{2}$ imply then that $U\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ satisfies the required properties. The uniqueness of $U$ with these properties follows from Theorem 1.1. ∎
A useful consequence of Theorem 11.4 is the following corollary.
Corollary 11.5.
Let the assumptions of Theorem 1.1 hold. Then, for every $m\in\mathbb{Z}_{+}$, there exists a constant ${\mathcal{C}}_{m}\in(0,\infty)$ such that, for any $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$ and any $\boldsymbol{h}\in H^{m-1}(\Omega;TM)$, the non-homogeneous Dirichlet problem (11.2) has a unique solution $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ and this solution satisfies
$\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}+\|p\|_{H^{m}(\Omega)}\leq{\mathcal{C}}_{m}\left(\|\boldsymbol{h}\|_{H^{m-1}(\Omega;TM)}+\|\boldsymbol{f}\|_{H^{m+1/2}(\Gamma;TM)}\right)\,.$
Let $U$ is the unique solution of Problem (11.2) for the given data $(\boldsymbol{h}\ \ \boldsymbol{f})^{\top}$ and let ${\mathfrak{A}}(\boldsymbol{h}\ \ \boldsymbol{f})^{\top}:=U$. In particular, the map
${\mathfrak{A}}:H^{m-1}(\Omega;TM)\oplus H^{m+1/2}(\Gamma;TM)\to H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega),$
is linear and continuous.
11.3. Dirichlet-to-Neumann operator and boundary behavior of $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}$
Let assumptions of Theorem 1.1 hold. Let $\mathcal{N}_{\rm{ST}}:H^{3/2}(\partial\Omega;TM)\to H^{1/2}(\partial\Omega;TM)$ be the Dirichlet-to-Neumann operator defined as follows (see [131, p.37] in the case of the Laplace operator on a compact manifold). For $\boldsymbol{f}\in H^{3/2}(\partial\Omega;TM)$ arbitrary, let $U:=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\Omega;TM)\oplus H^{1}(\Omega)$ be the unique solution of the Dirichlet problem (11.1) (that is, $\boldsymbol{\Xi}_{V,V_{0}}U\,=\,0$ in $\Omega$ and $\boldsymbol{u}|_{\Gamma}\,=\,\boldsymbol{f}$ on $\Gamma:=\partial\Omega$), see Theorem 1.1. Then we define
(11.3) $\mathcal{N}_{\rm{ST}}\boldsymbol{f}\mathrel{\,:=\,}[\boldsymbol{T}_{\boldsymbol{\nu}}U]_{+}\ \mbox{ on }\Gamma\,$,
the limit being evaluated from $\Omega$.
The next result shows that there is no jump of the conormal derivative $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})$ across $\Gamma:=\partial\Omega$ (see [131, Proposition 11.4] in the case of the Laplace operator).
Theorem 11.6.
Under the assumptions of Theorem 1.1, there is no jump across $\Gamma$ of $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})$. More precisely, for $\boldsymbol{f}\in H^{3/2}(\Gamma;TM)$,
$\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})]_{+}\,=\,\boldsymbol{T}_{\boldsymbol{\nu}}[\mathcal{D}_{\rm{ST}}(\boldsymbol{f})]_{-}\,=\,\left(\frac{1}{2}+{\boldsymbol{K}}^{*}\right)\mathcal{N}_{\rm{ST}}\boldsymbol{f}\,.$
Proof.
The second relation of Proposition 8.11 implies that
[…]
where $U$ is the unique solution of the Dirichlet problem (11.1) with $\boldsymbol{f}\in H^{3/2}(\Gamma;TM)$.
Now considering the vector part of identity (11.4), taking the limit of (11.4) on $\Gamma$ from $\Omega$, and using Theorem 8.14 and Theorem 8.16, we obtain
$\left(\frac{1}{2}+\boldsymbol{K}\right)\boldsymbol{f}-\boldsymbol{S}(\mathcal{N}_{\rm{ST}}\boldsymbol{f})=\boldsymbol{u}|_{\Gamma}=\boldsymbol{f}\,$,
and hence
$\boldsymbol{S}(\mathcal{N}_{\rm{ST}}\boldsymbol{f})\,=\,\left(-\frac{1}{2}+\boldsymbol{K}\right)\boldsymbol{f}\,.$
Thus we obtain the identity
$\boldsymbol{S}\mathcal{N}_{\rm{ST}}\,=\,-\frac{1}{2}+\boldsymbol{K}\,.$
(The same identity follows if we take the limit of the vector part of (11.4) on $\Gamma$ from $\Omega_{-}:=M\setminus\overline{\Omega}$ and use again Theorem 8.14 and Theorem 8.16.) This identity and the ellipticity of the operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ (see also Theorem 8.14 and Theorem 8.16) imply that $\mathcal{N}_{\rm{ST}}$ is elliptic as well.
Next we apply the operator $\boldsymbol{T}_{\boldsymbol{\nu}}$ to both sides of identity (11.4). Evaluating it first from $\Omega$, we obtain
(11.5) $[\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})]_{+}-[\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\rm{ST}}(\mathcal{N}_{\rm{ST}}\boldsymbol{f})]_{+}=[\boldsymbol{T}_{\boldsymbol{\nu}}U]_{+}=\mathcal{N}_{\rm{ST}}\boldsymbol{f}\,$,
while evaluating from $M\setminus\overline{\Omega}$ implies
(11.6) $[\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})]_{-}-[\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\rm{ST}}(\mathcal{N}_{\rm{ST}}\boldsymbol{f})]_{-}=0\,.$
Since both limits $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{S}_{\rm{ST}}(\mathcal{N}_{\rm{ST}}\boldsymbol{f})]_{\pm}$ exist, by Theorem 8.16, formulas (11.5) and (11.6) show that the limits $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})_{\pm}$ exist as well, and they are given by
$[\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})]_{+}\,=\,\mathcal{N}_{\rm{ST}}\boldsymbol{f}+\left(-\frac{1}{2}+{\boldsymbol{K}}^{*}\right)\mathcal{N}_{\rm{ST}}\boldsymbol{f}\,$,
$[\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})]_{-}\,=\,\left(\frac{1}{2}+{\boldsymbol{K}}^{*}\right)\mathcal{N}_{\rm{ST}}\boldsymbol{f}\,$,
and hence
$\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})]_{\pm}\,=\,\left(\frac{1}{2}+{\boldsymbol{K}}^{*}\right)\mathcal{N}_{\rm{ST}}\boldsymbol{f}\,.$
This completes the proof. ∎
12. Applications and extensions
We now consider an application and an extension of our results so far.
12.1. The Dirichlet problem for the generalized Navier-Stokes system
The first part is of this section is devoted to the Dirichlet problem for a Navier-Stokes type system on a manifold with cylindrical ends. We analyze the Dirichlet problem when the norms of the given data (the right hand side of the Navier-Stokes equation and the Dirichlet datum) are restricted to satisfy a boundedness condition which ensures the well-posedness of the Dirichlet problem.
A key role in the analysis of the nonlinear Dirichlet problem for the Navier-Stokes system on a domain with cylindrical ends is played by estimates of the product of two functions in suitable Sobolev spaces. An example is the following result, which is well known in $\mathbb{R}^{n}$ (see e.g. [17, Theorem 7.3, Theorem 8.2], [114, §4.5.2, Theorem], and [90, Lemma 8.1]).
Lemma 12.1.
Let $m\in\mathbb{Z}_{+}$ and $M$ be as in Assumption 8.3. Assume that $2\leq\operatorname{dim}\,M=n<2(m+2)$. Then there exists $C_{m}>0$ with the following property: for all $u\in H^{m+1}(M)$ and $v\in H^{m}(M)$, we have $uv\in H^{m-1}(M)$ and
$\|uv\|_{H^{m-1}(M)}\leq C_{m}\|u\|_{H^{m+1}(M)}\|v\|_{H^{m}(M)}\,.$
Proof.
This follows from the analogous result on $\mathbb{R}^{n}$ just mentioned above using a dyadic partition of unity (see, for instance [63]). ∎
We notice that the same argument works for manifolds with bounded geometry. Also, notice that the result of Lemma 12.1 extends to Sobolev spaces on sections of vector bundles over $M$ (see e.g. [114, §4.2.2, Proposition] in the Euclidean setting), and the next lemma is a consequence of this result (see [39, Lemma 7.5] in the case of a Lipschitz domain on a compact manifold).
Lemma 12.2.
Let $M$ be as in Assumption 8.3 and $\Omega\subset M$ be as in Theorem 1.1, that is, $\Omega$ is a smooth domain with straight cylindrical ends in the connected manifold with straight cylindrical ends $M$. Assume that $m\in\mathbb{Z}_{+}$ and that ${\rm{dim}}\,M=n<2(m+2)$. Then there exists a constant $C>0$, such that
(12.1) $\|\nabla_{\boldsymbol{u}}\boldsymbol{v}\|_{H^{m-1}(\Omega;TM)}\leq C\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}\|\boldsymbol{v}\|_{H^{m+1}(\Omega;TM)},$
(12.2) $\|\nabla_{\boldsymbol{u}}\boldsymbol{u}-\nabla_{\boldsymbol{v}}\boldsymbol{v}\|_{H^{m-1}(\Omega;TM)}$
$\hskip 20.00003pt\leq C\left(\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}+\|\boldsymbol{v}\|_{H^{m+1}(\Omega;TM)}\right)\|\boldsymbol{u}-\boldsymbol{v}\|_{H^{m+1}(\Omega;TM)},$
for all $\boldsymbol{u},\boldsymbol{v}\in H^{m+1}(\Omega;TM)$.
Proof.
First, we notice that $\Omega$ is a manifold with bounded geometry. Indeed, $\Omega$ is a smooth domain with straight cylindrical ends in the smooth manifold with straight cylindrical ends $M$, and hence both $M$ and $\Omega$ have bounded geometry. Hence the result of Lemma 12.1 holds also on $\Omega$. Thus, for each $m\in\mathbb{Z}_{+}$ such that $n<2(m+2)$, there exists a constant $c_{m}>0$ (depending only on $n$ and the bounded geometry of $M$) such that the estimate (12.1) holds for all $u\in H^{m+1}(\Omega)$ and $v\in H^{m}(\Omega)$. Thus, we have $uv\in H^{m-1}(\Omega)$ and
$\|uv\|_{H^{m-1}(\Omega)}\leq c_{m}\|u\|_{H^{m+1}(\Omega)}\|v\|_{H^{m}(\Omega)}\,.$
(see [114, §4.2.2, Proposition] in $\mathbb{R}^{n}$). This result extends also to vector bundles and, thus, there exists a constant $C>0$ (depending only on $n$, $m$ and the bounded geometry of $M$) such that the estimate (12.1) holds for all $\boldsymbol{u},\boldsymbol{v}\in H^{m+1}(\Omega;TM)$.
Moreover, the linearity of the covariant derivative $\nabla_{\boldsymbol{u}}\boldsymbol{v}$ with respect to $\boldsymbol{u}$ and the additivity with respect to $\boldsymbol{v}$, and estimate (12.1) give
$\|\nabla_{\boldsymbol{u}}\boldsymbol{u}-\nabla_{\boldsymbol{v}}\boldsymbol{v}\|_{H^{m-1}(\Omega;TM)}$ $\leq\|\nabla_{\boldsymbol{u}-\boldsymbol{v}}\boldsymbol{u}\|_{H^{m-1}(\Omega;TM)}+\|\nabla_{\boldsymbol{v}}(\boldsymbol{u}-\boldsymbol{v})\|_{H^{m-1}(\Omega;TM)}$
$\leq C\left(\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}+\|\boldsymbol{v}\|_{H^{m+1}(\Omega;TM)}\right)\|\boldsymbol{u}-\boldsymbol{v}\|_{H^{m+1}(\Omega;TM)}$
for all $\boldsymbol{u},\boldsymbol{v}\in H^{m+1}(\Omega;TM)$, that is, the estimate (12.2). ∎
Let us now prove the well-posedness of the Dirichlet problem for the Navier-Stokes system under the assumption of small data (see Theorem 7.3 in [39] in the case of a $C^{1}$, or even Lipschitz domain on a compact Riemannian manifold).
Theorem 12.3.
Let $M$, $V$, and $V_{0}$ satisfy Assumption 8.3 and also the assumptions in Theorem 1.1. Let $\Omega\subset M$ be a smooth domain with straight cylindrical ends in the smooth connected manifold with straight cylindrical ends $M$. Assume also that $\Omega$ is the interior of $\overline{\Omega}$. If $m\in\mathbb{Z}_{+}$ and $2\leq{\rm{dim}}\,M=n<2(m+2)$, then there exist two constants $\zeta,\eta\in(0,\infty)$ such that for all given data $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$ and $\boldsymbol{h}\in H^{m-1}(\Omega;TM)$ satisfying the condition
(12.3) $\|\boldsymbol{h}\|_{H^{m-1}(\Omega;TM)}+\|\boldsymbol{f}\|_{H^{m+1/2}(\Gamma;TM)}\leq\zeta,$
the Dirichlet problem for the Navier-Stokes system
(12.4) $\boldsymbol{\Xi}U+\left(\begin{array}[]{c}{\nabla_{\boldsymbol{u}}\boldsymbol{u}}\\ {0}\end{array}\right)\,=\,\left(\begin{array}[]{c}{\boldsymbol{h}}\\ {0}\end{array}\right)\ \mbox{ in }\Omega\ \mbox{ and }\ \boldsymbol{u}|_{\Gamma}\,=\,\boldsymbol{f}$
has a unique solution $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ such that
(12.5) $\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}\leq\eta\,.$
Moreover, there exists a constant $C_{0}\in(0,\infty)$ such that
$\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}+\|p\|_{H^{m}(\Omega)}\leq C_{0}\left(\|\boldsymbol{h}\|_{H^{m-1}(\Omega;TM)}+\|\boldsymbol{f}\|_{H^{m+1/2}(\Gamma;TM)}\right)\,.$
Proof.
Let $\left(\boldsymbol{h}\ \ {\boldsymbol{f}}\right)^{\top}\in H^{m-1}(\Omega;TM)\oplus H^{m+1/2}(\Gamma;TM)$ be the given data that will be kept fixed in all arguments of the proof. Also, for a fixed $\boldsymbol{v}\in H^{m+1}(\Omega;TM)$, we consider the linear Dirichlet problem for the generalized Stokes system
(12.6) $\begin{cases}\boldsymbol{\Xi}U^{0}\,=\,\left(\begin{array}[]{c}{\boldsymbol{h}}\\ {0}\end{array}\right)-\left(\begin{array}[]{c}{\nabla_{\boldsymbol{v}}\boldsymbol{v}}\\ {0}\end{array}\right)\ \mbox{ in }\Omega\,,\\ \boldsymbol{u}^{0}|_{\Gamma}\,=\,\boldsymbol{f}\end{cases}$
with the unknown $U^{0}=(\boldsymbol{u}^{0}\ \ p^{0})^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$.
Lemma 12.2 gives that $\nabla_{\boldsymbol{v}}\boldsymbol{v}\in H^{m-1}(\Omega;TM)$. Corollary 11.5 then implies that the problem (12.6) with the given data $\left(\boldsymbol{h}-\nabla_{\boldsymbol{v}}\boldsymbol{v}\ \ \boldsymbol{f}\right)^{\top}\in H^{m-1}(\Omega;TM)\oplus H^{m+1/2}(\Gamma;TM)$ has a unique solution $U^{0}=(\boldsymbol{u}^{0}\ \ p^{0})^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$, which can be expressed in the form
(12.7) $\left(\boldsymbol{u}^{0}\ \ p^{0}\right)^{\top}={\mathfrak{A}}\left(\boldsymbol{h}-\nabla_{\boldsymbol{v}}\boldsymbol{v}\ \ \boldsymbol{f}\right)^{\top}=:\left({\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{v})\ \ {\mathcal{P}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{v})\right)^{\top},$
where ${\mathfrak{A}}:H^{m-1}(\Omega;TM)\oplus H^{m+1/2}(\Gamma;TM)\to H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ is the solution operator given by Corollary 11.5. This is a linear and continuous operator. Hence the nonlinear operator
(12.8) $\left({\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}\ \ {\mathcal{P}}_{\boldsymbol{h},{\boldsymbol{f}}}\right)^{\top}:H^{m+1}(\Omega;TM)\to H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$
satisfies the estimate
$\|{\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{v})\|_{H^{m+1}(\Omega;TM)}+\|{\mathcal{P}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{v})\|_{H^{m}(\Omega)}$
$\hskip 30.00005pt\leq{\mathcal{C}}_{m}\|\left(\boldsymbol{h}-\nabla_{\boldsymbol{v}}\boldsymbol{v}\ \ \boldsymbol{f}\right)^{\top}\|_{H^{m-1}(\Omega;TM)\oplus H^{m+1/2}(\Gamma;TM)}$
$\hskip 30.00005pt\leq{\mathcal{C}}_{m}\left(\|\left(\boldsymbol{h}\ \ \boldsymbol{f}\right)^{\top}\|_{H^{m-1}(\Omega;TM)\oplus H^{m+1/2}(\Gamma;TM)}+\|\nabla_{\boldsymbol{v}}\boldsymbol{v}\|_{H^{m-1}(\Omega;TM)}\right)$
(12.9) $\hskip 30.00005pt\leq{\mathcal{C}}_{m}\|\left(\boldsymbol{h}\ \ \boldsymbol{f}\right)^{\top}\|_{H^{m-1}(\Omega;TM)\oplus H^{m+1/2}(\Gamma;TM)}+{\mathcal{C}}_{m}C\|\boldsymbol{v}\|_{H^{m+1}(\Omega;TM)}^{2}\,$,
where ${\mathcal{C}}_{m}$ and $C$ are the constants provided by Corollary 11.5 and Lemma 12.2. Thus, the operator $\left({\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}\ \ {\mathcal{P}}_{\boldsymbol{h},{\boldsymbol{f}}}\right)^{\top}:H^{m+1}(\Omega;TM)\to H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ is continuous. In addition, by (12.6) and (12.7) we have
$\begin{cases}\boldsymbol{\Xi}\left(\begin{array}[]{c}{{\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{v})}\\ {{\mathcal{P}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{v})}\end{array}\right)\,=\,\left(\begin{array}[]{c}{\boldsymbol{h}}\\ {0}\end{array}\right)-\left(\begin{array}[]{c}{\nabla_{\boldsymbol{v}}\boldsymbol{v}}\\ {0}\end{array}\right)\ \mbox{ in }\Omega\,,\\ {\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{v})|_{\Gamma}\,=\,\boldsymbol{f}\,.\end{cases}$
Therefore, if we show that the nonlinear operator ${\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}$ has a fixed point $\boldsymbol{u}\in H^{m+1}(\Omega;TM)$, i.e., ${\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{u})=\boldsymbol{u}$, then $\boldsymbol{u}$ and the pressure $p={\mathcal{P}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{u})$ will determine a solution of the nonlinear problem (12.4) in $H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$. To prove this claim, we consider the constants
(12.10) $\zeta:=\frac{3}{16C{\mathcal{C}}_{m}^{2}}>0\ \ \mbox{ and }\ \ \eta:=\frac{1}{4C{\mathcal{C}}_{m}}>0$
and the closed ball of the space $H^{m+1}(\Omega;TM)$
(12.11) ${\bf B}_{\eta}:=\left\{\boldsymbol{w}\in H^{m+1}(\Omega;TM):\|\boldsymbol{w}\|_{H^{m+1}(\Omega;TM)}\leq\eta\right\},$
and we assume that the given data $\left(\boldsymbol{h}\ \ {\boldsymbol{f}}\right)^{\top}$ satisfy the inequality (12.3), that is,
$\|\left(\boldsymbol{h}\ \ {\boldsymbol{f}}\right)^{\top}\|_{H^{m-1}(\Omega;TM)\oplus H^{m+1/2}(\Gamma;TM)}=\|\boldsymbol{h}\|_{H^{m-1}(\Omega;TM)}+\|\boldsymbol{f}\|_{H^{m+1/2}(\Gamma;TM)}\leq\zeta\,.$
In view of (12.3), (12.1), (12.10), and (12.11) it follows that
(12.12) $\left\|\left({\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{w})\ \ {\mathcal{P}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{w})\right)^{\top}\right\|_{H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)}\leq{\frac{1}{4C\mathcal{C}}_{m}}=\eta\,,\quad\forall\,\boldsymbol{w}\in\mathbf{B}_{\eta}\,.$
Hence the nonlinear operator ${\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}$ maps ${\bf B}_{\eta}$ into ${\bf B}_{\eta}$.
Next we prove that ${\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}$ is a contraction on ${\bf B}_{\eta}$. Indeed, the expression of ${\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}$ in (12.7), the linearity and continuity of the operator $\mathfrak{A}$}, the estimate of Corollary 11.5, and inequality (12.1) of Lemma 12.2 give
$\left\|{\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{v})-{\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{w})\right\|_{H^{m+1}(\Omega;TM)}\leq\big\|{\mathfrak{A}}\left(\nabla_{\boldsymbol{v}}\boldsymbol{v}-\nabla_{\boldsymbol{w}}\boldsymbol{w}\ \ 0\right)^{\top}\big\|_{H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)}\\ \leq{\mathcal{C}}_{m}\|\nabla_{\boldsymbol{v}}\boldsymbol{v}-\nabla_{\boldsymbol{w}}\boldsymbol{w}\|_{H^{m-1}(\Omega;TM)}\\ \leq C{\mathcal{C}}_{m}\Big(\|\boldsymbol{v}\|_{H^{m+1}(\Omega;TM)}+\|\boldsymbol{w}\|_{H^{m+1}(\Omega;TM)}\Big)\|\boldsymbol{v}-\boldsymbol{w}\|_{H^{m+1}(\Omega;TM)}\\ \leq 2\eta C{\mathcal{C}}_{m}\|\boldsymbol{v}-\boldsymbol{w}\|_{H^{m+1}(\Omega;TM)}\,=\,\frac{1}{2}\|\boldsymbol{v}-\boldsymbol{w}\|_{H^{m+1}(\Omega;TM)}\,$,
for all $\boldsymbol{v},\boldsymbol{w}\in\mathbf{B}_{\eta}$. The Banach-Picard Fixed Point Theorem (see e.g. Theorem A.1.17 in [63]) gives that there exists a unique fixed point $\boldsymbol{u}\in\mathbf{B}_{\eta}$ of ${\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}$, i.e., ${\mathcal{U}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{u})=\boldsymbol{u}$. Moreover, $\boldsymbol{u}$ and the pressure function $p={\mathcal{P}}_{\boldsymbol{h},{\boldsymbol{f}}}(\boldsymbol{u})$ (where ${\mathcal{P}}_{\boldsymbol{h},{\boldsymbol{f}}}$ is given by (12.7)) determine a solution of the nonlinear problem (12.4) in the space $H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$. The Banach-Picard Fixed Point Theorem implies also the uniqueness of the solution under the assumptions (12.3) and (12.5).
The condition $\boldsymbol{u}\in{\bf B}_{\eta}$ and inequality (12.1) imply that
$\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}$ $+\|p\|_{H^{m}(\Omega)}$
$\leq{\mathcal{C}}_{m}\big\|\left(\boldsymbol{h}\ \ {\boldsymbol{f}}\right)^{\top}\big\|_{H^{m-1}(\Omega;TM)\oplus H^{m+1/2}(\Gamma;TM)}+\frac{1}{4}\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}$
and hence the estimate
$\|\boldsymbol{u}\|_{H^{m+1}(\Omega;TM)}+\|p\|_{H^{m}(\Omega)}\leq\frac{4}{3}{\mathcal{C}}_{m}\big\|\left(\boldsymbol{h}\ \ {\boldsymbol{f}}\right)^{\top}\big\|_{H^{m-1}(\Omega;TM)\oplus H^{m+1/2}(\Gamma;TM)}\,$,
as asserted. This completes the proof. ∎
12.2. The case $M$ closed
We now consider the extension of our results to the case of a closed manifold (unlike most of the other sections of this paper, in which we considered, for the most part, manifolds with straight cylindrical ends). Recall that a closed manifold is one that is smooth, compact and does not have a boundary. Almost all the results on manifolds with straight cylindrical ends hold also for closed manifolds (when they make sense).
Here are the main modifications:
  1. (1)
    There are no restrictions “or conditions” at infinity on our operators, functions, and vector bundles. For instance, $\mathcal{C}^{\infty}_{\operatorname{inv}}$ is replaced with $\mathcal{C}^{\infty}$, all open subsets and all vector bundles are compatible, and $\Psi_{\operatorname{ess}}^{m}$ is replaced with $\Psi^{m}$.
  2. (2)
    The restuls about the limit at infinity $\mathcal{R}_{\infty}$} do not make sense anymore (and are not needed). They have to be ignored in the closed case.
  3. (3)
    The same is true for the conditions involving $\mathcal{R}_{\infty}$}; they are to be removed, as they are no longer needed.
The case $M$ closed was considered in great detail in [69], so we will not go through the exercice of performing all the modifications abour. Instead, we recall here only two results from that paper. The first result describes the kernel of $\boldsymbol{\Xi}$ and does not follow from the results of this paper (directly):
Theorem 12.4.
Let us assume that $M$ is a closed manifold, that $V,V_{0}$ are smooth and non-negative and that $M$ is a smooth, compact manifold without boundary (i.e., a closed manifold). Then $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}:H^{1}(M;TM)\oplus L^{2}(M)\to H^{-1}(M;TM)\oplus L^{2}(M)$ is a self-adjoint Fredholm operator. Let $\mathcal{N}\subset\mathcal{C}^{\infty}(M;TM\oplus\mathbb{C})$ be defined by
  1. (1)
    $\mathcal{N}:=\{(\boldsymbol{u},p)\mid\operatorname{Def}\,\boldsymbol{u}=0\,,\ \nabla p=0\}$ if $V=0$ and $V_{0}=0$ on $M$;
  2. (2)
    $\mathcal{N}:=\{(\boldsymbol{u},0)\mid\operatorname{Def}\,\boldsymbol{u}=0\}$, if $V=0$ and $V_{0}\not\equiv 0$ on $M$;
  3. (3)
    $\mathcal{N}:=\{(0,p)\mid\nabla p=0\}$, if $V_{0}=0$ on $M$ and either $V\not\equiv 0$ on $M$ or $M$ does not have non-zero Killing vector fields;
  4. (4)
    $\mathcal{N}:=\{0\}$, if $V_{0}\not\equiv 0$ on $M$ and either $V\not\equiv 0$ on $M$ or $M$ does not have non-zero Killing vector fields.
Then the kernel of $\boldsymbol{\Xi}_{V,V_{0}}$ is given by $\ker\boldsymbol{\Xi}_{V,V_{0}}=\mathcal{N}$. Moreover, $\boldsymbol{\Xi}$ has a (unique) Moore-Penrose pseudoinverse $\boldsymbol{\Xi}^{(-1)}\in\Psi_{\operatorname{cl}}^{-\boldsymbol{s}-\boldsymbol{t}}(M;TM\oplus\mathbb{C})$, with $\boldsymbol{s}=\boldsymbol{t}=(1,0)$, whose image is $\mathcal{N}^{\perp}$.
The well-posedness result of the Dirichlet problem becomes
Theorem 12.5.
Let $\overline{\Omega}$ be a smooth, compact, connected manifold with boundary and let $V:\overline{\Omega}\to[0,\infty)$ be a smooth function that can be extended to a smooth, non-negative function on the double of $\overline{\Omega}$. We consider the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,0}$ (thus $V_{0}=0$ in $\Omega$). Let $\Gamma:=\partial\overline{\Omega}$, as before. Then, for every $m\in\mathbb{Z}_{+}$ and for any $\boldsymbol{f}\in H^{m+1/2}(\Gamma;T\overline{\Omega})$ such that $(\boldsymbol{f},\boldsymbol{\nu})_{\Gamma}=0$, there exists a solution $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{m+1}(\Omega;T\overline{\Omega})\oplus H^{m}(\Omega)$ of the Dirichlet problem
(12.13) $\boldsymbol{\Xi}U\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,0}U\,=\,0\ \mbox{ in }\Omega\ \mbox{ and }\ \boldsymbol{u}|_{\Gamma}\,=\,\boldsymbol{f}\ \mbox{ on }\partial\Omega\,.$
Any two solutions $U_{1}$ and $U_{2}$ of this Dirichlet problem differ by a constant scalar field: $U_{2}-U_{1}=(0\ \ \ c)^{\top}$, $c\in\mathbb{C}$. Also, any of the following two formulas provides a solution of this Dirichlet problem:
$U_{1}\mathrel{\,:=\,}\mathcal{D}_{\rm{ST}}\Big(\Big(\frac{1}{2}+\boldsymbol{K}\Big)^{(-1)}\boldsymbol{f}\Big)\quad\mbox{or}\quad\mbox{U}_{2}\mathrel{\,:=\,}\mathcal{S}_{\rm{ST}}\left(\boldsymbol{S}^{(-1)}\boldsymbol{f}\right)\,.$
Moreover, there exists a constant $C_{m}\geq 0$, independent of $\boldsymbol{h}$, such that all solutions $U=(\boldsymbol{u}\ \ p)^{\top}$ satisfy
$\|\boldsymbol{u}\|_{H^{m+1}(\Omega;T\overline{\Omega})}+\Big\|p-\frac{1}{\operatorname{vol}(\Omega)}\int_{\Omega}p\,\,d\operatorname{vol}\Big\|_{H^{m}(\Omega)}\leq C_{m}\|\boldsymbol{f}\|_{H^{m+1/2}(\Gamma;T\overline{\Omega})}\,.$

### 28 May 2026 v2

@@
 Under further positivity assumptions at infinity, we prove that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are also Fredholm.
-Under slightly stronger assumptions on $V$ and $V_{0}$ (including positivity everywhere), we prove the invertibility of the operators $\boldsymbol{\Xi}$, $\boldsymbol{S}$, and $\frac{1}{2}+\boldsymbol{K}$.
+Under slightly stronger assumptions on $V$ and $V_{0}$ (including non-negativity everywhere, i.e., $V,V_{0}\geq 0$), we prove the invertibility of the operators $\boldsymbol{\Xi}$, $\boldsymbol{S}$, and $\frac{1}{2}+\boldsymbol{K}$.
 As it is well-known, the invertibility of these operators leads to well-posedness results for the associated (linear) Stokes boundary value problem with Dirichlet boundary conditions on domains with straight cylindrical ends.
@@
 1. Introduction
-We prove the well-posedness of the Dirichlet boundary value problem for a generalized Stokes system (1.4) on a domain with straight cylindrical ends $\Omega$ using the method of layer potentials.
-In the process, we obtain many other results of independent interest on the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (1.3), on related differential operators, and on their associated layer potentials.
+We prove the well-posedness of the Dirichlet boundary value problem for a generalized Stokes system
+[…]
+for a domain with straight cylindrical ends $\Omega$ using the method of layer potentials.
+(Here $\nabla$ is the gradient, $\boldsymbol{L}$ is the deformation Laplacian, and $V$ and $V_{0}$ are potentials, see the next subsection.
+See also Equation (1.5) for a reformulation of this system using the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ (1.4).) In the process, we obtain many other results of independent interest on the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$, on related differential operators, and on their associated layer potentials.
 We carefully develop the needed concepts and methods and prove in detail our results.
-Before stating in more detail our main results, let us introduce the setting of our paper.
-1.1. The setting of our well-posedness results
-We will work on manifolds and domains with straight cylindrical ends (see the next figure, Figure 1, as well as Figure 2 at the beginning of Section 6).
+Before stating our main results, let us first introduce the setting of our paper.
+1.1. The setting of our well-posedness results: Domains and manifolds with straight cylindrical ends
+We will work on manifolds $M$ and domains $\Omega$ with straight cylindrical ends (see the next figure, Figure 1, as well as Figure 2 at the beginning of Section 6 for pictorial description of our setting).
 Figure 1. A boundaryless manifold with straight cylindrical ends.
-More precisely, by a manifold with straight cylindrical ends $M$ we will mean a non-compact Riemannian manifold without boundary (i.e., boundaryless) isometric to one of the form
+1.1.1. Manifolds with straight cylindrical ends
+Intuitively, a manifold with straight cylindrical ends is one that, outside a bounded set, it is a (half-)infinite cylinder $M^{\prime}\times(-\infty,R]$ with a product metric, where $M^{\prime}$ is a closed manifold (i.e.
+a smooth, compact manifold without boundary).
+More precisely, by a manifold with straight cylindrical ends $M$ we will mean a non-compact Riemannian manifold without boundary (i.e., “boundaryless”) isometric to one of the form
 (1.2) $M\,=\,M_{0}\cup\big[M^{\prime}\times(-\infty,R_{M}]\big]\,$,
 where
-  • $(M_{0},g_{0})$ is a compact Riemannian manifold,
+  • $(M_{0},g_{0})$ is a compact Riemannian manifold with boundary $\partial M_{0}$,
   • $M^{\prime}:=\partial M_{0}\neq\emptyset$ has induced metric $g_{\partial}$ and is identified with $M^{\prime}\times\{0\}$, and
   • the cylindrical end $M^{\prime}\times(-\infty,R_{M})$ is endowed with the product metric $g=g_{\partial}+(dt)^{2}$, $t\in(-\infty,R_{M})$,
-(see, for instance, [63, 67, 68]).
+(where we followed the conventions in [83, 87, 88]; but see also [23, 27, 29, 30, 54, 56, 94], where manifolds with cylindrical ends arise in the setting of “wave-guides”).
 There is no loss of generality to assume that $M$ is connected, so we will do that throughout the paper, in order to simplify our statements.
 (Also, we may assume $R_{M}=0$, which we will often do.)
-Second, $\Omega$ will be a smooth domain with straight cylindrical ends in $M$, in the sense that
+1.1.2. Domains with straight cylindrical ends
+In fact, we are more interested in the smooth domain with straight cylindrical ends $\Omega$, on which our boundary value problem (1.5) is formulated.
+Often $\Omega$ will be contained naturally in a manifold with straight cylindrical ends $M$ in a compatible way, in the sense that
 (1.3) $\Omega\cap\big[M^{\prime}\times(-\infty,R_{\Omega})\big]\,=\,\Omega^{\prime}\times(-\infty,R_{\Omega})$
@@
 This implies that $\Omega^{\prime}$ is also on one side of its boundary $\Gamma^{\prime}:=\partial\Omega^{\prime}\subset M^{\prime}$.
-The generalized Stokes operators $\boldsymbol{\Xi}_{V,V_{0}}$ are defined as follows (see Section 2 for more details, as well as for the concepts not recalled here). Let
+1.1.3. The Dirichlet boundary value problem for the generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$
+Our main interest is, in fact, in the Dirichlet boundary value problem for a generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$, this problem is formulated below, see Equation (1.5).
+The generalized Stokes operators $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ are defined as follows.
+Let
 $\operatorname{Def}\,:\mathcal{C}^{\infty}(M;TM)\to\mathcal{C}^{\infty}(M;T^{*}M\otimes T^{*}M)$
-be the deformation operator, (see formula (2.1) and the discussion around it for more details), and $\boldsymbol{L}:=2\operatorname{Def}^{*}\,\operatorname{Def}\,$ be the “Deformation Laplacian.” Let $\operatorname{End}(E)$ be the set of linear maps $E\to E$ (the endomorphisms of $E$).
-For suitable “potentials” $V\in\mathcal{C}^{\infty}(M;\operatorname{End}(TM))$ and $V_{0}\in\mathcal{C}^{\infty}(M)$, we consider the operator
+be the deformation operator (see formula (2.1) for the definition of $\operatorname{Def}\,$ and, in general, see Section 2 for more details).
+Let then $\boldsymbol{L}:=2\operatorname{Def}^{*}\,\operatorname{Def}\,$ be the “Deformation Laplacian.” Let $\operatorname{End}(E)$ be the set of linear maps $E\to E$ (the endomorphisms of $E$).
+For suitable “potentials” $V\in\mathcal{C}^{\infty}(M;\operatorname{End}(TM))$ and $V_{0}\in\mathcal{C}^{\infty}(M)$ that are translation invariant in a neighborhood of infinity (Definition 3.14), we consider the operator
 (1.4) $\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}\mathrel{\,:=\,}\left(\begin{array}[]{ccc}\boldsymbol{L}+V&\nabla\\ \nabla^{*}&-V_{0}\end{array}\right):\mathcal{C}^{\infty}(M;TM\oplus\mathbb{C})\to\mathcal{C}^{\infty}(M;TM\oplus\mathbb{C})\,.$
-Let
+This operator extends to Sobolev spaces in distribution sense. Let
 $U\mathrel{\,:=\,}\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)\in L^{2}(\Omega;TM\oplus\mathbb{C})\,$,
-so that $\boldsymbol{u}$ is the “vector part” of $U$. We then consider the Dirichlet boundary value problem
+so that $\boldsymbol{u}$ is the “vector part” of $U$. We then consider the Dirichlet boundary value problem, Equation (1.1), which, in this notation, becomes
 (1.5) $\begin{cases}\boldsymbol{\Xi}U\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}U\,=\,0&\mbox{ in }\Omega\\ \boldsymbol{u}\,=\,\boldsymbol{h}&\mbox{ on }\partial\Omega\,.\end{cases}$
@@
 1.2. Main results
-Our main results are for manifolds with straight cylindrical ends $M$ and for compatible open subsets $\Omega\subset M$ (with boundary $\Gamma$, as above) or for closed manifolds (i.e., compact, smooth, boundaryless manifolds).
-The case of closed manifolds was treated in great detail in [69], so we will largely ignore this case in this paper (but we will recall the results for the closed case in Subsection 12.2 following [69]).
+Our main results are for manifolds with straight cylindrical ends $M$ and for compatible open subsets $\Omega\subset M$ with boundary $\Gamma:=\partial\Omega$, as above.
+Our main results also hold for closed manifolds (i.e., compact, smooth, boundaryless manifolds), see [89] (some of these results are recalled in Subsection 12.2).
 Let us thus assume in this introduction that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a manifold with straight cylindrical ends (see 1.1 or Definition 3.11); this will be the case for the most part of our paper.
-Our first main results are to prove the invertibility on $\Gamma$ of the single layer and double layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$, associated to suitable generalizations $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ of the Stokes operator $\boldsymbol{\Xi}_{0,0}$, which will be defined shortly, where $V$ and $V_{0}$ are suitable non-negative functions, see Theorems 10.4 and 10.8.
+Our first main results are to prove the invertibility on $\Gamma$ of the single layer and double layer potential operators $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$, associated to suitable generalizations $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ of the Stokes operator $\boldsymbol{\Xi}_{0,0}$, which will be defined shortly, where $V$ and $V_{0}$ are suitable non-negative potentials (as in 1.1), see Theorems 10.4 and 10.8.
 (We usually denote the identity operator $\mathbb{I}$ with 1, we thus write $\frac{1}{2}+\boldsymbol{K}$ instead of $\frac{1}{2}\mathbb{I}+\boldsymbol{K}$.) We also obtain, for suitable potentials $V$ and $V_{0}$ satisfying slightly weaker assumptions, that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are Fredholm.
@@
 We use these results to define the single and double layer potential operators $\mathcal{S}_{\rm{ST}}$, $\boldsymbol{S}$, $\mathcal{D}_{\rm{ST}}$, and $\boldsymbol{K}$ associated to the Dirichlet boundary value problem for $\boldsymbol{\Xi}$.
-We prove that they satisfy the usual mapping and “jump” properties and that they have the usual principal symbols (Theorem 8.16; but in order to obtain this result, we develop an extensive theory: energy and representation formulas, geometric representations and formulas for our operators, a “took-kit” for proving jump relations, and many other results of independent interest).
+We prove that they satisfy the usual mapping and “jump” properties and that they have the usual principal symbols (Theorem 8.16; but in order to obtain this result, we develop an extensive theory: energy and representation formulas, geometric representations and formulas for our operators, the definition and suitable classes of integral operators, a “took-kit” for proving jump relations, and many other results of independent interest).
 Then, using the jump relations and principal symbol properties of the layer potential operators, we prove that $\boldsymbol{S}$ and $\frac{1}{2}+\boldsymbol{K}$ are invertible.
-As is well known, this result implies the solvability of the Dirichlet problem (1.4), Theorem 1.1 next, which is the main result of this paper.
+As is well known, this result implies the solvability of the Dirichlet problem (1.5) (see Equation (1.5) for a more explicit form of this equation), Theorem 1.1 next, which is the main result of this paper.
 We write $v\succ 0$ for a non-negative function that does not vanish identically on any connected component of its domain (see Definition 2.13).
-The space $\mathcal{C}^{\infty}_{\operatorname{inv}}(M)$ consists of functions (or sections of a vector bundle in case of $\mathcal{C}^{\infty}_{\operatorname{inv}}(M;E)$) that are translation invariant in a neighborhood of invinity (Definition 3.14)
+The space $\mathcal{C}^{\infty}_{\operatorname{inv}}(M)$ consists of functions (or sections of a vector bundle in case of $\mathcal{C}^{\infty}_{\operatorname{inv}}(M;E)$) that are translation invariant in a neighborhood of infinity (Definition 3.14).
 Theorem 1.1.
@@
   3. (iii)
-    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, and
+    $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, and
   4. (iv)
-    if $\Omega_{0}$ is a connected component of $M^{\prime}$ contained in $\Omega^{\prime}$, then $\color[rgb]{1,0,1}\widehat{V}\not\equiv 0$ on $\Omega_{0}$.
+    if $\Omega_{0}$ is a connected component of $M^{\prime}$ contained in $\Omega^{\prime}$, then $\widehat{V}\not\equiv 0$ on $\Omega_{0}$.
 Then, there exists $C_{m}\geq 0$ such that, for any $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$, the homogeneous Dirichlet problem (11.1) has a unique solution $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ and this solution satisfies
@@
 The proof of this theorem is provided in Subsection 11.1.
-At this time, a proof in general of this theorem using energy methods (i.e., the Riesz representation formula) seems out of reach.
+At this time, a proof in general of this theorem using energy (or variational) methods (i.e., the Riesz representation formula) seems out of reach, in the sense that it does not follow from the Riesz representation formula using standard methods.
 We also obtain applications of this well-posedness results to the Dirichlet-to-Neumann operator and to the solvability of the stationary Navier-Stokes system on domains with straight cylindrical ends of small dimension.
 We have outlined here the most important results.
-However, in this paper, we obtain many other results of independent interest that lead to the main results.
+However, in this paper, we obtain many other results of independent interest that lead to the main results (they were mentioned briefly above).
+We also include a lot of background material, which is, hence not entirely new.
+However, our presentation is original and it gathers many results that are difficult to find in earlier papers.
+Moreover, most of our background material is presented in great detail, so that to make our main result more accessible (a notable exception is, though, the background material on pseudodifferential operators, for which many other well-developed works exist).
 These other results are discussed in much more detail in the part on the contents of the paper, which is next.
 1.3. Contents of the paper
-Our paper consists of 12.1 sections (including an introduction), each divided into subsections.
+Our paper consists of 12 sections (including an introduction), each divided into subsections.
 The assumptions of this paper (especially on the manifold $M$ on which we work) change from section to section.
 These assumptions are always stated at the beginning of that section and are reminded regularly, especially at the beginning of the subsections and in the statement of the most important results.
-(Note, however, that, beginning with Section 5, we work almost exclusively on manifolds with straight cylindrical ends.)
+(Note, however, that, beginning with Section 5, we work almost exclusively on manifolds with straight cylindrical ends.) The first six sections are devoted mostly to background material, but the presentation is original (and, except for the part on pseudodifferential operators, quite detailed, including details and even results that are hard to find in the literature).
 The first section is this introduction.
@@
 We also explain what modifications need to be brought to the statements of our results for manifolds with straight cylindrical ends so that they are valid also in the case of closed manifolds.
-These resuls were proved in [69] (for closed manifolds), so in this paper we content ourselves with reminding a few of these results.
+These results were proved in [89] (for closed manifolds), so in this paper we content ourselves with reminding a few of these results.
 1.4. Overview of the existing literature
-The layer potentials method is a classical method for solving elliptic boundary value problems [32, 35, 37, 39, 45, 54, 55, 59, 63, 76, 87, 96, 117, 134].
+The layer potentials method is a classical method for solving elliptic boundary value problems [39, 42, 47, 49, 59, 71, 72, 73, 79, 83, 96, 108, 118, 140, 158].
 It plays a crucial role in the analysis of elliptic boundary value problems in various function spaces, their main advantage consisting in the fact that they provide explicit formulas for the the solutions and that they provide a convenient way of extending the results to $L^{p}$-type Sobolev spaces.
@@
 An example is provided by our well-posedness theorem, Theorem 1.1.
-More on the history of the method of layer potentials can be found, for example, in [2, 43, 45, 54, 59, 76, 87, 115, 135].
+More on the history of the method of layer potentials can be found, for example, in [2, 57, 59, 71, 74, 79, 96, 108, 138, 159].
+See also [51, 60, 75].
 Let us now review some of the more specialized results in the subject, most of them from research papers devoted to the analysis of the Stokes and Navier-Stokes equations on domains, in Euclidean spaces, or on Riemannian manifolds.
@@
 Girault and Sequeira [61] used a variational approach to prove the well-posedness of the exterior Dirichlet problem for the Stokes system on exterior Lipschitz domains in $\mathbb{R}^{n}$, $n=2,3$.
-Galdi [45, Chapters 5, 11, 12] used variational approaches in the analysis of steady Stokes and Navier-Stokes flows in exterior domains.
+Galdi [59, Chapters 5, 11, 12] used variational approaches in the analysis of steady Stokes and Navier-Stokes flows in exterior domains.
 Medková [111] used an integral equation method to analyze the Dirichlet problem for the Brinkman and Darcy-Forchheimer-Brinkman systems in Sobolev spaces on unbounded domains with compact Lipschitz boundary in $\mathbb{R}^{n}$.
-Exterior Dirichlet and transmission problems for anisotropic Stokes and Navier-Stokes systems with $L_{\infty}$ variable coefficients have been studied in [65] and [63, Chapters 5-7] using a variational approach and in [92].
+Exterior Dirichlet and transmission problems for anisotropic Stokes and Navier-Stokes systems with $L_{\infty}$ variable coefficients have been studied in [85] and [83, Chapters 5-7] using a variational approach and in [113].
 Other boundary problems for the Stokes system in Sobolev spaces on Lipschitz domains via layer potential theoretical methods have also been studied in [109, 110, 112, 138, 141, 160].
-An important paper is Costabel’s contribution [35] that establishes many very useful properties of layer potentials using elementary methods.
 Chandler-Wilde et al.
 [38, 39] studied boundary integral equations on locally-dilation-invariant Lipschitz domains and fractal screens.
-Buffa, Hiptmair, Petersdorff, and Schwab [28] studied Maxwell transmission problems in Lipschitz domains by using boundary element methods.
-Labarca-Figueroa and Hiptmair [75] used coupled boundary and volume integral equations in the study of electromagnetic scattering.
+Cakoni, Hsio, and Wendland have studied a mixed boundary value problem for the biharmonic equation using boundary integral equations.
+Haller-Dintelmann, R.
+and Jonsson, A.
+and Knees, D.
+and Rehberg have studied the regularity of a mixed, second order boundary value problem in divergence form [67].
 Singularly perturbed boundary value problems based on a functional analytic approach have been studied by M.
@@
 The theory of Fredholm and layer potential operators in Euclidean spaces and in Riemannian manifolds plays an important role in their book [118] (and also in our book [83]).
+See also [34, 35, 43].
+Layer potentials play an important role in applications, for instance in the study of numerical methods.
+In this direction, in an important paper, Costabel [42] has establishes many useful properties of layer potentials using elementary methods.
+Buffa, Hiptmair, Petersdorff, and Schwab [33] studied numerical methods for Maxwell’s transmission problems in Lipschitz domains by using boundary element methods.
+Labarca-Figueroa and Hiptmair [95] used coupled boundary and volume integral equations in the study of electromagnetic scattering.
+Steinbach and Wendland [152] have obtained resulsts on a Neumann series method for solving boundary integral equations.
+Costabel, Stephan, and Wendland have studied boundary integral equations of the first kind for the bi-Laplacian in a polygonal plane domain [45].
+See also [19, 44, 53, 123, 124, 127].
 The layer potential methods are also very useful in the analysis of elliptic boundary value problems on manifolds.
@@
 In this regard, we refer to the contributions of Ammann, Lauter and Nistor [13, 14], Mazzeo and Melrose [105], Melrose [114], Melrose and Mendoza [116], Nistor [125], Nistor, Weinstein, and Xu [126] Schrohe [143, 144], Schulze [145, 146], Schulze and Wong [147], Shubin [151], to mention just a few among very many others.
+For instance, some results relevant for boundary value problems were obtained by Grieser [58, 63], who used the Calderon projection in the framework of the cusp-edge calculus [105] for related boundary value problems for (classically) elliptic operators.
+The cusp-edge calculus is a completion of a groupoid calculus [126] corresponding to a structural Lie algebra of vector fields that is, in fact, more general than the one appearing in our setting of manifolds with straight cylindrical ends.
 Another important issue in the analysis on non-compact manifolds is finding the “right definition” of the relevant function spaces.
@@
 Lie manifolds are special manifolds with bounded geometry [12, 13], which allow the implementation of the layer potential methods.
-See also [101, 102].
 The layer potential theory for the Laplace operator on manifolds with cylindrical ends was initiated by Lewis and Parenti [102] and further extended by Mitrea and Nistor [120].
@@
 Manifolds with cylindrical ends appear in may applications that involve “wave-guides” or “layers” (see e.g.
-[19, 20, 21, 22, 23, 24, 25, 42, 63, 68, 74, 137]).
+[23, 24, 25, 26, 27, 28, 29, 30, 52, 54, 56, 83, 88, 94, 161]).
 Let us finally add that the main operator in the structure of the Stokes and Navier-Stokes equations on Riemannian manifolds is the deformation Laplacian $\boldsymbol{L}:=2\operatorname{Def}\,^{*}\operatorname{Def}\,$.
@@
 Our generalized Stokes operator is closely related to the Brinkman operator.
-The Brinkman flow was studied by Karageorghis, Lesnic, and Marin in [57, 58], and Bacuta, Hassell, Hsiao, and Sayas in [26].
+The Brinkman flow was studied by Karageorghis, Lesnic, and Marin in [77, 78], and Băcuţă, Hassell, Hsiao, and Sayas in [31].
 A class of solutions to the Navier-Stokes equations for flows in plane-parallel channels and related singular perturbation problems were studied by Mazzucato and Taylor [107].
@@
 We thank Massimo Lanza de Cristoforis, Dorina, Irina, and Marius Mitrea, Sergey Mikhailov, and Mihai Putinar for useful discussions.
+We also thank Daniel Grieser and Karl-Mikael Perfekt for useful suggestions.
 2. Green formulas for the Stokes operator
@@
   1. (i)
-    or any vector field $X$, we have [107, page 49]
+    or any vector field $X$, we have [130, page 49]
     (2.7) $\mathcal{L}_{X}\,d\operatorname{vol}_{g}\,=\,\operatorname{div}(X)\,d\operatorname{vol}_{g}\,$,
@@
 in which case ${\mathcal{R}_{\infty}}\nabla^{LC}=\nabla^{LC}$.
+The choice of manifolds (i.e., with straight cylindrical ends) and of the associated functions and differential operators is motivated by practical questions, since these are the objects that are the most likely to appear in practical applications.
+However, our results extend to the more general case of manifolds with cylindrical ends (not necessarily straigh, but asymptotically straight).
+One just needs to slightly enlarge the class of functions, vector bundles, differential operators, and, especially, pseudodifferential operators.
+These questions will be considered in another paper.
 For the rest of this section, $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ will be a manifold with straight cylindrical ends and $E\to M$ will be a fixed compatible Hermitian vector bundle.
@@
 Proof.
-The result follows by applying Lebesgue’s dominated convergence theorem. See [41, Appendix E]. ∎
+The result follows by applying Lebesgue’s dominated convergence theorem. See [55, Appendix E]. ∎
 If $N=0$ (hence $\alpha=0$), then this theorem simply states that $F(x)$ is defined for all $x\in\mathbb{R}^{k}$, that it is continuous in $x$, and that it is bounded by $|F(x)|\leq\int_{\mathfrak{M}}G_{0}(\omega)\,d\mu(\omega).$ An immediate consequence of this theorem is the following strong operator convergence result.
@@
 Also, in the following, $E\to M$ will be a fixed compatible Hermitian vector bundle.
-We now define two needed classes of pseudodifferential operators on $M$, the $\operatorname{inv}$-calculus” and the $\operatorname{ess}$-calculus.” For $M$ closed, these definitions are trivial (they recover the usual pseudodifferential operators, so the give nothing new), so the only interesting case is the one when $M$ is not compact.
-Including the case $M$ compact is useful, however, in order to formulate our results in a more general setting, as they are true for both closed manifolds and for manifolds with straight cylindrical ends.
-The resulting classes of pseudodifferential operators are very closely related to the $b$- and $c$- calculi of Melrose [93, 84].
+We now define two needed classes of pseudodifferential operators on $M$, the $\operatorname{inv}$-calculus” and the $\operatorname{ess}$-calculus.” Our presentation is complete, but short (unlike the previous sections).
+The reason is that there are many works on related classes of pseudodifferential operators, beginning with the $b$- and $c$- calculi of Melrose [114, 105].
 Related calculi have also been considered by Schulze and Schrohe [144, 145, 147].
 See [87] for more references and for a discussion of the relation between our calculi and the ones of Mazzeo, Melrose, Schrohe, and Schulze.
-The results of this section are close to those in [63, 67].
-The presentation is, however, different and, occasionally, it provides more general results.
+See also [62, 100, 101, 117, 147].
+The results of this section are close to those in [83, 87], but the presentation is different and, occasionally, it provides more general results.
 For instance, in [83], we work only with classical pseudodifferential operators, whereas here we allow the Kohn-Nirenberg classes of symbols recalled in Definition 3.5.
@@
   3. (iii)
-    If $M=M^{\prime}\times\mathbb{R}$ and $P$ is translation invariant (that is, $P\in\Psi_{\operatorname{inv}}^{m}(M^{\prime}\times\mathbb{R};E,F)^{\mathbb{R}}$), then ${\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}=P$.
+    If $M=M^{\prime}\times\mathbb{R}$ and $P$ is translation invariant (that is, $P\in\Psi_{\operatorname{inv}}^{m}(M^{\prime}\times\mathbb{R};E,F)^{\mathbb{R}}$), then ${\mathcal{R}_{\infty}}(P)=P$.
 Definition 5.12.
 Let $P\in\Psi_{\operatorname{inv}}^{m}(M;E,F)$ and
-${\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)\mathrel{\,:=\,}\Psi^{m}(M\times\mathbb{R};E,F)^{\mathbb{R}}$
-be as in Lemma 5.11. Then $\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)$} is called the limit operator associated to $P$. The resulting map
-${\mathcal{R}_{\infty}}:\Psi_{\operatorname{inv}}^{m}(M;E,F)\ni P\longmapsto{\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)$
+${\mathcal{R}_{\infty}}(P)\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)\mathrel{\,:=\,}\Psi^{m}(M\times\mathbb{R};E,F)^{\mathbb{R}}$
+be as in Lemma 5.11. Then ${\mathcal{R}_{\infty}}(P)$ is called the limit operator associated to $P$. The resulting map
+${\mathcal{R}_{\infty}}:\Psi_{\operatorname{inv}}^{m}(M;E,F)\ni P\longmapsto{\mathcal{R}_{\infty}}(P)\in\psi_{\operatorname{sus}}^{m}(M^{\prime};E,F)$
 will be called the limit operator map.
@@
     $\sigma_{m}({\mathcal{R}_{\infty}}(T))\,=\,{\mathcal{R}_{\infty}}(\sigma_{m}(T))\in\mathcal{C}^{\infty}(S^{*}(M^{\prime}\times\mathbb{R});\operatorname{Hom}(E;F))^{\mathbb{R}}\,.$
-We also have the following characterization of Sobolev spaces [67], [63].
+We also have the following characterization of Sobolev spaces [87, 83].
 Corollary 5.16.
@@
 If $P\in\Psi_{\operatorname{ess}}^{m}(M;E,F)$ and $u\in\mathcal{C}^{\infty}_{c}(M^{\prime}\times\mathbb{R};E)$, then, for $s$ large enough, $P\Phi_{s}(u)$ is defined and we have the following norm convergence in $L^{2}(M^{\prime}\times\mathbb{R};F)$,
-${\color[rgb]{1,0,1}{\mathcal{R}_{\infty}}(P)}(u)\mathrel{\,:=\,}\lim_{s\to\infty}\Phi_{-s}\big[P\Phi_{s}(u)|_{\{t<0\}}\big]\,.$
+${\mathcal{R}_{\infty}}(P)(u)\mathrel{\,:=\,}\lim_{s\to\infty}\Phi_{-s}\big[P\Phi_{s}(u)|_{\{t<0\}}\big]\,.$
 In addition, the resulting operator satisfies ${\mathcal{R}_{\infty}}(P)\in\Psi_{\operatorname{sus}}^{m}(M^{\prime};E)$ and ${\mathcal{R}_{\infty}}\circ s_{0}(P_{0})=P_{0}$ for $P_{0}\in\Psi_{\operatorname{sus}}^{m}(M^{\prime};E)$.
@@
 Unless stated otherwise, our definition of $\widehat{T}(\tau)$ will be given by regarding the operator $T$ as an operator in $\Psi_{\operatorname{sus}}^{m}(\mathfrak{S};E,F)$ (i.e., as an invariant operator, which gives us “fewer” indicial operators $\widehat{T}(\tau)$).
-6. Normal lateral limits on manifolds with straight cylindrical ends
-In this section, we adapt the results from Section 4 on normal lateral limits on half-spaces to smooth, open domains in manifolds with cylindrical ends.
+6. Main properties of the $\operatorname{inv}$ and $\operatorname{ess}$ calculi
+In this section, we establish some of the main properties of the operators in the $\operatorname{inv}$” and $\operatorname{ess}$” calculi of the previous section.
+We concentrate on the properties needed in the study of layer potentials.
+We first adapt the results from Section 4 on normal lateral limits on half-spaces to smooth, open domains in manifolds with cylindrical ends and our calculi of the previous section ($\operatorname{inv}$ and $\operatorname{ess}$).
+Then we recall Douglis-Nirenberg-operators and their Fredholm, (elliptic) regularity, and spectral invariance properties.
 We thus continue to assume that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a smooth Riemannian manifold with straight cylindrical ends (Definition 3.11) and that $E,F\to M$ are two compatible vector bundle (see Definition 3.13).
@@
 ($\Gamma=\partial\Omega$) is well defined and is a diffeomorphism onto its image, then we shall say that $\Gamma$ has an $\epsilon$-normal tubular neighborhood.
-It is well-known that if $\Gamma$ is compact or if $M$ has straight cylindrical ends and $\Omega$ is compatible with this structure, then $\Gamma$ will have an $\epsilon$-normal tubular neighborhood, for some $\epsilon>0$ small enough, see [107, Corollary 5.5.3].
+It is well-known that if $\Gamma$ is compact or if $M$ has straight cylindrical ends and $\Omega$ is compatible with this structure, then $\Gamma$ will have an $\epsilon$-normal tubular neighborhood, for some $\epsilon>0$ small enough, see [130, Corollary 5.5.3].
 The curves $t\mapsto\exp(t\boldsymbol{\nu}(x))$, $x\in\partial\Omega$, will be called the normal geodesics to $\partial\Omega$.
@@
 ∎
+By extending the above mapping property to Besov spaces, one can extend our well-posedness result to $L^{p}$-type Sobolev and Besov spaces.
+This will be discussed in a future publication.
 Recall from Notation 6.2 that $\boldsymbol{\nu}$ is a fixed vector field on $M$ that is the outer unit normal vector to $\Gamma:=\partial\Omega$.
@@
 Proof.
-The given formula for principal symbol of $\boldsymbol{\Xi}$ is a consequence of [130, Chapter 5 §12, (12.20)] (but see also Corollary 2.6 for the principal symbol of $\operatorname{Def}^{*}\,\operatorname{Def}\,$ and Remark 2.8(iii) for the principal symbol of $\nabla$).
+The given formula for principal symbol of $\boldsymbol{\Xi}$ is a consequence of [154, Chapter 5 §12, (12.20)] (but see also Corollary 2.6 for the principal symbol of $\operatorname{Def}^{*}\,\operatorname{Def}\,$ and Remark 2.8(iii) for the principal symbol of $\nabla$).
 The result then follows by verifying that the product of these two matrices is the identity.
 ∎
-See also [68, Theorem 5.11 and Remark 5.12] and [36]. We notice that the formula for the inverse of the principal symbol was obtained using Schur’s complement formula.
+See also [88, Theorem 5.11 and Remark 5.12] and [46]. We notice that the formula for the inverse of the principal symbol was obtained using Schur’s complement formula.
 We now return to the case when $M$ is a manifold with straight cylindrical ends (Definition 3.11) and the vector bundles are compatible (Definition 3.13).
@@
 To that end, we shall use the following characterization of the Fredholm property in the essentially translation invariant calculus.
-(This result was discussed and proved in great detail in [63, Theorem 15.4.19] and [67, Theorem 7.12], so we will not include a proof here.
+(This result was discussed and proved in great detail in [83, Theorem 15.4.19] and [87, Theorem 7.12], so we will not include a proof here.
 It builds on earlier results of many people, including Kondratiev, Melrose, and Schulze, see the above quoted works for references.) Recall the notation of the last remark, Remark 6.14.
@@
 (See also [36, 58, 76, 88, 83, 91, 116, 144] and the references therein.)
+6.4. Further properties: regularity and spectral invariance
+We shall repeatedly make use of the following spectral invariance and elliptic regularity results from [83, 87].
+We have the following generalization of the classical elliptic regularity result (see Theorem 15.3.30 of [83] or [87, Proposition 6.5], see also [115, 145]).
+We use the notation of Definition 6.11.
+Thus $E_{j},F_{i}\to M$, $i,j=1,\ldots,k$, are compatible vector bundles, $E=\oplus_{j=1}^{k}E_{j}$, and $F=\oplus_{i=1}^{k}F_{i}$, and $\boldsymbol{s}=(s_{1},\ldots,s_{k})$ and $\boldsymbol{t}=(t_{1},\ldots,t_{k})$, $s_{i},t_{j}\in\mathbb{R}$.
+Theorem 6.16.
+Let $P\in\Psi_{\operatorname{ess}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E,F)$ be Douglis-Nirenberg-elliptic.
+Let also $m,m^{\prime}\in\mathbb{R}$.
+If $u\in H^{m^{\prime}+[\boldsymbol{t}]}(M;E)$ is such that $Pu\in H^{m-[\boldsymbol{s}]}(M;F)$, then $u\in H^{m+[\boldsymbol{t}]}(M;E)$.
+We have the following related spectral invariance property ([83, Theorem 15.4.11], see also [20, 87, 120]).
+Theorem 6.17.
+Let $s_{i},t_{j}\geq 0$, $m\in\mathbb{R}$, and $T\in\Psi_{\operatorname{ess}}^{[\boldsymbol{s}+\boldsymbol{t}]}(M;E,F)$. Let
+$\mathcal{D}(T_{\rm{max}})\mathrel{\,:=\,}\{\,\xi\in H^{m-[\boldsymbol{s}]}(M;E):\ T\xi\in H^{m-[\boldsymbol{s}]}(M;F)\,\}$
+and assume that $T:\mathcal{D}(T_{\rm{max}})\to H^{m-[\boldsymbol{s}]}(M;E)$ is Fredholm with pseudo-inverse denoted $T^{(-1)}$.
+If one of the $s_{i}$ or $t_{i}$ is (strictly) positive (i.e., >0>0), we also assume that $T$ is $(\boldsymbol{s},\boldsymbol{t})$-Douglis-Nirenberg-elliptic.
+Then its Moore-Penrose pseudo-inverse $T^{(-1)}\in\Psi_{\operatorname{ess}}^{[\mathbf{-s-t}]}(M;F,E)$.
+This result was proved in case $T$ is actually invertible in the above mentioned publications, so let us include a proof for the statement in case $T$ is Fredholm, using the corresponding statement for “$T$ invertible”.
+Proof.
+We can assume that $E=F$ and $\boldsymbol{s}$ and $\boldsymbol{t}=0$ are constant, so we work with the usual order (this is the only case used in this paper).
+Let us notice that $T$ is elliptic (if it is of order zero, this follows from the Fredholm property, Theorem 6.15).
+Elliptic regularity, Theorem 6.16 gives that its kernel $\mathcal{N}$ consists of functions in $H^{\infty}(M)$ and hence the orthogonal projection onto $p_{\mathcal{N}}\in\Psi_{\operatorname{ess}}^{-\infty}(M;E)$.
+Then $p_{\mathcal{N}}+T^{*}T$ is invertible and hence $(p_{\mathcal{N}}+T^{*}T)^{-1}\in\Psi_{\operatorname{ess}}^{[-2\boldsymbol{s}]}(M;E)$.
+The desired result follows from $T^{(-1)}=(p_{\mathcal{N}}+T^{*}T)^{-1}T^{*}\in\Psi_{\operatorname{ess}}^{[-2\boldsymbol{s}]}(M;E)$.
+∎
+These results are true (and were known for a long time) if $M$ closed, in which case we replace $\Psi_{\operatorname{ess}}^{m}(M;E)$ with $\Psi^{m}(M;E)$.
 7. Green formulas and invertibility of $\boldsymbol{\Xi}$ on cylinders
@@
 We thus assume in this section that $M:=\mathfrak{S}\times\mathbb{R}$, where $\mathfrak{S}$ is a closed manifold (i.e., smooth, compact, without boundary).
-Moreover, all other objects used in this section will be translation invariant, thus $\Omega=\Omega^{\prime}\times\mathbb{R}$ and $V$ and $V_{0}$ descend to functions on $\mathfrak{S}$, denoted, respectively, $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ (see Equation (5.6) for a justification of this notation).
-Explicitly, we have $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$ for $x^{\prime}\in\mathfrak{S}$ and $t\in\mathbb{R}$.
+Moreover, all other objects used in this section will be translation invariant, thus $\Omega=\Omega^{\prime}\times\mathbb{R}$ and $V$ and $V_{0}$ descend to functions on $\mathfrak{S}$, denoted, respectively, $\widehat{V}$ and $\widehat{V}_{0}$ (see Equation (5.6) for a justification of this notation).
+Explicitly, we have $V(x^{\prime},t)=\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\widehat{V}_{0}(x^{\prime})$ for $x^{\prime}\in\mathfrak{S}$ and $t\in\mathbb{R}$.
 The results of this section will be used for $\mathfrak{S}$ possibly disconnected, so we will not assume that $\mathfrak{S}$ is connected, although the proofs of the results in this section can be reduced immediately to the connected case.
@@
 The results of this section will be used for $\mathfrak{S}$ not necessarily connected, so we will not assume that $\mathfrak{S}$ is connected, although the general case can immediately be reduced to the connected case.
-Also, we let $\color[rgb]{1,0,1}\widehat{V}(x^{\prime})={\mathcal{R}_{\infty}}V(x^{\prime},t)=V(x^{\prime},t)\geq 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})={\mathcal{R}_{\infty}}V_{0}(x^{\prime},t)=V_{0}(x^{\prime},t)\geq 0$ be smooth, translation invariant functions (this notation is justified by Equation (5.6)).
+Also, we let $\widehat{V}(x^{\prime})={\mathcal{R}_{\infty}}V(x^{\prime},t)=V(x^{\prime},t)\geq 0$ and $\widehat{V}_{0}(x^{\prime})={\mathcal{R}_{\infty}}V_{0}(x^{\prime},t)=V_{0}(x^{\prime},t)\geq 0$ be smooth, translation invariant functions (this notation is justified by Equation (5.6)).
 Again, this choice of $V$ and $V_{0}$ satisfies all our previous assumptions.
-(Note that in some of our previous papers, the notation $\tilde{V}$ and $\tilde{V}_{0}$ was used instead of $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$.
+(Note that in some of our previous papers, the notation $\tilde{V}$ and $\tilde{V}_{0}$ was used instead of $\widehat{V}$ and $\widehat{V}_{0}$.
 However, in this paper, the notation $\tilde{u}$ means $\tilde{u}(x):=u(-x)$.)
@@
 Our results are formulated in general, but there is no loss of generality to assume that $\Omega^{\prime}$ is connected (that is, a smooth domain).
-Also, recall that, in this section, $\color[rgb]{1,0,1}\widehat{V}(x^{\prime})=V(x^{\prime},t)$ and $\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})=V_{0}(x^{\prime},t)$, as motivated by Equation (5.6).
-In particular, $V=\color[rgb]{1,0,1}\widehat{V}$ and $V_{0}=\color[rgb]{1,0,1}\widehat{V}_{0}$ are translation invariant.
+Also, recall that, in this section, $\widehat{V}(x^{\prime})=V(x^{\prime},t)$ and $\widehat{V}_{0}(x^{\prime})=V_{0}(x^{\prime},t)$, as motivated by Equation (5.6).
+In particular, $V=\widehat{V}$ and $V_{0}=\widehat{V}_{0}$ are translation invariant.
 The main equation of Notation 2.10 becomes.
-(7.12) $\begin{gathered}U\mathrel{\,:=\,}(\boldsymbol{u}\ \ p)^{\top}\,,\ W\mathrel{\,:=\,}(\boldsymbol{w}\ \ q)^{\top}\in H^{1}(\Omega^{\prime};TM)\oplus L^{2}(\Omega^{\prime})\,,\\ \qquad\mathfrak{v}\mathrel{\,:=\,}({\color[rgb]{1,0,1}\widehat{V}}\boldsymbol{u},\boldsymbol{w})_{\Omega^{\prime}}-({\color[rgb]{1,0,1}\widehat{V}_{0}}p,q)_{\Omega^{\prime}}\,,\quad\mbox{and}\\ B_{\Omega^{\prime},\tau}(U,W)\mathrel{\,:=\,}2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{w})_{\Omega^{\prime}}+(\widehat{\nabla}^{*}(\tau)\boldsymbol{u},q)_{\Omega^{\prime}}+(p,\widehat{\nabla}^{*}(\tau)\boldsymbol{w})_{\Omega^{\prime}}+\mathfrak{v}\,.\end{gathered}$
+(7.12) $\begin{gathered}U\mathrel{\,:=\,}(\boldsymbol{u}\ \ p)^{\top}\,,\ W\mathrel{\,:=\,}(\boldsymbol{w}\ \ q)^{\top}\in H^{1}(\Omega^{\prime};TM)\oplus L^{2}(\Omega^{\prime})\,,\\ \qquad\mathfrak{v}\mathrel{\,:=\,}({\widehat{V}}\boldsymbol{u},\boldsymbol{w})_{\Omega^{\prime}}-({\widehat{V}_{0}}p,q)_{\Omega^{\prime}}\,,\quad\mbox{and}\\ B_{\Omega^{\prime},\tau}(U,W)\mathrel{\,:=\,}2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{w})_{\Omega^{\prime}}+(\widehat{\nabla}^{*}(\tau)\boldsymbol{u},q)_{\Omega^{\prime}}+(p,\widehat{\nabla}^{*}(\tau)\boldsymbol{w})_{\Omega^{\prime}}+\mathfrak{v}\,.\end{gathered}$
 Also, in this subsection, we let $(\ ,\ )^{\prime}:=(\ ,\ )_{\partial\Omega^{\prime}},$ the scalar product on the boundary $\Gamma^{\prime}=\partial\Omega^{\prime}$ of $\Omega^{\prime}$.
@@
 It follows from the definition of indicial operators (Definition 5.26) that
-(7.13) $\widehat{\boldsymbol{\Xi}}(\tau)\,=\,\left(\begin{array}[]{cc}2\widehat{\operatorname{Def}}^{*}\,(\tau)\widehat{\operatorname{Def}}\,(\tau)+\color[rgb]{1,0,1}\widehat{V}&\widehat{\nabla}(\tau)\\ \widehat{\nabla}^{*}(\tau)&-\color[rgb]{1,0,1}\widehat{V}_{0}\end{array}\right)\,$,
-where, we recall, we are identifying $\color[rgb]{1,0,1}\widehat{V}=\widehat{V}(\tau)$ and $\color[rgb]{1,0,1}\widehat{V}_{0}=\widehat{V}_{0}(\tau)$.
+(7.13) $\widehat{\boldsymbol{\Xi}}(\tau)\,=\,\left(\begin{array}[]{cc}2\widehat{\operatorname{Def}}^{*}\,(\tau)\widehat{\operatorname{Def}}\,(\tau)+\widehat{V}&\widehat{\nabla}(\tau)\\ \widehat{\nabla}^{*}(\tau)&-\widehat{V}_{0}\end{array}\right)\,$,
+where, we recall, we are identifying $\widehat{V}=\widehat{V}(\tau)$ and $\widehat{V}_{0}=\widehat{V}_{0}(\tau)$.
 (The formula (7.13) is valid for any manifold with straight cylindrical ends, not just for our manifold $M:=\mathfrak{S}\times\mathbb{R}$.) The last proposition gives then the following lemma analogous to Corollary 2.14.
@@
 Lemma 7.6.
-Let $\tau\in\mathbb{R}$, let $\color[rgb]{1,0,1}\widehat{V}\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R};\operatorname{End}(TM))^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(\mathfrak{S};TM)$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R})^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(\mathfrak{S})$ be non-negative (i.e., $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$), and let
+Let $\tau\in\mathbb{R}$, let $\widehat{V}\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R};\operatorname{End}(TM))^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(\mathfrak{S};TM)$ and $\widehat{V}_{0}\in\mathcal{C}^{\infty}(\mathfrak{S}\times\mathbb{R})^{\mathbb{R}}\simeq\mathcal{C}^{\infty}(\mathfrak{S})$ be non-negative (i.e., $\widehat{V},\widehat{V}_{0}\geq 0$), and let
 $U=\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$
@@
 Then the following vanishing relations hold in $\Omega^{\prime}$:
-$\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}\,=\,0\,,\quad\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u}\,=\,0\,,\quad\widehat{\nabla}^{*}(\tau)\boldsymbol{u}\,=\,0\,,\quad\widehat{\nabla}(\tau)p\,=\,0\,,\quad\mbox{and}\quad\color[rgb]{1,0,1}\widehat{V}_{0}p\,=\,0\,.$
+$\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}\,=\,0\,,\quad\widehat{V}\boldsymbol{u}\,=\,0\,,\quad\widehat{\nabla}^{*}(\tau)\boldsymbol{u}\,=\,0\,,\quad\widehat{\nabla}(\tau)p\,=\,0\,,\quad\mbox{and}\quad\widehat{V}_{0}p\,=\,0\,.$
 Proof.
@@
 Then $\operatorname{Re}\big[-(\widehat{\nabla}^{*}(\tau)\boldsymbol{u},p)_{\Omega^{\prime}}+(p,\widehat{\nabla}^{*}(\tau)\boldsymbol{u})_{\Omega^{\prime}}\big]=0$.
-Let $\mathfrak{v}:=(\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u},\boldsymbol{u})_{\Omega^{\prime}}+(\color[rgb]{1,0,1}\widehat{V}_{0}p,p)_{\Omega^{\prime}}$, as before.
+Let $\mathfrak{v}:=(\widehat{V}\boldsymbol{u},\boldsymbol{u})_{\Omega^{\prime}}+(\widehat{V}_{0}p,p)_{\Omega^{\prime}}$, as before.
 Together with the definition of $B_{\Omega^{\prime},\tau}$ in Equation (7.12), this gives
@@
 = $\,=\,2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u})_{\Omega^{\prime}}+\mathfrak{v}$
-= $\,=\,2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u})_{\Omega^{\prime}}+(\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u},\boldsymbol{u})_{\Omega^{\prime}}+(\color[rgb]{1,0,1}\widehat{V}_{0}p,p)_{\Omega^{\prime}}\,.$
+= $\,=\,2(\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u},\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u})_{\Omega^{\prime}}+(\widehat{V}\boldsymbol{u},\boldsymbol{u})_{\Omega^{\prime}}+(\widehat{V}_{0}p,p)_{\Omega^{\prime}}\,.$
 All three terms in the last sum are non-negative, so each of them equals zero.
-Therefore $\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}=0$, $\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u}=0$, and $\color[rgb]{1,0,1}\widehat{V}_{0}p=0$.
+Therefore $\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}=0$, $\widehat{V}\boldsymbol{u}=0$, and $\widehat{V}_{0}p=0$.
 We also have
-$0\,=\,\widehat{\boldsymbol{\Xi}}(\tau)U\,=\,\left(\begin{array}[]{c}{2\widehat{\operatorname{Def}}^{*}\,(\tau)\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}+\color[rgb]{1,0,1}\widehat{V}\boldsymbol{u}+\widehat{\nabla}(\tau)p}\\ {\widehat{\nabla}^{*}(\tau)\boldsymbol{u}-\color[rgb]{1,0,1}\widehat{V}_{0}p}\end{array}\right)\,=\,\left(\begin{array}[]{c}{\widehat{\nabla}(\tau)p}\\ {\widehat{\nabla}^{*}(\tau)u}\end{array}\right)\,.$
+$0\,=\,\widehat{\boldsymbol{\Xi}}(\tau)U\,=\,\left(\begin{array}[]{c}{2\widehat{\operatorname{Def}}^{*}\,(\tau)\widehat{\operatorname{Def}}\,(\tau)\boldsymbol{u}+\widehat{V}\boldsymbol{u}+\widehat{\nabla}(\tau)p}\\ {\widehat{\nabla}^{*}(\tau)\boldsymbol{u}-\widehat{V}_{0}p}\end{array}\right)\,=\,\left(\begin{array}[]{c}{\widehat{\nabla}(\tau)p}\\ {\widehat{\nabla}^{*}(\tau)u}\end{array}\right)\,.$
 Hence $\widehat{\nabla}^{*}(\tau)\boldsymbol{u}=0$ and $\widehat{\nabla}(\tau)p=0$, as claimed. This completes the proof. ∎
@@
 Corollary 7.7.
-Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$ and let $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$ satisfy $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega^{\prime}$ and $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(0)U,\boldsymbol{u})_{\Gamma^{\prime}}=0$ (as in Lemma 7.6, but for $\tau=0$).
+Let $\widehat{V},\widehat{V}_{0}\geq 0$ and let $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$ satisfy $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\Omega^{\prime}$ and $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(0)U,\boldsymbol{u})_{\Gamma^{\prime}}=0$ (as in Lemma 7.6, but for $\tau=0$).
   1. (i)
-    If $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$ (Definition 2.13), then $p=0$ on $\Omega^{\prime}$.
+    If $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$ (Definition 2.13), then $p=0$ on $\Omega^{\prime}$.
   2. (ii)
@@
     2. (b)
-      $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega^{\prime}$,
+      $\widehat{V}\succ 0$ on $\Omega^{\prime}$,
     then $\boldsymbol{u}=0$ in $\Omega^{\prime}$.
@@
 Lemma 7.6 gives $\widehat{\nabla}(0)p=0$ in $\Omega^{\prime}$.
-The assumption that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ in $\Omega$ together with Proposition 7.3 (stating the $L^{2}$-unique continuation property of $\widehat{\nabla}(0)$ in $\Omega^{\prime}$) then give $p=0$.
+The assumption that $\widehat{V}_{0}\succ 0$ in $\Omega$ together with Proposition 7.3 (stating the $L^{2}$-unique continuation property of $\widehat{\nabla}(0)$ in $\Omega^{\prime}$) then give $p=0$.
 Similarly, Lemma 7.6 gives $\widehat{\operatorname{Def}}\,(0)\boldsymbol{u}=0$.
@@
 Corollary 7.8.
-Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$ and let $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\mathfrak{S};TM)\oplus H^{1}(\mathfrak{S})$ satisfy $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\mathfrak{S}$.
+Let $\widehat{V},\widehat{V}_{0}\geq 0$ and let $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\mathfrak{S};TM)\oplus H^{1}(\mathfrak{S})$ satisfy $\widehat{\boldsymbol{\Xi}}(0)U=0$ in $\mathfrak{S}$.
   1. (i)
-    If $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$, then $p=0$ on $\mathfrak{S}$.
+    If $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$, then $p=0$ on $\mathfrak{S}$.
   2. (ii)
-    Similarly, if $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\mathfrak{S}$, then $\boldsymbol{u}=0$ in $\mathfrak{S}$.
+    Similarly, if $\widehat{V}\succ 0$ on $\mathfrak{S}$, then $\boldsymbol{u}=0$ in $\mathfrak{S}$.
 The case $\tau\neq 0$ is very similar, in fact, even easier than the case $\tau=0$, because of the injectivity of the relevant operators for $\tau\neq 0$.
 Corollary 7.9.
-Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$, $\tau\neq 0$, and $U^{\top}=(\boldsymbol{u}\ \ p)\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$ satisfy $\widehat{\boldsymbol{\Xi}}(\tau)U=0$ in $\Omega^{\prime}$ and $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)U,\boldsymbol{u})^{\prime}=0$ (the same assumptions as in Lemma 7.6, except that $\tau\neq 0$ in this corollary).
+Let $\widehat{V},\widehat{V}_{0}\geq 0$, $\tau\neq 0$, and $U^{\top}=(\boldsymbol{u}\ \ p)\in H^{2}(\Omega^{\prime};TM)\oplus H^{1}(\Omega^{\prime})$ satisfy $\widehat{\boldsymbol{\Xi}}(\tau)U=0$ in $\Omega^{\prime}$ and $(\widehat{\boldsymbol{T}}_{\boldsymbol{\nu}}(\tau)U,\boldsymbol{u})^{\prime}=0$ (the same assumptions as in Lemma 7.6, except that $\tau\neq 0$ in this corollary).
 Then $U=0$.
@@
 Corollary 7.10.
-Let $\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$, $\tau\neq 0$, and $U^{\top}=(\boldsymbol{u}\ \ p)\in H^{2}(\mathfrak{S};TM)\oplus H^{1}(\mathfrak{S})$ satisfy $\widehat{\boldsymbol{\Xi}}(\tau)U=0$.
+Let $\widehat{V},\widehat{V}_{0}\geq 0$, $\tau\neq 0$, and $U^{\top}=(\boldsymbol{u}\ \ p)\in H^{2}(\mathfrak{S};TM)\oplus H^{1}(\mathfrak{S})$ satisfy $\widehat{\boldsymbol{\Xi}}(\tau)U=0$.
 Then $U=0$.
@@
 Recall from Theorem 6.15 that the Fredholm property of an operator depends on the invertibility of its limit operator, which acts on the cylinder associated to $M$.
-The limit operator of $\boldsymbol{\Xi}_{V,V_{0}}$ is $\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$.
-It is the purpose of this subsection to establish the invertibility of $\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$ acting on the associated cylinder under the assumption that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
+The limit operator of $\boldsymbol{\Xi}_{V,V_{0}}$ is $\boldsymbol{\Xi}_{\widehat{V},\widehat{V}_{0}}$.
+It is the purpose of this subsection to establish the invertibility of $\boldsymbol{\Xi}_{\widehat{V},\widehat{V}_{0}}$ acting on the associated cylinder under the assumption that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
 We thus continue to assume, in this subsection, that $M$ is the cylinder $M=\mathfrak{S}\times\mathbb{R}$ with $\mathfrak{S}$ a closed manifold and with a product metric.
-We also assume that $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and that $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$, and hence $V$ and $V_{0}$ are thus translation invariant.
+We also assume that $V(x^{\prime},t)=\widehat{V}(x^{\prime})$ and that $V_{0}(x^{\prime},t)=\widehat{V}_{0}(x^{\prime})$, and hence $V$ and $V_{0}$ are thus translation invariant.
 We also assume that they are both $\succ 0$ on $M$ (see Definition 2.13).
-This condition is equivalent to $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
+This condition is equivalent to $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
 Let us notice that, since $M:=\mathfrak{S}\times\mathbb{R}$ is endowed with the product metric, this metric is translation invariant, and hence $\boldsymbol{\Xi}$ is also translation invariant.
@@
 Proposition 7.11.
-Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
+Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder and that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
 Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$.
@@
 Let then $U=\left(\begin{array}[]{c}{\boldsymbol{u}}\\ {p}\end{array}\right)=(\,\boldsymbol{u}\ \ p\,)^{\top}\in L^{2}(\mathfrak{S};TM\oplus\mathbb{C})$ be such that $\widehat{\boldsymbol{\Xi}}(0)U=0$.
-The ellipticity of $\widehat{\boldsymbol{\Xi}}(0)$ guarantees then that $U\in H^{\infty}(\mathfrak{S};TM\oplus\mathbb{C})$, by elliptic regularity (see, for example, Theorem 15.3.30 in [63]).
+The ellipticity of $\widehat{\boldsymbol{\Xi}}(0)$ guarantees then that $U\in H^{\infty}(\mathfrak{S};TM\oplus\mathbb{C})$, by elliptic regularity (see Theorem 6.16 or [83, Theorem 15.3.30] or [87, Proposition 6.5]).
 Assumption 8.3 allow us to use Corollary 7.8 to conclude that $U=0$.
@@
 ∎
-The case $\tau\neq 0$ is completely similar, but only requires the weaker assumptions $\color[rgb]{1,0,1}\widehat{V}:={\mathcal{R}_{\infty}}(V)\geq 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}:={\mathcal{R}_{\infty}}(V_{0})\geq 0$.
+The case $\tau\neq 0$ is completely similar, but only requires the weaker assumptions $\widehat{V}:={\mathcal{R}_{\infty}}(V)\geq 0$ and $\widehat{V}_{0}:={\mathcal{R}_{\infty}}(V_{0})\geq 0$.
 Proposition 7.12.
-Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $\tau\neq 0$, and that $\color[rgb]{1,0,1}\widehat{V}_{0},\color[rgb]{1,0,1}\widehat{V}\geq 0$.
+Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $\tau\neq 0$, and that $\widehat{V}_{0},\widehat{V}\geq 0$.
 Let $\boldsymbol{s}=\boldsymbol{t}=(1,0)$.
@@
 The proof is almost the same as that of Proposition 7.11 (but easier).
-The only difference is that, instead of Corollary 7.8, we use Corollary 7.10, which only requires that $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ be non-negative.
+The only difference is that, instead of Corollary 7.8, we use Corollary 7.10, which only requires that $\widehat{V}$ and $\widehat{V}_{0}$ be non-negative.
 ∎
@@
 Remark 7.13.
-Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ with $\mathfrak{S}$ compact and that $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ are non-negative.
+Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ with $\mathfrak{S}$ compact and that $\widehat{V}$ and $\widehat{V}_{0}$ are non-negative.
 Then, for any $\tau,r\in\mathbb{R}$, $\tau\neq 0$, we have an isomorphism
@@
 Theorem 7.14.
-Let us assume that $M=\mathfrak{S}\times\mathbb{R}$, that $V$ and $V_{0}$ are translation invariant and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
-Then the operator
+Let us assume that $M=\mathfrak{S}\times\mathbb{R}$, that $V$ and $V_{0}$ are translation invariant and that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$. Then the operator
 $\boldsymbol{\Xi}\mathrel{\,:=\,}\boldsymbol{\Xi}_{V,V_{0}}:H^{1}(M;TM)\oplus L^{2}(M)\to H^{-1}(M;TM)\oplus L^{2}(M)$
@@
 We shall say that $V$ and $V_{0}$ satisfy the positivity assumption at infinity on $M$ if
-$\color[rgb]{1,0,1}\widehat{V}\mathrel{\,:=\,}{\mathcal{R}_{\infty}}V\succ 0\quad\mbox{and}\quad\color[rgb]{1,0,1}\widehat{V}_{0}\mathrel{\,:=\,}{\mathcal{R}_{\infty}}V_{0}\succ 0\quad\mbox{on }M\,.$
+$\widehat{V}\mathrel{\,:=\,}{\mathcal{R}_{\infty}}V\succ 0\quad\mbox{and}\quad\widehat{V}_{0}\mathrel{\,:=\,}{\mathcal{R}_{\infty}}V_{0}\succ 0\quad\mbox{on }M\,.$
 Notice that we are not assuming that, globally, $V$ and $V_{0}$ are non-negative.
 This will be done in Assumption 8.3, which is, hence slightly stronger than Assumption 8.1.
-We also recall that, if $a$ is a multiplication operator, then we use also the notation $\color[rgb]{1,0,1}\widehat{a}$ for ${\mathcal{R}_{\infty}}(a)$ (see Equation (5.6) for a justification of this notation).
+We also recall that, if $a$ is a multiplication operator, then we use also the notation $\widehat{a}$ for ${\mathcal{R}_{\infty}}(a)$ (see Equation (5.6) for a justification of this notation).
 We also recall (see Definition 2.13) that the explicit forms of the last two assumptions amount to the following:
   1. (i)
-    $\color[rgb]{1,0,1}\widehat{V}\geq 0$ and, in every connected component of $M^{\prime}$, there is a point $x$ (and hence a whole open set containing it) where $\color[rgb]{1,0,1}\widehat{V}(x)$ is invertible (i.e., $V_{0}(x)>0$ in $M_{n}(\mathbb{C})$).
+    $\widehat{V}\geq 0$ and, in every connected component of $M^{\prime}$, there is a point $x$ (and hence a whole open set containing it) where $\widehat{V}(x)$ is invertible (i.e., $V_{0}(x)>0$ in $M_{n}(\mathbb{C})$).
   2. (ii)
-    $\color[rgb]{1,0,1}\widehat{V}_{0}\geq 0$ and, in every connected component of $M^{\prime}$, there is a point $x$ (and hence a whole open set containing it) where $\color[rgb]{1,0,1}\widehat{V}_{0}(x)$ is invertible (i.e., $V_{0}(x)\in(0,\infty)$).
+    $\widehat{V}_{0}\geq 0$ and, in every connected component of $M^{\prime}$, there is a point $x$ (and hence a whole open set containing it) where $\widehat{V}_{0}(x)$ is invertible (i.e., $V_{0}(x)\in(0,\infty)$).
 Recall that by the statement $\phi\not\equiv 0$ on AA,” we mean that there exists $a$ in the domain of $\phi$ such that $\phi(a)\not=0$.
@@
 Let $\mathcal{N}$ be the kernel of $\boldsymbol{\Xi}:H^{1}(M;TM)\oplus L^{2}(M)\to H^{-1}(M;TM)\oplus L^{2}(M)$.
-The elliptic regularity of $\boldsymbol{\Xi}$ implies that $\mathcal{N}$ will consist of smooth sections.
+The classical elliptic regularity of $\boldsymbol{\Xi}$ implies that $\mathcal{N}$ will consist of smooth sections.
 Assume that $\mathcal{N}\subset L^{2}(M;TM\oplus\mathbb{C})$ and that it is finite dimensional and let $p_{\mathcal{N}}$ be the $L^{2}$–orthogonal projection onto $\mathcal{N}$.
@@
 Theorem 8.2.
-Let $M$ be a manifold with straight cylindrical ends and assume that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ (Assumption 8.1).
-Then the following properties hold.
+Let $M$ be a manifold with straight cylindrical ends and assume that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ (Assumption 8.1). Then the following properties hold.
   1. (i)
@@
 Proof.
-Assumption 8.1 implies that ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})=\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$ satisfies the assumptions of Theorem 7.14, which then yields right away the first point.
+Assumption 8.1 implies that ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})=\boldsymbol{\Xi}_{\widehat{V},\widehat{V}_{0}}$ satisfies the assumptions of Theorem 7.14, which then yields right away the first point.
 The generalized Stokes operator $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}$ is elliptic by Proposition 6.13.
 Since ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ is invertible, Theorem 6.15 implies that $\boldsymbol{\Xi}$ is Fredholm.
-The last point follows from the Beals-type result proved in [67, Theorem 7.4] (this result can also be found in [63]).
+The last point follows from the Beals-type result proved in [87, Theorem 7.4] (this result can also be found in [83]).
 ∎
@@
 Let $U\in H^{1}(M;TM)\oplus L^{2}(M)$ be such that $\boldsymbol{\Xi}U=0$ on $M$.
-Then $U\in H^{2}(M;TM)\oplus H^{1}(M)$, by elliptic regularity.
+Then $U\in H^{2}(M;TM)\oplus H^{1}(M)$, by elliptic regularity (see Theorem 6.16 or [83, Theorem 15.3.30] or [87, Proposition 6.5]).
 Let us notice that the fact that $V$ and $V_{0}$ satisfy the non-vanishing assumption on $M$ (Assumption 8.3) and that $M$ is connected implies that $V_{0}\not\equiv 0$ and $V\succ 0$ on $M$.
@@
 Moreover, the condition that $V$ does not vanish identically on $M$ can be replaced with the condition that $M$ does not have non-vanishing Killing vector fields.
-These results were proved in [69, Theorem 6.1(4)].
+These results were proved in [89, Theorem 6.1(4)].
 In relation to this result, the reader might have noticed that we are not considering the case when $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ has straight cylindrical ends but does not have non-vanishing Killing vector fields.
@@
 The proof is complete. ∎
-For the Stokes operator $\boldsymbol{\Xi}_{0,0}$, see also[39], [62, Lemma 3.1], [70, (6.1), (6.2)], [99, Theorem 3.1], [100], [134, Lemma 1.3].
+For the Stokes operator $\boldsymbol{\Xi}_{0,0}$, see also[49], [82, Lemma 3.1], [90, (6.1), (6.2)], [121, Theorem 3.1], [122], [158, Lemma 1.3].
 Note that $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ and $\frac{1}{2}+\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$.
@@
 Let $\Gamma^{\prime}:=\partial\Omega^{\prime}$.
-In this section, in addition to $V$ and $V\geq 0$ smooth, we assume that they are translation invariant, and hence $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$ identify to sections on $\mathfrak{S}$.
-Also, in this section, we use Assumption 8.1, that is that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$ (that is, they have points where they are strictly positive in every connected component of the space of definition, see Definition 2.13).
+In this section, in addition to $V$ and $V\geq 0$ smooth, we assume that they are translation invariant, and hence $V(x^{\prime},t)=\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\widehat{V}_{0}(x^{\prime})$ identify to sections on $\mathfrak{S}$.
+Also, in this section, we use Assumption 8.1, that is that $\widehat{V}\succ 0$ and that $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$ (that is, they have points where they are strictly positive in every connected component of the space of definition, see Definition 2.13).
 Also, recall that $\Omega$ is on one side of its boundary (Assumption 2.7), and hence $\Omega^{\prime}$ is also on one side of its boundary.
@@
 Definition 9.2.
-Assume that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$ (Assumption 8.1).
+Assume that $\widehat{V}\succ 0$ and that $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$ (Assumption 8.1).
 Let $\boldsymbol{h}\in H^{s}(\Gamma^{\prime};TM)$, $s\in\mathbb{R}$.
@@
 Proposition 9.6.
-Let $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$, $\Gamma^{\prime}:=\partial\Omega^{\prime}$, $\color[rgb]{1,0,1}\widehat{\mathfrak{f}}:=\frac{\color[rgb]{1,0,1}\widehat{V}_{0}+1}{2\color[rgb]{1,0,1}\widehat{V}_{0}+1}$ $\color[rgb]{1,0,1}\widehat{\mathfrak{g}}:=\frac{1}{2\color[rgb]{1,0,1}\widehat{V}_{0}+1}$.
+Let $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$, $\Gamma^{\prime}:=\partial\Omega^{\prime}$, $\widehat{\mathfrak{f}}:=\frac{\widehat{V}_{0}+1}{2\widehat{V}_{0}+1}$ $\widehat{\mathfrak{g}}:=\frac{1}{2\widehat{V}_{0}+1}$.
 Let $\boldsymbol{K}:=\boldsymbol{P}_{0}\in\Psi^{0}(\Gamma^{\prime};TM)$ be as in Equation (8.6).
@@
   1. (i)
-    $\mathcal{V}_{\tau}(\boldsymbol{h}^{\prime})_{\pm}\,=\,\widehat{\mathcal{A}}_{0}(\tau)\boldsymbol{h}^{\prime}$, and $\widehat{\mathcal{A}}_{0}(\tau)$ is elliptic with self-adjoint symbol $\sigma_{-1}(\mathcal{A}_{0}(\tau);\xi^{\prime})\,=\,\frac{1}{4|\xi^{\prime}|}(2-\color[rgb]{1,0,1}\widehat{\mathfrak{f}}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp}-\color[rgb]{1,0,1}\widehat{\mathfrak{f}}\eta^{\sharp}\otimes\eta)\,$, where $\eta:=|\xi^{\prime}|^{-1}\xi^{\prime}$.
+    $\mathcal{V}_{\tau}(\boldsymbol{h}^{\prime})_{\pm}\,=\,\widehat{\mathcal{A}}_{0}(\tau)\boldsymbol{h}^{\prime}$, and $\widehat{\mathcal{A}}_{0}(\tau)$ is elliptic with self-adjoint symbol $\sigma_{-1}(\mathcal{A}_{0}(\tau);\xi^{\prime})\,=\,\frac{1}{4|\xi^{\prime}|}(2-\widehat{\mathfrak{f}}\boldsymbol{\nu}\otimes\boldsymbol{\nu}^{\sharp}-\widehat{\mathfrak{f}}\eta^{\sharp}\otimes\eta)\,$, where $\eta:=|\xi^{\prime}|^{-1}\xi^{\prime}$.
   2. (ii)
-    $\mathcal{P}_{\tau}(\boldsymbol{h}^{\prime})_{\pm}=\big(\mp\frac{\color[rgb]{1,0,1}\widehat{\mathfrak{g}}}{2}\boldsymbol{\nu}^{\sharp}+\widehat{\mathcal{C}}_{0}(\tau)\big)\boldsymbol{h}^{\prime}$.
+    $\mathcal{P}_{\tau}(\boldsymbol{h}^{\prime})_{\pm}=\big(\mp\frac{\widehat{\mathfrak{g}}}{2}\boldsymbol{\nu}^{\sharp}+\widehat{\mathcal{C}}_{0}(\tau)\big)\boldsymbol{h}^{\prime}$.
   3. (iii)
@@
   • $M=\mathfrak{S}\times\mathbb{R}$,
-  • $V$ and $V\geq 0$ are smooth, translation invariant, and hence $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}_{0}(x^{\prime})$ identify to sections on $\mathfrak{S}$.
-  • $V$ and $V_{0}$ satisfy the positivity assumption at infinity (Assumption 8.1, that is, $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$).
+  • $V$ and $V\geq 0$ are smooth, translation invariant, and hence $V(x^{\prime},t)=\widehat{V}(x^{\prime})$ and $V_{0}(x^{\prime},t)=\widehat{V}_{0}(x^{\prime})$ identify to sections on $\mathfrak{S}$.
+  • $V$ and $V_{0}$ satisfy the positivity assumption at infinity (Assumption 8.1, that is, $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$).
 (Recall that $\succ$ was introduced in Definition 2.13.) We also keep the notation and assumptions of equations (9.2), (9.1), and (9.3).
@@
   1. (i)
-    $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\mathfrak{S}$ and
+    $\widehat{V}\succ 0$ on $\mathfrak{S}$ and
   2. (ii)
-    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}\smallsetminus\Gamma^{\prime}$.
+    $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}\smallsetminus\Gamma^{\prime}$.
 Let $\widehat{\boldsymbol{S}\,}(0)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16.
@@
 Our assumptions imply that $\widehat{\boldsymbol{\Xi}}(0)$ is invertible (Proposition 7.11), and hence the indicial single layer potential $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{0}(\boldsymbol{h}^{\prime})$ is defined (see Definition 9.2).
-Since $\widehat{\boldsymbol{S}\,}(0)$ is elliptic and $\Gamma^{\prime}$ is smooth and compact, we have that $\boldsymbol{h}^{\prime}\in H^{s}(\Gamma^{\prime};TM)$ for all $s\in\mathbb{R}$, by elliptic regularity.
+Since $\widehat{\boldsymbol{S}\,}(0)$ is elliptic and $\Gamma^{\prime}$ is smooth and compact, we have that $\boldsymbol{h}^{\prime}\in H^{s}(\Gamma^{\prime};TM)$ for all $s\in\mathbb{R}$, by classical elliptic regularity.
 Proposition 9.5 for $s=1/2$ then gives that the restrictions of $U$ to $\Omega_{+}^{\prime}:=\Omega^{\prime}$ and to $\Omega_{-}^{\prime}:=\mathfrak{S}\smallsetminus\overline{\Omega}{\,}^{\prime}$ satisfy
@@
 The same corollary gives that $p$ is constant on each connected component of $\mathfrak{S}\smallsetminus\Gamma^{\prime}$.
-Because $\color[rgb]{1,0,1}\widehat{V}_{0}$ is not identically equal to zero on any connected component of $\mathfrak{S}\smallsetminus\Gamma^{\prime}$, we obtain that, in fact, $p=0$.
+Because $\widehat{V}_{0}$ is not identically equal to zero on any connected component of $\mathfrak{S}\smallsetminus\Gamma^{\prime}$, we obtain that, in fact, $p=0$.
 We have thus obtained that $U=0$ and hence $\widehat{\boldsymbol{S}\,}(0)$ is injective.
 Finally, the fact that $\widehat{\boldsymbol{S}\,}(0)^{-1}\in\Psi^{1}(\Gamma^{\prime};TM)$ follows from Beals’ theorem (see, for instance, Theorem 7.4 in [87]). ∎
-The same result holds for $\tau\neq 0$ under the weaker hypothesis that $V\succ 0$ and $V_{0}\succ 0$ on $M$ (equivalently, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$).
+The same result holds for $\tau\neq 0$ under the weaker hypothesis that $V\succ 0$ and $V_{0}\succ 0$ on $M$ (equivalently, that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$).
 Proposition 9.8.
-Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder and that $V$ and $V_{0}$ are smooth, translation invariant, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
+Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder and that $V$ and $V_{0}$ are smooth, translation invariant, and $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
 We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$.
@@
 Proof.
-The proof is exactly the same as that of Proposition 9.7, except that we are using Corollary 7.9 (which only requires $\color[rgb]{1,0,1}\widehat{V}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}$ to be non-negative) instead of Corollary 7.7.
+The proof is exactly the same as that of Proposition 9.7, except that we are using Corollary 7.9 (which only requires $\widehat{V}$ and $\widehat{V}_{0}$ to be non-negative) instead of Corollary 7.7.
 ∎
@@
 Theorem 9.9.
-Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\mathfrak{S}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}\smallsetminus\Gamma^{\prime}$.
+Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and $\widehat{V}\succ 0$ on $\mathfrak{S}$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}\smallsetminus\Gamma^{\prime}$.
 We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$.
@@
 In this subsection, we keep the assumptions of the previous two subsections.
-In particular, we assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$.
+In particular, we assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$.
 In the following, we will need the operators $\boldsymbol{P}:=-2\mathcal{A}\boldsymbol{D}_{\boldsymbol{\nu}}^{*}+\mathcal{B}\boldsymbol{\nu}^{\sharp}$ and $\boldsymbol{K}\mathrel{\,:=\,}\boldsymbol{P}_{0}$.
@@
 Proposition 9.10.
-Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega_{+}^{\prime}:=\Omega^{\prime}$.
+Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and that $\widehat{V}_{0}\succ 0$ on $\Omega_{+}^{\prime}:=\Omega^{\prime}$.
 We furthermore assume that $\Omega^{\prime}$ is on one side of its boundary, as usual (i.e., $\Omega^{\prime}$ is the interior of $\overline{\Omega}^{\prime}$, see Assumption 2.7).
@@
 Recall that by the statement $\phi\not\equiv 0$ on AA,” we mean that there exists $a$ in the domain of $\phi$ such that $\phi(a)\not=0$.
-To negate this statement, we shall write $\phi=0$ in AA.” Of course, in view of our conventions, the condition $\color[rgb]{1,0,1}\widehat{V}_{0}\not\equiv 0$ on $\Omega^{\prime}$ implies the condition $V_{0}\not\equiv 0$ on $\Omega$ (since $V$ and $V_{0}$ are translation invariant at infinity).
+To negate this statement, we shall write $\phi=0$ in AA.” Of course, in view of our conventions, the condition $\widehat{V}_{0}\not\equiv 0$ on $\Omega^{\prime}$ implies the condition $V_{0}\not\equiv 0$ on $\Omega$ (since $V$ and $V_{0}$ are translation invariant at infinity).
 Proof.
@@
 Let then $\boldsymbol{h}^{\prime}\in L^{2}(\Gamma^{\prime};TM)$ be such that $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)^{*}\big]\boldsymbol{h}^{\prime}=0$.
-Since $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)^{*}$ is elliptic and $\Gamma^{\prime}$ is smooth and compact, we have that $\boldsymbol{h}^{\prime}\in H^{s}(\Gamma^{\prime};TM)$ for all $s\in\mathbb{R}$, by elliptic regularity.
+Since $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)^{*}$ is elliptic and $\Gamma^{\prime}$ is smooth and compact, we have that $\boldsymbol{h}^{\prime}\in H^{s}(\Gamma^{\prime};TM)$ for all $s\in\mathbb{R}$, by classical elliptic regularity.
 Let us consider the indicial single layer potential $U:=\mathcal{S}_{0}\boldsymbol{h}^{\prime}$ (see Definition 9.2 with $\tau=0$).
@@
 Let next $\Omega_{0}$ be a connected component of $\Omega_{-}^{\prime}$.
-Our assumptions give that $\color[rgb]{1,0,1}\widehat{V}\not\equiv 0$ on $\Omega_{0}$.
+Our assumptions give that $\widehat{V}\not\equiv 0$ on $\Omega_{0}$.
 Hence we obtain that $\boldsymbol{u}=0$ on $\Omega_{0}$, by Corollary 7.7.
@@
 Corollary 7.7 then gives that $\boldsymbol{u}=0$ on $\Omega^{\prime}$.
-Because $\color[rgb]{1,0,1}\widehat{V}_{0}\not\equiv 0$ on every connected component of $\Omega^{\prime}$, the same corollary gives that $p=0$ on $\Omega^{\prime}$.
+Because $\widehat{V}_{0}\not\equiv 0$ on every connected component of $\Omega^{\prime}$, the same corollary gives that $p=0$ on $\Omega^{\prime}$.
 We have thus obtained that $U=0$ on both $\Omega_{\pm}^{\prime}$, and hence $U=0$ on $\mathfrak{S}$.
@@
 Therefore $\widehat{\boldsymbol{S}}(0)$ is injective, and hence it is invertible.
-The fact that $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$ is also a classical result on pseudodifferential operators [16] (a proof can also be found in [63], Theorem 15.4.11, and in [67], Theorem 7.4).
+The fact that $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$ is also a classical result on pseudodifferential operators [20] (see also Theorem 6.17; a proof can also be found in [83], Theorem 15.4.11, and in [87], Theorem 7.4).
 This completes the proof.
@@
 Proposition 9.11.
-Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder $V$ and $V_{0}$ are smooth, translation invariant, and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
+Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is a cylinder $V$ and $V_{0}$ are smooth, translation invariant, and that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $\mathfrak{S}$.
 We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$.
 Let $\widehat{\boldsymbol{K}}(\tau)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ of Theorem 8.14 (see also Equation (6.1)).
-If $\tau\neq 0$, then $\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
+If $\tau\neq 0$, then $\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$, see Theorem 6.17.
 Proof.
@@
 Theorem 9.12.
-Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega_{+}^{\prime}:=\Omega^{\prime}$.
+Let us assume that $M=\mathfrak{S}\times\mathbb{R}$ is cylinder, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $\widehat{V}_{0}\succ 0$ on $\Omega_{+}^{\prime}:=\Omega^{\prime}$.
 We furthermore assume that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$, as usual.
@@
 Proposition 10.1.
-Let us assume that $M$ is a manifold with straight cylindrical ends and that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$.
+Let us assume that $M$ is a manifold with straight cylindrical ends and that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\widehat{V}\succ 0$ on $M^{\prime}$ and $\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$.
 Let $\widehat{\boldsymbol{S}\,}(0)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16.
@@
 ∎
-The same result holds for $\tau\neq 0$ under the weaker hypothesis that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$.
+The same result holds for $\tau\neq 0$ under the weaker hypothesis that $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$.
 Proposition 10.2.
-Let us assume that $M$ is a manifold with straight cylindrical ends and that $V$ and $V_{0}$ are smooth, asymptotically translation invariant at infinity, and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}$.
+Let us assume that $M$ is a manifold with straight cylindrical ends and that $V$ and $V_{0}$ are smooth, asymptotically translation invariant at infinity, and $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $M^{\prime}$.
 Let $\widehat{\boldsymbol{S}\,}(\tau)\in\Psi^{-1}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi^{-1}(\Gamma;TM)$ of Theorem 8.16.
@@
 The proof is the same as that of Proposition 10.1, but using Proposition 9.8 instead of Proposition 9.7.
-(As in that proof, the assumptions $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}$ are needed for $\boldsymbol{\Xi}$ to be Fredholm, so that the layer potentials are defined.) ∎
+(As in that proof, the assumptions $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $M^{\prime}$ are needed for $\boldsymbol{\Xi}$ to be Fredholm, so that the layer potentials are defined.) ∎
 The above propositions lead to the following useful corollary.
@@
 Let us assume that $M=M_{0}\cup\big[M^{\prime}\times(-\infty,0)\big]$ is a manifold with straight cylindrical ends, that $V$ and $V_{0}$ are smooth, translation invariant at infinity, and that $\Omega^{\prime}$ is the interior of $\overline{\Omega}^{\prime}$.
-We also assume that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$.
+We also assume that $\widehat{V}\succ 0$ on $M^{\prime}$ and $\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$.
 Then $\boldsymbol{S}$ is an unbounded Fredholm operator on $L^{2}(\Gamma;TM)$.
@@
 Proof.
-The operator ${\mathcal{R}_{\infty}}(\boldsymbol{S})$ is the single layer potential operator associated to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi}_{V,V_{0}})=\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$, by Theorem 6.10.
+The operator ${\mathcal{R}_{\infty}}(\boldsymbol{S})$ is the single layer potential operator associated to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi}_{V,V_{0}})=\boldsymbol{\Xi}_{\widehat{V},\widehat{V}_{0}}$, by Theorem 6.10.
 The assumptions of the corollary give that the assumptions of Theorem 9.9 are therefore satisfied by ${\mathcal{R}_{\infty}}(\boldsymbol{S})$, which is hence invertible.
@@
 Theorem 10.4.
-Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions 8.3 (hence, in particular, $M$ is with straight cylindrical ends and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$).
+Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions 8.3 (hence, in particular, $M$ is with straight cylindrical ends and $\widehat{V}\succ 0$ on $M^{\prime}$).
 Let $\boldsymbol{S}:=\mathcal{A}_{0}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ be as in Theorem 8.16.
@@
   1. (i)
-    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$.
+    $\widehat{V}_{0}\succ 0$ on $M^{\prime}\smallsetminus\Gamma^{\prime}$.
   2. (ii)
@@
 To this end, let $\boldsymbol{h}\in H^{-1/2}(\Gamma;TM)$ be such that $\boldsymbol{S}\boldsymbol{h}=0$ and let $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$ be the corresponding single layer potential.
-The ellipticity of $\boldsymbol{S}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$ and the elliptic regularity property of Theorem 8.16 assure that $\boldsymbol{h}\in H^{s}(\Gamma;TM)$ for all $s\in\mathbb{R}$.
+The ellipticity of $\boldsymbol{S}\in\Psi_{\operatorname{ess}}^{-1}(\Gamma;TM)$, the ellipticity property of Theorem 8.16, and elliptic regularity (see Theorem 6.16 or [83, Theorem 15.3.30] or [87, Proposition 6.5]) assure that $\boldsymbol{h}\in H^{s}(\Gamma;TM)$ for all $s\in\mathbb{R}$.
 Then, by Proposition 8.9, the restrictions of $U$ to $\Omega_{+}:=\Omega$ and to $\Omega_{-}$ satisfy the property
@@
 Consequently, $\boldsymbol{S}$ is injective and, in view of the above arguments, it is invertible.
-Finally, the property that $\boldsymbol{S}^{-1}\in\Psi_{\operatorname{ess}}^{1}(\Gamma;TM)$ follows from Theorem 7.4 in [67] (see also Theorem 15.4.11 in [63]).
+Finally, the property that $\boldsymbol{S}^{-1}\in\Psi_{\operatorname{ess}}^{1}(\Gamma;TM)$ follows from Theorem 6.17 (see Theorem 7.4 in [87] or Theorem 15.4.11 in [83]).
 ∎
@@
 This case needs additional assumptions and arguments.
-Recall that we have $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime},t)$ for $(x^{\prime},t)\in M^{\prime}\times(-\infty,0)$ and $t\ll 0$ and that $\color[rgb]{1,0,1}\widehat{V}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ is independent of $t$.
-The function $\color[rgb]{1,0,1}\widehat{V}_{0}$ on $M^{\prime}$ is defined similarly.
-The functions $\color[rgb]{1,0,1}\widehat{V}_{0}$ and $\color[rgb]{1,0,1}\widehat{V}$ correspond to the limit operators of the operators defined by $V_{0}$ and $V$.
+Recall that we have $V(x^{\prime},t)=\widehat{V}(x^{\prime},t)$ for $(x^{\prime},t)\in M^{\prime}\times(-\infty,0)$ and $t\ll 0$ and that $\widehat{V}(x^{\prime},t)=\widehat{V}(x^{\prime})$ is independent of $t$.
+The function $\widehat{V}_{0}$ on $M^{\prime}$ is defined similarly.
+The functions $\widehat{V}_{0}$ and $\widehat{V}$ correspond to the limit operators of the operators defined by $V_{0}$ and $V$.
 Proposition 10.5.
 Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions (Assumption 8.3) and let $\widehat{\boldsymbol{K}}(0)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ given in Theorem 8.14 (see also Equation (10.2)).
-We furthermore assume that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and that $\Omega^{\prime}$ on one side of its boundary $\Gamma^{\prime}:=\overline{\Omega^{\prime}}$.
+We furthermore assume that $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and that $\Omega^{\prime}$ on one side of its boundary $\Gamma^{\prime}:=\overline{\Omega^{\prime}}$.
 Then $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
-Of course, in view of our conventions, the condition $\color[rgb]{1,0,1}\widehat{V}_{0}\not\equiv 0$ on $\Omega^{\prime}$ implies the condition $V_{0}\not\equiv 0$ on $\Omega$ (since $V$ and $V_{0}$ are translation invariant at infinity).
+Of course, in view of our conventions, the condition $\widehat{V}_{0}\not\equiv 0$ on $\Omega^{\prime}$ implies the condition $V_{0}\not\equiv 0$ on $\Omega$ (since $V$ and $V_{0}$ are translation invariant at infinity).
 Proof.
@@
 The result then follows from Proposition 9.10 applied to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ acting on $M^{\prime}\times\mathbb{R}$ and the fact that layer potentials commute with limit operators, Theorem 6.10.
+The last point follows from Theorem 6.17.
 ∎
 Proposition 10.6.
-Let us assume that $V,V_{0}\geq 0$ satisfy Assumption 8.3 (so $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}$)and let $\widehat{\boldsymbol{K}}(\tau)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ of Theorem 8.14 (see also Equation (10.2)).
+Let us assume that $V,V_{0}\geq 0$ satisfy Assumption 8.3 (so $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $M^{\prime}$)and let $\widehat{\boldsymbol{K}}(\tau)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ of Theorem 8.14 (see also Equation (10.2)).
 If $\tau\neq 0$, then $\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
@@
 Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions 8.3 and let $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ be as in Theorem 8.14 (as before, see also Equation (10.2)).
-We furthermore assume that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$.
-Then $\frac{1}{2}+\boldsymbol{K}$ is a Fredholm operator on $L^{2}(\Gamma;TM)$.
+We furthermore assume that $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$.
+Then $\frac{1}{2}+\boldsymbol{K}$ is a Fredholm operator on $L^{2}(\Gamma;TM)$ and its pseudo-inverse $(\frac{1}{2}+\boldsymbol{K})^{(-1)}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$.
 Proof.
 The proof is similar to that of Corollary 10.3, but using Theorem 9.12 instead of Theorem 9.9.
-Indeed, the operator ${\mathcal{R}_{\infty}}(\frac{1}{2}+\boldsymbol{K})$ is the double layer potential operator associated to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi}_{V,V_{0}})=\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$, by Theorem 6.10.
+Indeed, the operator ${\mathcal{R}_{\infty}}(\frac{1}{2}+\boldsymbol{K})$ is the double layer potential operator associated to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi}_{V,V_{0}})=\boldsymbol{\Xi}_{\widehat{V},\widehat{V}_{0}}$, by Theorem 6.10.
 The assumptions of the corollary give that the assumptions of Theorem 9.12 are therefore satisfied by ${\mathcal{R}_{\infty}}(\frac{1}{2}+\boldsymbol{K})$, which is hence invertible.
 Because $\frac{1}{2}+\boldsymbol{K}$ is elliptic, Theorem 6.15 gives that $\frac{1}{2}+\boldsymbol{K}$ is Fredholm.
+The last point follows from Theorem 6.17.
 ∎
@@
 Theorem 10.8.
-Let us assume that $M,V$, and $V_{0}$ satisfy Assumption 8.3 (so, in particular $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$), that $\Omega$ is connected, and that $\Omega$ is the interior of $\overline{\Omega}$.
+Let us assume that $M,V$, and $V_{0}$ satisfy Assumption 8.3 (so, in particular $\widehat{V}\succ 0$ on $M^{\prime}$), that $\Omega$ is connected, and that $\Omega$ is the interior of $\overline{\Omega}$.
 We furthermore assume the following:
   1. (i)
-    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$,
+    $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$,
   2. (ii)
-    $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and
+    $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and
   3. (iii)
@@
 To this end, let $\boldsymbol{h}\in L^{2}(\Gamma;TM)$ be such that $(\frac{1}{2}+\boldsymbol{K}^{*})\boldsymbol{h}=0$.
-The ellipticity of $\frac{1}{2}+\boldsymbol{K}^{*}$, the fact that $\Gamma$ has straight cylindrical ends, and the elliptic regularity property ([67, Proposition 6.5]) imply that $\boldsymbol{h}\in H^{s}(\Gamma;TM)$ for all $s\in\mathbb{R}$.
+The ellipticity of $\frac{1}{2}+\boldsymbol{K}^{*}$, the fact that $\Gamma$ has straight cylindrical ends, and the elliptic regularity property (see Theorem 6.16 or [83, Theorem 15.3.30] or [87, Proposition 6.5]) imply that $\boldsymbol{h}\in H^{s}(\Gamma;TM)$ for all $s\in\mathbb{R}$.
 Let now the single layer potential $\mathcal{S}_{\rm{ST}}\boldsymbol{h}$.
@@
 The fact that $\boldsymbol{u}_{+}=0$ at $\Gamma$ allows us to use Corollary 2.14 that gives then that $\boldsymbol{u}=0$ in $\Omega$.
-Because $\Omega$ is connected and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, we also have $V_{0}\succ 0$ on $\Omega$.
+Because $\Omega$ is connected and $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, we also have $V_{0}\succ 0$ on $\Omega$.
 Therefore, the same corollary implies that $p$ is constant on $\Omega$.
@@
 Therefore the operator $\frac{1}{2}+\boldsymbol{K}^{*}:L^{2}(\Gamma;TM)\to L^{2}(\Gamma;TM)$ is injective implying that the Fredholm operator of index zero $\frac{1}{2}+\boldsymbol{K}:L^{2}(\Gamma;TM)\to L^{2}(\Gamma;TM)$ is onto and hence an isomorphism, as asserted.
-The fact that $(\frac{1}{2}+\boldsymbol{K})^{-1}\in\Psi^{0}(\Gamma;TM)$ follows from Theorem 5.4.11 in [63] (see also Theorem 7.4 in [67]). This completes the proof. ∎
+The fact that $(\frac{1}{2}+\boldsymbol{K})^{-1}\in\Psi^{0}(\Gamma;TM)$ follows from Theorem 6.17 (see Theorem 7.4 in [87] or Theorem 15.4.11 in [83]). This completes the proof. ∎
 In all the results above, the assumptions that $M$ and $\Omega$ are connected do not really decrease the generality.
@@
 11.1. The homogeneous Stokes system
-The invertibility of each of the operators $\frac{1}{2}+\boldsymbol{K}$ and $\boldsymbol{S}$ gives the following well-posedness result for the modified Stokes system with Dirichlet boundary condition, problem (1.4).
+The invertibility of each of the operators $\frac{1}{2}+\boldsymbol{K}$ and $\boldsymbol{S}$ gives the following well-posedness result for the modified Stokes system with Dirichlet boundary condition, problem (1.5) (see Equation (1.1) for a more explicit form of this problem).
 This result is known for the usual Stokes operator (i.e.
-when $V=0$ and $V_{0}=0$, see [100, Proposition 10.5.1, Theorem 10.6.2] in the case of a bounded Lipschitz domain in $\mathbb{R}^{n}$, $n\geq 2$, and [39, Theorem 5.1] in the case of a $C^{1}$ domain on a compact manifold, see also [63] for further related applications).
+when $V=0$ and $V_{0}=0$, see [122, Proposition 10.5.1, Theorem 10.6.2] in the case of a bounded Lipschitz domain in $\mathbb{R}^{n}$, $n\geq 2$, and [49, Theorem 5.1] in the case of a $C^{1}$ domain on a compact manifold, see also [83] for further related applications).
 We consider the homogeneous Dirichlet problem
@@
 Let $M$, $V$, and $V_{0}$ satisfy Assumption 8.3 (so, in particular, $M$ is has straight cylindrical ends) and let $\Omega\subset M$ be a compatible smooth domain (so $\Omega$ also has straight cylindrical ends, see Equation (1.3)).
-We assume that $V_{0}\succ 0$ on $\Omega$, that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, and that $\Omega$ is on one side of its boundary $\Gamma:=\partial{\Omega}$.
+We assume that $V_{0}\succ 0$ on $\Omega$, that $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, and that $\Omega$ is on one side of its boundary $\Gamma:=\partial{\Omega}$.
 We further assume one of the following conditions:
   1. (a)
-    $V\succ 0$ on $\Omega_{-}$ and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ or
+    $V\succ 0$ on $\Omega_{-}$ and $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ or
   2. (b)
-    $V_{0}\succ 0$ on $\Omega_{-}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega_{-}^{\prime}$.
+    $V_{0}\succ 0$ on $\Omega_{-}$ and $\widehat{V}_{0}\succ 0$ on $\Omega_{-}^{\prime}$.
 Then, for every $m\in\mathbb{Z}_{+}$ and any $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$, the homogeneous Diriechlet problem (11.1) has a unique solution $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$.
@@
 Because $V_{0}\not\equiv 0$ on $\Omega$ and $\boldsymbol{u}_{0}=0$ on $\partial\Omega$, Corollary 2.14 then ensures that $U_{0}=0$ in $\Omega$, which implies the desired uniqueness.
-Our assumption (a) implies that the operator $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi^{0}(\Gamma;TM)$, by Theorem 10.8.
-Moreover, the elliptic regularity of the operator $\frac{1}{2}+\boldsymbol{K}$ (Theorem 8.14) implies that $\boldsymbol{h}:=(\frac{1}{2}+\boldsymbol{K})^{-1}\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$.
+Our assumption (a) implies that the operator $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$, by Theorem 10.8.
+The mapping properties of the $\operatorname{ess}$-calculus” then gives that $\boldsymbol{h}:=(\frac{1}{2}+\boldsymbol{K})^{-1}\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$.
 Then $U:=\mathcal{D}_{\rm{ST}}(\boldsymbol{h})$ satisfies the required properties, including its boundedness with respect to the given datum $\boldsymbol{f}$, with a constant $C_{m}$ given by $\|\mathcal{D}_{\rm{ST}}(\frac{1}{2}+\boldsymbol{K})^{-1}\|$.
-If, on the other hand, condition (b) is satisfied, then the operator $\boldsymbol{S}$ is invertible on $L^{2}(\Gamma;TM)$ and $\boldsymbol{S}^{-1}\in\Psi^{1}(\Gamma;TM)$, by Theorem 10.4.
-Moreover, the elliptic regularity of the operator $\boldsymbol{S}:=\mathcal{A}_{0}$ (Theorem 8.16) implies that $\boldsymbol{h}:=\boldsymbol{S}^{-1}\boldsymbol{f}\in H^{m-1/2}(\Gamma;TM)$.
+If, on the other hand, condition (b) is satisfied, then the operator $\boldsymbol{S}$ is invertible on $L^{2}(\Gamma;TM)$ and $\boldsymbol{S}^{-1}\in\Psi_{\operatorname{ess}}^{1}(\Gamma;TM)$, by Theorem 10.4.
+The mapping properties of the $\operatorname{ess}$-calculus” then imply that $\boldsymbol{h}:=\boldsymbol{S}^{-1}\boldsymbol{f}\in H^{m-1/2}(\Gamma;TM)$.
 Theorem 6.6 then gives
@@
 Next we prove the existence part of Theorem 1.1.
-Given the potential $V$ satisfying the assumptions of this theorem, we can modify it on $\Omega_{-}$ such that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $V\succ 0$ on $\Omega_{-}$.
+Given the potential $V$ satisfying the assumptions of this theorem, we can modify it on $\Omega_{-}$ such that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $V\succ 0$ on $\Omega_{-}$.
 Then our assumptions imply that the operator $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi^{0}(\Gamma;TM)$, by Theorem 10.8.
@@
 We notice that the solution $U$ of the above theorem can be represented either by a single or as a double layer potential, depending on how we modify $V$ and $V_{0}$ on $\Omega_{-}$.
-Indeed, modifying $V$ on $\Omega_{-}$ such that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $V\succ 0$ on $\Omega_{-}$, then the assumptions of Theorem 10.8 are satisfied and the solution $U$ can be represented as a double layer potential
+Indeed, modifying $V$ on $\Omega_{-}$ such that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $V\succ 0$ on $\Omega_{-}$, then the assumptions of Theorem 10.8 are satisfied and the solution $U$ can be represented as a double layer potential
 $U\mathrel{\,:=\,}\mathcal{D}_{\rm{ST}}\Big(\Big(\frac{1}{2}+\boldsymbol{K}\Big)^{-1}\boldsymbol{f}\Big)\,.$
@@
 Indeed, it depends only on the norms of the operators $\mathcal{D}_{\rm{ST}}$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}$ and can be chosen the product of them.
+See also [3] and [4] for other results related to the Stokes operator on unbounded domains.
 11.2. The non-homogeneous Stokes system
@@
 ∎
-A useful consequence of Theorem 11.4 is the following corollary.
+Let us now consider the case $r=0$ in Equation (11.2). Then a consequence of Theorem 11.4 that will be used in what follows is the following corollary.
 Corollary 11.5.
-Let the assumptions of Theorem 1.1 hold.
+Let us further assume in Theorem 11.4 that $r=0$.
 Then, for every $m\in\mathbb{Z}_{+}$, there exists a constant ${\mathcal{C}}_{m}\in(0,\infty)$ such that, for any $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$ and any $\boldsymbol{h}\in H^{m-1}(\Omega;TM)$, the non-homogeneous Dirichlet problem (11.2) has a unique solution $U=(\boldsymbol{u}\ \ p)^{\top}\in H^{m+1}(\Omega;TM)\oplus H^{m}(\Omega)$ and this solution satisfies
@@
 Let assumptions of Theorem 1.1 hold.
-Let $\mathcal{N}_{\rm{ST}}:H^{3/2}(\partial\Omega;TM)\to H^{1/2}(\partial\Omega;TM)$ be the Dirichlet-to-Neumann operator defined as follows (see [131, p.37] in the case of the Laplace operator on a compact manifold).
+Let $\mathcal{N}_{\rm{ST}}:H^{3/2}(\partial\Omega;TM)\to H^{1/2}(\partial\Omega;TM)$ be the Dirichlet-to-Neumann operator defined as follows (see [155, p.37] in the case of the Laplace operator on a compact manifold).
 For $\boldsymbol{f}\in H^{3/2}(\partial\Omega;TM)$ arbitrary, let $U:=(\boldsymbol{u}\ \ p)^{\top}\in H^{2}(\Omega;TM)\oplus H^{1}(\Omega)$ be the unique solution of the Dirichlet problem (11.1) (that is, $\boldsymbol{\Xi}_{V,V_{0}}U\,=\,0$ in $\Omega$ and $\boldsymbol{u}|_{\Gamma}\,=\,\boldsymbol{f}$ on $\Gamma:=\partial\Omega$), see Theorem 1.1.
@@
 the limit being evaluated from $\Omega$.
-The next result shows that there is no jump of the conormal derivative $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})$ across $\Gamma:=\partial\Omega$ (see [131, Proposition 11.4] in the case of the Laplace operator).
+The next result shows that there is no jump of the conormal derivative $\boldsymbol{T}_{\boldsymbol{\nu}}\mathcal{D}_{\rm{ST}}(\boldsymbol{f})$ across $\Gamma:=\partial\Omega$ (see [155, Proposition 11.4] in the case of the Laplace operator).
 Theorem 11.6.
@@
 An example is the following result, which is well known in $\mathbb{R}^{n}$ (see e.g.
-[17, Theorem 7.3, Theorem 8.2], [114, §4.5.2, Theorem], and [90, Lemma 8.1]).
+[21, Theorem 7.3, Theorem 8.2], [137, §4.5.2, Theorem], and [111, Lemma 8.1]).
 Lemma 12.1.
@@
 Also, notice that the result of Lemma 12.1 extends to Sobolev spaces on sections of vector bundles over $M$ (see e.g.
-[114, §4.2.2, Proposition] in the Euclidean setting), and the next lemma is a consequence of this result (see [39, Lemma 7.5] in the case of a Lipschitz domain on a compact manifold).
+[137, §4.2.2, Proposition] in the Euclidean setting), and the next lemma is a consequence of this result (see [49, Lemma 7.5] in the case of a Lipschitz domain on a compact manifold).
 Lemma 12.2.
@@
 $\|uv\|_{H^{m-1}(\Omega)}\leq c_{m}\|u\|_{H^{m+1}(\Omega)}\|v\|_{H^{m}(\Omega)}\,.$
-(see [114, §4.2.2, Proposition] in $\mathbb{R}^{n}$).
+(see [137, §4.2.2, Proposition] in $\mathbb{R}^{n}$).
 This result extends also to vector bundles and, thus, there exists a constant $C>0$ (depending only on $n$, $m$ and the bounded geometry of $M$) such that the estimate (12.1) holds for all $\boldsymbol{u},\boldsymbol{v}\in H^{m+1}(\Omega;TM)$.
@@
 for all $\boldsymbol{u},\boldsymbol{v}\in H^{m+1}(\Omega;TM)$, that is, the estimate (12.2). ∎
-Let us now prove the well-posedness of the Dirichlet problem for the Navier-Stokes system under the assumption of small data (see Theorem 7.3 in [39] in the case of a $C^{1}$, or even Lipschitz domain on a compact Riemannian manifold).
+Let us now prove the well-posedness of the Dirichlet problem for the Navier-Stokes system under the assumption of small data (see also [17, 49, 83, 111] for results in bounded Lipschitz domains).
 Theorem 12.3.
@@
     There are no restrictions “or conditions” at infinity on our operators, functions, and vector bundles.
-For instance, $\mathcal{C}^{\infty}_{\operatorname{inv}}$ is replaced with $\mathcal{C}^{\infty}$, all open subsets and all vector bundles are compatible, and $\Psi_{\operatorname{ess}}^{m}$ is replaced with $\Psi^{m}$.
+For instance, $\mathcal{C}^{\infty}_{\operatorname{inv}}$ is replaced with $\mathcal{C}^{\infty}$, all open subsets and all vector bundles are compatible, and $\Psi_{\operatorname{inv}}^{m}$ and $\Psi_{\operatorname{ess}}^{m}$ are replaced with $\Psi^{m}$, $m\in\mathbb{Z}\cup\{\pm\infty\}$.
   2. (2)
-    The restuls about the limit at infinity $\mathcal{R}_{\infty}$} do not make sense anymore (and are not needed). They have to be ignored in the closed case.
+    The results about the limit at infinity $\mathcal{R}_{\infty}$} do not make sense anymore (and are not needed). They have to be ignored in the closed case.
   3. (3)
     The same is true for the conditions involving $\mathcal{R}_{\infty}$}; they are to be removed, as they are no longer needed.
-The case $M$ closed was considered in great detail in [69], so we will not go through the exercice of performing all the modifications abour.
-Instead, we recall here only two results from that paper.
-The first result describes the kernel of $\boldsymbol{\Xi}$ and does not follow from the results of this paper (directly):
+The case $M$ closed was considered in great detail in [89], so we will not go through the exercice of performing all the modifications needed here.
+Instead, we recall here only two results from that paper that do not follow directly from the results of this paper.
+The first result describes the kernel of $\boldsymbol{\Xi}$.
+The second result formulates the well-posendess of the Stokes operator.
 Theorem 12.4.
-Let us assume that $M$ is a closed manifold, that $V,V_{0}$ are smooth and non-negative and that $M$ is a smooth, compact manifold without boundary (i.e., a closed manifold).
+Let us assume that $V,V_{0}$ are smooth and non-negative and that $M$ is a smooth, compact manifold without boundary (i.e., a closed manifold).
 Then $\boldsymbol{\Xi}:=\boldsymbol{\Xi}_{V,V_{0}}:H^{1}(M;TM)\oplus L^{2}(M)\to H^{-1}(M;TM)\oplus L^{2}(M)$ is a self-adjoint Fredholm operator.
@@
     $\mathcal{N}:=\{0\}$, if $V_{0}\not\equiv 0$ on $M$ and either $V\not\equiv 0$ on $M$ or $M$ does not have non-zero Killing vector fields.
-Then the kernel of $\boldsymbol{\Xi}_{V,V_{0}}$ is given by $\ker\boldsymbol{\Xi}_{V,V_{0}}=\mathcal{N}$.
+Then $\mathcal{N}$ is the kernel of $\boldsymbol{\Xi}_{V,V_{0}}$ (i.e., $\ker\boldsymbol{\Xi}_{V,V_{0}}=\mathcal{N}$).
 Moreover, $\boldsymbol{\Xi}$ has a (unique) Moore-Penrose pseudoinverse $\boldsymbol{\Xi}^{(-1)}\in\Psi_{\operatorname{cl}}^{-\boldsymbol{s}-\boldsymbol{t}}(M;TM\oplus\mathbb{C})$, with $\boldsymbol{s}=\boldsymbol{t}=(1,0)$, whose image is $\mathcal{N}^{\perp}$.
@@
 Any two solutions $U_{1}$ and $U_{2}$ of this Dirichlet problem differ by a constant scalar field: $U_{2}-U_{1}=(0\ \ \ c)^{\top}$, $c\in\mathbb{C}$.
-Also, any of the following two formulas provides a solution of this Dirichlet problem:
-$U_{1}\mathrel{\,:=\,}\mathcal{D}_{\rm{ST}}\Big(\Big(\frac{1}{2}+\boldsymbol{K}\Big)^{(-1)}\boldsymbol{f}\Big)\quad\mbox{or}\quad\mbox{U}_{2}\mathrel{\,:=\,}\mathcal{S}_{\rm{ST}}\left(\boldsymbol{S}^{(-1)}\boldsymbol{f}\right)\,.$
 Moreover, there exists a constant $C_{m}\geq 0$, independent of $\boldsymbol{h}$, such that all solutions $U=(\boldsymbol{u}\ \ p)^{\top}$ satisfy
 $\|\boldsymbol{u}\|_{H^{m+1}(\Omega;T\overline{\Omega})}+\Big\|p-\frac{1}{\operatorname{vol}(\Omega)}\int_{\Omega}p\,\,d\operatorname{vol}\Big\|_{H^{m}(\Omega)}\leq C_{m}\|\boldsymbol{f}\|_{H^{m+1/2}(\Gamma;T\overline{\Omega})}\,.$
+Moreover, using layer potentials, we obtain explicit formulas for the solution.
+Indeed, any of the following two formulas provides a solution of the Dirichlet problem for $\boldsymbol{\Xi}_{V,0}$ in the above theorem:
+$U_{1}\mathrel{\,:=\,}\mathcal{D}_{\rm{ST}}\Big(\Big(\frac{1}{2}+\boldsymbol{K}\Big)^{(-1)}\boldsymbol{f}\Big)\quad\mbox{or}\quad\mbox{U}_{2}\mathrel{\,:=\,}\mathcal{S}_{\rm{ST}}\left(\boldsymbol{S}^{(-1)}\boldsymbol{f}\right)\,.$
+If, however, $V_{0}\not\equiv 0$ on $\Omega$ in the above theorem, we obtain the uniqueness of the solution $U$, as in this paper.